Home
        2007 as pdf
         Contents
1.                                                                 File  http   ttfinfo desy de CALICEelog data 2007 16 18 04 2007 04 18T11 11 02 00 gif    2007 04 18 18 19 noise development    CALICE eLogbook   Results for search of  20070401 Page 9 of 14                                                       File  http   ttfinfo desy de CALICEelog data 2007 16 18 04 2007 04 18T18 19 02 00 gif    2007 04 18 18 19 current development    CALICE eLogbook   Results for search of  20070401 Page 10 of 14      SEG EHQ 8605 Detail N                Modue 28A 82 4V    Modue 288 77 7V    EEEEREER  BAS                      SCH  File  http   ttfinfo desy de CALICEelog data 2007 16 18 04 2007 04 18T18 19 05 00 gif  2007 04 19 18 22 Sebastian CRC SERO001 FE5A is broken  Schaetzel    EDIT  Benjamin found on Tuesday  24 4 07  that just the SCSI cable was broken     2007 04 20 17 03 Sebastian Modules 25  27 in DESY testbeam    Schaetzel  crate 1  slot 12       hold from runs 202391  202392  ahcBeamHoldScan  v 26   Voltages and currents as of 20 April 2007  17h using  DACvalue mod25 A40 9 B44 2 txt  DACvalue mod27 A68 2 B45 8 txt    EDIT  same currents on 23 April  8h    gt  stable  no increase   Vision  1  1  Main HCAL pni  E rd     SEG EHQ 8605 Detail No                   moeitoritw  p Swited On   Off  Voltage  V  Current  wA  Set  V  bestial  V          1 10    10 00 EA      44 46    10 00 el      as 40    10 00 Tree   NEN n Io e el   4 pee 0 00     0 00 j  aa ho     0 00 H sei ail    fi 0 00 R
2.                     2007 07 30 21 00 Erika loading of files    A possible explanation for the too high rate is the sequence of commands executed to change  from pions to electrons and vice versa  It is advisable to first insert the absorber leaving  the target in  as to maximum screan the detector   After this one can load collimators only  from a file by de selecting the magnets when the program asks    Finally the magnets can be set and the target can be removed  This is a very safe procedure  which is normally not necessary when switching between files of various energy     2007 07 30 21 42 Currents in HV10 20 during time of the radiation alarm    CALICE eLogbook   Results for search of  20070701 Page 13 of 20          022019 ri ai        500      400     300       200     100       19 56 19 56 19 57  30  OF  SS 3D  07  2007 30  Ui  ZOD     File  htt   Ittfinfo desy  de CALICEelog data 2007 31 30  07 2007 07 30T21 42 02 00 jpg    2007 07 30 22 05 Alex Catherine 15 GeV electron run    Ok  the scalers where all off  after a second access they are on again now  and we see a rate of  2k pps  The trackers also work again  we did some positioning  so far everything seems  to be fine  As Roman just mentioned ECAL is also stable now        2007 07 30 22 08 Roman Cogri Scripts restarted   gt  Copy to desy    I have realized that the copy to desy did get stuck while copying run  331172  I don t know why         I have stopped and restarted the scripts  with fingers crossed    It see
3.              Ee E    i    rFTTTTTETTT3    shes es F   mE         a       i M bo pe Db    oP     i   12 14 d   10 7E NC  x  cm  A right   gt  inden I  cm        E 1 2 H 1  yt    emp    File  http   ttfinfo desy de CALICEelog data 2007 29 21 07 2007 07 21T13 03 02 00 gif    2007 07 21 13 52    with 6 CMB recovered after manual re init    6 CMB were found off since about run 330732  please keep monitoring the online plot for LED amplitude  if more than 5 CMB are below 2000 ADC counts click on manual re init on the CMB slowcontrol pannel    2007 07 21 14 23 Morin  Beni   Kieffer    330770 JoeamData  10x10    CALICE eLogbook   Results for search of  20070701 Page 10 of 19       Pb pressure  changing  stopped with 38kev  change TRIM5 from  3 3 to  1 to better center beam in y direction    2007 07 21 14 30 Morin  Beni   Kieffer    330771 beamData  10x10   20GeV pi  523 4mm air 0 76 SES  Pb  settings  in  Itransitico   table    cherenkov press  s  better centered   6 8k particles per spill  DAQ rate 84 Hz       2007 07 21 15 30 Morin  Beni  18 GeV pions  cherencov in transition  Kieffer       330772 lbeamData  10x10  18GeV loi  522 4mm air N A  33 cherenko  Pb Still un  transitio    trigger 3 8k  DAQ 80 Hz    2007 07 21 15 39 Morin  Beni  18 GeV pions  Kieffer    Pb    trigger 3 8k  DAQ 80 Hz       2007 07 21 16 28 Oliver Allister Henpien run    Pb    trigger rate  2 9 kpps  DAQ rate  81 Hz       no beam for 15 min  run stopped after 166kEvents    A 2007 07 21 17 08 Oliver Allister H
4.              no debugging symbols found   Using host libthread db library   lib64 tls libthread db so 1    Attaching to program   proc 6595 exe  process 6595  shared object read from target memory  has disappeared  keeping its symbols    no debugging symbols found    done    no debugging symbols found    done    no debugging symbols found    done    no debugging symbols found    done    no debugging symbols found    done    no debugging symbols found    done    Thread debugging using libthread db enabled    New Thread 4129945280  LWP 6595     no debugging symbols found    done    no debugging symbols found    done    no debugging symbols found    done    no debugging symbols found    done    no debugging symbols found    done    no debugging symbols found    done    no debugging symbols found    done    no debugging symbols found    done    no debugging symbols found    done    no debugging symbols found    done    no debugging symbols found    done    no debugging symbols found    done    no debugging symbols found    done    no debugging symbols found    done    no debugging symbols found    done    no debugging symbols found    done    no debugging symbols found    done    no debugging symbols found    done   Oxffffe410 in   kernel vsyscall     1  1 0x00766c63 in   waitpid nocancel    from  lib tls libc so 6  2  2 0x007113f9 in do system    from  lib tls libc so 6  3  3 0x0082c83d in system    from  lib tls libpthread so 0  4  4 Oxf7ac7985 in TUnixSystem  Exec       from  
5.         AJAST  ta x       File  http   ttfinfo desy de CALICEelog data 2007 28 09 07 2007 07 09T07 48 02 00 jpg    2007 07 09 07 55 Michal  Cristina   ECAL linearity  RunId     Runid id  330215 4560 MIPs  330212 6790_MIPs  302137   302087     330213 6803 MIPS  330208 EN STE      weighting scheme for ECAL 1 2 3       2007 07 09 08 21 Michal  Beni  100 GeV pion run  Cristina    CALICE eLogbook   Results for search of  20070701 Page 3 of 25    o MONOS  i 516 ER          11 6 pps in 10x10 trigger coincidence  Average data taking rate   75 Hz    2007 07 09 09 30 Michal  Beni  ahcGain  Cristina  m 2007 07 09 09 48 Beni  Alex wrong LED was switched off    During the last gain run a to high amplitude for some channels in Module 4 showed up  A crosscheck  with Sebastians website showed that the wrong LED was switched off on this board  The   broken LED is no 2  counting from 0  and not 1  After disabeling the right LED the signal   looks fine for this Module     EDIT  2007 07 09  12 32 Due to a misunderstood mapping  wrong LEDs got switched off  Alexander  is fixing this now  New Settings are shown below     Ref  http   www desy de  sschaetz hcal cmbstatus cmbstatus xhtml    mainCMB  V2 pnl   2   5 Ka  p Pulser LED enable disable  e123 48 6786 3 nN       SS A A  7 os of    E  pv En pn p qe JO  11333    u ERNE SS oc ocn  un     MN A  Y  sg                   File  http   ttfinfo desy de CALICEelog data 2007 28 09 07 2007 07 09T09 48 39 03 png    2007 07 09 09 56 Michal  Beni   Cristina 
6.        Nicola  Hengne  Manqi CHANGE of AHC cfg    HCAL crate slot fe 5 7 moved to 5 1    Logbook entry   CALICEelog data 2007 30 23 07  23 07 2007    23 19 Hengne ECal Linearity Check  Low Energy     data of stage position   6  0  added    when stage position is at  0  0  and   6 0   linearity seems to be better  for beam    energy lagger  than 18 GeV     27 09 2007 11 39    CALICE eLogbook http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log       ECal Linearity Check  Low Energy         Logbook entry   CALICEelog data 2007 30 23 07    sy MM 2m Beni changed data disk    changed data link on calice00  updated runNumber txt  changed data link in online monitor    Roman will change the cogri script as soon as all runs are copied  He will tell the shift  when this  task is finished and monitoring should continue     Logbook entry   CALICEelog data 2007 30 23 07  4 me Nicola  Hengne  Manqi    EECHER EIERE        DAQ program  TestBeamControl  seems to be dead at 22 20 23 07 2007      Expert is called      The disk is full              we are changing disk from data01 to data00          Logbook entry   CALICEelog data 2007 30 23 07    ay ces Nicola  Hengne  Manqi    EE    330880 JUNK RUN         Logbook entry   CALICEelog data 2007 30 23 07    EN E Nicola  Hengne  Mango     EECHER    Junk Run          Logbook entry   CALICEelog data 2007 30 23 07       2 of 28 27 09 2007 11 39    CALICE eLogbook http   ttfinfo desy de elog results jsp docroot  export web htdoc
7.       330252 lahcPmNois 0x10    100GeV fpi  516 air 4mm off 20 k  Pb levts    2007 07 09 10 24 Michal  Beni  muon run  tuned HCAL HV layer per layer    CALICE eLogbook   Results for search of  20070701 Page 4 of 25       Cristina    po qr  Pb   pq  Ph    EN NEN LUE    pese  Pb  poe  Pb  ee  Pb    reinstalled front muon counter         E 2007 07 09 11 06 Beni broken trigger cable    During Cherenkov comissioning I found the cable from LVDS converter to backplane input for channel  14 broken  Therfore     CHERENKOV input moved to channel 13       2007 07 09 13 09 Michal  Cristina  ECAL  wrong mapping in the online display  Benjamin  Clearly seen in the attached image  Concerns only the layers of bottom wafers   ECAL   cas   Tour   tracker   Pesci  Do   ocres  sinus   zegsa   Pepe J Hi astenie   TEAT Epia   Costera         Tite  S045  16481 Men Jul 9 2007    Run 330254 0 Event 121050 His  24 Esergy  27 7271             File  http   ttfinfo desy de CALICEelog data 2007 28 09 07 2007 07 09T13 09 37 00 jpg    2007 07 09 14 01 Cristina HCAL linearity  pi      CALICE eLogbook   Results for search of  20070701 Page 5 of 25    30210 74 65  30239 1212    30238 1848  30250  100GeV 2274    24   1 4   34   37 MIP GeV taking into account the e pi ratio       2007 07 09 14 45    No beam for at least 10 minutes     3300259 ahcGain spill sync switched off   used for CMB response  tests    checked following CMBs which did not show proper signal in hstDisplay  E     be switched on manually  gt  
8.       em    10 20  Layer ID    File  http   ttfinfo desy de CALICEelog data 2007 12 25 03 2007 03 26T01 00 03 01 jpg    2007 03 26 09 26 S  Uozumi    Runs 400796 400801 are over night precise 9x9 scan with 3rd configuration     2007 03 26 10 27 S  Uozumi beam runs finished    We have finished all the beam runs        2007 03 26 13 09 S  Uozumi saturation curve measurement    Now entering to the saturation curve measurement  with lots of help from Beni and Erika       2007 03 26 18 01 S  Uozumi Response curve  Response curve of WLSF readout strip with AHCAL electronics   Thanks to Beni  Erika and Peter    Horizontal axis is not yet calibrated to number of fired pixels     CALICE eLogbook   Results for search of  20070301 Page 70 of 76    Q3   e  e  ce  e    N  o  e  e  o    1    MPPC response  ADC counts   3  S       0 100 200 300 400  PMT response  ADC counts     File  http   ttfinfo desy de CALICEelog data 2007 13 26 03 2007 03 26T18 01 35 01 jpg    2007 03 27 07 45 S  Uozumi response map    This is the response map of extruded scintillator  which shows uniformity for the MIPs       300       File  http   ttfinfo desy de CALICEelog data 2007 13 26 03 2007 03 27T07 45 12 00 jpg    CALICE eLogbook   Results for search of  20070301 Page 71 of 76       2007 03 28 08 47 S  Uozumi and AllThe SCECAL beam test has been finished  Thanks  CALICE DESY     Mesurement of saturation curve  and shipping are done   Analysis will be followed soon     We really appreciate lots of people who s
9.      2007 07 11 20 37 ECAL heats first layers of HCAL    Today s temperature development of layer 1  At aprroximately 12 00 the ECAL was switched off     CALICE eLogbook   Results for search of  20070701 Page 18 of 25     c  rara en woe  SEE             CEDER       y  peperere rien qen TS TERRENT Y T eS prr revue vereor ever AS  3 00 22 00 01 00 wie 07 00 00 13 00 16 00 19 00 22 00 01 00 04 00 onto Gel  17 2007 10 07 9007 11 07 2007 11 07 2007 11 07 2007  rm 11 07 2007 121 07 2007 11 07 2007 11 07 2007 12 07 2007 12 07 2007 12 07 2007       File  http   ttfinfo desy de CALICEelog data 2007 28 11 07 2007 07 11T20 37 02 00 jpg       the effect can still be seen on layer 2    07 12 2007 04 03 06  8 395   07 12 2007 04 03 06 CH    07 12 2007 04 03 06 me          DEER          y  pepe  air enn a A S S TERT ad AA O A PH V I ven  3 00 22 00 01 00 04 00 07 00 00 13 00 16 00 19 00 22 00 01 00 04 00 E Gel  17 2007 10 07 900  11 07 2007 11 07 2007 11 07 2007  rm 11 07 2007 121 07 2007 11 07 2007 11 07 2007 12 07 2007 12 07 2007 12 07 2007    File  http   ttfinfo desy de CALICEelog data 2007 28 11 07 2007 07 11T20 38 02 00 jpg    Layer 38 shows a different behaviour  So the effect seen bevor is not from a temperature change  im the hall     CALICE eLogbook   Results for search of  20070701 Page 19 of 25           e Pos 38   Module 38   CMB 36 PHEL to det    Tene    1    C       07 12 2007 04 03 06 44 395                     2  2  S  cf HR             File  http   ttfinfo desy de CALICEel
10.      Re  NET   weit rpegr  NNNNNE    Moe  Seed  Aa we    Deag F1  21 0807    File  http   ttfinfo desy de CALICEelog data 2007 25 21  06 2007 06 21T 13 41  35 00 pdf    2007 06 23 00 25 Ecal Installation    Dear Colleagues     today a team of LLR LAL with support from the DESY team at CERN has  installed the SiW Ecal in front of the  AHcal  The installation went very smooth  nothing particular to report     The Ecal assembly consists of the the central 15 slabs   30 layers  and   6 lower slabs  The r o cables were connected to the   PCBs  Cables for three more slabs expected to be commissioned very soon  are in place  The final routing of the cables   to the rack housing the CRC boards will be done  or has already been   done  by K  Gadow    The Ecal is ready for the electronis tests which will start monday morning   Cheers     Roman    P S   A table with the current cabling is attached to this posting    CALICE eLogbook   Results for search of  20070601    LETT cep tu A Qm des ds dl id dl E    Ja  s    28  m  ij  aT  28  dE  28    File  http   ttfinfo desy de CALICEelog data 2007 25 23 06 2007 06 23T00 25 02 00 pdf    2007 06 23 09 18    Alexander     Pen CS calm d    Tc  uec  4 c      c  ac  a  Le  a c  a c  m    He   4 c  me   c  UE C Undated     c    ra n    La  um  none oom    Be    E  fake    Sebastian    CAN bus cable was disconnected at HV supply      i    i        i    f  wech  d    E  EF    BESESSREGRRESAATHRRESRRSS RACE BREBHEY    HV does not switch on    ILC DLL T
11.      nx      e Fcat   pecas   rener f mex   Parcino Coupe   Ces Chen   sho   Settings    i A rr Become   reee Zezecpe    L  I               gt     2  H a E  A VELLE  pr  D  pene    e EA Tom   LI d    z 2  f les e       em  At  pm       N      SS TT H D   wi w     er Fe            See 14228 404 009 a Av 16 ENT  Se ng teres ome             2          ie  Tm X Monitor B caceondcabcesen  he      File  http   ttfinfo desy de CALICEelog data 2007 28 14 07 2007 07 14T14 49 42 00 png    2007 07 14 14 50 caliceonOcaliceherescpMstivih   CALICE user     moved stage 2cm down  beam hits ECAL x centered and y between the wafers  target position  10 15     HCAL Motor control panel    actual position target position    x   10 01 bo ox   yf u s bs cora  Y     O lost X reference point       lost Y reference point       O lost 2 reference point    Q stage in motion  O manual operation get Pos    ficus 1 slow 1 pulsezfree   extern STOP   ref 24   304 968 304  068        Offset   dE EI a KREE  File  http   ttfinfo desy de CALICEelog data 2007 28 14 07 2007 07 14T14 50 02 00 jpg  2007 07 14 15 48 hold scans    Pee   ahcPmLedHqspStan  sync on       iol ahcPmLedHdspasd  n Y   sync off vcalib  settings    CALICE eLogbook   Results for search of  20070701 Page 19 of 20    330443 ahcPmLedHqspStan changed  sync on vcalib  settings    330444 ahcPmLedHdsast  n changed  sync on vcalib   settings   330445 laahcPmLedHgsgSct    n changed  sync on vcalib   settings    no beam       2007 07 14 18 12 George  Tao
12.     2007 07 05 19 44 Francois Jean Yvd amp Stasiughrilum Fab    330098 lbeamData 100x100 mu  Combined HCAL TCMT  voltages 0 0    Beam file loaded  H6B 226  120 GeV secondary  beam focused in H6a     Some collimators  settings have been changed   C3 from 10 to 15 in JAW1 2  C1 from 10 to 15 in JAW1 2  C10 from 5 to 10 in JAW1 2    Average rate  1000 pps on 10x10  2000 pps on 20x20  8500 pps on 100x100   Run stopped at 83K events   Average DAQ rate 84Hz    2007 07 05 19 44 Francois Jean Yvaniiusnauftrika Fab    330099 lbeamData 100x100 mu  Combined HCAL TCMT  voltages 0 0    Beam file loaded  H6B 226  120 GeV secondary  beam focused in H6a     Some collimators  settings have been changed   C3 from 10 to 15 in JAW1 2  C1 from 10 to 15 in JAW1 2  C10 from 5 to 10 in JAW1 2  C11 from 5 to 10 in JAW1 2    Average rate  1800 pps on 10x10  4000 pps on 20x20  14000 pps on 100x100    beam stopped     run stopped    2007 07 05 19 45 Francois Jean Yv  uftlas  n Erika Fab    330100 lbeamData 100x100 mu  Combined HCAL TCMT  voltages 0 0    run stops automatically after 2000 events     gt  junk run    wFkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkzk    I am not sure what the above entry means  The run on disk has 28613 events  as is easily shown using   icalice00   count 330100  It seems to have been stopped by the run operator  by hand  as the log file  Log1183650014 0ut icalice00  shows the next run and sequence starts immediately  If the run had stopped  itsel
13.     CALICE eLogbook   Results for search of  20070201 Page 19 of 21       to 1 4  this change happens when the SIPMs are turned on  because the additional voltage drop over  the 100k0hm can only happen when the circuitry is closed  which means when the SIPMs  are ON       to 1 3  Multimeter correction is needed and is 1  of the measured voltage    U SIPM    U 1 3     100kOhm 10MOhm   U 1 3      39 2V   0 392V    39 6V   SUMMARY         1   cable HV drop  100mV from HV distribution box to VFE electronics  HBAB input   2   if SIPM operates and draws a current of 300 500nA     gt  U drop   100k0hm   0 03   0 05 V   if SIPM has long discharge   gt  1muA  U_drop  gt  100mV    measured with SIPM 1648  U_drop   0 15V  tile tester current 1 84muA     gt  SIPM with high current drops the HV      gt  we have an average total U_drop of  200mV       File  http   ttfinfo desy de CALICEelog data doc CERN docu 2007 02 27T16 49 14 02 pdf    a 2007 02 27 21 18 Marius Simple Binary Analysis  SBA  zipfile with code    Manual and documentation can be found in the next elog entry     CALICE eLogbook   Results for search of  20070201 Page 20 of 21       File  http   ttfinfo desy de CALICEelog data doc CERN docu 2007 02 27T21 52 58 00 zip    la 2007 02 27 21 19 Marius Simple Binary Analysis  SBA     The enclosed zip file  next elog entry  contains small pieces of DAQ analysis code for the AHCAL     If you have the DAQ code you can recreate a daquser area with the following links to the DAQ code    
14.     RE daa  V U40i   9  ES                    LL fal possible    l  dead  V U401   0       2007 07 09 15 27 Michael  Beni   Cristina    lead    Run stopped to allow the preparation of a lower energy beam             2007 07 09 15 33 Beni DAQ work 8  ECAL stage re integration    I tried to switch off all switches and timer that could bring the ECAL stage laptop to any other  mode than alive     all configuration files should be synchronized again   Changed calice00   online src runnerDefine icc to connect to ECAL stage  buildAll on calice00   buildAll on calice02    restart of GUI    REMARK  on calice01 runnerDefine icc already had the ECAL stage in  skipped build all on this  machine    2007 07 09 16 31 Mary Cruz HengneNiefsrun       330261 lbeamData  100x100   80GeV mu  555 8mm air on  Pb calibratipn   limitatiohs    disk is changed to data01 from data00 and DAQ restarted at 17 20 09 07 2007    2007 07 09 17 17 Mary Cruz Hengne Niels    Pb       Pb    2007 07 09 18 28 Mary Cruz Hengn   tdintsmuon trigger removed from beam line  We ll move on to electrons and removed the 100x100 muon trigger in front of the experiment   2007 07 09 19 06 Mary Cruz Hengn amp taigs moved  Stage position moved to   6 0     2007 07 09 19 47 Fab DC trip    CALICE eLogbook   Results for search of  20070701 Page 6 of 25       All three DC tripped after shutting down muon chamber  Mapping of DC HV cables  in case needed to turn them on after a trip    DC1   gt  HV channel 36  DC2   gt  HV channel 37  
15.     TA   2007 03 14 17 02 S  Uozumi LED run   00143      EmMoise Pedestal              JEiber directp  v         I                 1  2007 03 14 17 03 S  Uozumi LED run    400145 Embed     TEp                              Eiber directp 5v   gt   mea 7     CALICE eLogbook   Results for search of  20070301 Page 6 of 76       to see  lp e  peak  at L6S14    2007 03 14 17 03 S  Uozumi LED run    fiber direct  2 5V            2007 03 14 17 04 E Uozumi  400445 CmLed ptun  direct 2 5V CmLed  to see lpe  peak around  L6S14  2007 03 14 17 04 SE Uozumi  400445 CmLed II direct 2 5V CmLed  to see lpe  peak around  L6S14  2007 03 14 17 05 S  Uozumi LED run  440447 CmLed LED fiber direct 2 5v CmLed  to see lpe  peak at  L11S5  2007 03 14 17 05 S  Uozumi LED run  440447 CmLed LED fiber direct  2 5V CmLed  to see lpe  peak at  L11S5  2007 03 14 17 05 S  Uozumi LED run  400448 CmLed LED fiber direct  2 5V CmLed  to see lpe  peak around  L11515  2007 03 14 17 05 S  Uozumi LED run  400448 CmLed LED fiber direct  2 5V CmLed  to see lpe  peak around  GLISIS  2007 03 14 17 06 S  Uozumi LED run  fiber direct   2 5V CmLed  for  inter calibr  tion  2007 03 14 17 06 S  Uozumi LED run  400449 CmLed LED fiber direct  2 5V CmLed  for  inter calibr  tion  2007 03 14 17 17 S  Uozumi LED run    400450 PmLED LED fiber direct  2 5V PmLed  run for  inter calibr  tion    CALICE eLogbook   Results for search of  20070301 Page 7 of 76       2007 03 14 17 17 S  Uozumi LED run       2007 03 14 17 17 Run summary of 
16.     enc DEP Switch SO  WI  lower LES 17 504 Y CR  a        S kass u mg D  REF 1 2894  3 235 v E p n 12345678  Well call  6x3 v    PAEL to elos         File  http   ttfinfo desy de CALICEelog data 2007 28 11 07 2007 07 11T21 13 02 00 jpg    2007 07 11 21 14 All green after check    CALICE eLogbook   Results for search of  20070701 Page 23 of 25       File  http   ttfinfo desy de CALICEelog data 2007 28 11 07 2007 07 11T21 14 02 00 jpg       All temperature sensors of module 20 show exactely the same temperature    ps iv Pos 18  Module 20   CMB 31                   WA kal MA Va Met Ute niih id ajii A HHG AN MU OA y  BE 20 00 00 00 06 00 14 00 20 00 40 00 08 00 14 00 20 00 02 00 08 00 14 00 20 00 ee   107 2007 06 07  2007 09 07  007 48 07  2007 09 07  2007 09 07 2007 10 07 2007 10 07  2007 10 07  200  10 07 2007 11 07  2007 11 07  2007 11 07 2002 14 07 2007 12 07        File  http   ttfinfo desy de CALICEelog data 2007 28 11 07 2007 07 11T21 23 01 00 jpg    CALICE eLogbook   Results for search of  20070701 Page 24 of 25    CEDER             REH CC Y  14 00 20 00 00 00 080 14 00 20 00 40 00 08 00 14 00 20 00 02 00 08 00 14 00 20 00  0   07 2007 06 07 2007 09 07 2007 09 07 2007 09 07 2007 09 07 2007 10 07 2007 10 07 2007 10 07 2007 10 07 2007 11 07 2007 11 07 2007 11 07 2007 11 07 2007 12 07           File  http   ttfinfo desy de CALICEelog data 2007 28 11 07 2007 07 11T21 28 02 00 jpg       This shows the typical temperature behaviour  Only remarkable feature are the big spikes i
17.     m 1  HE  mE iri  Trach tz wae   m     DO  FE  sE  in  Le   E       E ELLEN A A    File  http   ttfinfo desy de CALICEelog data 2007 27 08 07 2007 07 08T 19 32 31 04 eps    2007 07 08 19 54 Niels  Jean Yves   Hengne       330228 beamData  10x10   50GeV e  531 air 6mm off 250K  Pb events    DAQ mean rate  73 Hz    2007 07 08 21 05 Niels  Jean Yves   Hengne    330229 ahcGain 4 mm air off 0 0 11  330230 lahcPmNoi zk 4 mm air off 0 0 50 k  pee Lobo     ka P Bo 1  o ves               2007 07 08 21 39 Niels  Jean Yves  Voltage changed  Hengne    CALICE eLogbook   Results for search of  20070701 Page 24 of 24       HCAL voltage changes for some modules     3  0 3V  4  0 3V  5  0 3V  6  0 3V  7  0 3V  8  0 5V  9  0 3V  10  0 3V  11  40 3V  12  0 3V  14  0 5V  34  0 3V  2007 07 08 21 43 Niels  Jean Yves     Hengne    Pee    air 250k  Events    tl  Pb Events    Modified voltage for some HCAL modules   before these  runs          CALICE eLogbook   Results for search of  20070701 Page 1 of 22             2007 07 01 05 21 Roman cogri scripts and software release    Dear Colleagues   I have launched the copy  and grid    therefore cogri from now on      scripts     The scripts can be found on caliceserv under  scratch home caliceon cp xxxx xxxx ecal hcal comb    The scripts copy the binary files into the desy dcache and immediately launch the  conversion on the grid  As usual for the starttup phase I had to remove some bugs in  the scripts  In combination with other problems  see below 
18.    E NEN     330077 ahcPmNoise combined HCAL TCMT 0 0  nominal 0 3V    40 0  TCMT        330078 ahcGain LED combined HCAL TCMT 0 0  nominal 0 3V    40 0  TCMT     Run 330076   new settings loaded  file CERN 030707  0 3V txt    ECAL   same features  a bit better  First 150 events  layer 0 noisy    600 MIPS  and total energy  up to 900 MIPS     then it decreases to total energy 100 MIPS     Some peaks in energy  not in coincidence with HCAL though  see below         CALICE eLogbook   Results for search of  20070701 Page 18 of 22                L     A hEcalEnergqy v sTime      jaa ee    wr    ES  Sat s  E roa K     1  d en  eg  a  ad  Pa  ica  4 r  ET  bei    i hHealEnergy Temo M  1 ve e         E isa r i     UU  E emim a    rri UT   nd aom DATE  E Ha beier ee Abde kaal i i  mi  a     H  a  Iron eee    ieee gs   EN T S   a  E 1   i  B erg i  3 ET mri A  a LE   E aa  i SR  y hz de y Ema z TE     a CA bk    T  e e   TE   eeneg der Dn  File  http   ttfinfo desy de CALICEelog data 2007 27 04 07 2007 07 04T01 07 19 03 eps  2007 07 04 02 11 Anne Marie  Heng imp hifi 0 3      B    il o  Do  e  ahcGain LED combined nominal1 0 3   0 0   for  ECAL HCAL    40 for TCMT          CALICE eLogbook   Results for search of  20070701 Page 19 of 22       330081 is test run for ECAL in same conditions    There are several chips in layer 0  middle raw  left column in  x y  display  see below  being  really   noisy for 2 minutes  nearly exactly 120s  see plot in next entry  and 12000 beam events    s
19.    mr j  ele   ask                Bm     Eee ol j  M Crate OFF    SREL we    recover all el LE      Si Wi values to file   dm HAL valves  position                  File  http   ttfinfo desy de CALICEelog data 2007 29 22 07 2007 07 22T19 52 02 00 jpg    o 2007 07 22 21 25 Hengne Allister ECal Linearity Check  High Energy   Seems to be alright    for different stage position  0 0    0 43     3  3     3  3     6 0       data chosen of stage position  0 0  are Run 4 330428 to 330433   taken on 14  07  2007 night shift      no error considered since they are neglectable    thanks to Allister for the fits                   Hengne  ECal Linearity Check  High Energy     26500  E6000  o i  3   5500  c   5000  z D  4500 j  g  4000  D  3500  3000   Y  2500 oe  P  2000 e   KEE AA AAA Ree EE  Loo O 20 25 30 35 40 45 50  Beam Energy  GeV   File  http   ttfinfo desy de CALICEelog data 2007 29 22 07 2007 07 22T22 00 25 00 png        2007 07 22 22 16 Hengne Allister ECal Linearity Check  Low Energy     Seems to be alright  splitted a bit when energy lagger than 18 GeV    for different stage position  0 0    0 43     3  3     3  3       Since Run330717  Run330719  Run330720  which correspond to energy 15GeV and 18GeV at stage  position     0   3   were not online Analysed  these two points on the plot of stage position  0 0   are left blank     no error considered since they are neglectable      thanks to Allister for the fits    Hengne    CALICE eLogbook   Results for search of  20070701 Page
20.    the conversion is not fully in phase with the datataking    Ecal and combined runs are nearly fully converted  the hcal runs are tumbling in    The ecal and combined binaries are fully at desy the hcal runs are twenty runs behind     Therewere two problems this night  a file which was already copied by hand blocked  the continue of the data transfer and therefore also the conversion  On the other hand  there was a grid communication problem with desy between 3 h and 4 h this night     Now the situation looks better and runs are tumbling in  both binaries and converted files   When you wake upon sunday  the processsing should be in phase     Shift crews should check whether the scripts are running by  typing  ps aux   grep cp     in any X window on caliceserv    You should see the three perl scripts active      The status of copying and conversion can be most easily checked by typing    lfc ls  1  grid calice tb cern native dat   grep Run3n0 where n 3 4 5   for the binaries and   lfc ls  1  grid calice tb cern raw conv  v0402   grep Run3n0 where n 3 4 5  for the converted files    The shift crew should check that the files displayed there are not way behind the data taking   five runs at most       The latter action can be done from any grid ui in the world        will try to make the control of the data trasnfer processing more convenient for the  shift crews in the next days     Cheers   Roman  P S   Please do also have a look on my annoucnement on the software release in
21.   07    22 07 2007  16 30    I had modified the record in elog about absorber and target for run 330478 330508   for point  out        Manqi Re  Mistake in elog    Logbook entry   CALICEelog data 2007 29 22 07  22 07 2007  15 27    5  5  435 1  435 2  435 1  1   3   6   8   2    Brient Morin Lutz    3993  283   282   275   222   221 43  178 58                Logbook entry   CALICEelog data 2007 29 22 07    22 07 2007 caliceon  calicehcalsc0O2 cern ch  CALICE  13 14 user     channel 28 HV   module 16 B   slot fe 17 0 current dropped after DAC setting reload  current limit for channel 28 changed from 740 to 350       module 16 B noisy          Thanks    27 09 2007 11 39    CALICE eLogbook    22 of 28    http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log               M 40 00  MT 10 0 10 00 M fume See   OFF  hs A asaz  42 60 rJ SE    BS 2 42 4  0 2        07 00 03 00 11 00 13 00   gt  2 07   2007 22 07 2007 22 07 2007 22 07 2007       Zosen  0 2  ada    AS A 41 0 90 2  i  So a Mi    HY settings   file actions      cenible   disable ALL channels     PRA ee             E ae Rowe   OFF TEXTEN  A e o       10 00 M  Bee Paso   OFF   OC M6 Crate oF    FEET AS  EAT 0 We      10 00 W Ramo   OFF Ja T2 A 46 64 0 2  recover abl rel LW3   Ws j m3 LY dean   T 10 00 W Rane   OFF E dae Wi volus te File   _dinp REL values  position       Jj                Logbook entry   CALICEelog data 2007 29 22 07    22 07 2007  10 12    330798  beamData 10x10  18GeV  pi  522 2 air 0 
22.   10GeV e  520 air air in 12k  CT TT peoe TT D P Pt Eau events    H popa rep E E E E p pes    events    e  beam slightly contaminated by pions         2007 07 11 06 51 Kurt    330348 lbeamData  10x10  8GeV e  525 air air 3 0005 200k  events          e  beam slightly contaminated by pions  17     ECAL layer 12 has factor  2 too many hits  judging from neighbour layers   We have not seen this  earlier tonight     QUAD11 is also unstable now  just as QUAD1 has been for all of the shift  Neither deviates much  from nominal values  but they are frequently off value     A 2007 07 11 07 34 Trygve HCAL Channel 50   Current in HCal channel 50 has been rising slowly throughout the shift  The rate is approximately  luA hour  The channel is still not far above threshold  but since it has a persisiting  trend this should be looked into     CALICE eLogbook   Results for search of  20070701 Page 16 of 25             2  5  F  11 88  Z     ifaw    A  Jaso    470     460  450  440    18700 22300 02700 CHA  1 07 2007 10 07 2007 11 07 2007 11 07 2007       File  http   ttfinfo desy de CALICEelog data 2007 28 1 1 07 2007 07 11T07 34 01 00 jpg    2007 07 11 10 18 Erika optical fiber swap    the optical fibers of calice02 and calice01 have been swapped  according to Paul request          2007 07 11 11 53 Beni temperature development in our detector   The plot shows the average temperature  calculated from the first temperature record in each run  taken since 4th of July  The start of the records co
23.   118 320A 149 W RIA Ea npea  Sen  jaman      a aan X sa a    384  565   w SISA  60   042    144 RES    63 T8     EM mo n    w 338 j a 146 W HESE   205 wrae E x    WA 2 170 TSA       17     Fan t 151 W  004   iso  In    2      ss      125    ejan i z   199 w opoe  319 7e EN 5     109 maisa    7   103 wW ipia    3 1  ws Y mme DESS   45     103 monse Ra ae MW ope zi pe 781999 EE o   Ce So  E  DLW j CAT   He Crate ON   EN Crate OFF    recover  KILL wy    HEA   o   87 1 19     GE 228251 es  et  recover all tripe   403   Bre   ma   Lv detail     due H   values to file       dump AL valuse  position          X sa x 46 X  X       Experts       HY settings   file actions enable   disable AL chanele  sw   Lom   ur   enable f disable               File  http   ttfinfo desy de CALICEelog data 2007 31 30 07 2007 07 30T08 20 02 00 jpg    2007 07 30 08 21 Erika  Ariane    Pb       2007 07 30 09 15 wobbling change  change wobblings to  80GeV    2007 07 30 09 36 Ariane  Erika  2007 07 30 09 43 Ariane  Beni  Muon Calibration run  Nanda    deg    muon calibration  beam dump closed  20x20 trigger rate 5000 pps  DAQ rate 90 Hz  200 kevt    no beam     stop run to load new vcalib settings for       2007 07 30 10 25 Ariane  Beni  Muon Calibration run  Nanda    331158 lbeamData  20x20  80GeV mu  555 Air Air 20  deg  50 kevt    LED signal in layer 25 is fine now    layer 28 29 35 still don t show any LED signal   few layers show LED signals above 18000 ADC channels  try to lower these   2000  2000  
24.   2007 07 27 03 28 Nanda  Roman e  run  per qeu ms bees eye  ur e soper se c   d  Pb  200 kevt  2007 07 27 04 14 Nanda  Roman  e run  PRA TA  a pepe AAA  Pb  230 kevt  2007 07 27 05 05 Nanda  Roman e  run       Pb    200 kevt    CALICE eLogbook   Results for search of  20070701 Page 2 of 16       2007 07 27 05 49 Nanda  Roman  e  run    Pb    200 kEvents                             2007 07 27 06 49 Nanda  Roman calibration runs  30597  30998  30999  31000  31001  2007 07 27 09 26 Erika No beam  from 8 00 to 12 00 no beam to SPS  2007 07 27 11 03 Erika  331002 ahcGain  331003 ahcPmLedVcalibscan  331004 ahcCmLedVcalibscan  2007 07 27 11 49 Erika 20x20 trigger config    setup the DAQ for 20x20 trigger    2007 07 27 13 08 Erika Fabrizio Mar20x20 muons    331005 lbeamData  20x20  80GeV mu  555 Air Air 10deg  10k  events    DAQ rate  80 Hz  6600 pps    2007 07 27 13 27 Erika Fabrizio Mar20x20 muons    331006 lbeamData  20x20  80GeV mu  555 Air Air 10deg  190k  events    DAQ rate  90 Hz  6600 pps    2007 07 27 13 57 Erika  Fab 20x20 muon at 10deg  muons at 10deg go straight through ECAL and HCAL  we did a good job of stuggering   HCAL has 7 noise hits and 42 hits for the muon peak   7   35  not bad approximation of 38 layers     ECAL has   1 noise hit and 30 hits for the muon peak    CALICE eLogbook   Results for search of  20070701 Page 3 of 16       nos d  Zelt rightsas   denk      hHealfBarc ende rSlay eraX Y ka          Spe ENN        4 6 H i   S   i    y icm     File  http   tt
25.   400302  point0 35    5Gev 400296  point0 35   400310  point0 35   400311  point0 35   400314  point0 35   6GeV 400285 400291  point0 35    center part scan  9 points    1GeV     400309  point36 44    2GeV 400308  point36 44    3GeV     400305  point36 44    4GeV     400302 point36 44   400313  point36 44   5GeV 400296  point36 44   400310  point36 44   400311  point36 44   400314  point36 44   6GeV 400264  Run400285 400291  point36 44   Center injection  1 points x many times    1GeV     400317 90k ev    2Gev 400318 90k ev    3GeV     400319 90k ev    4GeV     400320 90k ev    5GeV 400314  point45 54  400321 90k ev    6GeV 400322 90k ev     end    of  list  configuration Ist pound esse SoS Ue REPE    LED  Fun  with  st configuration   only 3  layers  eelere    Pedestal runs    400382     CmNoise   400398     PmNoise   for Layer7 Strip15    400406     CmLed for inter calibration   400407     PmLed for inter calibration   400385     CmLed to see lpe peak   400386     CmLed to see lpe peak with over voltage  1V  3 5V     for Layer8 Strip15        400393     CmLed for inter calibration   400396     PmLed for inter calibration   400394     CmHoldScan   400395     PmHoldScan   400401     CmLed to see lpe peak  100k ev    400402     CmLed to see lpe peak  100k ev  but with slightly higher light to see adjacent strips  400419     CmLed to see lpe peak  100k ev    400420     CmLed to see lpe peak  with over voltage  1V  3 5V     for Layerl4 Strip13                    400411   CmLED 
26.   5   MIA 2 w MA    100 m8   100 98158  ou MW NA 117 W SoA    10   114 w 4008  w TARTA      115 SIA   34 wA      1x10 3138   34 paie    ous W E d 35 KUER      117 M NIB    35 amp  w 44p2B    102 MASA    38 mIa    99 wr OSB   w 4638    100 MA 157 MIRAR 27  W TADA    90  E CANO 107 AGA ru Ww spe E DS  gesch MV  ramp UP DOWN Lv am ofr Grey HY OFF Oey LY OFF  COo MET las e EE   m  recover    KILL Hy    Jede     HY settings   File actione erable   disable ALL chanel  AL two   oi   enable   disable         w TIROA  w 701588  w ITRIA      Ww 7398    Js helle lol    mg          Nw  recover all tripa   H3   we J w9 J LY detal   duae Hi values to file   dap AL values  position                File  http   ttfinfo desy de CALICEelog data 2007 29 21 07 2007 07 21T09 06 02 00 jpg    CALICE eLogbook   Results for search of  20070701 Page 8 of 19       2007 07 21 09 30 Morin  Lutz   Kieffer       330760 lbeamData  10x10  20GeV e  523 air air 0 76  3   3 position  tuning  run    trigger rate 8K       2007 07 21 09 37 Morin  Lutz   _Kieffer  330761  Jesambata     x10    206ev e 23 fir opie    De L3  T     run is after ERIS k events to account for the high pion contamination   3 1     w r t electrons  estimated 100 k ev good electrons    2007 07 21 10 53 Morin  Lutz   Kieffer   230762  beamData  10x10     25GeV fe        526     fair fair jo  5 tT    6300 particles per spill  DAQ rate 85 Hz                      2007 07 21 12 22 new position  moved to   3  3  which is  50 10  in stage postion   
27.   5 001 V WE upper LED 17 423 V    intct dimisi  UIP Switch S002  Power LON   32 440 V MIO lover LED    434 y    ET P ET  PEESI   5 235 v Ventum pu 12345678  A be  ca    PAEL to elos       File  http   ttfinfo desy de CALICEelog data 2007 28 11 07 2007 07 11T21 09 02 00 jpg    2007 07 11 21 10 problematic CMB    CALICE eLogbook   Results for search of  20070701 Page 21 of 25    Module 12   CMBO  Pulsar Settings  Apel DAC  Vidth Dat pe     gt  A RR  Jael Vealib  Width  351  J52     J feel DAC  a  351  352  fiidh DAC    ho     feel Vealib  Width DAC  r Enable LEDs  0123456789104       Hlpoc EXI CARa U EE OEE    Tupper D  I eem  C Wi uper LED  ass y  1    a   manual RE INIT    V calib at Uti   0 005 v WE pper LED 17 471 V      PDIP Switch S009    Power AN  iso v MMWLlowerLE  7 508 V   bass P DH  REF 1 2894  3 235 v coo ER 12348678  VLL upper  LES LECH Vocal bat 1051 110 005 X    PAEL to elos         File  http   ttfinfo desy de CALICEelog data 2007 28 11 07 2007 07 11T21 10 02 00 jpg    2007 07 11 21 13 problematic CMB    CALICE eLogbook   Results for search of  20070701 Page 22 of 25    Module 34   CMB 37   Pulser Settings    J rel DAC  Von DAC pH    pa jue  Jal Vealib  Width  351  J52     Jael DAC  Width  351  352   HIEN DAC        20     feel Vealib  Width DAC    y Enable LEDs  012 789104       TA R O   C 121217 000  c Tihi   t   14  317 000 e  19 217 000 at    Tapper 018  57 555  e Wi uper LED   gt  ass y  L ben i manual RE INIT  V calib at  Jl   bo Y MIC upper LED   7 530 Y
28.   9points  plus central  injection  lpointxl0times     2007 03 13 08 41 S  Uozumi 3GeV stage scan       3GeV fiber 2 5V scanned overnight  direct scan    with detector scan  36point    plus central part scan  9points  plus central  injection  lpointxl0times     2007 03 13 13 07 S  Uozumi et al     We have no beam this morning  doing analysis and working on LED monitoring system     2007 03 13 13 11 S  Uozumi LED run  INN SLE  gt  PAE E A  run  2007 03 13 13 16 S  Uozumi LED run  pow AA pee e  run  2007 03 13 13 16 S  Uozumi LED run  ue  uw peru AAA Jn MERA  run  2007 03 13 13 17 S  Uozumi LED run  pecu MO AA  2007 03 13 13 17 S  Uozumi LED run        400391 LED LED fiber direct  2 5V  events for  inter calibr  tion    2007 03  ios 13 19 i Uozumi LED run    400393 fiber direct  2 5V CmLED  with 50k  events for  inter calibr  tion    same light yield with to check stability of the LED light pulse    2007 03 14 09 22 daniel new calibrations in online monitoring  i updated the calibrations in the online monitoring with values from run 400262    before they were a mixture of rough estimates and guesses        these calibrations are now applied to all of our runs     400000    CALICE eLogbook   Results for search of  20070301 Page 4 of 76       2007 03 14 10 57 S  Uozumi mapping for direct fiber configuration  Here you will find a mapping for 2nd configuration  direct fiber    http   azusa shinshu u ac jp  satoru direct WLSF mapping v6 dat    2007 03 14 11 24 S  Uozumi 6GeV run  36
29.   Calibration constant  ADC counts     10  Layer   p 0    File  http   ttfinfo desy de CALICEelog data 2007 12 25 03 2007 03 25T23 23 38 00 jpg    2007 03 25 23 24 S  Uozumi    This is a distribution of calibration constant for fiber readout module     3 dead channels are not filled in this plot     CALICE eLogbook   Results for search of  20070301 Page 63 of 76    Fiber readout strips    calibdist1    Entries 231    y  e    Mean 245    RMS 49 32    Number of strips          400 200 300 400    Calibration constant  ADC counts     File  http   ttfinfo desy de CALICEelog data 2007 12 25 03 2007 03 25T23 24 34 00 jpg    2007 03 25 23 25 S  Uozumi  And this is a calibration constant distribution for direct readout module   2 dead channels are not filled     I  e    Direct readout strips    calibdist2    Entries 232    Co  e    Mean 110 6    RMS 10 51    Number of strips  N  O    e  e         100 200 300 400    Calibration constant  ADC counts     CALICE eLogbook   Results for search of  20070301 Page 64 of 76       File  http   ttfinfo desy de CALICEelog data 2007 12 25 03 2007 03 25T23 25 17 00 jpg    2007 03 25 23 25 S  Uozumi    If you want to see more distributions  you can look into attached root file           File  http   ttfinfo desy de CALICEelog data 2007 12 25 03 2007 03 25T23 25 56 00 root    2007 03 25 23 26 S  Uozumi    Table of the calibration constant from MIP runs with 1st configuration after tungsten removal   Format   layerID stripID calibconst calibconst err f
30.   Cheers    Roman   I should add that root can be found via  ROOTSYS and LCIO via  LCIO   The  bin files are visible via  mnt data00  01           2007 06 20 11 02 Beni calice00 login problems    The reason for the calice00 login problems was most probably a full root partition  Someone saved  hughe software archives in   and  root  Please  use a different location for software  installation in future  I moved some data from this places to  srv calice data01 root bk    If you miss something look there first     I still have some problems to get the system back online  I will edit this entry if the system is  back     EDIT  Computer is back online  For security reasons the root password was changed     2007 06 21 13 41 Doerte David assembling status    All cassettes are assembled    SiPMs for module   24 and 36 are exchanged    Module 2 has one half with new the other half with old SiPMs tiles    Module 1 has a pure mix of old SiPMs  All were tested  all are working but not with best conditions   Hopefully it will be ok as long as necessary     Shipment to CERN will be next Wednesday by our colleagues     CALICE eLogbook   Results for search of  20070601 Page 4 of 19       cassettes work in progress  D    cassettes  nearly  ready for transport  1  2  24  36  cassettes at CERN  E  status of production  25c   07 cassette assem bing HAL   FLO    assembling steps staff  3 4 D D D 9  base frame SPMs FRA plate pets cabing fibres     tempaensors cover plate no of persons       H     
31.   Hakan       350285 beamData  10x10  50GeV Git 612 4mm Air 0 5    6 0 20deg   Pb 166k    CALICE eLogbook   Results for search of  20070801 Page 12 of 22          A events         2007 08 20 02 47 Satoru  Allister   Hakan  Pb 157k  events   amp  2007 08 20 03 47 Satoru  Allister  data taking stopped due to unstable magnet  Hakan    When we loaded beam file 614  one magnet  BENDO3  fails to adjust to the new reference value and  is also rather unstable  Reloading the beam file doesn t solve the problem  We have called   the Accelerator Control Room to ask for support        Accelerator guys told that it is due to a problem of magnet power supply    Experts will fix it in 30 minutes    Take calibrations meanwhile        2007 08 20 03 57 Satoru  Allister  calibrations  Hakan  50287 lahcGain TT  50288  50289  50290 lahcCmLedVcalibScan CO U U U O  50291 jahcPmLedVcalibScan SO   2007 08 20 04 47 Satoru  Allister   Hakan    Call from control room   Magnet power supply problem has been fixed   Continue data taking with beam     2007 08 20 04 48 Satoru  Allister   Hakan    g WI     oit 614 Air 0 5  Pb  MINO  Pb  2007 08 20 04 55 Satoru  Allister     Hakan    The problem on magnet BENDO3 is appeared again   Called to accelerator control room        2007 08 20 05 32 Satoru  Allister     Hakan  BENDO3 is back   Start run 350293     2007 08 20 05 39 Satoru  Allister     Hakan  BEND03 sometimes drop off from the right value   But continue with run 350293 anyway     2007 08 20 06 18    BEND
32.   Ix    i T  1 f s 2 x F a     la z I    am s     a  E K   a a    E E Di   pu mn ipa ew Fe  ua  File  http   ttfinfo desy de CALICEelog data 2007 27 06 07 2007 07 06T01  21 39 02 eps  2007 07 06 02 12 Alex  eshi SH muon run  HCAL TCMT voltage scan    230115 lbeamData 100x100 combined HCAL amp TCMT  250k events  nominal 0 3V    Average event rate in Run is  90Hz   Ecal three noisy layers  41  415 and 417    2007 07 06 02 59 Alex  Yoshi  AM muon run  HCAL TCMT voltage scan    330116 lbeamData 100x100 mu  combined HCAL amp TCMT  250k events  nominal 0 3V       CALICE eLogbook   Results for search of  20070701 Page 3 of 24       TCMT layer  7 is noisy   TCMT1 dark current distribution shows noisy spikes as below   HV Main Panel    monitoring Switch On   Off  Voltage  V  Current Lat Set  V  initiat  v     74 00  60   10  oo    73 2    1 Trey  01 00  0 00    67 2007 06 07 2007 06     Wi settings   file actions erable   disable RL churrels    Sve we wmr   enable disable    E  ONLY Ww OFF ONLY Lv OFF  W Crate ON   W Crate OFF  rense  KILL wl AJ 44   w7 j wi     ma   ws   wo    recover all trips    W3   mwe   mg tY daii   dap Hi values to File dune BL values  position                    File  http   ttfinfo desy de CALICEelog data 2007 27 06 07 2007 07 06T02 59 34 03 png    2007 07 06 03 48 Alex  osm  EM muon run  HCAL TCMT voltage scan    330120 ahcGain 100x100 combined HCAL amp TCMT  0 0 11  events  pee el ahcPmNoise 100x100 EE Lo HCAL amp TCMT  eer events  pu gere A ur O Pre    330
33.   Op 8  oo SP Of Of O  Zell SJ of FJ do  cof cof 5  oo    Fit performance in region  gt  sigma to avoid wafer gap effect and low energy noise    CALICE eLogbook   Results for search of  20070701 Page 17 of 20          2007 07 14 13 40 Allister  Beni  Beam Run  Mangi  poene eere TAE AA ce   2 2007 07 14 13 59 Beni updated AHC cfg  10x10  amp  20x20  hold and VCALIB  values    I tried to synchronize the LED hold and vcalib values in the 10x10 and 20x20 AHC cfg   Several vcalib settings were at a unreasonable low value for physics mode     TO DO   Test if new vcalib settings are appropiate  synchronize with 100x100 file  complete PM mode hold scans  old scans miss a lot of the modules due to the wrong vcalib settings                                                  AHC cfg  10x10 trigger      5 0 AHCAL8 38 38 36 38 19 17 33 61 28000 30500  5 3 AHCAL8 37 37 40 37 18 19 35 64 28000 30500  5 4 AHCAL8 35 36 39 35 17 13 34 60 28000 30500  5 5 AHCAL8 34 35 37 34 16 16 34 60 28000 30500  5 6 AHCAL8 33 34 28 33 15 16 33 61 28000 30500  5 7 AHCAL8 32 33 35 32 14 14 34 59 28000 30500  7 0 AHCAL8 31 32 34 31 13 15 34 60 28000 30500  9 0 AHCAL8 36 31 33 36 37 14 35 60 28000 30500  7 3 AHCAL 24 30 3024 36 15 33 60 27750 30500  7 4 AHCAL 2 29 6 2 35 14 35 60 28200 30500  12 7 AHCAL 1 28 14 1 34 20 33 60 28000 30500  7 7 AHCAL 30 27 23 30 33 15 34 60 28140 30500  9 1 AHCAL 29 26 26 29 32 13 33 59 28140 30500  9 2 AHCAL 28 25 10 28 12 15 34 62 42500 48000  9 3 AHCAL 27 24 38 27 31 17 33 61 2790
34.   Slot number   12  CRC Component   10 BE Trg    Label 0x01   0500000001   write  inet_pton status    DuplexSocket connect failed  retry 0 100  inet_pton status    DuplexSocket connect failed  retry 1 100  inet_pton status    DuplexSocket connect failed  retry 2 100  inet_pton status    DuplexSocket connect failed  retry 3 100  inet_pton status    DuplexSocket connect failed  retry 4 100  inet_pton status    DuplexSocket connect failed  retry 5 100  inet_pton status                        DuplexSocket connect failed  retry 99 100  inet_pton status          and then silence         CALICE eLogbook   Results for search of  20070701 Page 10 of 22       extracted from Log1183429747 out_icalice00    Since I have no idea at all what kind of communication cannot   be established I have decided to switch the Ecal off and leave the  System sunning with  non recorded  Crc Noise runs on both calice01 and  calice02  I ve got rates on both individual DAQs     Cheers   Roman    BTW   Although the DAQ was clearly at run 330050 the Gui display still said 330048     e Ck Ck Ck Ck Ck Ck Ck Ck CC CC CC C CC CC CC CC CC CC CC Ck Ck CC C CC CC Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck Ck ck ck k kx A amp     For future reference  the socket connection failed error is to one of the other   PCs in the system  It looks from the above like the first one tried failed  which  is the one to the ECAL  i e  calice01  Hence  i
35.   Uozumi LED run  400597 CmLed KNU fiber  peak  2007 03 21 21 35 e Uozumi  400599 CmLed deg     peak  2007 03 21 21 35 SC Uozumi LED run  400599 CmLed KNU fiber  peak  2007 03 21 21 35 S  Uozumi LED run  pee qe pev de  a  inter calibr  tion  2007 03 21 21 35 S  Uozumi LED run  CT ree cans  inter calibr  tion  2007 03 21 21 36 S  Uozumi LED run  ee isterz eatibel  inter calibr  tion  2007 03 21 21 36 S  Uozumi LED run  400603 PmLed LED KNU fiber 2 5V for  DUM AA oco wm TE ls  2007 03 21 21 36 SS Uozumi LED run  400605 CmLed KNU fiber  peak  2007 03 21 21 36 Ls  Uozumi  400605 CmLed deg      peak  2007 03 21 21 36 ES Uozumi LED run  400607 CmLed KNU fiber  peak  2007 03 21 21 36 S  Uozumi LED run    400607 CmLed KNU fiber  see lpe  peak    CALICE eLogbook   Results for search of  20070301 Page 33 of 76       2007 03 21 21 37 S  Uozumi LED run  400609 CmLed LED KNU fiber 2 5V for  peu ees AO AA E  2007 03 21 21 37 S  Uozumi LED run  400609 CmLed LED KNU fiber 2 25V for  E EE E listezestibejeion  2007 03 21 21 37 S  Uozumi LED run  400610 PmLed LED KNU fiber 2 5V for  CT E EL listezestisejeien  2007 03 21 21 37 S  Uozumi LED run  400610 PmLed LED KNU fiber Zi  OV  for  RR PA nes con PARA TL ls  2007 03 21 21 38 S  Uozumi LED run  400612 CmLed LED KNU fiber 2 5V to  peak  2007 03 21 21 38 S  Uozumi LED run  400612 CmLed LED KNU fiber 25V to  peak  2007 03 21 21 38 S  Uozumi LED run  400614 CmLed LED KNU fiber 25V to  peak  2007 03 21 21 38 S  Uozumi LED run  400614 CmLed LED KNU f
36.   We looked up in the 3 dim display that the run is with muons and it fills nicely the HCAL   defocussed    muons      The 2 dim display of the ECAL is attached     CALICE eLogbook   Results for search of  20070701 Page 11 of 19          ECAL leet   veut   chu   Panow   Dipin30   Cremichec   zen zogen   Response   Pesestan Te roectons                   1       i       4     i  f           i  ul       1  1       i     n  i  d Ld    hEcalYX    i  i       ipn    1    w  nehme Kiom   hicaba    nm                   Ka Le  e             ee 03 09 23 907 202 tors A 14 2007 br 130529 000   d       File  http   ttfinfo desy de CALICEelog data 2007 29 16 07 2007 07 16T05 06 23 00 jpg    2007 07 16 05 09 Oliver Hakan Valemuon run    ae  lll   lll          2007 07 16 08 34 Allister  Imad new shift  New shift   2007 07 16 08 35 Allister  Imad muon run    NN       NN    CALICE eLogbook   Results for search of  20070701 Page 12 of 19    completed  since  there s  no beam   0 42995 of run 0529     Beam stopped at  09  r event          2007 07 16 11 03 Allister  Imad calibration runs  EEIEEI ENTE O      3  330531 ahcCmNoise actually reached more  than 100Kevts since  the  split beamRun   flag wasn t turned off  3    30532 EncEmNoise                         s 19000 AA  330533 BhcCmLedVcalibScan                     M       2007 07 16 12 00 Allister  Imad muon run       stopped at 19Kevts due to beam preparation    2007 07 16 12 02 Allister  Imad calibration runs  D30535                   RhePmte
37.   boseframe tios __SPMs Digg      amp     pa  s   cabing         fibres  tempsensors cover plate no of persons          tour  Tc as vt D Daid F1  25 04 07    File  http   ttfinfo desy de CALICEelog data 2007 1 7 26 04 2007 04 26T08 52 43 00 pdf    2007 04 26 15 53 Sebastian modules 25 27 29 31 installed  Schaetzel    crate 1  slot 12       CALICE eLogbook   Results for search of  20070401 Page 13 of 14       An TU ER    voltage files   flchcaldaq01  caliceon cfg ahcal DACvalue mod25 A40 9 B44 2 txt  DACvalue mod27 A68 2 B45 8 txt   DACvalue mod29 A78 8 B77 3 txt   DACvalue mod31 A79 4 B79 5 txt    Da 2007 04 26 22 27 Beni module calibration    202401 ahcBeamStagefoar  6 e  modules scan module  only 25A shows   25 27 29 31 no current       The trigger rate is 230 Hz at the beginning of the run  This will change during HERA injection   which just started  I requested that the machine control restores the electron settings  after the injection   positions  cfg StagePositions dat  Run stopped while scanning position x 120  y 240 because local disk is full    Request sent to Roman to implement in copy script  test validity of dCache file after copy to dCache  and removal of local file    A 2007 04 26 22 29 Beni module 25 A has no current  When I set up the overnight module scan  I noticed that    module 25A shows no current    The powerline seems to be ok  at least I could measure the set high voltage at the connector of  the base board     For the moment I started the scan  never
38.   my guess is that it has  failed during the complete october run period     see also the enclosed pdf file    Detailed bug description        Paul is using the received mapping information from the slowControl     AhcSlowRunDataV1  print      Timestamp   1161820663   Thu Oct 26 01 57 43 2006   Modules   4 5 6 8 9 11 12 13 14 15 17 18 19 20 21 22 23016070100302010000000  0 0    this means  module  4 is in layer 1 and so on    But Sven thought that the DAQ asks for all 28 layer positions and not for the filtered list  based  on the above information  In the slowControl mapping he has only 23 entries  which are   standing for the active layer number  but no empty slots for the not equipped layers    Therefore his list has only 23 entries  which is different from the one the DAQ expect     Sven SC mapping   mb  112       WOWDIHDUONS A    14  10  ES  ceri   ES     LS    16     23    24  D    21  20    19  ld    VO OO    1 OY O1 4S Q 0 DO EB  OO t0 OO  12 0 O1 4   Q0 PO ES          NNN  DEISI       N  LA  l    Therefore  we have a mismatch in the mappings which leads to the result that we have not read the  module information for module 18  20 and 22  because the DAQ skips these values     In the first run period we have not run into this problem  because we have not left the empty slots  in the slowControl mapping free  which was also not 100  correct  But since it was homogenous  over the complete AHCAL weh have not thought about it     CALICE eLogbook   Results for search of  2
39.   new position    27 09 2007 11 39    CALICE eLogbook http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log       bas      e     EA       Logbook entry   CALICEelog data 2007 30 23 07    Ed oem Allister  Robert  Riccardo calibration runs    330830 shcCmNoise  e 20000  330832 ahcCmLedVcalibScan       Logbook entry   CALICEelog data 2007 30 23 07    A BEN Allister  Robert  Riccardo position check             Logbook entry   CALICEelog data 2007 30 23 07       E E Allister  Robert  Riccardo stability of ECAL response  update   gaussian fit mean of hEcalEnergy vs Run Number  i e  time   error bars are actually the    fit sigma     13 of 28 27 09 2007 11 39    CALICE eLogbook    14 of 28    http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log       Stability of ECAL Response         rJ  2  E       E      E  z     D       Run Number       Logbook entry   CALICEelog data 2007 30 23 07  23 07 2007  02 38    Allister  Robert  Riccardo calibration runs    Logbook entry   CALICEelog data 2007 30 23 07  23 07 2007  02 29    Rund    Guassian Fit for recent electron    Allister  Robert  Riccardo  runs     18 2263 150 6 6 65                Logbook entry   CALICEelog data 2007 30 23 07    23 07 2007  02 25       Allister  Robert  Riccardo       trigger rate 4 7K    Logbook entry   CALICEelog data 2007 30 23 07    23 07 2007 slow conversion of data files from  02 22 bin to slcio  We notice a significant delay in the conversion of bin files into lcio 
40.   nohup   p_comb pl  before   the command  less nohup out   After disscused with Fab  the latest id for the     cp comb pl is killed  After that we checked the data transfer of bin files to the grid   All files seems to be transfered properly  We will check agein after having started a  new run     2007 07 14 06 37 Hakan Oliver Wenti30GeV       2007 07 14 07 36 Hakan Oliver Went 25 GeV       TRIM6 is off    2007 07 14 07 38 Hakan Oliver Wenb  agy data to the GRID    Due to the start of a second copy script  as mentioned before  the original copy script was not  able to sent files to the GRID  because the logical filenames were already registered   Therfore  we deleted the files Run330428 000 bin to Run330428 004 bin from the GRID by  using the command    lcg del   vo calice  a lfn  grid calice tb cern native dat Run330428 000 bin  up to Run330428 004     After that the files are copied again by the original copy script  The output in the file nohup out  shows now copying of the files Run330428 xxx bin again  An lfc 1s on the directory shows  all the Run330428 xxx bin files in  grid calice tb cern native dat     The files Run330429 xxx bin are also being copied to the GRID now     2007 07 14 08 03 Hakan Oliver Wenbanversion of data files from bin to slcio    CALICE eLogbook   Results for search of  20070701 Page 16 of 20       Due to the problems with the copying of the bin files  see below  the converter of data files from  bin to slcio might not work properlt anymore  We che
41.   pion run  Nanda       Pb    DAQ rate  70 Hz   3100 pps in 10x10 trigger coincidence    stopped since there is no beam    2007 07 14 18 35 Nanda hold scans    330447 ahcPmLedHqsast  n changed  sync on vcalib  settings       preliminary results so far     no signal at all in modules   1 2 12 20 23 24 25 26 27 28 29 30 31 34 36      will try to reanimate CMBs    successfully determined holdvalues   module    Id value  2    ES ES ES    ES ES ES ES    ON ooo of  gt     Beo wo  To E    N       E    Of OF da EN  Hs  OM  DPD  O    o o NrermrrrTrqmre   CPS  oO  RP nie  Of cof BS  a  a  oc  K  OY DL DT DY OY  DOJ Pop dof co       3  2007 07 14 18 50 George  Tao  pion run  Nanda       Pb    DAQ rate  80 Hz   8500 pps in 10x10 trigger coincidence    mean  1220 mip  sigma  97 mip    2007 07 14 19 40 George  Tao  pion run  Nanda       Pb    DAQ rate  87 Hz   10k pps in 10x10 trigger coincidence    CALICE eLogbook   Results for search of  20070701 Page 20 of 20       mean  1891 mip  sigma  131 mip          2007 07 14 20 00 Beni cleanup of data00    I removed runs 330000 to 330261 from data00 with Roman s check and cleanup scripts     2007 07 14 20 21 Beni changed data disk    I changed the disk to data00       changed data link on calice00     updated runNumber txt     changed data link in online monitor    changed cp comb pl    2007 07 14 20 33 George  Tao  pion run  Nanda    Pb       DAQ rate  80 Hz   6 9k pps in 10x10 trigger coincidence    mean  2294 mip  Sigma  165 mip    2007 07 14 
42.   rather than  v 21 as probably  intended  This can be easily verified using  e g     icalice00   example 330415    which prints out the DaqRunStart information  This means the trigger required  only a single hit on input 13 rather than the 10x10 coincidence  I don t recall  what input 13 is but I assume it is not beam related  in which case no hits would  have been seen     Both runs are therefore junk   I guess this happened because  v 13 is the default  I have just recompiled    runStart to have  v 21 as the default so this should not happen again   PaulD    2007 07 13 22 13 Marie Cruz  Tao   Nanda    CALICE eLogbook   Results for search of  20070701 Page 13 of 20       DAQ rate  80 Hz    4000 pps in 10x10 trigger coincidence    stage at 60 0    2007 07 13 22 24 rising current    Current in HV channel 25  amp  16   module 15a  amp  15b is slowly  but steadiliy increasing since  the afternoon                  200    es xn d    16 00 EH 2 QU 22300  H  2007 13  OF  20  13 07 2007 13  OF  2007    File  http    ttfinfo  den delCALICFelog data 2007 28 134 07 2007 07 13T22 24 03 00 jpg    2007 07 13 23 03 Marie Cruz  Tao  electron run  Nanda       Pb    DAQ rate  50 Hz    940 pps in 10x10 trigger coincidence    stage at 60 0    2007 07 14 00 33 Hakan Oliver Went 25GeV    8 MM LL  Pb    330422 beamData  10x10  25GeV 6mm 205k  Pb events    Stopped intial run due to high looking current values  channel 28  up to 800uA   Seemed to rectify  itself        2007 07 14 01 35 SES    330423
43.   the bin files   to the grid seems to be going well nevertheless    Current bin file uploaded  330826 004   Current lcio file uploaded  330820       Allister  Robert  Riccardo    The upload of    27 09 2007 11 39    CALICE eLogbook    15 of 28    http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log       Logbook entry   CALICEelog data 2007 30 23 07    23 07 2007  01 03    See low channels noisy in slot 17 for HCAL  Data  12GeV e  run       Allister  Robert  Riccardo       H        D   D           1       Logbook entry   CALICEelog data 2007 30 23 07    23 07 2007  01 00    RMS for pedestals in TCMT looks better compared to the reference map       Allister  Robert  Riccardo    Data  12GeV e  run    27 09 2007 11 39    CALICE eLogbook    16 of 28    http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log          LJ  LI  ae  LI  cx  E  D  er  RH        Logbook entry   CALICEelog data 2007 30 23 07    23 07 2007  00 58    330826 beamData 10x10  10GeV e  520  air air 2 29   6 0 252kE vents    10x10 trigger rate  9 65 kpps   DAQ rate  92 Hz   Collimator settings   COLLO1 COLLO5 COLLO6 COLL10 COLL11     5mm  COLL02 COLL03 COLL08     5mm       Allister  Robert  Riccardo electron run                                     Logbook entry   CALICEelog data 2007 30 23 07  23 07 2007  00 51    Beam centered in Y  see zoomed boxes  but not in X  where we observe  average values of 40 50 units    Allister  Robert  Riccardo    Data  12GeV e  run    27 
44.   the current in HV channel 62   module 36B is constantly rising during the last 8 hours   from  60 to  85   module 36A was stable    CALICE eLogbook   Results for search of  20070701 Page 5 of 22       AR ee e   LP    40  oe E TT om    nap NER ri Prout TE zT i Ti d  M NT  13700 15 00 17300 13 00  POF 2007 17 07  2007 l  07 200  17 07 2007    File  http   ttfinfo desy de CALICEelog data 2007 29 17 07 2007 07 17T20 02 02 00 jpg    aee ez pa c edt    2007 07 18 00 03 Nanda Manqi Valemarbtakan Nicola George    i   i ef oP fF oP oP  i   8 dq O oP oP  i   B ef fF fF PoP op    beamData  100x100   80Gev mu 555 air air N A 15k  events                         2007 07 18 02 09 Hakan Nicola GeoatieGain  330618 ahcGain  2007 07 18 02 35 Hakan Nicola Geomaeon run    eamData  100x100   80Gev mu 555 air air N A split  beam  file in  250 kevt    D   beamData  100x100   80Gev mu 555 air air N A split  beam  file in  250 kevt   ppp po    2007 07 18 03 11 how to check the CMB power supply  The colors of the CMB LV button may be misleading  therfore the CMB power has to be checked in the  CMB panel  Go to the slowcontrol  open the CMB menue  switch to the detailed view of  a random CMB and check if the Power_12V is larger than 12V  If the voltage is lower   especially if it is between 6 to 7 V please call the HCAL expert  16 5554        CALICE eLogbook   Results for search of  20070701 Page 6 of 22    CME Main 1234506789 10  11 12 13 14 15 16 17 18 19 20 21 22    Module 7   CMB 5     Pulser  Set
45.   y   9       0 25 50 100  oen E GeV       0 25 50 100  oil E GeV    File  http   ttfinfo desy de CALICEelog data 2007 31 30 07 2007 07 30T17 33 54 00 pdf    non linearity    2007 07 30 18 21 Alex Catherine 45 GeV pion run    CALICE eLogbook   Results for search of  20070701 Page 12 of 20    331175 lbeamData  10x10    45GeV pi  515 4mm Air 0 1 20 deg   Pb changed   trim06  from 15   to 10       331174 lbeamData  10x10  45GeV loi  515 4mm Air EL 20 deg   Pb junk    10x10 trigger rate  1 4 kpps  DAQ rate  65 Hz    2007 07 30 19 16 Alex Catherine   80 GeV    331176 lbeamData  10x10   80GeV pi  510 20 deg   ES changed   trim06  from 15   to 10    Noisy ECAL layer        2007 07 30 20 04 GEES 15 GeV electron RI    lbeamData  10x10 15GeV 536 Air 20 deg   Es junk  k  331180  em pem    Somehow we managed to initiate a radiation alarm by loading file 536  the previous was 510   Three  bending magnets were off  With the friendly help of Roman we managed to reset it  by   calling the control room    But now having loaded file 536 and we do not see anything in the tracker in the online monitor   Are we totally off position   We played a bit with the settings for TRIMO5 and TRIMO6  also loading yesterday s settings  but  with no effect     ECAL was off before run 331182  turned it on     Is the Tracker off  Yes Tracker had tripped  Ramped up Channels 36 37 38   Roman takes over the ECAL slow control via VNC to keep an eye on them  since currents are still  creeping up                     
46.  03 13 08 27 S  Uozumi    400368 LED LED fiber direct  LED  run for  L1 L2 with  physics mode    same LED pulse with previous run for inter calibration             2007 03 13 08 27 S  Uozumi LED run    400368 LED LED fiber direct  LED  run for  L1 L2 with  physics mode    same LED pulse with previous run for inter calibration                2007 03 13 08 27 S  Uozumi LED run  AA A AA E AA  run  2007 03 13 08 27 S  Uozumi LED run  pepe E ZAPPA s   run  2007 03 13 08 29 S  Uozumi 3GeV stage scan  400370 beam 3GeV fiber 2 5V scanned overnight  direct scan  with detector scan  36point  central part scan  9points  central injection lpointx10times   2007 03 13 08 29 S  Uozumi 3GeV stage scan       400370 beam 3GeV fiber 2 5V scanned overnight  direct scan    with detector scan  36point  central part scan  9points  central injection lpointx10times     CALICE eLogbook   Results for search of  20070301 Page 3 of 76       2007 03 13 08 40 S  Uozumi 3GeV stage scan    400371 beam 3GeV fiber 2  5V scanned overnight  direct scan  Q       with detector scan 6point  plus central part scan points  plus central  EE    2007 03 13 08 40 S  Uozumi 3GeV stage scan    400371 beam 3GeV fiber 2 5V scanned overnight  direct scan       with detector scan  36point    plus central part scan  9points  plus central  injection  lpointx10times     2007 03 13 08 41 S  Uozumi 3GeV stage scan    400372 beam 3GeV fiber 2 5V scanned overnight  direct scan       with detector scan  36point  plus central part scan
47.  03 20 20 41 S  Uozumi et al  comissioning run for 3rd configuration  KNU fiber  2 5V   15  42  comissioninp  position  run  of strip 1  on X layer  2007 03 20 20 41 S  Uozumi et al  comissioning run for 3rd configuration    KNU fiber  2 5v   15 42  comissioninp  position  run  of strip 1  on X layer    2007 03 20 21 04 S  Uozumi et al  comissioning run for 3rd configuration                KNU fiber  2 5V  2  20  comissioninp  position Wun  of strip 1  on Y layer    comissioning run for    S  Uozumi et al  3rd configuration    KNU fiber  2 5V  2  20  comissioning  position Wun  of strip 1  on Y layer    S  Uozumi et al  comissioning run for 3rd configuration    3GeV KNU fiber  2 5V scanned 36  points  detector  scan for  comissioninp    The voltage for KNU module is set to 1 2V above OV        2007 03 20 21 30 S  Uozumi et al  comissioning run for 3rd configuration       36   points  detector  scan for  comissioninp    The voltage for KNU module is set to 1 2V above OV   E 2007 03 21 00 40 RomanTrigg Trigger Info for ScEcal Running  Dear Colleagues    I have updated the software and the database such that the trigger  information can be extracted from the converted files  see also    entry in the calice s w list      Since it was not clear when the triggers were connected how I have  queried the database and created two files     1  One containing the run start and stop times    CALICE eLogbook   Results for search of  20070301 Page 26 of 76       2  All the  non zero  trigge
48.  10      100 M  METROS o A NEN   a iW      s0 wj EN   oc ees 715   tr    RES zen  EISEN aa e  mamm JD MY ramp UP DOWN   Vum om j  vum   mmm   ouer    umm zem 1   apra mas   ER GUI Z Vile alo T rua Y labia NU ale        sw j wm   mu   exe   diete  i                            File  http   ttfinfo desy de CALICEelog data 2007 30 29 07 2007 07 29T15 12 02 00 jpg    2007 07 29 15 24 Beni  Martin  junk run  Ariane       331106 lbeamData  10x10  12 GeV   pi  524 8mm Air 1 3 20deg  Pb junk    beam slightly off  for a change    CALICE eLogbook   Results for search of  20070701 Page 13 of 16       jg u nk    2007 07 29 15 33 Beni  Martin  pi  12 GeV  Ariane       Ph    1600 pps  DAQ rate  60 Hz  150k events    2007 07 29 16 34 Catherine  Tom   15 GeV pi   Alex    Data  10x10  15 GeV   pi  521 8mm Air 20deg   Pb trim05 5   trim06  8   beamData  15 GeV   pi  521 8mm Air 20deg    Pb junk  beamDat  15 GeV   pi  8mm A 20deg   E ES GEM RENTE  beamData  15 GeV   pi  8mm A 20deg   pe rou   at 15     20deg   ET    ata 15   A       P    31108 EE                               2007 07 29 17 37 Catherine  Tom   20GeV pi  run  Alex    ME REX pee qe      Nu TT EN LO RH a  Pb    noisy ECAL layer 9    20deg   trim05 0   trim06  3       2007 07 29 17 55 Alex Run Monitor Pedestal Update    I updated the pedestal data for the run monitor for HCAL and TCMT  Pedestal data was taken from  ahcPmNoise run 331108    2007 07 29 18 09 Catherine  Tom  AHCAL Led Ped    3 HCAL DAC  settings  reload    31117  AHCAL 
49.  10GeV e  626 Air 6mm  2 0 0 20deg   Pb 100k    2007 08 19 00 10 daniel electron analysis    similar plots for e  runs at different angles  amp  positions                       File  http   ttfinfo desy de CALICEelog data 2007 33 16 08 2007 08 19T00 10 50 00 ps    CALICE eLogbook   Results for search of  20070801 Page 2 of 22       2007 08 19 02 53 Allister HCAL pion response  combined runs   We plot the response of the HCAL to pions  obtained by gaussian fit of hHcalEnergy  as a function  of beam energy for the combined ECAL HCAL runs  33xxxx series   The linear fit is constrained  to the origin  Some points deviate from the linear fit due to changes in HV settings     The errors bars are actually the gaussian fit sigma     HCAL Response       Fit Mean  mips        0 20 40 60 80 100 120 140 160 180  Beam Energy  GeV     File  http   ttfinfo desy de CALICEelog data 2007 33 19 08 2007 08 19T02 53 51 00 png    2007 08 19 02 58 Allister HCAL pion response stability  combined runs   We plot the HCAL response to pions  obtained from gaussian fit of hHcalEnergy  vs the run number   i e  time   The error bars are actually the gaussian fit sigma   Stability of HCAL Response        Fit Mean  mips   3  S       KA  e  o   Xx KH aei Les wi    200 400 600 800 1000 1200 1400 1600  Run Number  File  http   ttfinfo desy de CALICEelog data 2007 33 19 08 2007 08 19T02 58 26 00 png    2007 08 19 03 37 Allister ECAL response for different beam positions angle  We plot the HCAL response to pions 
50.  13 18 30 Marie Cruz  Tao  pion run  Nanda       Pb    DAQ rate  70 Hz  43000 pps in 10x10 trigger coincidence    stage at 60 0    Gaus fit in online monitoring panel     mean   2015 mip   sigma   109 mip     gt  sigma mean   5 4    2007 07 13 19 42 Marie Cruz  Tao  pion run    Nanda       Pb    DAQ rate  80 Hz    3700 pps in 10x10 trigger coincidence    stage at 60 0    Gaus fit in online monitoring panel     mean   986 mip   sigma   91 mip       sigma mean   9 2    2007 07 13 20 47 Marie Cruz  Tao  electron run    Nanda       Pb    DAQ rate  60 Hz    3400 pps in 10x10 trigger coincidence    CALICE eLogbook   Results for search of  20070701 Page 12 of 20       stage at 60 0    A 2007 07 13 21 07 caliceon calicehcaisn  atpeaksiin HCAL   CALICE user     HCAL currents have been instable for almost all channels during this evening           2  9 c c ra      sm          z052 21300 21708 21 16  P zou  13  W   200  13 07  S   13  E 2007    File  http   Ittfinfo  Soen delCALICFelog data 2007 28 13  07 2007 07 13T21 07  02  00 jpg          2007 07 13 21 44 Marie Cruz  Tao   Nanda    Pb    Pb       DAQ rate  75 Hz    4100 pps in 10x10 trigger coincidence    stage at 60 0    Do not see anything in the online monitoring       Panic  close DAQ  amp  start again    e Ck ck ck ck ck ck ck ck ck ck ck ckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckckck ck    These two runs were taken with beamData  v 13
51.  17 of 19         ECal Linearity Check  Low Ener          24000   E E  Stage Position   33500    00     E    0  3   Se    3 5      3000 Ss DAL  t  2500     i  En s  2000   s  1500        E e  1000               8 10 12 14 16 18 20 22 24 26  Beam Energy  GeV     File  http   ttfinfo desy de CALICEelog data 2007 29 22 07 2007 07 22T22 16 10 04 png    2007 07 22 22 32 Hengne Oliver Maelgctron run      load beam file 522  set cherenkov pressure to 0 79 bar        close collinators to   COLLO1 COLLO5 COLLO6 COLL10 COLL11    5 mm     trigger rate reduced to   7 8kpps    eee eer A qp    Trigger rate 7 8 kpps  DAQ rate  90Hz             2007 07 22 22 55 Oliver HCAL Response    The plot shows the response of the HCAL in MIPs for pions beams of different energies  The HCAL    response is given by the mean of a gaussian fit of hEcalEnergy spectrum  The error bars  indicate the standard deviation of this fit     Compared to the plot form July 20th 2007 a few more data samples are added        HCAL response   a 2400    2200 Lu   gt  2000 070716  E mt mE  ke 1600 070719   o 1400 070719 2  070720   3 1200    gt  1000   800   600   400   200   0    0 20 40 60 80 100  beam energy   GeV    File  http   ttfinfo desy de CALICEelog data 2007 29 20 07 2007 07 22T22 55 27 00 png    2007 07 22 23 08 Hengne Oliver Maelgctron run    AA PAE mec    Trigger rate 7 43 kpps  DAQ rate  90Hz       CALICE eLogbook   Results for search of  20070701 Page 18 of 19       2007 07 22 23 28 Oliver  Hengne  Noisy lay
52.  20 21 22  Pulser LED enable disable  012345678 981011    RSS    On 0 0 A yn                 li le    en bm Le    IS   ren es e    3          gt   i  ta                  ken    puto           E     LLLI  RIES    GEZ EAKASEKRCRKRHCRG CA    PANEL toelog      File  http   ttfinfo desy de CALICEelog data 2007 29 18 07 2007 07 18T03 17 02 00 jpg       2007 07 18 04 41 calibration runs    CMB 12V readings are all above 12V  all 13 13 5V            2007 07 18 05 51 Hakan Nicola Geomgeon run    WW TI TT  i     A II i i    2007 07 18 08 00 Hakan Nicola Geocgdibration runs    CMB 12V readings are all above 12V  all 13 13 5V               2007 07 18 08 12 Hakan Nicola Geo amp zdety run    ECAL safety run started    2007 07 18 10 15 Beni HV off    Sven will work on the CMB power alarm  I switched the HCAL HV to 10 V for this work     CALICE eLogbook   Results for search of  20070701 Page 8 of 22       2007 07 18 13 12 Sven CMB power LEDs    The problem with the power LEDs on the CMB panel is solved  Now one can exactly see  which power  supply is on or off     2007 07 18 13 14 Sven problems with eLogbook    The eLogbook group changed a router from 1 GBit to 10 GBit  Hence there were problems on the  ethernet   which should be gone now      amp  2007 07 18 13 19 set DAQ to 10x10 trigger    Switching back to e pi trigger     removed 100x100 front trigger plate   changed AHC cfg link to AHC CERNO7 beamhold 10x10 cfg  changed sequenceDelay in EmcConfiguration hh to 13  build runner   DAQ shu
53.  20070701 Page 2 of 19       Morin    330715     eamData   10x10    The rate was increasing so we close colll to  4 4 mm   trigger rate 7 8K          2007 07 20 10 35 kieffer  Fabbri   Morin   230715  beamData   10x10    10cev  e _____ 520__  air     fair _  2 6  0 3    9K        R t d th K t  un stopped wi  9K events  2007 07 20 10 36 kieffer  Fabbri   Morin  Events  2007 07 20 11 42 kieffer  Fabbri   Morin  KEvents   amp  2007 07 20 12 08 Roman Freeing Diskspace  Hi there     I am about to free  800 GByte diskspace on data0l     The action is ongoing and should be ready in  30 Min         data01 is available to store data  The link can be switched now           Cheers   Roman  2007 07 20 12 09 kieffer  Fabbri   Morin   2330718  beamData   10x10    i5Gev  e     521      air fair X  1 35  0 3            Garbage    2007 07 20 12 10 kieffer  Fabbri   Morin  pepe ue ee ce is  Events  2007 07 20 12 37 kieffer  Fabbri   Morin  ee ee E e ree  Events  2007 07 20 13 26 kieffer  Fabbri   Morin       330721 lbeamData   10x10  20GeV 523 air air 0 76 0 3 150k  Events    trigger rate 9 8K  many pion contamination    2007 07 20 14 21 kieffer  Fabbri   Morin   330723   LL1a8hecmNoise   ST EN   330724 l ahePmvoise Fe 20000 V TEN  2007 07 20 15 04 Beni changed data disk    changed data location to data01    changed link on calice00    changed link in online monitor    2007 07 20 15 14 Beni    330725 ahcBeamHo l   amp  amp 8n unknown  unknown  N A N A N A N A 0  3 summer  studenst    CALICE eL
54.  20070801 Page 14 of 22        Sleft right   gt    gt  index l  cm    File  http   ttfinfo desy de CALICEelog data 2007 34 20 08 2007 08 20T11 34 03 00 gif    2007 08 20 12 48 Jamie  Pavel   Beni  Erika    Compare of the HCAL energy distribution for 60GeV pion    hHcalEnergy hHcalEnergy   hHcalEnergy305    Entries 92572   Entries 141380  1239    Mean 1250  RMS 448 4                                  0 ae pant X Y N i  0 1000 2000 3000 4000 5000  energy per event  mip    File  http   ttfinfo desy de CALICEelog data 2007 34 20 08 2007 08 20T 12 48 41 01 jpg    2007 08 20 12 48 Jamie  Pavel  pion run  Beni       i  Pb    CALICE eLogbook   Results for search of  20070801 Page 15 of 22           Pb    2007 08 20 13 34 Jamie  Pavel   Ben    350307 lbeamData  10x10  30GeV Git 614 4 ram  Pb  P    air  b  Pb    2007 08 20 15 54 Jamie  Pavel  e   run  Beni    Pb  Pb    garbage   lost  TCMT dac  settings   reload       0 04   04    101k   deg       2007 08 20 16 19 Daniel  e  runs  Anne Marie  Petr    ui       110k  30  degs  INN ae          2007 08 20 19 44 Daniel  junk run  Anne Marie  Petr  2007 08 20 20 26 Daniel  e  run    Anne Marie  Petr    350319 was stopped because beam at  90 in x in tracking chamber 3   Move it back to 0 for 350320           2007 08 20 21 05 Daniel  Calibration run  Anne Marie  Petr   350321  JjaheGain   A  50324 jahcPmLedVcalibScan           50325 jahcCmLedVcalibScan         2007 08 20 21 51 Daniel  junk run  Anne Marie  Petr  2007 08 20 21 52 Daniel  pi     
55.  28 30  cont d   Schaetzel   2007 04 16 12 12 Sebastian MIP calibration modules 24 26 28 30  cont d   Schaetzel    like 202365 but with cfg StagePositions202368 dat    stopped at position 120  240  disk full    2007 04 16 13 17 Doerte David assembling status    Tiles for module 33 are tested but there are not enough good SiPMs for that module  Therefore the  tiles for the next module are now being tested     The problems with the new type of cable and the stripping machine were solved on Friday  We put    in new knifes  a new spring and wrote a new program  Assembling could continue again     Meanwhile we started to disassemble cassettes   1 and 2  All tiles with their current and in many  cases bad SiPMs will be shipped to Moscow where the colleagues will assort tiles with  in many cases new SiPMs for new modules   1 and 2  Later we will reassemble the first  cassettes at DESY     cassettes work in progress  35  2    cassettes ready  3o    CALICE eLogbook   Results for search of  20070401 Page 5 of 14       status of production  ew  cassette amen b  ng HAL   FLO           module assembling steps staff  1 2 3 4 s g 7 H 9    8 basetrame ities SN FRA plate   pa  s    cabing fibres  tempsensors cover piste no of persons    2    3 33          siot lc de vt D Daid F1  600107    File  http   ttfinfo desy de CALICEelog data 2007 16 16 04 2007 04 16T13 17 42 00 paf          2007 04 16 15 12 Sebastian MIP calibration modules 24 26 28 30  cont d   Schaetzel   202369 ahcBeamStageScan v 
56.  3   3 200k  Pb events       should be       330508    fboeamData 10x10  15GeV 536 air 6mm 1 34  3  3 200k  Pb events    Since    4mmPb absorber   air target   is the pi  setup  for e   it should be   air  absorber 6mmPb target       The runs with this error are from 330478   330508  taking on 15 07 2007     We checked that it is only a mistake in the elog  the paperbased logbook has no problem  and also  no problem of the data file     Please be careful of this mistake     2007 07 21 23 58 Oliver Allister Hengnergy resolutions electron runs    Electron Runs    pu EE  GeV  energy mean  mip  width of mean  mip    Sigma Mean   Es          30781 OTS OS SB  2007 07 22 01 00 Nicola  Riccardo  Fit result of ECAL energy per event spectrum  Robert    330784 Mean   1890 sigma   154  330785 Mean   2270 sigma   190    330786 Mean   2351 sigma   228  330787 Mean   2894 signa   248       2007 07 22 08 24 Morin  Brient  25 GeV pions    CALICE eLogbook   Results for search of  20070701 Page 14 of 19       Lutz    Pb       position checked  is ok  trigger rate 3 9k  DAQ rate 75 Hz    2007 07 22 10 27 Morin  Brient   Lutz       330799 beamData  10x10  18GeV 8mm 25K  Pb events    absorber change to 8mm Pb    2007 07 22 10 37 Morin  Brient  pion run 18 GeV  Lutz    330800 lbeamData  10x10           2007 07 22 10 42 Morin  Brient   Lutz    330801 beamData  10x10  A 2007 07 22 10 53 Morin  Brient  cherenkov pressure was to high  Lutz    We looked into the particle ID plots for the last run and re
57.  55 Erika Francois Sacha Jean Yves    330143 lbeamData  100x100  80GeV mu  510 Air air off good  muon  run  HV 0        DAQ rate  86 Hz  10x10 rate 9200 pps  100x100 rate 16500 pps    C8 set to  30  JAW1    27  JAW2  after 140000 events because 20x20 rate   13000 pps   gt  11000     now    10x10 rate 8200 pps  20x20 rate 11000 pps  100x100 rate 15300 pps    stopped after 170 000 events  No beam from CPS    2007 07 06 20 34 Erika Francois Sacha Jean Yves    330144 lbeamData  100x100  80GeV mu  510 Air air off good  muon  run  HV 0        DAQ rate  83 Hz   10x10 rate 8100 pps  20x20 rate 11100 pps  100x100 rate 16000 pps    Layers 1  15 and 17 of the ECAL are still noisy    CALICE eLogbook   Results for search of  20070701 Page 8 of 24       hEcalMitsPerLayor nmn mt  I Es E iam  a I mp a  E  oe          mm  H f a    u 1  m    ni  gt     e pop Lp EMEN T TER TL TEYSTIPENYS      PEcalEmemgyPerLayer          Se         t  A ui  Pe    al d         al  d D D   Bd aca BEL gsca c  i T   7    dd  File  http   ttfinfo desy de CALICEelog data 2007 27 06 07 2007 07 06T20 34 45 02 eps  2007 07 06 21 21 Francois Sacha Jean Yves       330145 lbeamData  100x100  80GeV mu  510 Air air off good  muon  run  HV 0     DAQ rate  86 Hz   10x10 rate 8800 pps  20x20 rate 11600 pps  100x100 rate 16800 pps    250k events    2007 07 06 22 09 Francois Sacha Jean Yves       330146 lbeamData  100x100  80GeV mu  510 Air air off good  muon  run  HV 0     DAQ rate  89 Hz   10x10 rate 8800 pps  20x20 rate 11500 
58.  9 10 scan  with tungsten  pure BEER Xe pec   direct  2007 03 14 11 24 S  Uozumi 6GeV run  36 9 10 scan  with tungsten  eee Pee Been PA e EL  direct  2007 03 14 12 36 S  Uozumi 5GeV run  36 9 10 scan  with tungsten  AAA A AAA  direct  2007 03 14 12 36 S  Uozumi 5GeV run  36 9 10 scan  with tungsten  pue ge qe A A A AA poo  direct  2007 03 14 13 56 S  Uozumi 4GeV run  36 9 10 scan  with tungsten  pee  gt  A Jeep E EE  direct  2007 03 14 13 56 S  Uozumi 4GeV run  36 9 10 scan  with tungsten  oben mo  D  E          direct  2007 03 14 13 56 S  Uozumi Run summary of 2nd round energy position scan with    fiber direct module  This is a run summary of 2nd round energy and position scan run with fiber direct module  after improving the drift chamber efficiency   All data contains data with 55 points   36 points for center of each strip of X and Y layers   9 points for central part scan   1 points  x 10 times for central injection    points have 10k events of beam data and 1k events of pedestal     n    1 GeV     Run400422  Run400423  2 GeV     Run400425  Run400426  3 GeV     Run400427  Run400428  4 GeV     Run400429  Run400438  5 GeV     Run400432  Run400437  6 GeV     Run400434  Run400435  2007 03 14 14 25 Doerte David assembling status    Tiles for cassettes no  33 to 36 were advertised to be send to DESY today  Usual time for shipping  and German custom amounts nearly 1 week  Then the tile tests at DESY will follow     Now the change from fine to coarse structure is well done   Du
59.  9 I Bam OFRI  oc 5 8 41 1 Hu etarga e icto ag    n ALL iners   sos T   e Ram WEE sue   uxo   com    _ enable   gea   gt  eer Soren na Experts  hee EN   E ro 46 395 DR 00 000  3 0  amp  Rand OFFI  HvCrae ON  HY Crate Ol     amp  Bamd OFF  AAN  TIEN ee   o W Bed ori   recover adios   HVI J  Hv6   Hv9  Vde  ME   Ame vda tom   ar ALL vitaes ote  _        File  http   ttfinfo desy de CALICEelog data 2007 20 14 05 2007 05 14T09 21 42 03 png    2007 05 14 10 01 Sebastian 30 HCAL modules in stack  Schaetzel    10 more modules are in the stack for a total of 30  Thanks to Uwe  Gert and colleagues     tack position      NN  oloni elol et oal oalon         f o NIe onon    NIN  OY          2007 05 15 08 32 Sebastian DAC loading broken  Schaetzel    We have a clear show stopper here  All LED work is delayed until the system is working     read back of DAC values loaded at startUp    CALICE eLogbook   Results for search of  20070501 Page 8 of 9       slot 5  FEO as an example                       Slot number   9  CRC Component   0   FEO  Label   1   0x01   0b00000001   write  AhcVfeStartUpData   print     Verification data   2863289685   0xaaaa5555   0b10101010 10101010 01010101 01010101  HAB 0 DACs   61  56 St 56 6l 56 30 1  30 25 30 30 30 25 40 35 35 20  HAB 1 DACs   25  Wb  5 Ee 087 SE SO 205  39  E 5 56 30 40 35 35 20 20  HAB 2 DACs   35 30 30 35 40 40 46 40 46 51 40 40 46 46 51 35 35 35  HAB 3 DACs   56 46 40 51 51 51 46 51 40 40 35 30 30 61 56 40 56 51  HAB 4 DACs   46 51 40 56 46 
60.  9 points to center area of detector   10 times of center  injection         This energy histograms is from 41th point which is center position among center area scan     CALICE eLogbook   Results for search of  20070301 Page 57 of 76       It means It is almost center injection           1149907  3  1    9 1378  2 175       z   TW M   MM 1000 1500 Lg       File  http   ttfinfo desy de CALICEelog data 2007 12 25 03 2007 03 25T11 00 36 00 JPG    2007 03 25 11 46 S  Uozumi MIP calibration    400784  beam TIL 3GeV KNU fiber  2 5V scanned 36  points  scan for  IP    2007 03 25 11 46 S  Uozumi MIP calibration          400784  beam TIL 3GeV KNU fiber  2 5V scanned 36  points  scan for  IP  calibration    2007 03 25 11 50 S  Uozumi ntuples for analysis     ve almost finished to copy beamtest ntuples to Japan    Temporary the data are put on my desktop at Shinshu  so if you are  interested in looking the data   please let me know    Then I will make your temporary account on my desktop    After flcscecalana arrives to Japan  it will be the data storage place and  analysis server      The data are found at   pcat102 shinshu u ac jp  BT_data newtrees v1_0  The root files contain ntuple which name of event_data   Here I will briefly write about available parameters   runNumber eventNumber     not necessary to explain   config     shows which scanning point we are    int  config 3 will give position number of the stage scan   starts from zero        date  time microsec     format of date
61.  AHCAL 23 21 21 23 21 32 33 59 28000 31000  1    2 4 PMT  1    5 1 AHCAL 22 20 32 22 20 32 33 59 28050 31000  5 2 AHCAL 21 19 24 21 19 32 34 59 28000 31000  5 3 AHCAL 20 18 31 20 18 32 33 59 27950 31000  5 4 AHCAL 19 17 27 19 17 32 34 59 28050 31000  5 6 AHCAL 18 16 15 18 16 32 34 59 42050 31000  5 7 AHCAL 17 15 25 17 15 32 33 59 28200 31000  1    7 0 AHCAL 16 14 29 16 17 32 33 59 28000 31000  7 1 AHCAL 15 13 3 15 17 32 40 56 42050 31000  7 2 AHCAL 14 12 8 14 17 32 31 56 41600 31000  7 4 AHCAL 13 11 18 13 17 32 35 56 41730 31000  7 5 AHCAL 12 10 0 12 17 32 34 56 41720 31000  7 6 AHCAL 11 9 13 11 17 32 34 56 41800 31000  7 7 AHCAL 10 8 2 14 17 32 36 56 42000 31000  1    9 4 AHCAL 9 719 9 17 32 33 56 42000 31000  S 1 AHCAL 8 6 11 12 17 32 32 56 41750 31000                         CALICE eLogbook   Results for search of  20070701 Page 5 of 22       19 2 AHCAL 7 5 5 717 32 33 56 41900 31000  19 3 AHCAL 6 4 7 6 17 32 33 56 41700 31000  19 5 AHCAL 5 3 1 517 32 38 56 41950 31000  19 6 AHCAL 4 2 4 4 17 32 32 56 41950 31000  19 7 AHCAL 3 1 12 11 17 32 100 56 41800 31000  2007 07 01 18 55 Michele Faucci    Giannelli    B UE E  un    r  ES MO TE  run            2007 07 01 20 51 S  Schaetzel    The Slow Control is painfully slow  CMB voltage has to be switched on by hand   Initialisation of CMB takes a couple of minutes     Module switch off  1 module HV on   9 5min  we had to switch HV off in the HV expert panel     2007 07 01 22 39 S  Schaetzel    green signal in scope picture is analo
62.  Additional  i detector  center  injection    more data        100k event    S  Uozumi  4GeV i   Additional   i center  injection       2007 03 18 20 46    2007 03 18 21 12 S  Uozumi 3 GeV run    400513 beam TIL 3GeV direct s    HN scanned one strip  fiber precise  scan       2007 03 18 21 12 S  Uozumi 3 GeV run  400513 beam TIL 3GeV direct 2 5V scanned one strip  fiber precise  scan    2007 03 18 22 22 S  Uozumi 3 GeV run       400514 beam TLL 3GeV direct 2 5V scanned 5x20  fiber points       CALICE eLogbook   Results for search of  20070301 Page 17 of 76            2007 03 18 22 22 3 GeV run       2 5V scanned 5x20  points  precise  2007 03 19 10 03 3 GeV run  beam direct 2 5V scanned  fiber points  precise  2007 03 19 10 03  UN  direct 2 5V scanned  fiber points  precise    3 GeV run    direct 2 5V scanned  fiber          3 GeV run    2007 03 19 10 04    2007 03 19 10 04 S  Uozumi       2007 03 19 14 56 S  Uozumi Tracking information    This plot shows incident position of the beam at detector surface extracted from drift chamber  information     Data is from run 400451  MIP calibration with fiber direct configuration    Black points are all the events  each color shows events where a certain strip has non pedestal  signal     Red     L1 S3 have  gt 50 ADC counts  Yellow     L1 S5  Blue sev clc  Pink     L1 S14    I expected to see sharp image of the strip  but the image is diluted   It means that optical cross talk of strips  or drift chamber position resolution is not ye
63.  Beam 10x10  20GeV 533  Data SE       CALICE eLogbook   Results for search of  20070701 Page 14 of 20          2007 07 14 02 28 Hakan Oliver Wenetakb  Pe LI CTT  Pb  Set collimators c2  c3  ue to low rate  2007 07 14 04 21 Hakan Oliver Weribest run       AAA AAA AA ASA MAA  Pb    Test run to check that ECAL is at y 10  Monitoring shows y  9 6  so will adjust some more     2007 07 14 04 35 new position    centered beam on ECAL for 50GeV electron beam  10   5        HCAL Motor control panel    actual position target position    x   10 01 bo GER  Y  50 I    ori   Y     O lost X reference point  O lost Y reference point    O lost 2 reference point    Q stage in motion  O manual operation get Pos    ficus 1 slow 1 pulsezfree   extern STOP   ref 24   204 556 204 958        OFF set   E  900 E 1290  File  http   ttfinto desy de CALICEelog data 2007 28 14 07 2007 07 14T04 35 02 00 jpg  2007 07 14 04 42 Hakan Oliver Went 40 GeV    ili   d   Ll LIII    Pb          Ep III     Pb    ECAL is now at correct position    CALICE eLogbook   Results for search of  20070701 Page 15 of 20       ww DN  ESAL    CAL   cur   Teacter  Padio   Dhglas22  Ce  Zeg  Stew   Zeg el  Aqu iio   Feier Z   rincon            ett KEN           AA Ten 390422 000   teg em    File  http   ttfinfo desy de CALICEelog data 2007 28 14 07 2007 07 14T04 42 48 01 eps          2007 07 14 05 38 Hakan Oliver Wenebiat   GeV       2007 07 14 06 09 wenbiao data transfer    When I check the data transfer  I make a mistake by typing
64.  C2 C3 C8 from   5 to   3 since H6B EXPTO1 went up to  100k  after change we  now have a rate of 6 3k       2007 07 26 22 11 Lukas  Alex beam data  EB call ONE E   SAI EU CENE   CIA EA E  Pb  We had to adjust collimators 1 5  mm from  6 to  5 mm to obtain desirable rate below  11 k ev   2007 07 26 22 46 Lukas  Alex beam data       Pb    We again had to change collimation settings   C1 4 5 C5 C6 C10 C11 5 0  Rate went from 11 3k ev to 7 5k ev     amp  2007 07 26 22 52 Roman Cogri Scripts restarted   gt  Copy to desy    I have restarted the cogri scripts and send the  files to desy again     Copying has resumed at run 330966     At the same time I have started the replication of  the files of runs 330871   330965 frm lyon to desy     I cannot however say whether the replication will be successful  Since apparently cc lyon has dcache problems at the moment           Cheers   Roman  2007 07 26 23 21 Lukas  Alex  Pb  2007 07 27 19 44 Lukas  me 10x10 SS  GCo C and Cll narrowed to  4  Ee GE screenshot prooves  the bean is way out of line    200k ev     NOTE  This run 331019 seems to be  15GeV e     Allister    CALICE eLogbook   Results for search of  20070701 Page 20 of 20       File  http   ttfinfo desy de CALICEelog data 2007 30 27 07 2007 07 27T19 44 34 01 png       2007 07 27 22 49 Lukas  Alex 10x10 pi   Pb   Collimators C1 C5 and C6 adjusted to   7  therefore same value like C10 and C11  to obtain desirable  pps  72130     Beam was off for few minutes during run    4150k ev   2
65.  CC CC CC CC CK CC CC CC CC CC CC CC Ck CC CK Ck C CK Ck Ck Ck Ck ck ck ckckckckckck    Thanks Alex  Just one point  the check for the CRCs being alive is only  done at DAQ startUp  i e  when starting from dead  and not each run  It   is really to check the CRCs have booted OK when the crate was last powered  on  Hence  your script does not need to loop as the log file won t change  unless the DAQ is closed down and restarted    Paul    2007 07 08 09 33 Evgeny Cristina Michal    b    e   330212 eamData  10x10  30GeV  e     33013 __Joeambata  0x19   30687 fe  52       CALICE eLogbook   Results for search of  20070701 Page 21 of 24       trying to get a broad electron beam in order to locate the interwafer gaps for ECAL  beam is hitting the left side of MWPC in x   change TRIM 6 to move beam in x direction    gt  run 330213 beam more centered    2007 07 08 11 09 Erika pion summary    summary of pion runs under following conditions   wobbling  80 GeV   AHCAL HV  nominal   0 3 V   Hodoscope still installed on beam line   10x10 trigger counters   100x100 front counter removed from beam line   Targ   Air  Abs   4 mm Pb    Eun f  E1007 LLL E 330198  330207  pg C P30171   330192  330204        run taken without absorber   higher electron contamination        amp  2007 07 08 12 19 hodoscope out    remove hodoscope from ECAL front    2007 07 08 12 40 Fab new beam file for 30 GeV e     H6B 530  e  at 30 GeV       Rate  1270 pps 10x10  1400 pps 20x20    2007 07 08 12 45 Evgueni  Michal  
66.  Cristina    330215  f boeamData 10x10  20GeV 533 6mm air Off 250k  Pb events  0 Hz    DAQ mean rate  5       330214 junk run     amp  2007 07 08 13 59 Evgueni  Michal   Cristina    Moved ECAL x position from 5150 to 4950     2007 07 08 14 04 Evgueni  Michal   Cristina    330217   beamData 10x10  20GeV 532 6mm air off 254k  Pb events       330218   beamData l0x10  30GeV  e  533 6mm air off 113k  Pb events    DAQ mean rate  50 Hz    2007 07 08 16 08 Niels  Jean Yves   Hengne    330219 ahcGain LED 533 6mm Pb fair Off only  one  configurbhtion    15k  events       330220 foeamData 10x10  30GeV 533 6mm air off 262k  Pb events     CALICE eLogbook   Results for search of  20070701 Page 22 of 24    DAQ mean  Em 56    330221 fbeamDatal10x10  40GeV 6mm Gees  Pb events    DAQ mean rate  8   Fab  Run 330221 Ge because beam not centered on the drift chambers   Changed current in Trim6 to steer beam  now at  10  saved in beam file   Sterted new run  330222   Will take same amount of events as previous run                     2007 07 08 17 00 Jean Yves Summary of electron runs  Summary of electron runs under following conditions    Hodoscope still installed on beam line   10x10 trigger counters   100x100 front counter removed from beam line   Targ   6 mm Pb  Abs   air    Electron Energy  Run f  330176  330179  330180  330188  330189  330194  330199  3  330191  330203  330210  330211  330212  330213    Run     330215  330217   330218 330220  330221 330222 330223 3330224  330228       2007 0
67.  DACs   0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0  HAB 8 DACs   0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0  HAB 9 DACs   0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0  HAB 10 DACs   0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0  HAB 11 DACs   0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0  FE2B is connected with module 5B which is cabled to channel 74 in the SlowControl HV panel   2007 05 11 15 57 Sebastian RMS mean  Schaeizel  slot 5 fe 0 is module 4  module 4A power seems broken  no current shown in the SC HV panel   module 4B is powered and CMB4 is connected  CAN address 4  appears at position 25 in SC CMB panel     CMB parameters  width 115 DAC  at CERNO6 this corresponded to 10ns   amplitude 115 DAC  Current in 4B at nominal HV is 231 muA  LEDs on or off makes no difference     Run ahcPmLed    See the attached plot for the amplitudes  Strange that the RMS is about the size of the  mean value  Online histogram obtained with   hstDisplay  c 172  s 5  w 1 HstCrcNoise    I wanted to look at the spectra of individual channels but  hstDisplay  c 172  s 5  f 0  p 11  n 17 HstCrcChannel  produces empty histograms  see next elog entry     CALICE eLogbook   Results for search of  20070501 Page 4 of 9        se fs  1 PEPE Chen       File  http   ttfinfo desy de CALICEelog data 2007 19 11 05 2007 05 11T15 57 03 00 ps    2007 05 11 16 08 Sebastian empty online histogram HstCrcChannel    Schaetzel  hstDisplay  c 172  s 5  f 0  p 11  n 17 HstCrcChannel    CALICE eLogbook   Results for search of  20070501 Page 5 of 9             
68.  Mi Crate ON Mi Crate OF  recover    KILL W  MA wa j wi j wi      recover all trips mi ma Wg J LY detail  Ae W values to file Amp RL values  position     AS    ala  ss          File  http   ttfinfo desy de CALICEelog data 2007 28 10 07 2007 07 10T19 07 01 00 jpg    2007 07 10 19 39 Mary Cruz Hengne Niels    330326 beamData  10x10  20GeV pi  523 4mm air 0 76 200k  Pb Events       330327 beamData  10x10  18GeV pi  522 4mm air 0 94 200k  Pb Events    2007 07 10 21 24 Mary Cruz Hengn     amp sts monitors removed    We removed the MWPCs and Sci counters for beam monitoring in order to minimize upstream material     2007 07 10 21 25 Mary Cruz Hengne Niels       330328 beamData  10x10 KE 35 0 200k  Events    QUADO1 shows slightly fluctuating current within     1A     2007 07 10 22 09 Mary Cruz Hengne Niels    330329 beamData  10x10  12GeV pi  524 4mm air in 0 0 20k  Pb transitiol Events       330330 beamData  10x10  12GeV pi  524 4mm air 2 11 0 0 300k  Pb Events    QUADO1 continues to misbehave   most probably a problem with the read back of the current  not with  the actual setting     2007 07 10 22 43 Mary Cruz Hengne Niels    The histogram dimensions for the online monitoring have been changed to some more reasonable values   The old settings are saved in HistoInit dat bak 20070710    2007 07 10 22 50 Mary Cruz Hengne Niels    Slightly high current  450muA  for module 28B  high voltage channel 50 25B  This is the current  limit   please monitor carefully  Call Beni  165554  when 
69.  Oliver Manqi electron run    load beam file 523  set cherenkov pressure to 0 688 bar   trigger rate   24kpps     close collinators to   COLLO1 COLLO5 COLLO6 COLL10 COLL11    4 mm     trigger rate reduced to   2 4kpps                30822 RETRATA AUTO RUE EEN    Trigger rate 2 4 kpps  DAQ rate  80Hz    Logbook entry   CALICEelog data 2007 29 22 07    ay TE Hengne Oliver Mangi electron run       30821 RENATA AO E  e Eze DEE    Trigger rate 7 25 kpps  DAQ rate  90Hz    Logbook entry   CALICEelog data 2007 29 22 07    ay T  Hengne Oliver Mangi electron run    330820 beamData 10x10  25GeV e  526 Air Air 0 4  6  O 1KEvents    Trigger rate 7 21 kpps  DAO rate  Hz          Stopped   adjusting cherenkov presure  goto 0 42 bar to enhance efficency for triggering  electrons     18 of 28 27 09 2007 11 39    CALICE eLogbook http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log       Logbook entry   CALICEelog data 2007 29 22 07      22 07 2007  o 19 02      Ceasar NetBeans is working again      loaded beam file 526  25GeV     due to high rate in 10 10 trigger  reduced collinator opennings   COLLO1 COLLO5 COLLO6 COLL10 COLL11   from 6 0 mm to 5 0mm  COLLO2 COLLO3 COLLO8   from 3 0mm to 2 5mm     10 10 trigger rate is now 7 0 kpps    Hengne Oliver Manqi Beni                            Logbook entry   CALICEelog data 2007 29 22 07    zy NU ICT CESAR problem solved       SPS  Sunday 22 Ju1 2007 Afternoon        o  Y    RO 2  Comment  Claude  Ron  created by spsop on cwo 
70.  TRIM0G  2A   Pb more centered now    trigger rate  2 5 kpps  DAQ rate  80 Hz                      NOTE  The beam was shifted in  y direction in run 330840  TRIMO5 is therefore changed  from 3A to     1A  Also in the same run  the beam was shifted in the  x direction  TRIMO6 is then  changed from  3A to  2A to center the beam     Logbook entry   CALICEelog data 2007 30 23 07      23 07 2007  SA 06 08    The second peak at larger energy correspond to the 8 points jump of the  energy mips  as visible in the picture below       Allister  Robert  Riccardo sudden jump in HCAL energy    1110144       mU               hEcalEnergyVsTime     gt t  i t  NEM     pin    d       or wr    hhHcalEnergyVsTime    race  e  E  D  ses  mas  LI  EI        m       Logbook entry   CALICEelog data 2007 30 23 07    SC A FAMAM  Allister  Robert  Riccardo sudden jump in HCAL energy    05 59  We observed a sudden jump in the HCAL energy in run 330839  We can t figure out the  reason yet   We didn t notice any problem with the beam at this point  And from the ECAL  we don t  See anything strange     EDIT  after some investigation we found the cause of the strange bump  no new discovery  unfortunately    just layer 14 of HCAL   module 16   slot fe 17 0 being very noisy because of DAC  settings   lost  please reset immediately HCAL DAC settings when you see something like this again     9 of 28 27 09 2007 11 39    CALICE eLogbook    10 of 28    E    23 07 2007  aco     http   ttfinfo desy de elog results j
71.  Y       actual position target position    Kei  deet wr     500     O lost X reference point   Q lost Y reference point   O lost 2 reference point   Q stage in motion     E get Pos          i mas Y Fast i pulsestree   extern STOP   reif    10 3 310 008    OFF set    E  0 H PEL to den          File  http   ttfinfo desy de CALICEelog data 2007 29 20 07 2007 07 20T20 04 02 00 jpg    2007 07 20 19 23 Oliver Allister Hengest run  stage has been moved to   3 3    2007 07 20 19 35 desy dcache unavailable  Hi there     according to the desy it news he desy mass storage is currently  not  available     CALICE eLogbook   Results for search of  20070701 Page 19 of 22       http   it news desy de  type trouble amp refresh 300 amp 1ang en   I don t know when the problem will be solved    I have stopped and restarted the copy scripts and redirected everything to lyon  The copying  and conversion has resumed at run 330725  sorry I have missed the switch to data01   this noon     This remains true until further notice    Control of the copy scripts handed back to shift crew    Unfortunately  this also affected the conversion of some runs which either   failed or which I had to abort due to the problem    Therefore there are some holes in the sequence of converted runs     Cheers     Roman    2007 07 20 19 47 Oliver Allister Henpien run       330736 beamData  10x10  25GeV pi  526 4mm air 0 49  3  3 a few k  Pb Events    10x10 trigger coincidence  5 1 kpps  DAQ rate    Hz    run stopped after
72.  a aa X W aa X a    Wang             due Hi values to file   dap AL values  position             File  http   ttfinfo desy de CALICEelog data 2007 34 21 08 2007 08 21T22 23 04 00 jpg    2007 08 21 22 36 Daniel   Anne Marie  Petr    350373 lbeamData  10x10  40GeV Git 613 4mm air 0 1 20k 30  Pb degs    adjust trim 06    350374 lbeamData  10x10  40GeV Git 613 4mm air 0 1 130k 30  Pb degs    2007 08 21 23 15 Daniel  Calibration run  Anne Marie  Petr    50375  50376    350381 lbeamData  10x10 i s 110k 30  pee  350382 lbeamData  10x10  15GeV i 30k 30  peg  p ue e          pi   eamData  10x10  15GeV pi  i 100k 30  degs  lbeamData  10x10  10GeV pi  520 i y 108k 30  degs    BhcGain E  Dunk rug  LI  BhcCcmLedVcalibScan          E   550380                     JahoPmbedVealibsean             2007 08 22 01 52 Satoru Hakan Allister    350385 lbeamData  10x10  10GeV e  520 air air 2 0 100k  30 degs  30    350386 lbeamData  10x10  15GeV e  521 air air 45k  degs   stopped  because    350387  peanpata meng  fisey    e          E21     r r pe fo from       CALICE eLogbook   Results for search of  20070801 Page 20 of 22          A O 8 sess    2007 08 22 02 25    Ch 59 and 56 seem to be moved     monitoring Switch On   Off cossents  Voltage  V  Current  uA  Set  V  Initial  V      79 60 oF H PE A 79 4   0 2   20 60 d    41 10    EH  ul    ofl    ww 2  E3 IJ  fad I  Lud    La  j    44 40    HY settings   file actions enable   disable ALL chamels  Sw   imb   twr   erable   disable        SSES    Ex
73.  a few 1000 events due to access to detector    2007 07 20 19 50 Oliver Allister Henpian run       330737 lbeamData  10x10   25GeV pi  526 4mm air 0 49 3   3 256k  Pb Events    10x10 trigger coincidence  5 1 kpps  DAQ rate  72 Hz    2007 07 20 20 09 caliceon 2calicesd osirarcCheck    position check in ECAL after changing of stage position to   3 3     CALICE eLogbook   Results for search of  20070701 Page 20 of 22       Feet ue a   char   tracker   artic  Deptt  Cooaicnecta   Sai   Sege   Rarporae   Peseta TePrtiecions      EE i Lech Aa  A oh E e EA a4  x  cm LE nah lee    Deene eelere       File  http   ttfinfo desy de CALICEelog data 2007 29 20 07 2007 07 20T20 09 02 00 png    2007 07 20 20 49 Oliver Allister Henpien run       330738 beamData  10x10  20GeV pi  523 4mm air 0 75  3  3 254k  Pb Events    10x10 trigger coincidence  5 4 kpps  DAQ rate  88 Hz    2007 07 20 21 44 ee run       330739 beamData  10x10  18GeV 522 0 936  3  3 255k  ES V    NOTE  Cherenkov went a bit above the desired threshold later in the run  actual value 0   desired   threshold 0 94   10x10 trigger coincidence  4 6 kpps   DAQ rate  86 Hz    2007 07 20 22 49 Oliver Allister Henpien run       330740 beamData  10x10  15GeV pi  521 4mm air 1 35  3423 254k  Pb Events    10x10 trigger coincidence  2 9 kpps  DAQ rate  80 Hz    2007 07 20 23 26 Oliver Allister Hengtability of ECAL response    We plotted the mean of the gaussian fit of hEcalEnergy for electron runs against the run number   i e  time   The er
74.  an    Ur an    AAA A       iin    ER D 74  KE  2     KC    Trend to   Los    las a 41 0  0 2   oo a L       s 34 41 1  0 2          r fe     H e  E Ce ip Wi settings   file actions   comable   disable ALL channels     d mg j ue   mr j  eei JI       a ATM   EER amer A RATA     Wi Grate tn   _ Wi Crate OFF     recover    KILL we    WA i j wi go   w2 J MS   recover all rel L3 1 ow  ea   LY detait     dme W voles to file   due  L values  position               B 72 4 40           Logbook entry   CALICEelog data 2007 30 23 07    23 07 2007  07 40          Trim05 5A   gt  1A       coll co105 co1l6 co110 co111   10       Logbook entry   CALICEelog data 2007 30 23 07    23 07 2007    07 22 HCAL    Channels 33 17A  1 26A and 2 31B showed seldomly  but continuously   a red alarm due to overcurrent  compared to maximum value loaded in config file        Aliste Robert Riccardo Comments on HV channels in    Logbook entry   CALICEelog data 2007 30 23 07       27 09 2007 11 39    CALICE eLogbook http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log         23 07 2007 i    0626 Allister  Robert  Riccardo    Amm  144kEvents  stopped because the beam  330840 beamData 10x10  18GeV  pi  523 Pb air 0 797 is not centered in both x and y in the  8 DCs    44kEvents  adjusting both TRIMO5 and  L0x10  18Gev  pi  523 4m  797 y eed 3       4mm 56kEvents  adjusting both TRIMO5 and  10x1   52 797  EE 06 magnets  getting there     330843 beamData 10x10  18Gev pi  523 4T a i r  o 297 o
75.  beamData  10x10  40GeV 150k  ES events    2007 07 26 14 47 Erika Fabrizio MarttieamData    330965   eampata  ioxio0  secev Je  35      Bir    em    pee po  good run        CALICE eLogbook   Results for search of  20070701 Page 17 of 20       DAQ rate  60 Hz          2000 pps  2007 07 26 17 50 all of us rotation to 10 deg  ROTATION of stage to 10deg  2007 07 26 17 51 Lukas  Alex  330967 lahcPmNoise  2007 07 26 18 25 caliceonOcalicehedttarifeacunrent threshold   CALICE user     set higher current threshold values for  ch  28   layer 14B   module 16   ch  62   layer 31B   module 36   ch  1   layer 1A   module 3      37 ATA   se      7  F 10 98 ya    ue E 310A    ss ow moe    se mima   95 Ramp    65 WJA        ge coo   ES 237 W  NEA    108 w 138 Tao w 4B 5 w 741378     14   F Dan 150   r w TSROA    n3 wea   f 169 p 7598    100 mansa   09   as e pA    57 w ose   mei   95     CA    120      126   W  maps       5  He Crate UN W Crate OFF  recover    KILL BY  MAS  A 1  sl    HY settings   File actions erable   disable  amp L chanel ws  recover we LY detail  sw   up   EMT     be   aab   mess all tipa  10033 mad    de Hi values to file dine HLL valuss  position   File  http   ttfinfo desy de CALICEelog data 2007 30 26 07 2007 07 26T18 25 02 00 jpg             AA  e fe                 2007 07 26 18 59 Erika junk runs  runs 330968   330974 optimization of ECAL   HCAL position w r t  beam   80GeV pi  is used but with changing conditions    2007 07 26 19 22 new position    We rotated to 1
76.  daquser   gt     code daquser   data  records   gt     code records    Then you create also a local inc and src and copy the files from the zip file in these files as  indicated in the zipfile structure     For the ahcBinTree tool also   mapping   and   calibration   directories are needed  Examples are in the zipfile  examples from the CERN testbeam 2006      Althought the newest DAQ BinReader can handle the DESY dCache file access the needed modified files  are also contained in the zipfile     CALICE eLogbook   Results for search of  20070201 Page 21 of 21       dcap h  dcap_errno h  RcdReaderBin hh    Take care  in this case you use a different Reader than the official one and you have to use the  build script provided by the zipfile    bin build      A documentation of the tools can be found in the attached pdf file        File  http   ttfinfo desy de CALICEelog data doc CERN docu 2007 02 27T21 19 00 04 pdf    la 2007 02 27 22 10 Marius Module01 had a wrong DAC file   The DAC file DAC_module01_031006 dat was wrong   it is the same like DAC_module02_031006 dat   Therefore module01 was not working properly during the october run period    It seems to be a copying failure or mine     The new corrected file copied to the calice diskarray is called   DAC module01 260207 dat    CALICE eLogbook   Results for search of  20070101 Page 1 of 5       2007 01 04 16 18 Sebastian Module 24 gain and pedestal runs  Schaetzel    202250 ahcGain 20k events per configuration  see  below for
77.  from 5 to 20  Now at 2 5 kpps      reduce collimator 9 from 35 to 20 to clean up beam profile  End this run after 35k Events and start new run     Pb          150 kevt    2007 07 28 02 07 Nanda  Roman pi  run    Pb    10x10 trigger rate  3 5 kpps    155 kevt    2007 07 28 02 08 Nanda  Roman pi  run       Pb    10x10 trigger rate  5 3 kpps    beam gone after  130 kevt    2007 07 28 02 46 Nanda  Homan calibration data       2007 07 28 05 18 Nanda  Roman e  run       closed collimators 1 5 6 from 6 to 4mm  and collimator 10 11 from 6 to 5mm       Now trigger rate in 10x10 coincidence  2 7 kpps    CALICE eLogbook   Results for search of  20070701 Page 6 of 16       DAQ rate  92 Hz  High pion contamination  but well seperated by cerenkov threshold    200kevt    2007 07 28 06 05 Nanda  Roman  e  run    o    331038  high triger rate    gt  closed collimator 1 amp 5 from 6 to 4mm  6 from 6 to 5mm  10811 from 7 to 5mm  not centered in DC  at 60 in Y       change trim05 from  3 to  6A  331039  wrong direction       change trim05 to  2A  still pion contamination visible      lower cerenkov pressure to 0 7  331040  10x10 trigger rate  7kpps  DAQ rate  90Hz  200kevt       2007 07 28 06 57 Nanda  Roman  e  run       10x10 trigger rate  1 6kpps  DAQ rate  70Hz  200kevt       2007 07 28 07 37 Nanda  Roman comb run  33  10    1042 lbeamData  10x10  6GeV e  pi  023 Air Air 2 98 10deg   good  10    x trigger rate   Akpps  DAQ rate  75Hz          2007 07 28 07 46 Roman cogri scripts stopped    
78.  gaussian fit mean of hHcalEnergy  for the combined ECAL HCAL  runs  i e  33xxxx run series  for different beam positions and rotation angles     CALICE eLogbook   Results for search of  20070801 Page 3 of 22             Zeg am geg ep EE  Dani bi 10  4 14 e Ei mma 2     mnw ewn  5 xy  330 E yor 400 H Eo DRETT ET vi  E E a 1   ge 1  g f  iz d    z  Sune dee    LG  M  00  p       mo  A EH   D  I  gt     Es 3  c   mi e g  A zi  S   DN    FX BEL Dm kl n e DN FL NEN Gg KE E  M    PUn a   VO PME Vafheshai s vi    2 se nm  3                                            Y nu    Dm oe DA   ae   Im KSE i 4 xy 04 19    ape  0 0 20  m  atl M E uo  l   E BE j  Ae  wal g  e  S g on end KI  zu B  m   s so    E 4 E  ege deeg  SSA rm JOE Y STEDT  D EOL PAR  87 cates GE   UU NUES TU URGUE  p t ne  n                       rf mau  se Bep S   ur     E yo ono EE yo  600  b  y ho  i  po   ge E cm  e o e  e  LEE  SE Jeep WWW KR E  m mmo  File  http   ttfinfo desy de CALICEelog data 2007 33 19 08 2007 08 19T03 37 55 00 png  2007 08 19 04 08 Hakan Allister Satoalibration runs       2007 08 19 05 08 Hakan Allister Sat8 GeV pion      Al AN    udi pm Md NE B    i               pi  350255 eamData  10x10  40GeV  i    dani HH GE     2007 08 19 09 11 Jamie  Pavel  pi   Beni    it    614  613  12          CALICE eLogbook   Results for search of  20070801 Page 4 of 22    350259 lbeamData  10x10 H50GeV  m  missing     Beam fail hamber appears to have failed  Investigation in  e A          3 3k trigger rate  80Hz 
79.  is mmdd  and format of time is hhmmss   temperature     not implemented yet   stagePositionX  stagePositionY     Position of the movable stage  Since    the stage position is already involved to the trackXPos and trackYPos   normally you will not need to use stage position information   trackXPos trackYPos     Position of the beam on surface of the SCECAL   calculated from tracking and stage position information    trackXSlope  trackYSlope     slope of a track   trackXHitMap  trackYHitMap     hitmap of the drift chamber X and Y wires   trackXChisq  trackYChisq     Chisquare of the tracking    ADC layerID 1   stripID 1      ADC counts for each channel  For the beam  data  pedestal is already subtracted    ADC xxx     ADC counts for trigger and veto counters    XXX T1R  T2R T3L V1U V1D V2D    TDC     raw TDC counts  normally not necessary     Since it does not contain temperature information and tracking is  necessary to be improved   there will be updates of the ntuples after the beam test     CALICE eLogbook   Results for search of  20070301 Page 58 of 76    But anyway for first analysis these data are already useful     2007 03 25 12 31 S  Uozumi uniformity of fiber readout module    This shows uniformity of light yield with a fiber readout strip  measured with MIP events     3003 due Entries 10000    Meanx 24 09  Mean y 35 21     RMSx 14 87    RMS y 20 56       File  http   ttfinfo desy de CALICEelog data 2007 12 25 03 2007 03 25T12 31 12 00 jpg    2007 03 25 12 34 S  Uozu
80.  lib libGui so  1  36 0xf6470685 in TGFrame  HandleEvent     from  scratch software i386 extern root lib libGui so  1  37 0xf644800c in TGClient  HandleEvent     from  scratch software i386 extern root lib libGui so  1  38 0xf6447c24 in TGClient  ProcessOneEvent     from  scratch software i386 extern root lib libGui so  1  39 0xf6447d43 in TGClient  HandleInput     from  scratch software i386 extern root lib libGui so  1  40 0xf64460ca in TGInputHandler  Notify     from  scratch software i386 extern root lib libGui so  1  41 Oxf7ac5029 in TUnixSystem  DispatchOneEvent     from  scratch software i386 extern root lib libCore so  1  42 Oxf79d4686 in TSystem  InnerLoop     from  scratch software i386 extern root lib libCore so  1  43 0xf79d4626 in TSystem  Run     from  scratch software i386 extern root lib libCore so  1  44 0xf794e054 in TApplication  Run     from  scratch software i386 extern root lib libCore so  1  45 0x080662e2 in main                   2007 07 29 10 41 Beni  Martin  junk run  Ariane       tracker slightly off  adjusting current  run stopped    2007 07 29 10 47 ECAL noisy    noisy ECAL FE in run 331094       2  5  Q  vo   amp   g  ki  uU  z 3  S a  Z ei  4  H                   uz  0 10 20 30 40 50 60 70 50 chip number  File  http   ttfinfo desy de CALICEelog data 2007 30 29 07 2007 07 29T10 47 02 00 gif  2007 07 29 10 54 Beni  Martin  pi  18 GeV  Ariane    N e ee E  EIE  Pb  2500 pps  DAQ rate  85 Hz  150k events  2007 07 29 11 25 Beni  Martin  pi  25 GeV  Ar
81.  mE  Pb    DAQ rate  90Hz  Spill  12kpps    2007 07 15 07 52 Valeria Wenbiao Qiosr run    Pb       DAQ rate  70 Hz  spill rate  6200pps          2007 07 15 08 00 Allister  new shift  Mary Cruz  Beni    2007 07 15 08 50 Allister  50 GeV electrons  Mary Cruz  Beni       pee  Pb    CALICE eLogbook   Results for search of  20070701 Page 4 of 19       2007 07 15 09 56 Allister   40 GeV electron run  Mary Cruz  Beni       weep Pep pp fe fm  Pb    2007 07 15 10 52 Allister   Mary Cruz  Beni    330472 lbeamData  10x10   30GeV e  534 air 6mm 0 3214  3  3  Pb  beam not  centered  330473 lbeamData  10x10   30GeV e  534 air 6mm 0 3214  3  3 Tunning  Pb the beam  nt nae E PA    2007 07 15 10 56 Allister   Mary Cruz  Beni    Pb          2007 07 15 11 54 Allister   Mary Cruz  Beni    Pb       2007 07 15 13 02 Allister   Mary Cruz  Beni    330476 beamData  10x10  20GeV 538 air 6mm 0 7184  3   3 Stopped  Pb Beam not  centered    2007 07 15 13 14 Allister   Mary Cruz  Beni    Pb          2007 07 15 13 30 Allister  restartet X server on caliceserv  Mary Cruz  Beni    X got stuck on caliceserv  restarted xserver  restarted cp_comb  cp_ecal  cp_hcal    2007 07 15 14 01 caliceonOcalicesenergy jimp in HCAL    This jump in energy was recognized during an electron run   No obvious reason found     All other detectors do not show this jump        hHcalEnergyVsTime Ene  7   mm        2 Man di  g ES ny  Ps ES   gt  e paja o  m integral 1908  9 14 E  vo    Guti IM M  e ch    8 H          o o 500 1000 150
82.  of the high current in module 27 B     We found a bend connecter pin on the baseboard module connection  The pin of chip 7 chan 4 was  touching   the neighboring channel 5  This means  that we should see no signal for the MIP runs   in 7 4  at least     But this was not the reason for the high current   The current is still 0 50mA   operation voltage  after repairing the bend pin     The test also showed that the current comes from the module and  not from the electronics     2007 04 26 08 52 Doerte David assembling status    All tiles for the last cassettes are at DESY  The last tests are started  Due to a lot of not good  SiPMs it was hard to find spare ones  But now we have complete lists for module 33 and  36  May be we will get some new spare later together with the SiPMs for  new   cassettes no  1 and 2     The stripping machine is working well again  But the workshop we share it with requires it more  and more     After a long time doing a very good job one colleague will leave us this week going back to Moscow   A new colleague from Moscow arrived already this week so he can be trained by two experienced  colleagues     The last assembled cassettes are not tested with electronic boards because the equipment is used  for the tile tests        cassettes ready  34   35  cassettes work in progress  33  cassettes ready  3     32  cassettes disassembled  i  2   status of production  Teo  cassette Imam bing HAL   FAC   module assembling steps staff   1 2 3 4 5   7 8 9  8   
83.  point       See     O lost Y reference point       Q lost Z reference point    Q stage in motion  o manual operation get Pos         Tem  s   slow   puise free   extern STOP   ref  8   280 0   280    OFF set    EI a Kl za PANEL toelog               J    File  http   ttfinfo desy de CALICEelog data 2007 28 10 07 2007 07 10T14 18 02 00 jpg       2007 07 10 14 26 Fab DC1  right wire  signal cable moved to different    discriminator s channel  Discriminator s channel where DC1  right wire  was plugged in found broken   Moved DC1  right  cable to different channel of same discriminator     2007 07 10 14 33 Beni  Michal   Cristina    330306 lbeamData  10x10  10Gev pi  520 4mm air 2 9 62K   20  el   30  muon   0 0      330305    eampata 0x0  roger fe 20 lair     hir     tt          abandonneh       2007 07 10 16 31 Mary Cruz Hengne Niels                            CALICE eLogbook   Results for search of  20070701 Page 11 of 25    P    tuning    not for  physics    B    N  E  E  o  O O O O O O O CH O O CH  C C Cn Cn C C C C Cn Cn  Ou    S      N  A  E    N  A  E    N   A  S8          UI UI GUN UI UI  NI  Ka  E  o    A A  Z   53    A  S         A  S                        2007 07 10 17 41  Fwd   calice vo support  Short Service Interruption  srm dcache desy de        Publication from   Christoph Wissing  lt christoph wissing desy de gt   DESY HH     This mail has been sent using the broadcasting tool available at http   cic gridops org       due an urgent reboot of the PNFS backup de
84.  positron beams in DESY II     Called to control room again at the morning  they told that their work with positron beam will take  some more hours       2007 03 23 07 45 S  Uozumi 3 GeV run       400658 beam T1L 3GeV KNU fiber  2 5V scanned  detector  scan    2007 03 23 07 45 S  Uozumi 3 GeV run    400658    em n         oev         ENU fiber 2 5V_      canned     Dx   7        CALICE eLogbook   Results for search of  20070301 Page 40 of 76    points    precise  detector  scan        with low intensity beam           2007 03 23 07 45 3 GeV run    KNU fiber  2 5V scanned 9x9  points  precise  detector    3 GeV run    Ki    S  Uozumi         S  Uozumi    i       with low intensity beam            2007 03 23 07 45    2007 03 23 10 02 S  Uozumi no beam today     Called to control room again  they told that there is a problem with HERA injection  no estimate  to rinish be e   I just asked them to give us a call when we can have the electron beam    Running the LED run meanwhile     2007 03 23 17 41 S  Uozumi LED run          a dd WE    7 tick is OK         2007 03 23 17 41    CmLedHoldScal KNU fiber 2 5V holdscan  to check  holdtime  with KNU  module    2007 03 23 17 41    Re e i a    tick is OK         2007 03 23 17 41    PmLedHoldScal KNU fiber 2 5V holdscan  to check  holdt ime  with KNU  module    2007 03 23 18 10 S  Uozumi fantastic     Control room people are still working on the HERA injection problem  and they have no estimation  to finish    But Erika and Beni kindly cal
85.  run    peu rne  inter calibr  tion  of L25515  2007 03 23 18 51 S  Uozumi LED run    400718 CmLed LED KNU fiber 2 5V to  see lpe  peak at    CALICE eLogbook   Results for search of  20070301 Page 45 of 76          odo  LL LLL LLL LL VETE    2007 03 23 18 51 S  Uozumi LED run  400718 CmLed LED KNU fiber 225V to  peak at  L25515  2007 03 23 18 52 S  Uozumi LED run  400719 CmLed LED KNU fiber 2 5V to  see lpe  L25514  2007 03 23 18 52 S  Uozumi LED run  400719 CmLed LED KNU fiber 25V to  see lpe  L25514  2007 03 23 18 52 S  Uozumi LED run         AAA e A  inter calibr  tion  of L26S15  2007 03 23 18 52 S  Uozumi LED run  KNU fiber   for  e D IL Fet  of L26S15  2007 03 23 18 53 S  Uozumi LED run  400723 CmLed LED KNU fiber 2 5V to  Er LP e EE sek    peak at  L26915  2007 03 23 18 53 S  Uozumi LED run  400723 CmLed LED KNU fiber 2 5V to  peak at  L26515  2007 03 23 18 53 S  Uozumi LED run  400724 CmLed LED KNU fiber KN to  TTT TT sg  peak at  L26914 16  2007 03 23 18 53 S  Uozumi LED run  400724 CmLed LED KNU fiber 2 5V to  peak at  L26514 16  2007 03 23 18 54 S  Uozumi LED run    400725 CmLed LED KNU fiber 2 5V to  see lpe  peak at  L14s15    2007 03 23 18 54 S  Uozumi LED run    CALICE eLogbook   Results for search of  20070301 Page 46 of 76    CmLed KNU fiber              2007 03 23 18 55 S  Uozumi LED run    CmLed KNU fiber   same  with  previous run  but with  increased  lover voltage    2007 03 23 18 55 LED run    CmLed KNU fiber   same  with  previous run  but with  increa
86.  scanned 9x9  fiber precise  scan       2007 03 20 10 13 overnight precise scan    Run400529 400533 are overnight precise scan with 9x9 points without tungsten     2007 03 20 10 14 S  Uozumi 3 GeV run    MIP run direct   scanned 3x10  fiber one strip    precise  scan       2007 03 20 10 14 S  Uozumi 3 GeV run  MIP run 3GeV direct d scanned 3x10  fiber one strip    precise  scan       2007 03 20 11 29 S  Uozumi    Final mapping file for 2nd configuration after fixing FPCB swapping     CALICE eLogbook   Results for search of  20070301 Page 23 of 76       File  http   ttfinfo desy de CALICEelog data 2007 12 19 03 2007 03 20T11 29 18 00 dat    2007 03 20 11 42 S  Uozumi et al  check on tracking information  Now the previous problem on tracking information is gone   I found that polarity of Y direction on tracking information is flipped   This shows response of Layer 1 strip 3 with various beam injection position  bug corrected   with a MIP calibration run  run400459    The image of the strip is still not very sharp  maybe due to tracking resolution and optical  cross talk     But anyway size of the strip is consistent with the real size  1x4 5cm      CALICE eLogbook   Results for search of  20070301 Page 24 of 76       Beam X position  mm   NO A ie           O       NO  O       20 0 20 40 60 680  Beam Y Position  mm     File  http   tttinfo desy de CALICEelog data 2007 12 20 03 2007 03 20T1 1 42 54 00 jpg    2007 03 20 16 04 S  Uozumi some runs    Runs 400536 400537 are precise s
87.  scheme    202251 EncPmNoise  20k events  202252 BhccmNoise  20k events    ahcGain scheme   nominal Vcalib  Vcalib in AHC cfg    configuration comment    E  pO oo  CMB     A peminal vealib 109  E                 ena vealib  LL EECHER          2007 01 04 16 40 Sebastian MIP calibration module 24 reverse orientation with  Schaetzel CMB   202253      ehcBeamStageScan__  v 13         OB GeV fe    stage positions  flchcaldaq01  home caliceon cfg StagePositions innercore dat  stopped on Jan 5 at 11h40 at position x  105  y 15 because no beam rate  machine problems         REVERSE ORIENTATION    gt  the chip channel data in the file correspond to the normal orientation     LEDs and temperature readout via CMB          5 ret d  1 Pre       i 1 1 L 1  12 14 16182022 Q 2 6  File  http   ttfinfo desy de CALICEelog data 2007 01 04 01 2007 01 04T16 40 23 03 gif  2007 01 05 11 45 Sebastian Module 24 gain and pedestal runs    Schaetzel    BneGain                        re 202250  lahcPmNoise                         Po events    jahcCmNoise                                         Ek events       2007 01 10 09 08 Doerte David assembling status    Due to xmas and holidays the production is slowly going on   All modules from CERN  1   23  are back at DESY   Others  24   26  are ready for tests     Last year we got tiles for module 27 and 28    Several SiPMs were not ok  defect or long discharge  but we had not enough spare tiles due to their  higher HV  Therefore we decided to combine 1 half of
88.  strip 14     For Layerll Stripl15                       400449   CmLed for inter calibration   400450 PmLed for inter calibration   400447     CmLed to see lpe peak   400448 CmLed to see lpe peak around strip 15     GNO GP LE Tune  Sep pS PSSS RS    MIP calibration with lst configuration after removing tungsten                  400451  400452  400453  400454  400455  400456  400457    end of MIP calibration run                                 Precise detector scan with 1st config after removing tungsten                  400459  81 points entire detector    400460  81 points entire detector    400462  81 points entire detector    400464  30 points  precise scan for X layer strip 5    end of precise scan                                         2nd configuration MIP calibration before putting tungsten                  400468   400469   400470   400471   end of MIP calibration                           2nd configuration precise detector scan before putting tungsten               00473  81 points entire detector              4  400474  81 points entire detector    400475  81 points entire detector    400477  30 points  precise scan for X layer strip 5    400478  30 points  precise scan for X layer strip 5    400479  30 points  precise scan for X layer strip 5    400480  30 points  precise scan for X layer strip 5    400481  30 points  precise scan for X layer strip 5    400482  30 points  precise scan for X layer strip 5    400483  30 points  precise scan for X layer strip 5  sto
89.  the calice sw  Jast   2007 07 01 06 53 Alex one CMB remaining for tuning  beamHoldScan started    run 330016  beamHoldScan  10x10 trigger  spill sync enabled  histogram output enabled  split beamRun  in 250k events     stopped ecal safety run   all CMBs are powered down  power supply switched off manually     2007 07 01 08 25 Beni Yoshi hold scan test    330018 BeamHoldScan   10cmX10cm muon combined Stopped  in a way       330017 BeamHoldScan  10cmX10cm muon combined Enable  histogram  output and  split  beamrun in  250k events     Average event rate  70Hz    Scalers  H6B_EXPT01 6 0E 03  H6B_EXPT02 6 2E 03  H6B EXPTO3 6 1E 02    2007 07 01 09 55 Beni Yoshi    All previous BeamHoldScan runs was not completed   Overwritten by mistake     330019 330022 are also junk run     CALICE eLogbook   Results for search of  20070701 Page 2 of 22       2007 07 01 10 21 Beni Yoshi hold scan   230023   BeamHoldScan  i0cmX10cm      muon combined TC   ____      Average event rate  65Hz  Scalers  H6B EXPTO1 6 0E10 03  10cmX10cm trigger   H6B EXPTO2 6 2E10 03  20cmX20cm trigger   H6B EXPTO3 6 1E10 02  ECAL layer  3 and  4 are noisy   ECAL slow control s  5V display is inactive because of network connection problem with the module    300k events were taken     GOOD run for determining hold values from beam in the center most cells covered by the 10x10 trigger    also allignementd of detectors should be possible from this file     2007 07 01 11 37 Beni Yoshi hold scan   230025  BeamHoldScan  i0cm
90.  the first set of modules       2    8233323234218       2222222255    Wu O8   Movement   160  Modul tefo      IER       mann   Pn j vro    A    OO OOO GON    16 00 17 00 18 00 19   07 065 200  07 08  2007 07 05  2007 07 0            10 00 jj  10 00 2 Fe   ere MY broken           ee AAA                   HI Crate ON EI oj   recover    KILL we   MD  Med 1  897   we      recover di tries  DWE S wwe J wg  e   dnp W vabas to File   dae ML values  position            10 00 meos              File  http   ttfinfo desy de CALICEelog data 2007 19 07 05 2007 05 07T18 53 01 00 jpg    Page 2 of 9                CALICE eLogbook   Results for search of  20070501 Page 3 of 9  2007 05 11 15 26 Sebastian Slot 5 FE2B no readback  Schaetzel  CrcLocationData  print    CrcLocation  print    Crate number   172   Oxac  Slot number   5  CRC Component   2   FE2  Label   2   0x02   0000000010   read  AhcVfeStartUpData  print    Verification data   21845   0x00005555   0500000000 00000000 01010101 01010101  HAB 0 DACs   199 179 194 194 199 179 30 46 51 30 46 35 20 20 51 30 46 25  HAB 1 DACs   51 25 51  46 40 26  30 35 46  51 199 199 56 15 20 20 40 51  HAB 2 DACs   25 51 40 30 46 81 107 143 174 61 128 163 138 138 122 15 46 35  HAB 3 DACs   143 158 122 112 138 61 87 117 148 128 56 20 30 138 117 117 133 138  HAB 4 DACs   117 112 66 102 61 76 179 46 20 35 87 61 76 97 76 97 66 61  HAB 5 DACs   71 71 20 46 35 87 61 71 71 61 122 102 148 138 168 40 35 56  HAB 6 DACs   0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0  HAB 7
91.  the tiles foreseen for module 27  and 1 half foreseen for module 28    Additional we had technical problems with the test system which were solved also in December    All     tiles for 27 and 28 are tested  Assembling of cassette 27 was started     cassettes work in progress  27  cassettes ready  24   26  1   23    CALICE eLogbook   Results for search of  20070101 Page 2 of 5          status of production Q2c   07 cassette assem b  ng HAL   FLO  module assembling steps staff  1 2 3 4 5 D H D 9  8     baseframe     tles         SiPMs   PRA plate pu  s cabing fibres     tempsensors cover plate no of persons       H          au de vt D  Deag F1  100107    File  http   ttfinfo desy de CALICEelog data 2007 02 10 01 2007 01 10T09 08 58 03 pdf    2007 01 10 13 15 Doerte David assembling status    planning     Tiles for module 29 and 30 are ready produced and tested in Moscow    Tiles for module 31 and 32 will be tested in Moscow soon    Sending all tiles including spare tiles for module 27 and 28    to DESY are planed for end of next  week     Our 2 Russian colleagues will come to DESY again end of January  beginning of February   With cassette no 31 we will start with an other granulation  coarse  instead of fine   Some questions about cabling  allocation SiPMs to front end electronic have to be answered     The machine  the program for measuring the position of SiPMs and drilling the FR4 plate has to be  readjusted     2007 01 15 19 14 Sebastian MIP calibration inner core module 
92.  there are lots of electrons in the Cherenkov filter    CALICE eLogbook   Results for search of  20070701 Page 8 of 20       hacha   Santas   Setegi          cn    11111151                See 011820 303 18 Er aa 13 2007    f 336090 006 Teen  am    File  http   ttfinfo desy de CALICEelog data 2007 28 13 07 2007 07 13T02 44 06 02 gif    2007 07 13 03 43 SEN rygveun 30 GeV    330384 beamData  10x10   30GeV 534 Air  n abandoned       330385 beamData  10x10   30GeV 6mm 250k  Pb events    Collimators c2  c3  c8 set to     7 0mm    2007 07 13 04 39 Hakan Valeria Trygveun 40 GeV       330386 beamData  10x10  40GeV e  531 Air 6mm 0 1  250k  Pb events    Cherenkov not much use    Most of the events do no fire Cherenkov  but it looks like a pure electron  beam anyway     2007 07 13 05 32 Hakan Valeria Trygveun 50 GeV    330387 beamData  10x10  50GeV e  530 Air 6mm zT 250k  Pb events    2007 07 13 06 27 Hakan Valeria Try ain Run  330388 feheGain  El            2007 07 13 06 29 A rypien run  40GeV    ili TI EE  Pb    330390 beamData  10x10  40GeV 4mm 250k  Pb events    Collimators c2  c3 and c8 set to     1 5mm  2 0mm and 2 0mm  respectively       2007 07 13 07 22 Hakan Valeria Try  vanmary of Electron Data  for this shift   By performing a Gaussian fit to LES main energy peak in ECAL we extract the following values     Beam Energy  GeV  San E  MIP        CALICE eLogbook   Results for search of  20070701 Page 9 of 20       Note  The distributions are quite far from being Gaussian since th
93.  type   Calibration Run w o Tungsten  Beam Energy   3GeV    overvoltage for KNU modules   3 7 V    Layer 1   Layer 13   gt  KNU  Layer 14   Layer 26   gt  Kuraray fiber coupling    odd number layer     Y layer  event number layer   gt  X layer    3rd_mipKNU   3rd mipKNU     Entries 234                            T TTTTTTTTTITTITT T       T       i f  i   2 f   SAA E E PIPA ME  0 100     200 300 40 50 600               File  http   ttfinfo desy de CALICEelog data 2007 12 25 03 2007 03 25T10 51 37 01 JPG    2007 03 25 10 53 S Chang et al  Responce of KNU modules   Calibration Run   3 7V     CALICE eLogbook   Results for search of  20070301 Page 56 of 76       3rd Config  500543  3 7V  X          i    50  MU WW       File  http   ttfinfo desy de CALICEelog data 2007 12 25 03 2007 03 25T10 53 23 00 JPG    2007 03 25 10 54 S Chang et al  Responce of KNU modules   Calibration Run   3 7V     3 d Config  500543  3 7V  Y       Sall    U    A TTT          File  http   ttfinfo desy de CALICEelog data 2007 12 25 03 2007 03 25T10 54 19 00 jpg  2007 03 25 11 00 S Chang etal   3rd config Report of results Before preliminary  t s a report about KNU modules  3rd configuration   linearity and resolution   Beam energies are 1GeV 2GeV 3GeV 4GeV 6GeV  all five  each beam run has two datas  two run number   t was results before getting 5GeV data   Even 4GeV run data is not complete data  center injection 10 times  missing among 55 points   Total 55 points       36 points to center of each strips  
94. 0 2000 2500   3000 3500  time elapsed  sec     File  http   ttfinfo desy de CALICEelog data 2007 28 15 07 2007 07 15T 14 01  02 00 gif    2007 07 15 14 33 Allister   Mary Cruz  Beni       330478 lbeamData  10x10  30GeV e  534 air 6mm 1 3309  Pb  because  beam not  centered    CALICE eLogbook   Results for search of  20070701 Page 5 of 19       330479 lbeamData  10x10 I 30GeV 534 air 6mm 1 3309  3 43 Beam  Pb tuning    A 2007 07 15 14 42 Beni UPS dead    One of the 3 UPS we installed last week is broken  It was already marked as not so good  I tested  if it can support the monitor in the computer room long enough to do a controlled manual  shut down  It switched off after approximately one minute   I removed this UPS and tried to build a test setup for the auto shutdown with this one  During a  power failure test it broke down completly     2007 07 15 14 47 Allister   Mary Cruz  Beni    o MN MONOS a  Pb    2007 07 15 16 08 George  Mangi  electron run  Nanda       Pb       10x10 trigger rate  1 6k pps  DAQ rate  60Hz    2007 07 15 17 30 George  Mangi  calibration runs  Nanda   230485 0 k bhetmnedvealibsceaa o o o E  2007 07 15 17 39 George  Mangi  checking stage position  Nanda       Pb    moved stage to   5  1 5  checking position with pions    2007 07 15 17 49 George  Manqi  electron run  Nanda    AAA a E up usu Je AA  Pb       10x10 trigger rate  1 8k pps  DAQ rate  50 Hz    2007 07 15 18 26 new stage position    moved to ECAL corner between wafers    CALICE eLogbook   Resul
95. 0 31000  9 4 AHCAL 26 23 22 26 11 16 33 60 28080 31000  9 6 AHCAL 25 22 16 25 30 16 33 59 41800 46000  9 7 AHCAL 23 21 21 23 10 15 33 61 28000 31000  5 1 AHCAL 22 20 32 22 29 15 33 61 28050 31000  5 2 AHCAL 21 19 24 21 9 16 33 61 28000 31000  5 3 AHCAL 20 18 31 18 28 14 33 62 27950 31000  5 4 AHCAL 19 17 27 19 8 16 33 63 28050 31000  5 6 AHCAL 18 16 15 20 27 16 33 70 42050 48000  5 7 AHCAL 17 15 25 17 7 15 33 61 28200 31000  7 0 AHCAL 16 14 29 16 26 16 33 61 28000 31000  7 1 AHCAL 15 13 3 15 6 17 39 69 42050 48000  7 2 AHCAL 14 12 8 14 25 19 33 62 41600 48000  7 4 AHCAL 13 11 18 13 5 19 35 68 41730 48000  7 5 AHCAL 12 10 0 12 24 19 34 67 41720 48000  7 6 AHCAL 11 9 13 11 4 17 34 65 41800 48000  7 7 AHCAL 10 8 2 10 23 19 35 67 42000 48000  9 4 AHCAL 9 719 9 3 20 33 66 42000 48000  9 1 AHCAL 8 611 8 22 20 33 65 41750 48000  9 2 AHCAL 7 5 5 7 218 32 56 41900 48000  9 3 AHCAL 6 4 7 6 21 16 34 64 41700 48000  9 5 AHCAL 5 3 1 5 116 38 68 41950 48000  9 6 AHCAL 4 2 4 4 2016 31 60 41950 48000  9 7 AHCAL 3 112 3 0 17 34 60 41800 48000  2 2 PIN 46 1   2 3 PIN 46 2   2 4 PMT O   2 5 TCMT DAC tcomtl dat 16 33 61 4500 11000   2 6 TCMT DAC tcmt2 dat 15 34 62 9000 16500       1  end of AHC cfg file    2007 07 14 14 47 4 hold scans       pees Pr EA AE ee  sync on  330439 lahcPmLedHgsgS amp ct  n only   mw  configurat ions    2007 07 14 14 49 Allister  Beni  Pion Beam Run after Stage Shift    CALICE eLogbook   Results for search of  20070701 Page 18 of 20    Mangi  o MONOS  Pb                
96. 0 40 so 60 70 80 90 100  beam setting  GeV  beam setting  GeV     resolution  E AE  H ss 1    E     at  x     sE         7   1E   amp 5 D D  6  a  t  5536 a CR O E   Een    File  http   ttfinfo desy de CALICEelog data 2007 31 31 07 2007 07 31T07 15 02 00 gif    2007 07 31 07 29 done    switch of HV  start ECAL saftey    CALICE eLogbook   Results for search of  20070701 Page 20 of 20    pee pata     pa pe p            File  http   ttfinfo desy de CALICEelog data 2007 31 31 07 2007 07 31T07 29 02 00 jpg       0 deg angle and 0 0 position    HCAL Motor control panel    rmt CNN TT    File  http   ttfinfo desy de CALICEelog data 2007 31 31 07 2007 07 31T16 42 02 00 jpg    CALICE eLogbook   Results for search of  20070701 Page 1 of 16       2007 07 27 00 07 Nanda  Roman Pion run       PD    150k events     2007 07 27 00 43 Nanda  Roman pion run       Pb    150k events        2007 07 27 01 17 Nanda  Roman e  run  fees E a IE spe 0 AN Nee er  200 k    2007 07 27 02 07 Nanda  Roman e  run    Pb     40k events        Stopped since we want to center the beam in x in the chambers      2007 07 27 02 18 Nanda  Roman       Pb    Testrun to adjust beam position   Switched trim 06 from 12 0 to 2 0A    2007 07 27 02 23 Nanda  Roman       Pb    Testrun to adjust beam position   Switched trim 06 to 22 0A    2007 07 27 02 26 Nanda  Roman electron run       Pb    switched trim 06 to 25 0A  200 kevt    2007 07 27 03 26 Nanda  Roman  e  run    Pb  Pb    Short runs to adjust beam position            
97. 0 9300        change collimator 1 to   3 3   collimator 5 to   5 5   Current trigger rate 9 5k    2007 07 19 07 18 Mangi  Petr  15 GeV electron run  Nicola       change collimator 1 and 5 to   3 3   trigger rate 4 2k    250k events    CALICE eLogbook   Results for search of  20070701 Page 12 of 22                      4  nes A   Bi cakceonP cabcesan X Moeitor O AXE io elogio 1 Mi cac P cler o Y O    ve    File  http   ttfinfo desy de CALICEelog data 2007 29 19 07 2007 07 19T07 18 38 02 png    2007 07 19 08 27 MORIN  Imad   330661   beambata   10x10      i2ceev  e         1524  air Jair     12 0977  9 0        Collimater 1 and 5 set to   2 2  and collimator 2 and 3 set to   5 5   trigger rate 10 5  few events taken because of beam stop    2007 07 19 08 52 MORIN  Imad  330662  beambata  10x10        126eV Je Joe   air    Jar     2 10  po       31  0 0  Same collimator setting than previous     Trigger rate 3K  Stopped at 50K event due to an increase of trigger rate  11 5K     2007 07 19 09 02 MORIN  Imad  B30663  beambata   10x10       izGev  e  32       Leir      ai     12 194  foo          Col 2 and 3 set to  3 3 mm    2007 07 19 09 51 Mangi  Petr   Nicola   230664  beambata   10x10 Lee Le        520    air    air _  2 59  00      colimator setting    col 1 i     2 2  cgl  53  col 3    lt 3 3  col 5 2 2  2    trigger rate 7 5K  245K events    2007 07 19 10 12 current limits raised    Raised current limit for following HV channels  14  28  amp  51    CALICE eLogbook   Results 
98. 00  29 99 0 00 00 28000 31000  28 99 0 00 00 28000 31000  27 99 0 00 00 28000 31000  26 99 0 00 00 28000 31000  25 99 0 00 00 28000 31000  23 99 0 00 00 28000 31000  22 99 0 00 00 28000 31000  21 99 0 00 00 28000 31000  0 0 00 0 0 0   19 99 0 00 00 28000 31000  18 99 0 00 00 28000 31000  17 99 0 00 00 28000 31000  16 99 0 00 00 28000 31000  3 99 0 00 00 28000 31000  14 99 0 00 00 28000 31000  0 0000 0 0    CALICE eLogbook   Results for search of  20070601 Page 11 of 19                7 5 AHCAL 12100 0 0 00 0 0 0  7 6 AHCAL 119 0 0 0 00 0 0 0  7 7 AHCAL 108 0 0 0 00 0 0 0  9 0 AHCAL 9 7 0 00000 0 0  9 1 AHCAL 8 6 0 0 0 00 0 0 0  9 2 AHCAL 7 5 0 0 0 00 0 0 0  9 3 AHCAL 6 4 0 0 0 00 0 0 0  9 5 AHCAL 5 3 0 00000 0 0  9 6 AHCAL 4 2 D 0 PDP D D 0  9 7 AHCAL 3 1 0 0 0 00 0 0 0  2 5 TCMT DAC_tcmtl dat 34 34 64 15000 15000  2 6 TCMT DAC tcmt2 dat 34 34 64 15000 15000    2007 06 28 21 17 pedestal check run    check dead channels  noisy HABs  etc     HV is on for most modules except for problematics ones with very high current  or non programmable FE     2007 06 29 13 56 A  M  Magnan   debugging run  330007 DEER ES EC ES  330008 beampata     EES muon                1                                   Comment  10x10 corresponds to option  v 21    m 1 option enabled   i e  taking buffering data during spill then transferring between  spills     Muon rate   4100 in 10x10  H6B_EXPTO2 in scaler status   4400 in 20x20  H6B_EXPTO1   DAQ average trigger rate   450 Hz in spill     Stopped after 
99. 0000             2007 07 30 11 21 Beni swap of data disk    changed   data link on calice00 to data01    CALICE eLogbook   Results for search of  20070701 Page 10 of 20       data link for online monitor on caliceserv to data01  diskSpaceMonitor to watch data01  changed cogri script to data01    2007 07 30 11 33 Ariane  Beni  Muon Calibration run  Nanda    TE    Paes  deg       muon calibration  beam dump closed  20x20 trigger rate 5000 pps  DAQ rate 90 Hz    2007 07 30 13 28 Ariane  Nanda 35 GeV pion    ph    10x10 trigger rate 1200 pps  DAQ rate 50 Hz  150 kevt    adjusted DAC setting for LEDs in layer 28  which are currently set manually from 110 to 130  Now we see light in this layer in the online monitoring      only 2 layers without signal left       2007 07 30 14 22 Ariane  Nanda pi  run    Pb  331165 lbeamData  10x10    45GeV pi  518 4mm Air 0 1 20 deg   Pb junk  p    20 deg    20 deg   junk       Pb    64 and 331165  beamtuning  offset in x  changed trim06 from 15  gt 10  gt 5 A       331166   10x10 trigger rate  2 kpps  DAQ rate  60 Hz          170 kevt  2007 07 30 15 10 Ariane  Nanda pi  run  preuve ES EA aie E  Pb  0x10 trigger rate   5 kpps  EA rate  65 Hz  160 kevt  2007 07 30 15 45 Ariane  Nanda pi  run       Ph    closed collimators 1 5  0811 from 12 to 5 mm  10x10 trigger rate  7  4 E  DAQ rate  80 Hz    2007 07 30 16 33 Alex Catherine    Ph  331170 lbeamData  10x10  50GeV 4mm  Pb MES    closed collimators 1 5  0 amp 11 from 6 to 5 mm  10x10 trigger rate  6  
100. 00437  6 GeV     400434  400435    end of 1st configuration 2nd round                                     LED run after 1st config 2nd round to see gain of Layer 6 strip 14                  Pedestal runs   400442     CmNoise    400443     PmNoise    for Layer6 Stripl4     400440     CmLed for inter calibration   400441 PmLed for inter calibration   400445     CmLed to see lpe peak   400446 CmLed to see lpe peak around strip 14     For Layerll Strip15                    400449   CmLed for inter calibration   400450     PmLed for inter calibration   400447     CmLed to see lpe peak   400448     CmLed to see lpe peak around strip 15     end oT  LED runs                                                                                         MIP calibration with 1st configuration after removing tungsten                  400451  400452  400453  400454  400455  400456  400457    end of MIP calibration run                                 Precise detector scan with 1st config after removing tungsten                  400459  81 points entire detector    400460  81 points entire detector    400462  81 points entire detector    400464  30 points  precise scan for X layer strip 5    end of precise scan                                         2nd configuration MIP calibration before putting tungsten                  400468   400469   400470   400471   end of MIP calibration                           2nd configuration precise detector scan before putting tungsten                 400473  81 
101. 007 07 19T02 04 43 01 png    2007 07 19 02 56 Mandi  Petr  25 GeV pion run  Nicola       Trigger rate 3 6k    2007 07 19 03 46 Mandi  Petr   Nicola  2007 07 19 03 49 Mandi  Petr   Nicola  2007 07 19 03 50 Mangi  Petr   Nicola  830053  ehecmnedvealibseaa E  2007 07 19 03 53 Mandi  Petr   Nicola  83005  ehePmnedvealibseaa LI OE    2007 07 19 03 56 Mangi  Petr   Nicola    CALICE eLogbook   Results for search of  20070701 Page 11 of 22        330655 o EEE EE  2007 07 19 04 41 Mandi  Petr  25 GeV electron run  Nicola    Pb    Trigger rate 2 9k    First set the target to be lead  But then read from monitor the pure pion signal  so decide to  remove  the lead     2007 07 19 04 56 Mandi  Petr  electron run 25GeV  Nicola    Changing the target to air  Then we get a mixed run with muon  pion and electrons  v wm AX  ECAR lues   mower   Reacher  Partes   Suerg   Comi Checa   Stats   sasagi                11111561                                   F E cech  er  eer E             teas ssh   MiadResdout eee     y as WE  B E                  ds  E    fe       E e  s   K   ir a E  See 10 827 000 078 Te A 12000    fuv 200082 000 Tee 1 me     ya    Ja x X Monitor D   e D ese  File  http   ttfinfo desy de CALICEelog data 2007 29 19 07 2007 07 19T04 56 59 02 png  2007 07 19 05 39 Mandi  Petr  electron run 20GeV  Nicola   230658  beamData  10x10 ice fe      523 Kr fir  0 7368  0 0 ESO    2007 07 19 06 30 Mangi  Petr  18 GeV electron run  Nicola   230659   beambata  10x10   Eier   fe    522 fi Ki  
102. 007 07 27 23 33 Lukas  Alex 10x10 pi        Pb     Collimators C1 C5 C6 C10 and Cll adjusted to  13 mm      1900 pps    Run cancelled because of X axis adjustment     CALICE eLogbook    1of28    http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log       1088 match es  for text or title   start date   01 07 2007  end date   31 07 2007       Previous 0 1234 56 7 8 9 Next       A    Logbook entry   CALICEelog data 2007 30 24 07    24 07 2007  23 18    30936 peaspaca  scev e zs  i 20200     Collimator 1 5 6 10 11 reduced to   4  rate 900 Hz       Nicola  Hengne  Manqi    during the run changed to   5  at 50k events   rate 2 5k    Logbook entry   CALICEelog data 2007 30 24 07  24 07 2007  22 09    530957 EE e paper bins C260    Nicola  Hengne  Manqi       collimators 6 and 11 put at  15  Rate 1 28 K  The beam looks very clean in the Veto little presence of double particle  neglectable     Logbook entry   CALICEelog data 2007 30 24 07    24 07 2007  21 45       Nicola  Hengne  Manqi Back to the electron runs       EE RUS zs pis EENEG    trigger rate too low   lt 300  we increase the collimator 6 and 11 from   10 10  to    20 20   Now  trrigger rate is 3 7k    Logbook entry   CALICEelog data 2007 30 24 07    24 07 2007  21 37    e changed sequenceDelay in EmcConfiguration to 13    e build runner    changed link of AHC cfg to AHC CERNO7 beamhold 10x10 cfg    Beni changed trigger setup to 10x10    Logbook entry   CALICEelog data 2007 30 24 07    24 07 2007  17 45
103. 007 29 19 07 2007 07 19T11 10 51 03 png  2007 07 19 12 11 MORIN  Imad    ahcCmNoise  e 20000  ahcPmNoise  e 20000    ahcCmLedVcalibScan  ahcPmLedVcalibScan       CALICE eLogbook   Results for search of  20070701 Page 14 of 22       2007 07 19 12 58 new position    moved stage 2 cm down    HCAL Motor control panel    actual position target position    x      8 31 20 ox   Y EE  do or  Y   5o    CH lost X reference point        Q lost Y reference point     O lost 2 reference point    Q stage in motion  o marras  operation get Pos    Sen 1 alow 1 pulse free   extern STOP   ref 8   239 3692399 959    OF ser    H   200 gi         File  http   ttfinfo desy de CALICEelog data 2007 29 19 07 2007 07 19T12 58 03 00 jpg    2007 07 19 12 59 MORIN  Imad   330674 Jpeambata   iox10    8GeV__ pi      1525   mPb  air  2 99  073  _     beam position control with pions   Used 525  obstacles changed   col setting       col 1 s  20 20   gol 2 t  3 3   col 3   s  3 3   col 5 GES 35  2007 07 19 13 00 MORIN  Imad   B30675 athe gat  2007 07 19 17 58              330680 10x10 Pes  sz  impo far porte       330681 fbeamData   10x10  8GeV pi  E   4mm Pb 1 6 2 5 0   3 170k  waiting  for  cherenko  pressure    2007 07 19 18 17 Oliver Allister Hengne       330682 lbeamData   10x10  8GeV pi  525 8mm Pb air 2 99 0 43 252k  Events    10x10 trigger coincidence  4 2 kpps  DAQ rate  75 Hz    Collimators COLLO2  COLLO3 and COLL08 have been increased to    10mm to increase trigger rate     2007 07 19 18 36 SiPM curr
104. 0070201 Page 18 of 21          File  http   ttfinfo desy de CALICEelog data doc CERN docu 2007 02 27T14 07 00 02 paf          2007 02 27 14 41 Marius DAQ online histogram reset    Usefull for the DAQ online histograms   to reset a histogram and the change the setting  use following commands     hstReset  1  hstReset  1  hstReset  1  hstReset  1  hstReset  1       reset fuer jede acquisition     reset fuer jede config    gt  reset fuer jedes event   gt    gt        never  reset fuer jeden run    k   OO b  A GA    to force  now  hstReset  r    a reset  use     la 2007 02 27 14 41 Marius DAQ trigger choice  v     v option is used for the trigger decision  in the meaning which trigger is used in the run     v 26   coincidence of 10x10 scint    v 28   coincidence of 3x3 scint    v 19   coincidence of 100x100 scint    v 2   single signal of 20x20 scint    v 9   external pulser for trigger checks  normally not used     this was valid for all beamRuns at CERN in 2006    la 2007 02 27 16 49 Marius HV drop on its way to the SIPM  We have to deal with a HV drop from the HV coming out of the HV distribution box on its way to the  SIPM     Measurement   HV set to 39 7 V Q SC panel     measurement point explanations in enclosed pdf   1 3 39 2V  multimeter correction   0 39V  HV   losses  1 4 OV  SIPM off  0 08 0 15V  SIPM ON  100kOhm voltage drop    39 4V   HV   DACoffset  no multimeter correction needed  because no 100kOhm in parallel     2 4  2 3 0 2V   DACoffset  DAC at 0 bit is not OV 
105. 0070701 Page 5 of 19    330383   Terum Sk    EE i ik i l    228 1k 5882 281 9    228 9k 4753 251 1   L  i layer  12    330119 B5             E28 7k HE 1593 ERA    14 07 2007    RunNum Energy Beam event Num Main Peak ain Sigma Mean Comments  position Peak Width    330220 EE EU 177 5 BN  330422 5      EES 5588             73 7             B                  L                    330423 5     EES 2458         pr              3                 L                      50k a  gt   7   E Ecal  layer    A SEEN EB m ie  330429 te 4761 245 2                   244 2k 3708 193 6    178k 3099 17 98    232 8k 2525 15573       233 7k 1905 138 1                15 07 2007   For all the runs this day    Beam  very low hadron components   Significant wafer gap effect     229 7k    peo          layer    40 228 4k          50 228 4       layer    50 228 7k   1 Slightly  noise in   layer 19  ecal     EN    EECH EE Ee  E3953          8         R5                330482 15 229 5k L 137 2 ER Slightly  noise in ecal  layer 2           330188 ES             153  9k 5721 511 8 DEE  330493 ERR EH DR EES Kl          CALICE eLogbook   Results for search of  20070701 Page 6 of 19    330199 59     EE BEE 103  9               L              4  330501 ES              i51 3k mu a E    330506 55             i51  5k    330508 15 200k     HT  3   ED      Noisy layer 2    16 07 2007 18 07 2007  No electron runs       19 07 2007  Many hadrons in the beam  layer 19 always have very large noise    RunNum Beam event Num M
106. 07 31 04 15 module 1A current    The current of module 1A is unsteady     CALICE eLogbook   Results for search of  20070701 Page 18 of 20                2  ccr a A    4 50  40  20   00  B    O O O CS O rrr O v n  200 00 00 02 00 04 00  OF 2007 31 07 2007 31 07  2007 31 07  2007  File  http   ttinfo desy de CALICEelog data 2007 31 31 07 2007 07 31T04 15 02 00 jpg  2007 07 31 04 20 Beni  Lukas  2007 07 31 04 24 Beni  Lukas    3312 11 beamData  10x10  50GeV e         1 8k pps   200k ev    2007 07 31 05 07    2007 07 31 05 41 Beni  Lukas         7 3k pp      pps    stopped after 11 5k ev   adjusted TRIMO5    8 and TRIMO6    8    2007 07 31 05 45 Beni  Lukas    Pb            stopped after 12k ev         further adjustments  TRIMO6    4 and TRIMO5    4  2007 07 31 05 49 Beni  Lukas  EPA PEA ES SMA PA  Pb    7 1k pps   150k ev  2007 07 31 06 16 Beni  Lukas       Pb     4 5k pps   aborted because of position corrections  TRIMO5    5 and TRIM06    5    CALICE eLogbook   Results for search of  20070701 Page 19 of 20                2007 07 31 06 23 Beni  Lukas  ESPA PE AA pepe  Pb   4 5k pps   150k ev  2007 07 31 06 50 Beni  Lukas  Er Jeep pee cce TE E pep pe  Pb   2 8k pps    stopped after 11k ev    position adjustments  TRIMO5    12 and TRIMO6    8  2007 07 31 06 58 Beni  Lukas  pepe qe uc  Pb   150k ev  2007 07 31 07 15 caliceon calices eststs for pion runs  fitted energies for pion runs    AN O              A                                               30 40 50 60 76  3  5 100 d 3
107. 09 2007 11 39    http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log                electron run    Allister  Robert  Riccardo    27 09 2007 11 39    CALICE eLogbook          o    gt    G   0   Uu   Q   O   o       v   y   v   ES   3   x    amp    d   EE   O      S  2   p  q     ex  A e  ES T  Q  0  S f    Q  Ze Oo   s  Kai on 3  a 2    E og    T ND oO  D oo  gt   O 07          o N  a  d    u o  Q go 2o  MA o Oo  po y  3  e D y  oc oo 4  S pp o  Mm E S H i  oO c oo 0    O   DOO   ch   argid     3337  o  o gt  m  5h o   8 H 3 Cm D CT Le sim tores  Q m   o     Aaleas          Logbook entry   CALICEelog data 2007 30 23 07    23 07 2007    17 of 28    CALICE eLogbook http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log       330825 beamData 10x10  12GeVle  524lairlair l 86  6 0 257kEvents    10x10 trigger rate  9 6 kpps   DAQ rate  97 Hz   Collimators reduced a bit just to reduce the 20x20 trigger rate which was slightly above  llkpps        New settings   COLLO1 COLLO5 COLLO6 COLL10 COLL11     5 5mm  COLLO2 COLLO3 COLLO8      4 0mm   New trigger rates    10x10 7 3 kpps   20x20 8 5 kpps                            Logbook entry   CALICEelog data 2007 30 23 07    ay vu ES Allister  Robert  Riccardo new shift       Logbook entry   CALICEelog data 2007 29 22 07      22 07 2007 i   Guassian Fit for electron run  SY 22 13    Em E ppn    Oliver  Hengne  Manqi 330821 330822          Logbook entry   CALICEelog data 2007 29 22 07     y GC Hengne
108. 0deg  X Displacement neccessary to get the beam back to ECAL  amp  HCAL center  This  is the new 0 0 position     HCAL Motor control panel    actual position target position    vlses  m mx  Y    19 97 ka    ori   Y     O lost X reference point         Q lost Y reference point     O lost 2 reference point    Q stage in motion  O mana  operation get Pos    Sp   Fast   pulsezfree   extern STOP   ref 24   594 390 6 4 932    OFF set     B  900 ge         CALICE eLogbook   Results for search of  20070701 Page 18 of 20       File  http   ttfinfo desy de CALICEelog data 2007 30 26 07 2007 07 26T19 22 02 00 jpg          2007 07 26 19 24 Lukas  Alex   mw ESP E T E FE E FR E  2007 07 26 19 38 check of new 0 0 position   Result of new 0 0 adjustment         hHealBarycenteflayergtY    is lll able ren rr porem per perennem       6    DD 12 14 b    zr Pal    File  http   ttfinfo desy de CALICEelog data 2007 30 26 07 2007 07 26T 19 38 02 00 gif       2007 07 26 20 47 Lukas  Alex    330977 beamData  10x10  45GeV loi  515 4mm Air 0 1 10deg   Pb 150k ev       2007 07 26 21 07 Lukas  Alex  330978 beamData  10x10   80GeV pi  510 4mm Air 0 1 10deg   Pb 150k ev    2007 07 26 21 19 Lukas  Alex    Pb          330979     peampata   oxi0   40Gev__pi      Bii     fme ir         i 0 0 ET    CALICE eLogbook   Results for search of  20070701 Page 19 of 20       Pb stopped  after  about  because  changes  during  this run    We had to adjust Collimator settings for C1 C5 C6 C10 and C11 from   12 to   6  and
109. 1 pore  DAQ rate  85 Hz       2007 07 30 17 13 Alex Catherine 100 GeV pion run       Pb    closed collimators 1 5  from 5 to  10x10 trigger rate  7  2 SEN  DAQ rate  88 Hz    2007 07 30 17 22 change current threshold    Change current threshold for module 15 B    CALICE eLogbook   Results for search of  20070701 Page 11 of 20    y senitoribg Switch On   Off  Voltage  V  Current  a     V   Initial D    SEN fi 10 00   ee   or f     110 00 M  fae   OFF   o    10 00 E fe   rr ff   d   10 00 fine Jer fh ie 2   1 1000 200 53 a     10 00 P fae OF m           1  t  i  1    Dn A add A  10 00 14 00 18 00 22 00  e 30 07 2007 30 07 2007 30 07 2007 20 02 2000    Trend to sog j     mo a Mi       S      r W settings   file actions     renable   disable AL chienels         set   cow   EMT     ele    dimble      M Crate OFF    a a a  SE   w2 Jwi we      Aer Wi volves to file   eee            10 00      10 00       File  http   ttfinfo desy de CALICEelog data 2007 31 30 07 2007 07 30T17 22 02 00 jpg    2007 07 30 17 33 Alex Catherine     35 GeV pion    Pb    10x10 trigger rate  1 5 kpps  DAQ rate  55 Hz       2007 07 30 17 33 Erika 20 deg HCAL energy    Summary of pion energy deposited in HCAL for 20deg angle  red dots  compared to 0 deg  blue angle      From the last muon run the muon peak is found at 50 mip with 20 deg angle  noise of 5 mip   while for the 0 deg angle it was more around 42 mip  same noise         u 25          225         2 2   E 0 25 50 100  ul   A E GeV   gt    g   O   Lu 
110. 10 trigger coincidence  2 2 kpps  DAQ rate  80 Hz    2007 07 18 21 37 Beni  Oliver  10 GeV pion run  Mangi    CALICE eLogbook   Results for search of  20070701 Page 9 of 22       330643 lbeamData  10x10  10Gev  oi  520 4mm air 2 613 250k  Pb events    2007 07 18 22 28 Beni  Oliver  12 GeV pion run  Mangi    330645 lbeamData  10x10  12GeV  oi  524 4mm air 2 613 250k  Pb events    2007 07 18 23 56 Mangi  Petr  Fabi 15 GeV pion run    Pb    Change the absorber to 8mm Pb to reduce the electron signal in ECAL  see attached plot          Change the collimator 2  3  8 to  6 0   6 0 to increase the trigger rate   beam off for about 15 min during this run     trigger rate now  1 70k       A em Gi  ELE  FEE   E P Pm  P  o2  EES SE  di     ss   s  E  Ss  DNA  a  8  zl  D  c  a    1000 1500    500             o    o    SE      T Ta       File  http   ttfinfo desy de CALICEelog data 2007 29 18 07 2007 07 18T23 56 52 04 ps    CALICE eLogbook   Results for search of  20070701 Page 10 of 22       2007 07 19 01 07 Mandi  Petr  18 GeV pion run  Nicola    TAPA Pep py fe p pap PA  Pb       Change Collimator 2  3 to   6 6   trigger rate 3 1k    2007 07 19 02 04 Mandi  Petr  20 GeV pion run  Nicola    Pb       Change Collimator 2  3 to   6 6   trigger rate 4 55k  ECAL noisy layer 2  19    ECAR lee   mower   Teacher  Parts   Oipisy30   Cronicas   Stn   atop   Peigonco   pecemen   Z  zsgcpeen                                       OE EXT    File  http   ttfinfo desy de CALICEelog data 2007 29 19 07 2
111. 1058    f  i                PITT  d          w   HN Crate  FF     RL aime peeable   disable  LL amis      N2   ws   ws    sw   two   BT     mele   gr  is      Fe  i e   due Hi values to file   _ damp AL values  position     File  http   ttfinfo desy de CALICEelog data 2007 33 19 08 2007 08 19T20 52 03 00 jpg             2007 08 19 21 24 caliceonOcaliceheiepdA berckannel slot17  fed   CALICE user     this one is not showing an error  red     hv channel 28  layer14  module 16    21259  r  i Z        em       emm O Leg       i   1o  dod o   poro rra CC RR ET CI g  g  13700  13700 ZU DU 21 00  J 08 2007 13 08 2007 13 08 2007 19 09  2007       File  http   ttfinfo desy de CALICEelog data 2007 33 19 08 2007 08 19T21 24 03 00 jpg    2007 08 19 21 43 Daniel  DAC reloaded  Anne Marie  Petr    2007 08 19 21 47 Daniel   Anne Marie  Petr    350276 lbeamData  10x10  10GeV e  626 air 6mm unknown 38k   Pb 20deg    CALICE eLogbook   Results for search of  20070801 Page 9 of 22             2007 08 19 21 52 Roman Shipping out slw  log and sum files    FYI    I have just started to ship out the sum  slw and log files for this years  cern running into the desy dcache    Thus you will find from now on a process called grid cp misc sh    I will monitor these transfers     Roman    2007 08 19 21 54 caliceon calicehedtstd2  fe0n ch   CALICE user     reloading dac settings brought this channel back  thanks Beni         2  2  S  n  i          adas        13 00 SH 21300 eet  13 08 2007 13 08  2007 13 0
112. 118 lbeamData only a few  minutes   330117 beamData 100x100 mu  combined HCAL amp TCMT  0 0 250k events   Sa pme AA PS basta r ees      2007 07 06 05 08 Alex  Yoshi  AM voltage  0 3  hominal 0  3V  nominal 0  3V  pas DT TT H   WT inma  nominal 0  3v  nominal40            lbeamData 00x100 mu  combined HCAL amp TCMT  22k events   Reload HCAL   330127 ahcPmNoise 00x100 mu  combined HCAL amp TCMT  20k events  ahcGain combined HCAL amp TCMT  e                                         HCAL LED has high level at layer  15   16   17  9  0  3   34   36   37 and  38     CMB Temperatures    Module 34   CMB 37   Pos 35  T2 T3 T4 T5 117   C   Module 1   CMB 14   Pos 28  T2 T3 T4 T5 117   C   Module 36   CMB 33   Pos 31  T5 117   C   Module 21   CMB 24   Pos 19  T5 117   C   The temperature on the other sensors on these Boards is between 25 and 30  C   All other CMB have temperatures between 25 and 30   C for all sensors    This was already noticed in the past and is due to wrong sensors     CALICE eLogbook   Results for search of  20070701 Page 4 of 24       Si Lekt      ease  Fuego a aria  PHASE ERR LC wegl Ta  tes  Ls H He    ed          ILLAE    hiHcalPedRMEParL ayer O C eege   e IT E La E el  I     Er co      per  r            e a   dai  E  H     E  File  http   ttfinfo desy de CALICEelog data 2007 27 06 07 2007 07 06T05 08 29 04 eps  2007 07 06 08 48 Michele     Evgueny  Hengne  e Det poxo pe DT Keck Kies    o  nominal 0 JM events  paa pr PTT TT Bamaepo fevents        nominal 0  5v ev
113. 13 14 15 and 16   new values for 9 10 11 and 12    2007 07 24 10 53 Beni  Petr    Laurent    330908 beamData  10x10  6GeV le  pion  23 air air 3 0 0 0 140k  Events       trigger rate 10x10  1 3k  DAQ  50 Hz   pion  rejection    99    e  detection    98     CALICE eLogbook   Results for search of  20070701 Page 6 of 20       ratio e  pion    9 1    2007 07 24 10 57 Beni  Petr    Laurent    330909 lbeamData  10x10  6GeV e  pion   23 air air 3 0 43k  events       no beam due to server rebooting    2007 07 24 11 11 Beni  Petr    Laurent    330910 lbeamData  10x10  6GeV e  pion   23 air air 3 0 159k  Events  Ek    trigger rate 10x10   DAQ  50 Hz  spike in hcal power line  run stopped for relod Hcal DAQ setting       2007 07 24 13 42 Beni  Petr    Laurent   2007 07 24 13 44 Beni  Petr    Laurent       330914 lbeamData  10x10  6GeV e  pion   23 air air 3 0 30k  Events   gt  1 4k    trigger rate 10x10        DAQ  50 Hz  2007 07 24 13 45 Beni  Petr    Laurent  30919 EE O O O  30918  ahcPmLedVcalibScan SO   30917 jancCmLedVcalibScan          30916  30915  2007 07 24 14 07 Beni  Petr   6 GeV run  Laurent        330920 lbeamData  10x10  6GeV e  pion   23 air air SU  afterwards    The summer students started without notifying us  last 20  of run are with unknown beam settings    trigger rate 10x10  1 4k  DAQ  50 Hz    2007 07 24 14 43 caliceon 2caliceh  nisi amp tseecuntent threshold   CALICE user     increase current threshold for channel 62  module 36 B from 200 to 210    CALICE eLogboo
114. 16 48 Daniel  Anna cerenkov pressure  Marie  Petr  Beni    Got call from control room     They are trying to contact an expert to solve our problem setting the cerenkov pressure  but having  problems getting a response   they will continue trying     2007 08 19 20 37 Daniel  15 GeV e    Anne Marie  Petr    pep          for both runs  the beam was unstable       Position seems at  10 0 more than  6 0   see plot below     hHcalBarycenter5layersXY    Er  a 65 Entries 68517      Dn Meanx 3575  60 Meany 47 03  d RMS x 2 012  2 108  255        25 30 5 40 45   lt  uleft right   gt  index I  cm     File  http   ttfinfo desy de CALICEelog data 2007 33 19 08 2007 08 19T20 37 36 07 eps    2007 08 19 20 52 2 hot hv channels    CALICE eLogbook   Results for search of  20070801 Page 8 of 22        ga  RIES ES MEL E nl    98 898   a w MIS    ep s 36 330      129 MANO aE rj E 116 W IRSA s SMA    101 w poe g 38 188 145 W ge PIT  zg Mia  31 pw MIA 192 MIRTA 23 35     106  Rap   5  13 w 98188   17 W ajere   mr osa     n ARA   12  33208   149 W SEA   10 4 w IRA    ou Epa   i22 waga   142 W gae      2 wa  yew aa   40   e RIA   TAS   12    Sms    1 NO mm y anne   149 XS   ias w rare    145 wean o   w GLA   155 M 900A   i175 SES  3 Kaap      19   m opo w 701588  103 SA   y Gs   103 APTA oe W TTRSA      3100 OSB   m4 sege 55 M 28   ies 70 998    102 DU 199 W 424A i25 MZA 354 W 7940A   o OS       SR     ER  ee     MY ramp  MM   vum ofr E mm    um   weer       1na  ane  HA  428  SpA  spe  THA  H  E   
115. 18 50 S  Uozumi 4 GeV tuns  400502 beam 4GeV direct scanned  fiber Po 10  stage scan  2007 03 17 18 50 S  Uozumi 4 GeV Dun   400502 beam 4GeV direct scanned  fiber E 10  stage scan  2007 03 17 22 34 S  Uozumi 6 GeV fun   400504 beam 6GeV direct scanned  fiber E 10  stage scan  2007 03 17 22 34 S  Uozumi 6 GeV tuns  400504 beam 6GeV direct scanned  fiber E 10  stage scan  2007 03 18 09 19 S  Uozumi 6 GeV fana  400505 beam 6GeV direct scanned  fiber n 10  stage scan  2007 03 18 09 19 S  Uozumi 6 GeV tuni  400505 beam 6GeV direct scanned  fiber us 10  stage scan  2007 03 18 09 20 S  Uozumi 6 GeV zn  400506 beam 6GeV direct scanned  fiber E 10  stage scan  2007 03 18 09 20 S  Uozumi 6 GeV SR  400506 beam 6GeV direct scanned  fiber E 10  stage scan    CALICE eLogbook   Results for search of  20070301 Page 16 of 76       2007 03 18 09 35 S  Uozumi 4 GeV run   p q  A qe qr qo Ems C   fiber 9 10  stage scan  2007 03 18 09 35 S  Uozumi 4 GeV tuno    400508 beam 4GeV direct scanned  fiber HE 10  stage scan    stopped at 50th point because beam is gone  due to HERA injection  will take  2 hrs     2007 03 18 10 53 S  Uozumi 5 GeV run  400509 beam T1L 5Gev direct 2 5V scanned 36  fiber 9 10  2007 03 18 10 53 S  Uozumi 5 GeV run  400509 beam T1L 5Gev direct 2 5V scanned 36  fiber 9 10  2007 03 18 17 37 S  Uozumi 5 GeV run  pq juu eu  fiber i 10  stage scan  2007 03 18 17 37 S  Uozumi 5 GeV run  p qe qp qe Jess   fiber 9 10  stage scan  2007 03 18 20 46 S  Uozumi 4 GeV run    beam    25 20 
116. 2 07 2007    m 0657 Ricardo Nicola Kieffer    330795 beamData 10x10  20GeV e  523 air air  0 76  3  43       260k  Events    Logbook entry   CALICEelog data 2007 29 22 07    ay n m Ricardo Nicola Kieffer          Events    Logbook entry   CALICEelog data 2007 29 22 07      22 07 2007  rd 05 42       Ricardo Nicola Kieffer    Logbook entry   CALICEelog data 2007 29 22 07  22 07 2007    zy 0530 Ricardo Nicola Kieffer    330792 ahcPmLedVcalibScan       24 of 28 27 09 2007 11 39    CALICE eLogbook    25 of 28    http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log       Logbook entry   CALICEelog data 2007 29 22 07    05 28 Ricardo Nicola Kieffer    330791 ahcCmLedVcalibSscan    Logbook entry   CALICEelog data 2007 29 22 07    22 07 2007        0526 5 26 Ricardo Nicola Kieffer       Logbook entry   CALICEelog data 2007 29 22 07  22 07 2007       05 25 Ricardo Nicola Kieffer  330785 fahccaNoise     Logbook entry   CALICEelog data 2007 29 22 07  22 07 2007  04 53       Ricardo Nicola Kieffer       Events    DAQ rate  94 Hz    Large contamination in e sample  Run repeated at 25GeV e      Logbook entry   CALICEelog data 2007 29 22 07    22 07 2007  03 20       Nicola  Riccardo  Robert       Events    DAQ rate  94 Hz    Logbook entry   CALICEelog data 2007 29 22 07    22 07 2007  01 56       Ricardo Nicola Kieffer    Events    DAQ rate  88 Hz       Logbook entry   CALICEelog data 2007 29 22 07  22 07 2007  01 13       Ricardo Nicola Kieffer       Events    DAQ ra
117. 2 beam TIL 1Gev direct 2 5V scanned 36  fiber 9 10  stage scan    2007 03 17 10 42 S Chang Detector Uniformity along the 36 Beam Positions   Calibration Run 400262    Calibration Run Detector Uniformity along the 36 Beam Positions    3GeV Calibration Run      2 100       3 GeV          90  80  70  60    ADC mean total energ    50  40  30  20       0 5 10 15 20 25 30 35  Beam Position    File  http   ttfinfo desy de CALICEelog data 2007 11 17 03 2007 03 17T10 42 26 00 jpg    2007 03 17 10 42 daniel tracking code    I made a first iteration of the tracking code and alignment   I have not yet tested it very carefully     you can find the code required on flcscecalana in  home jeans trackingCode  it runs on the slcio data     StraightTrackImpl is a class which stores useful data for our straight tracks  the usual class is  designed for helices        DriftChambertoTrack is the code which does the tracking and fills the track collection    these files can be put in the hcal v00 01 18 calice reco src directory  don t forget to include  them in the Makefile am     track steer shows which parameters to use for DriftChambertoTrack    at present it       considers only the first hit in a chamber in the case of multiple hits      ignores the y measurement of layer 4  nearest the ecal module  which behaves somewhat strangely       requires at least 3 hits in x and 3 in y    has a guess  a much inflated 1mm  as track point error  only used in chisq calculation       sometimes removes a hi
118. 20 000 events    Trigger configuration at the moment    the trigger we used is the coincidence of the backplane VME input 5 and 6   corresponding to the two 10x10 triggers   v 21     the 20x20 digital trigger is in input 9 of the back plane   the 4 plates of the veto counter are in trig  1 4   only one 100x100 muon counter is in the line  in the back of the TCMT  and connected to backplane  input 7     Still no analog readout for the 20x20 counter     2007 06 29 16 25    uon  A 2007 06 29 16 52 Nanda comparison between 10 and 50 GeV secondary beam    Comparison of HCAL stand alone electron data from august   september  06 from 10 and 50 GeV  secondary  beam  All runs are handled exactly the same concerning trigger selection  3cmx3cm trigger    on    100cmx100cm trigger of   calibration to mip scale  cut at 0 5 mip  saturation  correction   Used runs    320678   10 GeV from 50 GeV secondary beam       320671   15 GeV from 50 GeV secondary beam  320666   20 GeV from 50 GeV secondary beam  and   320603   10 GeV from 10 GeV secondary beam  320585   15 GeV from 10 GeV secondary beam    320608   20 GeV from 10 GeV secondary beam    The energy sums already show significant disagreement   The 10 GeV seconmdary beam runs always end up at a higher mean value  and a worse  resolution  The bigger the energy  the bigger the difference     The reason for this is shown in the number of hits above threshold plots   While they are in reasonable agreement for 10 GeV  the spectrum of the 20  Ge
119. 20deg    Pb set  wq pq qc d  a  331135 lbeamData  10x10  12 GeV fpi  524 Air 8mm  5 20deg     new noisy ECAL layer  20    CALICE eLogbook   Results for search of  20070701 Page 1 of 20             2007 07 24 01 29 Beni disk space monitoring    I installed a cronjob on caliceana that checks the free disk space   If the free space goes below 15  Roman and me will be warned by email  If the free space is less  than 10  a warning message will appear in the elog     The cronjob runs as caliceon on caliceana  The script is in  home caliceon diskSpaceMonitor     If the data partition is changed  the disk to be monitored has to be changed  too  Please edit the  line monitorDisk    in the checkDataDisk sh script     2007 07 24 01 52 Allister  Robert  electron run  Riccardo       330882 lbeamData  10x10  12GeV e  524 air air 1 87 56kEvent si   stopped  since  not  centered  in x   i i L 87 Een      330883     peanpata  10x10  12600               123      gt  dix     ES    10x10 trigger rate  5 4 kpps  DAQ rate  87 Hz  TRIMO6   0amp       2007 07 24 01 59 Allister  Robert  Guassian Fit for recent electron runs  Riccardo  Guassian Fit for recent electron runs  Run   E GeV  sean  mips   30874   98 075  30876 1329  30883 1587  30886 1928  2302  2007 07 24 02 38 Allister  Robert   Riccardo    a NM Pf   ff fd          330885 lbeamData  10x10  15GeV e  524 air air 0 98  tuning   just a  few  event    S  330886 lbeamData  10x10  15GeV 524 air air 0 98 beam  centered   250kEventis    TRIMO5 is 
120. 21       File  http   ttfinfo desy de CALICEelog data doc CERN docu 2007 02 27T09 30 59 00 zip    la 2007 02 27 08 00 Marius AHC cfg august run period    this files gives a overview about the used AHC cfg files during the august run period      a small comment in the first line gives an idea why a new file was needed     the list starts from the 25 08 06 since before only debugging runs have been taken     most of the changes are due to TCMT vcalib value changes  therefore no mapping change in the  readout      files  where only the TCMT frontends were activated a left out  but they are in the zip file in  the entry before    CALICE eLogbook   Results for search of  20070201 Page 9 of 21          File  http   ttfinfo desy de CALICEelog data doc CERN docu 2007 02 27T09 27 16 00 pdf    la 2007 02 27 09 00 Marius DAC files for august      enclosed the used DAC files for the august run period      belongs to the next entry  where explanations are given    CALICE eLogbook   Results for search of  20070201 Page 10 of 21       File  http   ttfinfo desy de CALICEelog data doc CERN docu 2007 02 27T09 16 21 00 zip    la 2007 02 27 09 03 Marius DAC files for august run period    DAC files for august      gives a overview about the HV settings     shows which DAC files for which module have been used      some comments    The mentioned DAC files are in the zip file in the entry below this one   A new DAC file have been used from the date  which is specified in the filename  on   DAC modu
121. 21 24 George  Tao  electron run  Nanda    Pb       DAQ rate  50 Hz   1 5k pps in 10x10 trigger coincidence    2007 07 14 22 42 George  Tao  gain run  Nanda   2330453       JeheGain 2 2  2007 07 14 22 42 George  Tao  electron run  Nanda       Pb    DAQ rate  50 Hz   1 5k pps in 10x10 trigger coincidence    CALICE eLogbook   Results for search of  20070701 Page 1 of 25       2007 07 08 00 22 Fab A new check for the shifters    Please add this new check to the checklist that needs to be performed during the shift    check that all the CRCs are properly read out by the DAQ    From calice00  open an xterm and find the last run in ECAL and AHcal      for ECAL  look in   data00 data cernEcalJul07 1og  and issue the command     gt  grep alive   data00 data cernEcalJul07 1log Logxxxxx out icalice01    check that the time stamp of the log file corresponds to the latest of the startUps     If everything is OK  you should see the following    CrcReadout  ctor   PCI card 0 crate Oxec slot 5 found alive  CrcReadout  ctor   PCI card 0 crate Oxec slot 7 found alive  CrcReadout  ctor   PCI card 0 crate Oxec slot 9 found alive  CrcReadout  ctor   PCI card 0 crate Oxec slot 15 found alive  CrcReadout  ctor   PCI card 0 crate Oxec slot 17 found alive  CrcReadout  ctor   PCI card 0 crate Oxec slot 19 found alive      for HCAL  look in   data00 data cernAhcalJul07 log    gt  grep alive   data00 data cernAhcalJul07 log Logxxxxx out_icalice02   check that the time stamp of the log file corresponds t
122. 24  Schaetzel    202260 ahcBeamStageScan  v 13 3 GeV StagePositions ig nercore dat  but flipped  Orientation    machine problems  data useless       The shutter closed automatically due to machine problems at some point since run start   Unfortunately the shutter does not open automatically when the machine is fixed       design error  DESY II and DESY III Tue J  DESY II e Current  mA  DESY III p          oe oA  ma  A    LILOLIL IISI  IL ee eee ee ee a  10 12 14 16 18       File  http   ttfinfo desy de CALICEelog data 2007 03 15 01 2007 01 15T19 14 13 02 gif    2007 01 16 09 33 Sebastian waiting for electrons  Schaetzel    With positrons the trigger rate at our experiment is 2Hz   With electrons the rate is around 50Hz  According to John Maidment e  and e  orbits are not  equal after switching magnet polarity due to permanent magnetic stray fields     CALICE eLogbook   Results for search of  20070101 Page 3 of 5           We cannot use positrons  Electrons cannot always be delivered due to other users  requests     2007 01 16 10 08 Sebastian MIP calibration module 24 reverse orientation with  Schaetzel CMB    202261 ahcBeamStageScan   v 13 3 GeV StagePositions  ijnercore dat  but flipped  Orientation    useless data  beam lost due to machine problems shortly after starting the run       2007 01 16 17 53 Sebastian MIP scan  Schaetzel    202266 ahcBeamStageScan  v 13 3 GeV StagePositions  ijnercore dat  but flipped  orientation    at around 18h00 switch over to positrons to fil
123. 25 40 40 46 46  HAB 10 DACs   76 112 107 87 102 46 97 51 61 46 66 107 81 87 51 87 107 133  HAB 11 DACs   28  25 VI 128   128 138 122  139 122 122  122 153  15 138 143 138  122  Slot number   5  CRC Component   6   FE6  Label   2   0x02   05000000010   read  AhcVfeStartUpData  print    Verification data   2863333375   Oxaaaaffff   0510101010 10101010 ts Ke ps i ge eget Ve Ta ag gt  HAB 0 DACs   38 155 255 255 138 138 138 122 133 127 255 250 153 143 143 143 250 250  HAB 1 DACs   26 122 235 215 230  115229  179 189 174 106 FIL 123 215 179 215 123 133  HAB 2 DACs   250 117 127 127 245 117 127 107 117 127 126 107 1 123 120 122 63 63  HAB 3 DACs   23  40 92 TOT 123  RP dd  1238 128 402 dd 1007  2228 2121 123 107 126 119  HAB 4 DACs   33 148 253 255 117 251 62  93 119 125 250 133 122 131123  122 2249  127  HAB 5 DACs   229 188 184 7 179 250  138 255 248 251 191 71 9 7 143 221 205 223  HAB 6 DACs   33 148 128 T 128  122 128117  112 122 6l  br Tl 20  1331339 1339 143  HAB 7 DACs   33 128 117 143 117 122  46  32 71  61 138 128 122 2 107 112 112 107  HAB 8 DACs   12 112   92  Si 891 LOT 102 107    99 T12  lIOT 107 2 2112 66 46 51  HAB 9 DACs   12 12 1107 7 117 112 107 107 112 107 46 97 76 25 40 32 46 46  HAB 10 DACs   76 112 107 87 102 46 97 51 61 42 66 107 81 87 51 87 107 129  HAB 11 DACs   28 25 117 7 128 128 138 122 133 122 122  122 153 5 138 143 138 122                                                                   CALICE eLogbook   Results for search of  20070501 Page 9 of 9    
124. 25V to  peak  2007 03 21 21 27 S  Uozumi LED run  400580 CmLed LED KNU fiber 25V to  peak  2007 03 21 21 28 S  Uozumi LED run    400586 CmLed LED KNU fiber 2 5V for  inter calibr  tion    CALICE eLogbook   Results for search of  20070301 Page 31 of 76       2007 03 21 21 28 S  Uozumi LED run  400586 CmLed LED KNU fiber 2  OV for  pes A EE EL listez estiejeion  2007 03 21 21 28 S  Uozumi LED run  Eee A ws O o qe pee  a  inter calibr  tion  2007 03 21 21 28 S  Uozumi LED run  400587 PmLed LED KNU fiber 2 5V for  Eee pee EA Ts  2007 03 21 21 29 S  Uozumi LED run  400590 CmLed LED KNU fiber 2 5V to  peak  2007 03 21 21 29 S  Uozumi LED run  400590 CmLed LED KNU fiber 25V to  peak  2007 03 21 21 29 S  Uozumi LED run  400592 CmLed LED KNU fiber 2 5V to  peak  2007 03 21 21 29 S  Uozumi LED run  CAS CO MM E s  see lpe  peak  2007 03 21 21 33 S  Uozumi LED run  pum eee E GE  inter calibr  tion  2007 03 21 21 33 S  Uozumi LED run  400594 CmLed LED KNU fiber 2 5V for  eee         listezestibejeion  2007 03 21 21 33 S  Uozumi LED run  poe quee cue quo RS dh A  inter calibr  tion  2007 03 21 21 33 S  Uozumi LED run  400595 PmLed LED KNU fiber 2 5V for  phs A EE E E  2007 03 21 21 34 S  Uozumi LED run  400597 CmLed LED KNU fiber 25V to  peak  2007 03 21 21 34 S  Uozumi LED run    400597 CmLed LED KNU fiber 2 5V to  see lpe    CALICE eLogbook   Results for search of  20070301 Page 32 of 76          YC      2007 03 21 21 34 Es Uozumi LED run  400597 CmLed KNU fiber  peak  2007 03 21 21 34 M
125. 26  like 202365 but with cfg StagePositions202369 dat   2007 04 17 18 57 Beni dead cells check  202370  202371  module 24  found dead channel  FE 1 chip 0 chan 5 hist  PM only   found dead channel  E 1 chip 5 chan 17 hist  found dead channel  FE 1 chip 8 chan 5 hist  found dead channel  E 1 chip 8 chan 12 hist  found dead channel  E 1 chip 9 chan 0 hist  found dead channel  E 1 chip 9 chan 3 hist  found dead channel  FE 1 chip 9 chan 5 hist  found dead channel  E 1 chip 10 chan 1 hist  found dead channel  E 1 chip 11 chan 6 hist  module 26  found dead channel  FE 2 chip 6 chan 9 hist  PM only   module 28  found dead channel  FE 4 chip 9 chan 2 hist                      2007 04 17 20 16  gt  current development    CALICE eLogbook   Results for search of  20070401                jroo    ARTI  RE RAA RI  201500 201600 201700 201800 20194    10 00 Ramo  OFAI   Mecue 20  ERR Mesue 208                      File  http   ttfinfo desy de CALICEelog data 2007 16 17 04 2007 04 17T20 16 02 00 gif    2007 04 17 20 34  gt  noise looks correct    Y ptaCrcNomeCrate   Sit JT u           nul       File  http   ttfinfo desy de CALICEelog data 2007 16 17 04 2007 04 17T20 34 02 00 gif    2007 04 18 11 10 current development          Page 6 of 14       CALICE eLogbook   Results for search of  20070401 Page 7 of 14                File  http   ttfinfo desy de CALICEelog data 2007 16 18 04 2007 04 18T11 10 02 00 gif    CALICE eLogbook   Results for search of  20070401 Page 8 of 14       A            
126. 3 330612  Nanda    Run NoE o         Mean tmips              piam Bigma Mean      330583 mg Oe El EE  30589 11 z    30590  30591  30592  30596  30600  30603  30604  30605  30606  330612 51 7    330613  5         EC    Run 330613 has only 80k events  After it we move the stage    n  LA       Wei xo  FR  00    ol      o  ot   w E    EIST ST STOT    WOT   f COF to  09  of   o     024    774    Hs  CO  SI  Go  Lol   Do  SP BP us  om Lo  Do   amp     Da  RH IL L            N  N  H H H W  oo ef   o   H Nf H N  EI EN A oyu O Wu  au   CO    ll  Wu Co  ado    ooo       o          2007 07 17 22 54 Nanda Mangi Valgsipn run       Pb    collimator C10 from 4 to 3  trigger rate  8kpps  DAQ rate  90 Hz     amp  2007 07 17 23 31 caliceon calicehansv  d stageh   CALICE user   Go back to muons     HCAL  amp  ECAL back to center of beam    moved 1lxlm trigger back into the beamline   changed AHC cfg file link to AHC_CERNO7_beamhold_100x100 cfg   change in EmcConfiguration hh sequenceDelay to 1   new beamsettings  file 555  absorber  amp  target   air    adjusted collimator settings to 14k events in 100x100 coincidence COLLO1   20 20   COLLO5 COLLO6    15  15     left beamdump closed    CALICE eLogbook   Results for search of  20070701 Page 19 of 19       File  http   tttinfo desy de CALICEelog data 2007 29 17 07 2007 07 17T23 31 02 00 jpg    CALICE eLogbook   Results for search of  20070701 Page 1 of 20       A 2007 07 11 03 12 Trygve HCAL high current    Channel 74  purple  slightly high  Tren
127. 38  5 3 AHCAL8 37  5 4 AHCAL8 35  5 5 AHCAL8 34  5 6 AHCAL8 33  5 7 AHCAL8 32  7 0 AHCAL8 31  9 0 AHCAL8 36  7 3 AHCAL 24  7 4 AHCAL 1  7 6 AHCAL 2  7 7 AHCAL 30  9 1 AHCAL 29             slot fe    LED Hold PM LED    slot fe    slot fe  slot fe    38 36 38  37 40 37  36 39 35  35 37 34  34 28 33  33 35 32    32 34 31  310 0  300 0  29 0 0  28 0 0  27 23 30    26 26 29       0 500 1000  RMS ADC     File  http   ttfinfo desy de CALICEelog data Analysis AHCAL 2007 06 30T14 09 30 00 eps    AHCAL   TCMT mapping    AHCAL moduleNo stackPos cmbId cmbCanAdr cmbPinId Hold ext  Vcalib CM Vcalib PM    TCMT dacFileName Hold ext Hold CM LED Hold PM LED Vcalib C  PIN hold PIN HBAB   PMT hold   99 15 33 61 28000 31000  99 15 35 63 28000 31000  99 15 35 00 28000 31000  99 15 34 00 28000 31000  99 15 34 00 28000 31000  99 15 35 00 28000 31000  99 15 00 00 28000 31000  SS 150 0 0 31000  0 50 0 0 0   0 50 0 D 0   0 50 0 D 0   99 15 00 00 28000 31000  99 15 00 00 28000 31000          CALICE eLogbook   Results for search of  20070601 Page 18 of 19                                                             9 2 AHCAL 28 25 10 28 99 15 00 00 28000 31000  9 3 AHCAL 27 24 38 27 99 15 00 00 28000 31000  9 4 AHCAL 26 23 22 26 99 15 00 00 28000 31000  9 6 AHCAL 25 22 16 25 99 15 00 00 28000 31000  9 7 AHCAL 23 21 21 23 99 15 00 00 28000 31000   2 4 PMT 14   5 1 AHCAL 22 20 32 22 99 15 00 00 28000 31000   5 2 AHCAL 21 19 24 21 99 15 00 00 28000 31000   5 3 AHCAL 20180 0 0 150 0 0 0   5 4 AHCAL 19 17 27 1
128. 4 15    400679     CmLed for inter calibration  400680     PmLed for inter calibration  400682     CmLed to see lpe peak   400683     Same with 400682 but with  1V increased base voltage   80V   400685     CmLed to see lpe peak at strip 14   For Layer4 strip14 15    400689     CmLed for inter calibration  400690     PmLed for inter calibration  400692     CmLed to see lpe peak   For Layer2 strip14 15    400697     CmLed for inter calibration  400698     PmLed for inter calibration                400700     CmLed to see lpe peak    CALICE eLogbook   Results for search of  20070301 Page 75 of 76       For Layer6 strip14 15     400706     CmLed for inter calibration  400707     PmLed for inter calibration  400709     CmLed to see lpe peak   For Layer24 stripl5    400710     CmLed for inter calibration  400711     PmLed for inter calibration    could not see lpe peak with this strip  due to too wide pedestal width     For Layer25 stripl5                       400715     CmLed for inter calibration   400716 PmLed for inter calibration   400718     CmLed to see lpe peak   400719 CmLed to see lpe peak around strip 15   For Layer26 stripl5    400721     CmLed for inter calibration   400722     PmLed for inter calibration   400723     CmLed to see lpe peak   400724     CmLed to see lpe peak around strip 15     Following runs are strips measured second time  so no inter calibration for this time     For Layerl4 stripl5     400725     CmLed to see lpe peak  400726     Same with 400725 b
129. 400   21 25 23    File  http   ttfinfo desy de CALICEelog data 2007 26 27 06 2007 06 27T21 34 46 00 png    G Mavromanolakisnonitor installed  a first version of the monitor code is installed at caliceon caliceserv  under  scratch software calice soft monitor caliceon monitor      start gui sh to open the gui panels      start sender sh to process data    more details and instructions can be found in the README txt or by pressing the   Help  button at the gui     The mapping cabling files for the ecal hcal tcmt are not available yet     2007 06 28 16 45  All AHCAL modules are installed and cabled   CMBs are installed for all modules with non zero entries     slot fe detector    HCA  HCA  HCA  CA  CA  CA    Dm Dm  004 GA 0    88   gt     BE EE EE  Aen LAACH  SEKR GEES SEKR  aa  Do  EEE          VO VO XO VO VO O    Ons Q0 2 E      Ons 00 0L    Oooo Oi Oi  pps PPPppp               o1        wWN PO    Li       L8  L8  L8  L8  L8  L8    L8  L8    E EES EE       E bets    38  37  35  34  33  32    31  36  24  T  2  30    29  28  27  26  25  23    22  21  20  19  18  17    38  37  36  35  34  33    32  31  30  29  28  27    26  25  24       new AHC cfg    module position cmb canbus pinId    36  40  39  37  28  35       38 99 0 33 61 28000 31000  37 99 0 35 63 28000 31000  35 99 0 00 00 28000 31000  34 99 0 00 00 28000 31000  33 99 0 00 00 28000 31000  32 99 0 00 00 28000 31000  31 99 0 00 00 28000 31000  0 0 00 0 0 0   0 0 00 0 0 0   0 0 00 0 0 0  0 0000 0 0   30 99 0 00 00 28000 310
130. 400744 CmLed LED KNU fiber   same  with  previous run  but with  increased  Over voltage        2007 03 23 19 05   Uozumi LED run    CmLed KNU fiber e same  with  previous run  but with  increased  lover voltage    2007 03 23 19 06    CmLed KNU fiber    2007 03 23 19 06   Uozumi LED run    a  2007 03 23 19 07    2007 03 23 19 07   Uozumi LED run    CmLed KNU fiber    2007 03 23 19 07    CmLed KNU fiber 3  with  previous run  but with  increased  Over voltage    2007 03 23 19 07   Uozumi LED run    CmLed KNU fiber S same  with  previous run  but with  increased  lover voltage                         2007 03 23 19 08 S  Uozumi LED run    8 a BEEN       CALICE eLogbook   Results for search of  20070301 Page 51 of 76          2007 03 23 19 08 S  Uozumi LED run       p  lle  2007 03 23 19 08 S  Uozumi LED run  DT EE J    r  8  see lpe  peak at  L21515  2007 03 23 19 08 S  Uozumi LED run  A dE ETT  see lpe  peak at  L21515  2007 03 23 19 09 S  Uozumi LED run    NI 7   1          2007 03 23 19 09    LED KNU fiber 3 5V  previous run  but with  increased  lover voltage    S  Uozumi       NI 7 rot    S  Uozumi    2007 03 23 19 09          2007 03 23 19 09          III   2007 03 23 19 10 S  Uozumi LED run  Pared dL XE  see lpe  peak at  L23s15  2007 03 23 19 10 S  Uozumi LED run    400753 CmLed LED KNU fiber 2 5V to  see lpe  peak at  L23515    CALICE eLogbook   Results for search of  20070301 Page 52 of 76       2007 03 23 19 11 S  Uozumi LED run    KNU fiber 3 5V same  with  previous 
131. 46 61 15 20 30 56 46 46 46 40 51 51 56  HAB 5 DACs   56 46 30 25 25 46 251 46 SL SIS 46 S   51 46 20 25 25  HAB 6 DACs   46 51 46 40 46 51 46 40 46 46 56 61 61 40 51 40 51 51  HAB 7 DACs   51 46 51 51 51 40 40 61 15 66 5 46 51 46 51 51 51 51  HAB 8 DACs   51 51 81 102 76 46 51 51 46 40 5 40 51 46 46 133 97 122  HAB 9 DACs   51 40 46 46 46 46 51 51 51 46 6 51 56 40 40 66 71 87  HAB 10 DACs   102 102 107 117 107 112 102 87 71 66 133 133 122 127 56 97 102 56  HAB 11 DACs   61 40 56 61 56 61 56 107 61 71 5 5 163 40 40 255 40 40  Slot number   5  CRC Component   0   FEO  Label   2   0x02   0b00000010   read  AhcVfeStartUpData   print     Verification data   43690   0x0000aaaa   0500000000 00000000 10101010 10101010  HAB 0 DACs   61 40 56 61 56 61 56 107 61 71 51 15 163 40 40 255 40 40  HAB 1 DACs   102 102 107 117 107 112 102 87 71 66 133 133 122 127 56 97 102 56  HAB 2 DACs   51 40 46 46 46 46 51 51 51 46 61 51 56 40 40 66 71 87  HAB 3 DACs   51 51 81102 76 46 51 51 46 40 51 40 51 46 46 133 97 122  HAB 4 DACs   51 46 51  1 51 40 40 61 15  66 S1  46 51 46 51 51 951 51  HAB 5 DACs   46 51 46 40 46 51 46 40 46 46 56 61 61 40 51 40 51 51  HAB 6 DACS   0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0  HAB 7 DACs   0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0  HAB 8 DACs   0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0  HAB 9 DACs   0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0  HAB 10 DACs   0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0  HAB 11 DACs   0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0                gt  values loaded as HAB 11 are rea
132. 5 kpps   DAQ rate  100 Hz    2007 07 17 00 22 Hakan Oliver Geopije 35GeV   330561  beamData  10x10    35GeV Ki 515 Wan fir ___  wa Cen BA      CALICE eLogbook   Results for search of  20070701 Page 15 of 19    Pb spill  rate  210k  events       2007 07 17 01 21 Hakan Oliver GeoNeise Runs  30562 lahcCmNoise  30563 lahcPmNoise  30564 lahcCmLedVcalibScan  30565 ahcPmLedVcalibScan    30566 ahcGain  30567 lahcCmNoise  30568 lahcPmNoise       2007 07 17 02 22 Hakan Oliver Geopje 35 GeV    33 0569 lbeamData  10x10   35GeV pi  OL  4mm Air N A 9k spill  Pb rate  200k  events       2007 07 17 02 57 Hakan Oliver Geopje 40 GeV    Mi pm T  Pb             2007 07 17 03 39 Hakan Oliver Geopje 45    LA DL d IB  Pb    2007 07 17 04 24 Hakan Oliver Geopje 50 GeV    330572 lbeamData  10x10   50GeV pi  512 4mm i R  Pb spill  Irate  200k  events       2007 07 17 05 12 Hakan Oliver GeoNeise Runs  30573 lahcCmNoise  30574 lahcPmNoise  30575 ahcCmLedVcalibScan    30576 ahcPmLedVcalibScan  30577 ahcGain       2007 07 17 06 19 Hakan Oliver Geopje 80GeV    330578 lbeamData  10x10   80GeV pi  510 4mm Air N A lbeamtune   Pb 10k  events  330579 beamData  10x10  80GeV pi  510 4mm Air N A Spill  Pb rate  7k260k  events    Spill rate was too high  so set collimators 2  3 and 8 from     2 0mm to     1 5mm   DCs not aligned with beam  changing TRIMO6 current from  13 3A to  7 0A   330579    Is aligned now        2007 07 17 07 18 Hakan Oliver Geopje 100 GeV    nil HN E  Pb    beamData  10x10  100GeV   pi  516 4
133. 53 CmLed LED KNU fiber 25V for   EC GR PASASE gos  2007 03 21 21 11 S  Uozumi LED run   400553 CmLed LED KNU fiber 2 5V for   pe MO NEM EMEN O aca E ege  2007 03 21 21 12 S  Uozumi LED run   400554 PmLed LED KNU fiber 2 5V for   pee ugue xe AA c E A  2007 03 21 21 12 S  Uozumi LED run   400554 PmLed LED KNU fiber 2 5V for   pee A cum   dee cp odes eis  2007 03 21 21 12 S  Uozumi pedestal run       400555 CmNoise pedestal KNU fiber Z 9V  pedestal  run for LED  calibration  2007 03 21 21 12 S  Uozumi pedestal run    400555 CmNoise  pedestal KNU fiber 2 5V  pedestal  run for LED  calibration       CALICE eLogbook   Results for search of  20070301 Page 29 of 76       2007 03 21 21 13 S  Uozumi pedestal run    400556 PmNoise pedestal KNU fiber 2 5V pedestal  run for LED  calibration       with LED system off    2007 03 21 21 13 S  Uozumi pedestal run  400556 PmNoise pedestal KNU fiber Aa pedestal  run for LED  calibration    2007 03 21 21 13 S  Uozumi pedestal run    400557 CmNoise pedestal KNU fiber 25V pedestal  run for LED  calibration    with LED system off                   2007 03 21 21 13 S  Uozumi pedestal run  400557 CmNoise pedestal KNU fiber 2 5V pedestal  run for LED  calibration  2007 03 21 21 14 S  Uozumi pedestal run  400558 PmNoise pedestal KNU fiber 2 5V pedestal  run for LED  calibration  2007 03 21 21 14 S  Uozumi pedestal run  400558 PmNoise pedestal KNU fiber PERI pedestal  run for LED  calibration  2007 03 21 21 15 S  Uozumi LED run  400562 CmLed LED KNU fibe
134. 6 3k pps in 10x10 trigger coincidence  Average data taking rate   85 Hz    2007 07 09 03 29 Kurt  Tao  Nanda pion run    330240 beamData  10x10  40GeV pi  514 air 4mm off 250k  Pb Events        10 7k pps in 10x10 trigger coincidence  Average data taking rate   80 Hz    2007 07 09 04 19 Kurt  Tao  Nanda pion run   230241  beamData  10x10    100GeV Ki 516 fai mm fof WI    CALICE eLogbook   Results for search of  20070701 Page 2 of 25          lo      pp    f  Events     11 6k pps in 10x10 trigger coincidence  Average data taking rate   80 Hz          2007 07 09 05 05 Kurt  Tao  Nanda gain run    330222 jaheGain TT  2007 07 09 05 15 Kurt  Tao  Nanda noise run    330223  ahcPmNoise          OK events    2007 07 09 05 36 Kurt  Tao  Nanda pion run    WW ud MIENNE NEN  Pb    330245 beamData  10x10  80GeV 4mm eee  Pb pcc ev        11 5k pps in 10x10 trigger coincidence  Average data taking rate   80 Hz       2007 07 09 06 37 Kurt  Tao  Nanda pion run       330246 lbeamData  10x10   50GeV  oi  512 air 4mm off 250k  Pb events     6k pps in 10x10 trigger coincidence  Average data taking rate   80 Hz    2007 07 09 07 25 Kurt  Tao  Nanda pion run       330247 lbeamData  10x10   40GeV pi  514 air 4mm Off 250k  Pb events     11 6 pps in 10x10 trigger coincidence  Average data taking rate   80 Hz       2007 07 09 07 48 high temperature   2 degree higher temperature in layer 10 and 18 compared to the average HCAL temperature   27  degree   b iv Pos 10   Module 12   CMBO PREL to den           
135. 7 08 18 16 e  run  40 Se    330222 lbeamData  10x10 a 531 6mm beam not  Pb  pointing  to centre  ot DC     Trim6  current increased to  15 to steer beam   Run Stopped after 112K events  Starting new run at sam energy to get a good  centered  point         2007 07 08 18 34 electron run    E illi MONO  Pb       Current in Trim6 up from 10 to 18 to center the beam on drift chambers     Please take 250K good events with this configuration and then move to 50 GeV e   250K evts   2007 07 08 18 36 Niels  Jean Yves   Hengne  2007 07 08 18 46 electron run       330224 lbeamData  10x10 ie 531 6mm beam  Pb centered  now  250K  events in  this  configurat ion    DAQ mean rate  87 Hz    2007 07 08 19 32 Niels  Jean Yves   Hengne    330225 lbeamData  10x10     50GeV e  531 air 6mm off 26k  Pb Events   dropped    330226 lbeamData  10x10   50GeV e  531 air 6mm off 36k  Pb Events   dropped   330227 lbeamData  10x10  50GeV e  531 air 6mm off ex  Pb Events   dropped       CALICE eLogbook   Results for search of  20070701 Page 23 of 24       dropped because beam is too board in the x direction   put collimators in x direction more closer   and then moved the beam to the right  x direction   TRIM 6 25  with respect to the drift chamber        A   e hTeack 0G IU      E E o Es  LN D K    s   runi      Tre  um i f  i     mm d  HE Wen  m Exc  Bee ee acd amb PE OR M XM LSC T GK SS  xe SE Sen  Track 8  A AA gio         Zeg   eg wi       e     m   E I E       A  im mum pm TET        aum  H   ves  Bi    
136. 7 25 07 26 07 27 07 28 07 29 07    time  File  http   ttfinfo desy de CALICEelog data 2007 31 30 07 2007 07 30T04 07 55 00 gif          2007 07 30 04 10 Beni Geneva temperature development    CALICE eLogbook   Results for search of  20070701       Tenperatur C C2  28    23 07    f wetter com   Daten aus Genf  Schweiz    t t t  4  3  3   gt  s s s E E  Ed S a S  amp  2    Legende  Durchschnitt    Maximum      Minimum    Zeitraum  23 07 2007   29 07 2007       File  http   ttfinfo desy de CALICEelog data 2007 31 30 07 2007 07 30T04 10 05 01 png        2007 07 30 04 12    Beni temperature development    detailed view of the temperature development       Trent  A                    D       File  http   ttfinfo desy de CALICEelog data 2007 31 30 07 2007 07 30T04 12 52 00 gif    2007 07 30 04 14    Beni temperature profile development    Page 7 of 20    This plots shows all temperature profiles available since the mapping correction overlayed  The    variations happen in parallel on all sensors      10 28    35   The system seems to be in good equilibrium     except the three layers with readout problems    CALICE eLogbook   Results for search of  20070701 Page 8 of 20    sum of temperature profiles      31c T    L f             average temperature     C              cook er VA A a AT CET 1 Lu f EE Ka Li   22  5 10 15 20 25  30 35 40  layer    File  http   ttfinfo desy de CALICEelog data 2007 31 30 07 2007 07 30T04 14 12 00 gif      A 2007 07 30 05 03 Roman  Beni 6 GeV run       Th
137. 8 2007 13 08     File  http   ttfinfo desy de CALICEelog data 2007 33 19 08 2007 08 19T21 54 04 00 jpg       2007 08 19 22 06 raised current limit    raise limits of HV channel 26 and 60       3233 w ad  iss  W SRSA  148  RER   1   36538   116  FER   4m      146 Fam  193 MIRTA    172 Maja  149 W RIA  141 W oss         3103 w ipia    ainu  E ciem  5 218     mm W AIRAA   5   w X   LESEN     gy 7        m iva ort J aw     mime Oy Lv OFF  al  BETE poe  SIE sn              PA    enable   disable PLL channels  a T PO TA       a   um   Qm j  ex   dele   reser all wpa   8035  wwe   wg wv aot  Zeg H   values to file J  See NL veluse  position                CALICE eLogbook   Results for search of  20070801 Page 10 of 22    File  http   ttfinfo desy de CALICEelog data 2007 33 19 08 2007 08 19T22 06 04 00 jpg    2007 08 19 22 18 Daniel   Anne Marie  Petr    350277 beamData  10x10  10GeV e  626 air 6mm unknown 5k   Pb 20deg    adjusted Trim06 to  1    2007 08 19 22 21 Daniel   Anne Marie  Petr    350278 lbeamData  10x10 H10GeV e  626 air 6mm unknown 60k    Pb 20deg    2007 08 19 22 59 Daniel  junk run  Anne Marie  Petr    350279 junk run          2007 08 19 23 12 caliceon 2calicehdalscbammelti 6 slots  fed   CALICE user     shows step                20 zu z100 22300 23300    13 08 2007 19 08 2007 19 08 2007 19 08  2007    File  http   ttfinfo desy de CALICEelog data 2007 33 19 08 2007 08 19T23 12 03 00 jpg    2007 08 19 23 12 Daniel  DAC reloaded  Anne Marie  Petr  2007 08 19 23 12 Danie
138. 8 set to  30  JAW1    27  JAW2     still no beam at the moment  Started ECAL safety run     2007 07 06 14 47 Fab key access requested    Key access requested by TCMT people to perform maintenance works   ECAL turned off since work is expected to last for at least one hour     2007 07 06 16 03 Fab Erika beam back    Key access finished  Beam back   SLOT 5 is ALIVE   Only FE 6 is connected  Slot 5 FE7 has been moved to SLOT 7 FE7 to prevent    problems  in case SLOT 5 disappears again     e Ck Ck Ck Ck Ck Ck Ck CC CK CC CC CC CC CK CC CK CC CC CC CC CC Ck Ck Ck Ck Ck Ck CC Ck CC Ck Ck Ck Ck Ck Ck Ck Ck kk kk ck ck kc kckckckckckckck    WARNING  ECAL slots 5 and 7 were both not booted properly and hence completely  missing for runs 330132 to 330134 inclusive     Paul   2007 07 06 16 35 Fab Erika selecting muon beam from 80 GeV file  Loaded 80 GeV file  H6B 510   Collimator C8 changed to select muons   C8 set to  30  JAW1    27  JAW2     015 and 016 turned off to have parallel beam in H6     One more beam file available  100 GeV   H6B 516    2007 07 06 17 15 Fab Erika Francois    330135 beam 100x100   80Gev mu  510 Air Air not  0 5HV defocus  data in  beam  8 0   27 mm    colimeter 8 set to     mainly pions in beam   10x10 trig rate   300 Hz   100x100 trig rate   300 Hz   DAQ rate 16 Hz   low intensity beam due to booster problems       2007 07 06 17 16 Fab Erika Francois    330136 beam 100x100   80Gev mu  510 Air Air not  0 5HV defocus  data in lbeam  8 5 mm    colimeter 8 set t
139. 86 extern root lib libHistPainter so  1  20 0xf712d913 in TH1  Paint     from  scratch software i386 extern root lib libHist so  1  21 0xf6d34411 in TPad  PaintModified     from  scratch software i386 extern root lib libGpad so  1  22 0xf6d3433d in TPad  PaintModified     from  scratch software i386 extern root lib libGpad so  1  23 Oxf6dl17adc in TCanvas  Update     from  scratch software i386 extern root lib libGpad so  1  24 Oxf6d32cla in TPad  Update     from  scratch software i386 extern root lib libGpad so  1  25 0x080783ae in MyMainFrame  HandleVetoInfo     2  26 0x08075988 in MyMainFrame  HandleList                                  CALICE eLogbook   Results for search of  20070701 Page 4 of 20       3  27 0x080742d3 in MyMainFrame   DoStart     4  28 0x08098073 in G LinkDict 722 0 10     5  29 0xf734da8b in Cint  G _CallFunc  Execute     from  scratch software i386 extern root lib libCint so  1  30 0xf79ac504 in TQConnection  ExecuteMethod       from  scratch software i386 extern root lib libCore so  1  31 0xf79b0233 in TQObject  Emit     from  scratch software i386 extern root lib libCore so  1  32 0xf643b304 in TGButton  Clicked     from  scratch software i386 extern root lib libGui so  1  33 0xf6435306 in TGButton  EmitSignals     from  scratch software i386 extern root lib libGui so  1  34 0xf6435060 in TGButton  SetState     from  scratch software i386 extern root lib libGui so  1  35 0xf64351f7 in TGButton  HandleButton    from  scratch software i386 extern root
140. 9 99 15 00 00 28000 31000   5 6 AHCAL 18 16 15 18 99 15 00 00 28000 31000   5 7 AHCAL 17 15 25 17 99 15 00 00 28000 31000   7 0 AHCAL 16 14 29 16 99 15 00 00 28000 31000   7 1 AHCAL 15 13 3 3 99 15 00 00 28000 31000   7 2 AHCAL 14 12 8 14 99 15 00 00 28000 31000   7 4 AHCAL 1311 0 0 0 150 0 0 0   7 5 AHCAL 1210 0 0 0 150 0 0 0   7 6 AHCAL 11 9 0 0 0 150 0 0 0   7 7 AHCAL 10 8 0 0 0 150 0 0 0   94 AHCAL 9 7 0 0 0 150 0 0 0   9 1 AHCAL 8 6 0 0 0 150 0 0 0   9 2 AHCAL 7 5 0 0 0 150 0 0 0   9 3 AHCAL 6 4 0 0 0 150 0 0 0   95 AHCAL 5 3 0 0 0 150 0 0 0   9 6 AHCAL 4 2 0 0 0 150 0 0 0   9 7 AHCAL 3 1 0 0 0 150 0 0 0   2 5 TCMT DAC tcmtl dat 46 32 27 10000 10000   2 6 TCMT DAC tcmt2 dat 46 33 27 5500 5500   2007 06 30 16 04 Erika new AHC cfg   5 0 AHCAL8 38 38 36 38 99 15 33 61 28000 31000  5 3 AHCAL8 37 37 40 37 99 15 35 63 28000 31000  5 4 AHCAL8 35 36 39 35 99 15 35 00 28000 31000  5 5 AHCAL8 34 35 37 34 99 15 34 00 28000 31000  5 6 AHCAL8 33 34 28 33 99 15 34 00 28000 31000  5 7 AHCAL8 32 33 35 32 99 15 35 00 28000 31000  7 0 AHCAL8 31 32 34 31 99 15 00 00 28000 31000  9 0 AHCAL8 36 310 0 99150 0 0 31000  7 3 AHCAL 24 300 0 0 150 0 0 0  7 4 AHCAL 1 290 0 0 150 0 0 0  7 6 AHCAL 2 280 0 0 150 0 0 0  7 7 AHCAL 30 27 23 30 99 15 00 00 28000 31000  9 1 AHCAL 29 26 26 29 99 15 00 00 28000 31000  9 2 AHCAL 28 25 10 28 99 15 00 00 28000 31000  9 3 AHCAL 27 24 38 27 99 15 00 00 28000 31000  9 4 AHCAL 26 23 22 26 99 15 00 00 28000 31000  9 6 AHCAL 25 22 16 25 99 15 00 00 28000 31000  9 7 A
141. 93  6 0 18 K events    trigger rate 3 6k  DAQ rate 80 Hz       Morien  Brient  Lutz 18 GeV pions       Logbook entry   CALICEelog data 2007 29 22 07    22 07 2007  09 17       Morin  Brient  Lutz       trigger rate 4 7 K  DAQ rate 83 Hz    Logbook entry   CALICEelog data 2007 29 22 07    22 07 2007  08 48    ECAL position   6 0        position check in 2D display       27 09 2007 11 39    CALICE eLogbook http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log       xi     ou  es 32    5       Logbook entry   CALICEelog data 2007 29 22 07    EZ 22 07 2007 caliceon G calicehcalsc02 cern ch  CALICE dw eositon  08 20 user  p    moved to   6 0  which is  80  10  in stage coordinates       otage co Cale nte    HCAL Motor control panel    actual position target position     o   A fp ox        Y    19 00 La ei    O lost X reference point  E    Q lost Y reference point    CH lost 2 reference point    O sem in motion  Q m operation get Pos    Ian Y Fast 1 pulsesfreo   extern STOP   eicht   819 305 819 998      OFF set      D    s00 i H Le          Logbook entry   CALICEelog data 2007 29 22 07    a O em Morin  Brient  Lutz new shift    23 of 28 27 09 2007 11 39    CALICE eLogbook http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log       Logbook entry   CALICEelog data 2007 29 22 07    Sy     ear Ricardo Nicola Kieffer Jump in RMS for ECAL ped    MEcalPedRMSCNpAverage sloti9  um EE      morem       Logbook entry   CALICEelog data 2007 29 22 07  2
142. AQ rate 88 Hz       A 2007 07 21 21 43 Oliver Allister HenMistakes in elog entries   There seems to be some mis assignment of target and absorber for some runs  For example  runs  330488 330508    supposedly electron runs  it is written in the elog that a  4mm Pb  absorber   was used and target is  air   The corresponding logbook entries seem to be   correct though  with target as  6mm Pb  and absorber is  air   The   analyses of the data seems to agree that these are indeed electron runs as well     CALICE eLogbook   Results for search of  20070701 Page 13 of 19       Maybe this means that we should reduce the use of  copy and paste  and be more careful  52        We suggest the shifters for these runs to re check and correct the e log entries for the  above mentioned          runs    2007 07 21 22 04 Oliver Allister Hengne   a eae EE GE da e E e  0 trigger rate  D kpps   E rate  86 Hz   Collimators reduced    COLLO1       5   COLLO2       2   COLLO3       2   COLLO5       5   COLLO6       10   COLLO8       2   COLL10       5   COLL11       10  2007 07 21 23 09 EE run    3       0    2  trigger rate  6 kpps  DAS rate  86 Hz    Collimators reduced   COLLO1       3       COLLO2       6  COLLO3       3  COLLO5       3  COLL06       3  COLL08       6  COLL10       4  COLL11       4  A 2007 07 21 23 26 Oliver Allister Henmistakes found in elog    We note that some entries in the elog have mistake in  absorber  and  target      e g   330508 fbeamDatal10x10  15GeV 536 4mm air 1 34 
143. Anne Marie  Petr    350327 lbeamData  10x10  80GeV Git 611 4mm air 154k  Pb degs    350328 lbeamData  10x10  60GeV Git 610 4mm air 0 3 0 0 151k  30  Pb degs  p 3 po         30       350323     peaupata  l0x10 150k     CALICE eLogbook   Results for search of  20070801 Page 16 of 22             pess         o yoo  zb j   pf    350330 lbeamData  10x10  40GeV Git 613 4mm air 0 1 40k  30  Pb degs    350331 lbeamData  10x10  40GeV Git 613 4mm air 0 1 15k  30  Pb degs   350332 lbeamData  10x10  40GeV Git 613 4mm air 0 1 155k  30  Pb degs    2007 08 21 00 08 Hakan Allister Sat amp tHCAL run    350333 lbeamData  10x10 H30GeV Git 614 4mm air 0 3 155k  gt   30  Pb degs    2007 08 21 01 01 Hakan Allister SatBositron Runs  Pb 30degrees   beamDat 0x10  40GeV e  dr  beamDat 0x10 e  k   30degrees  0x10 101k   ata  10x10    30degrees  beamD I e  625 air    air                                  2007 08 21 07 30 Hakan Allister Satoru  Stage moved from  0 0  to   6 0    Start taking data again from e  10 GeV     2007 08 21 07 31 Hakan Allister Satoru       2007 08 21 07 53  Shower profile of run 350346  10GeV e     Energy is dropped off at Layer 6   This has been appeared since current run  350346   after moving the stage to  6 0      CALICE eLogbook   Results for search of  20070801 Page 17 of 22          hHcalHitsPerLayer          2 s I    8  7  6  5  4  3  2  1  0  hHcalEnergyPerLayer      4  INTO TDT TEA a LS E mania PT EL  E  se 36  a  o 30  E  9 ap GE  20  15 A O O O O SEO  10  5  0 40  aye
144. CALICE eLogbook   Results for search of  20070901 Page 1 of 1       2007 09 17 11 00 Beni    CMBs 23  amp  14 reinstalled    2007 09 17 11 49 current   10V     Cem     REN     a gias   75 See     e Sa Fe SSES   0 Sms   o gomos 9 Sr   0 Rima 3 R ISA   1 ama 1   o SZ   0 NA   65 g DBA     e DEE   o raves   1 RA E E   o Sim   o Sms 1  e 2  o 748       RER      1         1 SOEER      2 gw     1 aroa         QONLYHVOFF  ONLY LVOFF    Hy Crate ON  HV Crate OFF     recover    KILL HIT Wal SHY ZO   Hv 10 1    HY2 1  Hv3  1 HEI   recover gue    ZHVI I Hv J    Hyg   LV detail   File  http   ttfinfo desy de CALICEelog data 2007 38 17 09 2007 09 17T11 49 02 00 jpg       d                            2disiidis                          fes 2007 09 17 14 54 Beni last measurement    Measurement of the difference between temperature sensor voltage and digitized value for pos 1  module    223 18 21 59  226 59 21 952  2 21 570    22 75    226 40 21 936  are       2 056       CALICE eLogbook   Results for search of  20070801 Page 1 of 22       2007 08 19 00 09 Hakan Allister SatbeamData    350243 beamData  10x10  25GeV e  623 Air 6mm 0 35 20deg   Pb 100k   350244 beamData  10x10  20GeV e  622 Air 6mm O 5 0 0 20deg   Pb 100k    350245 beamData  10x10 H15GeV e  625 Air 6mm 0 0 20deg   fp pe qe q  E  1    el  350246 beamData  10x10  10GeV e  626 Air 6mm 2 0 0 20deg    Pb 53k    350246 stopped early due to beam being off centre        Adjusted TRIMO6 current from 3 0A to OA    350247 beamData  10x10 
145. DACvalue mod2 A77 0 B79 9 txt  4  rw r  r   1 caliceon 3145 1749 Jun 28 20 40 DACvalue mod24 A68 4 B75 2 txt  4  rw r  r   1 caliceon 3145 1177 Jun 28 20 41 DACvalue mod36 A80 7 B73 3 txt  4  rw r  r   1 caliceon 3145 1667 Jun 28 20 42 DACvalue mod18 A41 0 B41 1 txt  all other files are from June 23    Apparently  the June 28 files were not copied to hcaldaq01     2007 07 02 00 28 S  Schaetzel summary of DAC file desaster  wrong LED settings for    Module 36 no DAC file      needs retuning    Module 2 wrong DAC file Vcalib increased by 200    Module 24 wrong DAC file Vcalib unchanged       For Module 36  the log files say   Log1183249096 0ut  error opening DAC module36 dat  Log1183249232 0ut  error opening DAC module36 dat  Log1183249323 out  error opening DAC module36 dat  It would be good to send this error to stdout     2007 07 02 02 27 S  Schaetzel no beam    CALICE eLogbook   Results for search of  20070701 Page 7 of 22       machine broken    monitor reads    CPS magnet problem   no piquet or expert available  No beam for the night   More news around 9 00 AM    2007 07 02 02 28 S  Schaetzel LED tuning today 14h 17h    We have to retune the LEDs for module 36 because the DAC value file was missing last time     2007 07 02 08 59 Yoshi No beam    No beam before 12 00 noon    When I arrived in shift room this morning  ECAL slow control was on without any DAQ running   REMIND  Please do ECAL SAFTY RUN when beam is out for long time   In order to avoid that slab is  kept in c
146. DAQ trigger average    The hodoscope has been removed   The discriminator on TDC 1  channel 5 has been exchanged            We went to the x position of last time 20 deg  From this position we measured the distances  driftchamber  to three points on the HCAL front plate     Top of drift chamber window closest to experiment to the incredibly precise and objective  middle   of the HCAL front plate     gt  170 Scm    Top of DC window to left hand HCAL edge  level with window       192cm    Top of DC window to right hand HCAL edge  level with window       164cm    Front plate of HCAL to rear plate  along beam direction       129cm       See p33 of logbook for a diagram        File  http   ttfinfo desy de CALICEelog data 2007 33 19 08 2007 08 19T 12 01  03 00 jpg        We moved 6 cm in direction Saleve to take a set of data in a different position    CALICE eLogbook   Results for search of  20070801 Page 5 of 22          gc ct es l i  HCAL Motor control panel    actual position target position    x    305 00 be oxi  Y   Aas H10 55     Y     Q lost X reference point       o lost Y reference point   O lost 2 reference point   O stage in motion   O marras  operation get Pos       feas   Fast V pulsesfeee   extern STOP   ref 57   595 2 495 000       OFF set    RES RS       File  http   ttfinfo desy de CALICEelog data 2007 33 19 08 2007 08 19T12 05 02 00 jpg    2007 08 19 12 14 Jamie  Pavel  beam back    Beni  The beam is back  but we have problems with BENDO3  It is oscillating aroun
147. DC3   gt  HV channel 38    2007 07 09 20 00 Mary Cruz Hengne Niels    330264 beam  steering  test    not for  physics    330265 beam  steering  test    not for  physics  beam  steering  test    not for  physics  beam  steering  test    not for  physics  beam  steering  test    not for  physics    beam  steering  test    not for  physics          2007 07 09 21 20 Roman cogri scripts restarted    I have restarted the cogri scripts       Copying and conversion resumed at Run 330259     Since there might be instabilities with the connection to desy  I will continue to watch them     Cheers     Roman    2007 07 09 21 25 Mary Cruz Hengne Niels    330270 lbeamData  10x10  50GeV  e  530 Air 6mm off 200 k  Pb evts    2007 07 09 21 27 Mary Cruz Hengne Niels  ron pepe E EE EE    Pb evts  HAC pmpa IO O IO e E D TEE    Pb evts       330273 lbeamData  10x10  50GeV  pi  512 4mm Air Off 200k  Pb evts   330274 lbeamData  10x10  40GeV  pi  512 4mm Air Off 200k  Pb evts       2007 07 10 00 07 Kurt  Wenbiao   Trygve    330275 ahcGain lead air Off 11  configurat ion    2007 07 10 00 18 caliceon calicehd GAQ2ctrrents   CALICE user     What are these spikes starting to appear at around 23 00     Plot shows HV 21 40     CALICE eLogbook   Results for search of  20070701 Page 7 of 25             bt       40     20    I I I I I I I I I m I I T I I I I I   I  22300 23 00 DUDU 01 00  03  OF  etl  09 07 2007 10  OF  2007 10  Hr  CH    File  ito   Ittfinfo desy  de CALICEelog data 2007 28 10  07 2007 07 10
148. EN          CO  00   pb  ON  al    CALICE eLogbook   Results for search of  20070701 Page 3 of 22       330558    50 E159 S             2007 07 17 13 33 HCAL currents    for reference    Em Ea CALEN   i02     292  Kaes  99  RI   F 0258   w 310A 4 114 W IRSA      w moe 144 F gae  Fima F RIA      me   53 M ajere  F ais f 1 M RBA    DATE  an   E F oes      i15 3A    175  Ra  w 338   1 w 9 238 Pe 748   F Dan   149 W  304   146 pr RRA    OW aia   73       195 w opoe   y 761568   Ww 315A    7   105 MBA   es w 7TRSA   e 01158 zo aa mae 72    Teja  PALLAS    150   w NXA   275 790A    o    m O     37 e     e     HEN  mmm JD M ramp  P DOWN wm en em  ONLY HY OFF crio    L ona    Ess     zen   en ced  recover    KILL HY   Wi settings   File actions      enable   disable ALL channels    w2   ws   ws   sw   vm   ET      eub       disable   mess ail tripe   3  Save   Sway Lv detail         S dae W   values to file   done ALL valise  position     File  http   ttfinfo desy de CALICEelog data 2007 29 17 07 2007 07 17T13 33 02 00 jpg                      A 2007 07 17 13 59 Beni one CMB poer supply failed    During CMB checks I discovered that one of the two CMB power supplies had failed  From the  slowcontrol  history it looks like this happened already 13th July between 14 15 and 14 30     The CMBs were supllied with 7 V instead of 14 V  Nevertheless we could see LED monitoring data in  the last runs     I reset the CMB power suplly  We now have 14 V again     2007 07 17 14 41 Allister  Imad  
149. Entry  graphs  15    i suggest to get the mppc temperature you average the values for the  monitors mounted directly on the layers  0  gt 5 and 8  gt 13          File  http   ttfinfo desy de CALICEelog data Analysis SCECAL 2007 03 16T09 15 36 00 txt    2007 03 16 09 15 S  Uozumi precise detector scan       400473 MIP run TIL 3GeV direct 2 5V scanned precise  fiber scan  9x9    2007 03 16 09 15 S  Uozumi precise detector scan    400473 MIP run TIL 3GeV direct 2 5V scanned precise  fiber scan  9x9       2007 03 16 09 16 S  Uozumi precise detector scan   400474  MIP run  rir een  direct       p 5v scanned  precise         CALICE eLogbook   Results for search of  20070301 Page 11 of 76        LLL LL ABE LL RARA     2007 03 16 09 16 S  Uozumi precise detector scan    400474 MIP run TIL 3GeV direct 2 5V scanned precise  fiber scan  9x9    2007 03 16 09 18 daniel temperature plotting script          File  http   ttfinfo desy de CALICEelog data Analysis SCECAL 2007 03 16T09 18 52 00 C  2007 03 16 11 35 daniel temperature data and macro  please see the following links     http   ttfinfo desy de CALICEelog show jsp dir  Analysis SCECAL amp pos 2007 03 16T09 15 36    http   ttfinfo desy de CALICEelog show jsp dir  Analysis SCECAL amp pos 2007 03 16T09 18 52    2007 03 16 14 25 S  Uozumi precise detector scan    CALICE eLogbook   Results for search of  20070301 Page 12 of 76    400475 MIP run T1L 3GeV direct fiber 2 5V scanned precise  detector  scan  81              2007 03 16 14 25 pre
150. HCAL 23 21 21 23 99 15 00 00 28000 31000   2 4 PMT 14   5 1 AHCAL 22 20 32 22 99 15 00 00 28000 31000   5 2 AHCAL 21 19 24 21 99 15 00 00 28000 31000   5 3 AHCAL 20180 0 0 150 0 0 0   5 4 AHCAL 19 17 27 19 99 15 00 00 28000 31000   5 6 AHCAL 18 16 15 18 99 15 00 00 28000 31000   5 7 AHCAL 17 15 25 17 99 15 00 00 28000 31000   7 0 AHCAL 16 14 29 16 99 15 00 00 28000 31000   7 1 AHCAL 15 13 3 3 99 15 00 00 28000 31000   7 2 AHCAL 14 12 8 14 99 15 00 00 28000 31000   7 4 AHCAL 13 11 0 0 0 150 0 0 0   7 5 AHCAL 12 10 0 0 0 150 0 0 0   7 6 AHCAL 11 9 0 0 0 150 0 0 0   7 7 AHCAL 10 8 0 0 0 150 0 0 0   9 4 AHCAL 9 7 0 0 0 150 0 0 0   9 1 AHCAL 8 6 0 0 0 150 0 0 0   9 2 AHCAL 7 5 0 0 0 150 0 0 0   93 AHCAL 6 4 0 0 0 150 0 0 0   9 5 AHCAL 5 3 0 0 0 150 0 0 0   9 6 AHCAL 4 2 0 0 0 150 0 0 0   9 7 ABCAL 3 1 0 0 0 150 0 0 0   2 5 TCMT DAC tcmtl dat 46 32 27 10000 10000   2 6 TCMT DAC tcmt2 dat 46 33 27 5500 5500    CALICE eLogbook   Results for search of  20070601 Page 19 of 19          2007 06 30 16 06 Erika  all runs not mentioned are garbage   50027  50028  50029  50032 jahcCmNoise o  o    2007 06 30 18 15 Shift crews hold scan   230014      BeamHoldScan iOxi0 muon      combined run                       y    2007 06 30 18 25 Shift crews hold scan    330015 BeamHoldScan I0emXi0cm  mon         combined           J   __     CALICE eLogbook   Results for search of  20070501 Page 1 of 9       2007 05 02 15 17 Sebastian Modules 27 29 31 deinstalled  Schaetzel   2007 05 04 14 37 Sebastian t
151. Have fun   Cheers   Roman    2007 03 13 08 15 Thanks     Thanks  Roman     Satoru    2007 03 13 08 22 S  Uozumi LED run  fiber direct  2 5V LED  run for    11 12 with  calibration  ode        light yield too low         LED run  fiber direct  2 5V  LED  run for  L1 L2 with       2007 03 13 08 22 S  Uozumi    calibration  mode    light yield too low          2007 03 13 08 23 S  Uozumi LED run  ee eee Par po far    2007 03 13 08 23 S  Uozumi LED run  eee DUE lo im    2007 03  Li 08 24 ES  Uozumi LED run    400365 fiber direct  2 5V LED  run for  L1 L2 with  calibration  mode       Small amplitude LED pulse for Ll    2007 03 13 08 24 S  Uozumi LED run  H00365       p      pe                          Eibez sirestp 5v            T        L9               CALICE eLogbook   Results for search of  20070301 Page 2 of 76    run for    L1 L2 with  calibration  mode        large light yield for L1         2007 03 13 08 25 S  Uozumi LED run       direct   LED  run for  L1 L2 with  calibration  mode        small amplitude LED pulse for L2          2007 03 13 08 25 S  Uozumi LED run  LED i direct   LED  run for  L1 L2 with  calibration  mode       small amplitude LED pulse for L2              2007 03 13 08 26 LED run       S  Uozumi  direct   LED  run for  L1 L2 with  calibration  mode        large LED pulse for inter calibration         2007 03 13 08 26 S  Uozumi LED run  LED i direct   LED  run for  L1 L2 with  calibration  mode           large LED pulse for inter calibration         2007
152. He Crate ON W Crate  FF  recover  KILL wy    HEA   o   87 1 19        EE wu   recover all trips je 3 wE wg   LV detail    so Hi values to file dump AL values  position     File  http   ttfinfo desy de CALICEelog data 2007 30 28 07 2007 07 28T 15 1 1 02 00 jpg    HY settings   File actions erable   disable  LL channels  SHE Lom   oi   enable   disable            2007 07 28 15 28 Ariane  Martin Catljarikarun       331058 beamData  10x10  15Gev pi  524 4mm Air 10deg  Pb junk    tracker off  run stopped    2007 07 28 15 36 Ariane Martin Cathunkeagain       331060 beamData  10x10  15Gev pi  524 4mm Air 10deg  Pb junk    tracker off  run again cancelled    2007 07 28 15 42 Ariane Martin Cathunkeagain again       331061 beamData  10x10    15GeV pi  524 4mm Air 10deg   Pb junk    surprise  again junk    2007 07 28 15 45 Ariane  Martin Catlpirid8 GeV    Pb    DAQ rate  85 Hz  3900 pps    2007 07 28 16 19 Lukas  Alex 10x10 pi        Pb    2007 07 28 16 20    beam stopped  we are rotating to 20 deg    2007 07 28 19 33 Erika Roman Beni Alex Lukas Fab  ECAL AHCAL rotated by 20 deg and re staggered    2007 07 28 19 34 Fab    After the rotation the laptop controlling the ECAL stage again lost contact with the stage     CALICE eLogbook   Results for search of  20070701 Page 9 of 16       The link between the laptop and the DAQ is established and the DAQ can communicate with the laptop  using the usual socket  but the slow control program controlling the movements of the  stage Cannot establi
153. I have stopped the cogri scripts around midnight due to  problems with a new ce at DESY     I took the opportunity to update the scripts such that they don t fail  on one unaccessible replica anymore     I will inform when the scripts will be restarted again    Cheers    Roman         2007 07 28 08 25 Roman Cogri Scripts restarted   gt  Copy to desy  I have restarted the cogri scripts     Copying and conversion resumed at run 331029   The output is sent to desy     The control of the cogri scripts has been handed over to the shift crew   Cheers     Roman    2007 07 28 09 14 Ariane  Martin Catloimb run finishing       DAQ rate  60 Hz  2100 pps    the beam went off from 400sec till 630sec  finished the last comb run   250k events    2007 07 28 10 17 Ariane  Martin Catlpiria amp  GeV    CALICE eLogbook   Results for search of  20070701 Page 7 of 16       Pb    DAQ rate  90 Hz  6000 pps  150k events    2007 07 28 10 55 Ariane Martin Cathpirin8 GeV    Pb    DAQ rate  70 Hz  3000 pps    layer no 20 noisy  changing the trim05 to  4A after 65k events  therefore we go for 200k events    2007 07 28 11 35 Ariane Martin CathpiriBeGeV    Pb       miserable rate  about 30 pps  we changed the collimators position according to general rule    DAQ rate  50 Hz  unstable beam from CPS reported  continuing run nevertheless    a big bug has attacked the crew  the situation is contained    2007 07 28 12 40 Ariane  Martin CatladYa Gain    no beam from CPS  no time estimate    2007 07 28 13 16 Arian
154. L           ZukEqeGRqmRMc          Tag ebed ele seta    u  T          LJ    AAA A E A Los Ae We De Gecko Yee oo    Page 5 of 19    CALICE eLogbook   Results for search of  20070601 Page 6 of 19       2007 06 24 15 22 S  Schaetzel PIN diode read out working again    The read out of the HCAL PIN diodes had suffered badly from the transport to CERN  It has been  rebuilt  and was checked today  All 432 channels are working     2007 06 24 17 41 first AHCAL run taken  gain comissioning run    50000 modules 37 4 38 connected  50001    these runs were without proper hold         2  2007 06 24 17 43 broken FEs   AHCAL slot 5 FE 1 bad pedestal and noise on one channel  Matt claimed he repaired this  but I still  see it    AHCAL slot 5 FE 2 B side doesn t receive SR anymore  this is new since DESY     2007 06 24 21 07 Beni holdscan    250002  hcCmLedHoid can   EMB was not properly set up  2007 06 24 21 10 Beni gain comissioning  2007 06 25 15 00 Beni trigger test run    330005 LLL Eeampata  2007 06 25 17 07 hadronic analysis note    This note contains preliminary CALICE results  and is for the use of membe    CALICE eLogbook   Results for search of  20070601 Page 7 of 19          File  http   ttfinfo desy de CALICEelog data Analysis AHCAL Hadronic shower analysis 2007 06 25T17 07 55 00 pdf          2007 06 25 18 07 Combined Pion Analysis note by NIU for LCWS 2007    Combined Pion Analysis note by NIU for LCWS 2007    This note contains preliminary CALICE results  and is for the use of m
155. L2S15  2007 03 23 18 46 S  Uozumi LED run  400697 CmLed LED KNU fiber 2 5V for  ml  E LOL I I  of L2S15  2007 03 23 18 46 S  Uozumi LED run  AR      inter calibr  tion  of L2S15  2007 03 23 18 46 S  Uozumi LED run  E PTE TL TA  inter calibr  tion  of L2S15  2007 03 23 18 47 S  Uozumi LED run  400700 CmLed LED KNU fiber 2 5V to  peak at  L2514 15  2007 03 23 18 47 S  Uozumi LED run  pA d 1L qa  See lpe  peak at  L2814 15  2007 03 23 18 48 S  Uozumi LED run       400706 CmLed LED KNU fiber 2 5V for  inter calibr  tion  of L6S14    CALICE eLogbook   Results for search of  20070301 Page 44 of 76                    2007 03 23 18 48 S  Uozumi LED run  p gea pce  inter calibr  tion  of L6S14  2007 03 23 18 48 S  Uozumi LED run  pg uec ap ue ec gp e  inter calibr  tion  of L6S14  2007 03 23 18 48 S  Uozumi LED run  EP a  inter calibr  tion  of L6S14  2007 03 23 18 49 S  Uozumi LED run  400709 CmLed LED KNU fiber 2 5V to  peak at  L6S14 15  2007 03 23 18 49 S  Uozumi LED run  400709 CmLed LED KNU fiber 25M  to  peak at  L6S14 15  2007 03 23 18 49 S  Uozumi LED run  KNU fiber 225V for  AAA E umero  of L24515  2007 03 23 18 49 S  Uozumi LED run  400710 CmLed LED KNU fiber 2 5V for  i A TI A res  of L24S15  2007 03 23 18 50 S  Uozumi LED run  p 3e qe qp PAT ene  inter calibr  tion  of L25S15  2007 03 23 18 50 S  Uozumi LED run  E qu qo dps ee  inter calibr  tion  of L25515  2007 03 23 18 51 S  Uozumi LED run  pup cp uer Oo ce  inter calibr  tion  of L25515  2007 03 23 18 51 S  Uozumi LED
156. LED pedestal for Layer25  2000  so we did a HCAL DAC settings reload       2007 07 29 18 17 Catherine  Tom   20 GeV e  run  Alex    E Sg LL LU due EK        m   D t IE a LA    There is a high number of pions in the beam  but cherenkov seems to work correct    CALICE eLogbook   Results for search of  20070701 Page 14 of 16    SERE    tan          File  http   tfinfo desy de CALICEelog data 2007 30 29 07 2007 07 29T 18 17 51  02 png    0 0 20deg   closed  Cl  05 06  E10    11  Ito 4 4   stopped  for pos     2007 07 29 19 01 Catherine  Tom   Alex       NN NN       II  331121 lbeamData  10x10    15 GeV 521 Air Air 0 0  pos corr   peccet    set    CALICE eLogbook   Results for search of  20070701    B NM    2007 07 29 19 36 Catherine  Tom  AHCAL LED pedestal  Alex    AHCAL LED pedestal for Layer25 still  2000    2007 07 29 20 19 Catherine  Tom   12 Gev electron run  Alex    8 MN  321125 be  331124    amData  10x10  12 GeV    TI  12 GeV    2007 07 29 21 11 Catherine   Alex    i NI i ep    set cl  c5 c6 c10 c11 to    5    2007 07 29 22 06 Catherine  Tom  20 GeV pion run  Alex    i 7  331130 beamData  10x10  20 GeV  A a EE    2007 07 29 22 50 Catherine  Tom  15 GeV pion run    Page 15 of 16       CALICE eLogbook   Results for search of  20070701 Page 16 of 16       Alex    i pi  Pb         E WI eit  e  Pb    331132 lbeamData  10x10  15 GeV   pi  20de9   E lpbositionihg     2007 07 29 23 27 Catherine  Tom  12 GeV pion run  Alex    331136 lbeamData  10x10  12 GeV   pi  524 Air 8mm BER 
157. LED run for Layer 6 and 11    Run summary of LED run performed today to see gain of Layer 6 strip 14     Pedestal runs    400442     CmNoise   400443     PmNoise   for Layer6 Stripl4    400440     CmLed for inter calibration   400441     PmLed for inter calibration   400445     CmLed to see lpe peak   400446     CmLed to see lpe peak around strip 14     For Layerll Stripl5              400449   CmLed for inter calibration  400450     PmLed for inter calibration  400447     CmLed to see lpe peak  400448     CmLed to see lpe peak around strip 15   2007 03 14 17 53 S  Uozumi et al  tungsten removed    Now we removed the tungsten plates from the SCECAL and take another calibration run  followed by  precise position scan with the MIP     2007 03 14 17 55 S  Uozumi calibration run without tungsten  MIP fiber 2 5V scanned 36  calibration direct points    scan for  IP       2007 03 14 17 55 S  Uozumi       calibration run without tungsten  MIP fiber scanned 36  calibration direct points  scan for   IP  calibration          S  Uozumi       2007 03 14 19 19    MIP fiber scanned 36  calibration direct points    calibration run without tungsten          2007 03 14 19 19 S  Uozumi calibration run without tungsten    MIP fiber scanned  calibration direct       calibration       2007 03 14 20 21 S  Uozumi calibration run without tungsten  MIP fiber scanned 36  calibration direct points   scan for       IP    2007 03 14 20 21 S  Uozumi calibration run without tungsten    MIP 3GeV fiber 2 5V 
158. LICEelogidata 2007 28 11 07 2007 07 11T07 30 07 00pg    2007 07 11 07 54 Trygve    New DAQ files for TCMT loaded  Making some calibration runs    30343  30343  30351    30352  30353 topped after 10k events  30354  30355       2007 07 11 08 40 Trygve    Machine day started  No beam  ECAL safety run until experts come             2007 07 12 11 30 PaulD In spill buffer limit reduced  2000   gt  1800    The figure shows an example high beam intensity run 330247   This is one of the beam data configurations  i e  not pedestal    or LED periods   Green is triggers  which only occur during  spills  and blue is event readout  which can be within and outside  SPILLS Js    There is a clear alternating structure  caused by taking more  triggers within a spill than can be read out before the next spill   Hence  the first event readout period eats into the time of the  second spill  causing it to be short and hence have a lower number  of events  The third spill does the same to the fourth  etc  This  reduces the overall average rate significantly  the gaps show the  full possible bandwidth is not being used     By limiting the number of events buffered per spill  then this  alternating structure can be smoothed out  Paradoxically  by  limiting to a lower number of events in a spill  the average rate  should be increased     I have changed the buffer limit from 2000 to 1800 events  Note   this is not the absolute limit of the number of triggers per  spill  it is the limit on the number wh
159. Mangi  calibration data  Nanda    330602    CALICE eLogbook   Results for search of  20070701 Page 17 of 19       2007 07 17 17 54 Valeria  Mangi  pion run  Nanda    Pb       trigger  8 2kpps  DAQ rate  69Hz          2007 07 17 17 59 Valeria  Manqi  LED check    Nanda  adjusted VCalib settings in AHC cfg     3000  3000    5000  3000    EE                LLL  3098   000  ES EIER B3000  EE                28000 SLI    0  000    Cot Gof Co  Co  Cop BS  Co  GO Co     LA     33000 IO  GEI   33000 LLL   B3000 S          performed DAQ shutDown  DAQ startUp  CMB manual re init  amp  CMB check   Now  online monitoring of Run330603 it looks much better   But we still don t see licht in layers 10 18 28 29 35    v   4X       ECAL HEAL   veut   rte   Padomi   Copan 00   Creech   Sio Sege   Responsa   Pedestus Less  une                              py ama                trigger  6 5kpps  DAQ rate  80Hz    CALICE eLogbook   Results for search of  20070701 Page 18 of 19       2007 07 17 19 30 Nanda Mangi Valgsipn run       Pb    collimator C10 from 4 to 3  trigger rate  10kpps  DAQ rate  90 Hz    2007 07 17 20 29 Valeria  Mangi  pion run  Nanda    Pb    trigger  4 3kpps  DAQ rate  85Hz          2007 07 17 21 16 Nanda Mangi Valezalibration runs  30607 ahcGain o e  30608  30609  30610 ahcCmLedvcalibScan    30611 ahcPmLedvcalibScan E   2007 07 17 22 04 Valeria  Manqi  pion run   Nanda       Pb    trigger  4 3kpps  DAQ rate  85Hz    2007 07 17 22 54 Valeria  Mangi  Guassian Fit for Pion run 33058
160. Marcel Roman    Please leve it untouched  unless there is a specific reason to do it   Please always document any changes of the ECAL position on the elog     2007 07 12 20 35 Marie Cruz  Tao  beam tuning  Nanda    330372 lbeamData  10x10  25GeV 535 air 6mm  887 check  Pb ECAL  position       ECAL back to centre of beam  1800 pps in 10x10 trigger coincidence  DAQ rate  60 Hz       A 2007 07 12 20 55 caliceonOcalicehdigblAceentch   CALICE user     Very high current spikes in HV channel 77   TCMT 1  HV is instable  too    CALICE eLogbook   Results for search of  20070701 Page 5 of 20          02299  na                         ia m A al A   apap ben Ge  oe m  mm    ef           _             m MG    P         enne RS SAO tease pe ek t seed        00 ZU7 20 202 40 21300   200  L207 200      12 07 2007 12 07  200    File  http   ttfinfo desy de CALICEelog data 2007 28 12 07 2007 07 12T20 55 01 00 jpg       2007 07 12 20 58 Marie Cruz  Tao  beam tuning  Nanda    330374 beamData  10x10  15Gev le  536 air 6mm 2 887 looks  Pb Crap    330375 beamData  10x10  15GeV e  536 air 6mm 2 887 also  Pb looks  crap       330376 beamData  10x10  15GeV e  536 air 6mm 2 887 seems  Pb Okay    530 pps in 10x10 trigger coincidence  DAQ rate  30 Hz          2007 07 12 21 26 caliceon caliceheusb BReckofiiCMT1   CALICE user     since HV channel 77 continued with very high current  up to 300 uA  we switched it off   Expert intervention is needed to check VFE electronics    CALICE eLogbook   Results for se
161. Marius DAQ   SC communication to CERN database    During the testbeam period  the DAQ was sometimes stucked  because the CERN database was offline     At a certain point we have used bit 7 from the DAQ runStart argument to start a run   v  to decide  if the DAQ should ask for the CERN database or not     set bit 7    v 128   for example    use trigger 100x100 with database   v 19   use trigger 100x100 without database   v 147    This is not needed anymore  since Sven has implemented a timeout in the SlowControl communication  to the CERN database     A 2007 02 27 14 07 Marius AhcSlowReadoutData  bug in temperature readout in  the october run period    A bug in the temperature   slow control information readout exists  which leads to the fact that  we have lost the slowReadout information for the following modules in the october run  period     Module number  18  20 and 22    Due to same bug you will get at the end of the slowReadout list information for module 25  27 and    CALICE eLogbook   Results for search of  20070201 Page 17 of 21       29  which are of course only filled with 0  because these modules did not exist   were  not connected     Also the part of the slowReadout information  which was read out over the very frontend side and  not over the CAN Bus CMB side has not worked always correctly  We have found some runs  in which this part of slowReadout information was for all modules the same  We do not  know what has happened and which runs are affected in detail  
162. N docu 2007 02 26T20 19 42 00 paf    la 2007 02 26 20 30 Marius DAC calculation procedure    All given DAC values in this documentation are calculated with the following formula   dac    hv hbab   U_sipm   dacoffset    voltagebin     hv hbab   voltage for the half module   U sipm   needed SiPM bias voltage   dacoffset   0 0  to give a common offset   not used  voltagebin   0 01953125     5V   256 bits    IMPORTANT    To avoid the non linear range of the DAC  we add to the maximun needed SiPM voltage for each module   hv hbab   0 3V  This leads to a file with no 0 entries for the DAC values    It is also important to avoid the DAC range from 4 7 V on    try to minimize DAC value entries   higher   than 240    Both should be possible for all produced modules  23  until now  except module 2  where even the  complete DAC range is not big enough  see DAC summary pdf in this elog folder      REMARK    It is not completly clear if the DAC voltage slope given by the voltagebin variable is correct   Since it is not easy to measure the DAC output without influencing it  there are three  different results on the market    LAL   Alexandre Karakash   Benjamin   Therefore we are using the conservative approach and are using the mathematical value      gt  should be checked again    la 2007 02 27 07 59 Marius AHC cfg files august    enclosed all used AHC cfg files during the august run period  more information in the next entry    CALICE eLogbook   Results for search of  20070201 Page 8 of 
163. O3 problem again     Stopped run 350293 and call to accelerator control room     2007 08 20 06 21    CALICE eLogbook   Results for search of  20070801 Page 13 of 22       Take calibrations meanwhile     2007 08 20 06 21  50294 BhcGain                             50255 BhcCmNoisg                                 SESS  50296 BhcPmNoise                                  50297 ahcCmLedVcalibScan  50298 ahcPmLedVcalibScan       2007 08 20 07 45    Problem on BENDO3 seems to be gone  but still need to keep eyes on it   Start another run anyway     2007 08 20 07 46  Pb          2007 08 20 10 34 Erika  Beni  rotation to 30 deg  Jamie  Pavel    We have rotated the HCAL to 28 3 degrees     2007 08 20 11 17 Erika  Beni  pion run  Jamie  Pavel    Pb       2007 08 20 11 23 caliceon calicehoatssittheasation movement   CALICE user     We move to where we guess the new 0 0 position is     Due to problems during the stage movement  the y position reading  is arbitrary                    meng carae j  HCAL Motor control panel  actual position target position  es  Tas TE  Q lost X reference point   Qo lost Y reference point  e lost 2 reference point  O ma osa  O ens ELT       ais 1 Fast V pulsestree   extern STOP   ref 6l   349 975 249 975    OFF set    H  900 H Loy PING  to elos          File  http   ttfinto desy de CALICEelog data 2007 34 20 08 2007 08 20T11 23 02 00 jpg    2007 08 20 11 34 beam position in HCAL    Our guess was good  We are centered     CALICE eLogbook   Results for search of 
164. Results for search of  20070701 Page 20 of 24       2007 07 08 02 19 Alex  Tao  Nanda ahcPmNoise run   2330200 fahcPmNoise_  CT S o S d S       30k events    checked CRCs and found all alive    2007 07 08 02 39 Alex  Tao  Nanda    i P LII  Pb    330202 lbeamData  10x10  40GeV  e  531 6mm  Pb   bid DEE KSE A AO ue  Pb                               2007 07 08 03 36 Alex  Tao  Nanda pion run  330204 lbeamData  10x10  80GeV  pi  510 Air 4mm Off 250k  Pb event  2007 07 08 04 28 Alex  Tao  Nanda pion run  330205 lbeamData  10x10  50GeV  pi  512 Air 4mm off 0 0 250  Pb even  2007 07 08 05 19 Alex  Tao  Nanda  330206 lbeamData  10x10  40GeV  pi  514 Air 4mm off 250  Pb even  2007 07 08 06 08 Alex  Tao  Nanda pion run  330207 lbeamData  10x10  100GeV  pi  516 Air 4mm off 270  Pb event  2007 07 08 07 00 Alex  Tao  Nanda electron run  330208 lbeamData  10x10  50GeV  e  530 6mm air off 25k  Pb events  2007 07 08 07 53 Alex  Tao  Nanda ahcPmNoise run  330209 ahcPmNoise 10k  events  2007 07 08 07 55 Alex  Tao  Nanda electron run    330210 lbeamData  10x10  40GeV  e  531 6mm air off 250k  Pb events       2007 07 08 07 58 Alex Script checking if CRCs are alive    I wrote a script  checking if ECAL and AHCAL CRCs are alive   Start a terminal on calice02 and type  ssh checkCrcReadout     One should cross check with the instructions given by Fabrizio below from time to time  since the  Script was written at night     e Ck Ck Ck Ck Ck CK CK CK CK CC CC CC C CC CK CC CC CC CC CC CC CC CC CC CC CC
165. Scan and  ahcPMLEDVcalibScan Runs    330457 jaahcCMNoise 10k  Run events  330458 jaahcPMNoise 10k  Sp Do  o po  oo Lies    330459 jaahcCMLEDVgalibScan 20k  Rp ITT Kies    330460 ahcPMLEDVqalibScan 43k  pa prep  gt  gt  pop    flv      Take AHCAL calibration       2007 07 15 03 01 Valeria Wenbiao  ii40 GeV       Pb    DAQ rate  82 Hz   4 0k pps in 10x10 trigger coincidence    beam is back    CALICE eLogbook   Results for search of  20070701 Page 2 of 19       Pb    DAQ rate  66 Hz  Spill rate 2 1k        Searching ECAL wafer edge     oops  wrong side    File  http   ttfinfo desy de CALICEelog data 2007 28 15 07 2007 07 15T05 04 02 00 jpg    ECAL p TAS 1 wafers    Stage control   Citrato      HCAL Motor control panel    File  http   ttfinfo desy de CALICEelog data 2007 28 15 07 2007 07 15T05 09 01 00 jpg        Test run to check movement of stage   40GeV Pions  Run330464    CALICE eLogbook   Results for search of  20070701 Page 3 of 19          Arpon   peores 2ePrjecton              Enecpyverime      E      E    aman o ausan nae oo a RO  pe SS                File  http   ttfinfo desy de CALICEelog data 2007 28 15 07 2007 07 15T04 55 08 00 jpg    2007 07 15 05 25 Valeria Wenbiao Qliosr run       Pb    junk run  the absorber and the targets were wrong    2007 07 15 05 28 Valeria Wenbiao Qliosr run       Pb    DAQ rate  90 Hz  Spill  10kpps    2007 07 15 06 17 Valeria Wenbiao Qliosr run       Pb    DAQ rate  101 Hz  Spill  5kpps    2007 07 15 07 06 Valeria Wenbiao Qliosr run      
166. T00 18 02 00 jpg    2007 07 10 00 39 Kurt  Wenbiao  beamData  Trygve    pre eq vent  CZ Im bens     pue 7  ERES  pe pues 6mm 200k  PPP  E ff tiene         E p p  20GeV 6mm  Pb  beam for  after  84k  events  in this  run    2007 07 10 02 32 Kurt  Wenbiao  ECAL energy  Trygve  Looking at the peak position for the energy depositied in ECAL when taking e  runs   40 GeV  Run330276      4800 MIP  30 GeV  Run330277    gt  2500 MIP  20 GeV  Run330278      2500 MIP    The  30 GeV run  looks more like 20 GeV       I checked the magnet setting files  bending magnets have same settings in files 532  amp  533 which  explains why Runs 330277 and 330278 appears to have same energy        2007 07 10 03 08 Roman Copy scripts stopped   Hi there   Since it seems as if there are persistent problems with the srm dcache I have  to stop the cogri scripts once more    Cheers     Roman    CALICE eLogbook   Results for search of  20070701 Page 8 of 25       2007 07 10 03 23 Beni moved to new position    reached  0  3cm     HCAL Motor control panel    O lost X reference point    O lost Y reference point       Q lost Z reference point   Q stage in motion   Qo manual operation get Pos     Seas   slow  puse free ememSTOP ref 8 319960 31995       OFF set  Biz za PANEL toeLog    File  http   ttfinfo desy de CALICEelog data 2007 28 10 07 2007 07 10T03 23 02 00 jpg  2007 07 10 03 36 Kurt  Wenbiao  Beam Data    Trygve    E   Gm  seg nad pem pum E  E mer 200k  mew   S    E  gt  E mE Efe  200  Pb   beam for   10
167. U fiber  2  scanned 9x9  points  precise  detector    S  Uozumi          with positron beam  DAQ rate is  120Hz             3 GeV run    KNU fiber  2  scanned    2007 03 24 09 21 S  Uozumi    detector  scan       3 GeV run    KNU fiber E  scanned 9x9  points  precise  detector    S  Uozumi       2007 03 24 09 21       3 GeV run    KNU fiber E  scanned    S  Uozumi       2007 03 24 09 22    detector  scan       3 GeV run   KNU fiber  2  scanned 9x9  points   precise    S  Uozumi       2007 03 24 09 22    detector    2007 03 24 09 22 S  Uozumi 3 GeV run    400770 beam T1L 3GeV KNU fiber E  scanned  detector  scan       CALICE eLogbook   Results for search of  20070301 Page 54 of 76          2007 03 24 09 22 S  Uozumi 3 GeV run  400770 beam T1L 3GeV KNU fiber  2 5V scanned 9x9  rit ta 3 mus  precise  detector  2007 03 24 09 37 S  Uozumi 4 GeV run  400772 beam T1L AGeV KNU fiber  2 5V scanned 36  9 10  pg x EA AA  2007 03 24 09 37 S  Uozumi 4 GeV run  400772 beam TLE 4GeV KNU fiber  2 5V scanned 36  9 10  SE EE rue  2007 03 24 12 47 S  Uozumi 4 GeV run  400773 beam T1L 4Gev KNU fiber  2 5V scanned 36  9 10  EE cal SESTINMLL  2007 03 24 12 47 S  Uozumi 4 GeV run  400773 beam T1L 4Gev KNU fiber  2 5V scanned 36  9 10  pace qp uq I  l  2007 03 24 13 30 S  Uozumi Electron beam  long time no see    We got a call from control room  they are still working on the HERA injectin problem    but for a moment they switched to the electron mode just to give us a nice beam    But they are not s
168. V tertiary beam resulting from a 10 GeV secondary beam seems to be  splitted into 3 peaks  The first peak might be muon contribution  although I  don t see any in the corresponding run from a 50 GeV secondary beam  But the  close by main peaks  indicate that it is not a good idea to extract a  tertiary beam from the tail of a secondary one     CALICE eLogbook   Results for search of  20070601 Page 12 of 19          File  http   ttfinfo desy de CALICEelog data Analysis AHCAL EM shower analysis 2007 06 29T 16 52 04 04 pdf    2007 06 29 17 18 Nanda em shower analysis note  This note explaeins the analysis procedure and contains preliminary CALICE results   It is for the use of members of the Calice Collaboration and others to whom permission has been  given     TITLE   First results from electron data with the CALICE tile HCAL prototype at the CERN test beam    CALICE eLogbook   Results for search of  20070601 Page 13 of 19       File  http   ttfinfo desy de CALICEelog data Analysis AHCAL EM shower analysis 2007 06 29T17 18 50 00 ps    2007 06 29 17 38 G Mavromanolakisnuon through ecal    CALICE eLogbook   Results for search of  20070601 Page 14 of 19       Time  06 45 56 071 972 Fri Jun 29 2007    Run 330008 0 Event 1020 Hits  44 Energy  52 7994 mips          sgets ee ba   aa aha sc    nme e     T           Time  06 45 56 071 972 Fri Jun 29 2007  Hits  54 Energy  54 7943 mips    eM    Den    Wan saat IM paa    Run 330008 0 Event 1020       CALICE eLogbook   Results for search o
169. We noticed afterwards a wrong beam file set     2007 08 19 15 54 jumpy current    HV channel 1  no lost DAC visible in data        150  ee H  100  re 50  m 0        L  I       L  I  L                  1 D    l        L       I I I    l i    15 20 15 40 16 00  139 08 2007 139 08 2007 13 08  2007  mum File  http  tfinfo desy de CALICEelog data 2007 33 19 08 2007 08 19T15 54 03 00  pg    2007 08 19 16 06 Jamie  Pavel  no beam  Beni  no beam  no Cherenkov  2007 08 19 16 12 Daniel  ahcGain run  Anne Marie  Petr   Beni  350268 fac     2350269 EEE      250270 JjahctmNoise               0    U U U O  2007 08 19 16 29 raised current limits    changed current limits for HV channels 25  amp  53    CALICE eLogbook   Results for search of  20070801 Page 7 of 22       w STA ES   is 83A   w Js a Re 1448 wm   w 35 185 X IRBA   434 TAR   PEL     145 21088   277 w esua   w IMIA NM WRIA w SRSA     113 m   in Apre   sos 70888  120 W SoA   140   SRA P w TIGA    121 SA 141 W eza cz we    E  e  ies M MSA   126 w TARTA   an wann Kan       197  ep   a  EEN   154 M poa   aw mA    197 O08   2   MENS    103 BIA    e me    60 E    95   pig   169 7 998   102 W  w XX  354 RI   7   e    o oe    KA j Ly LV OFF  J  He Crate ON j W Crate OFF                HN settings   file actione  erable   disable ML channels    sw   we   on   eble   gr  is          2    t dup Wi values to file dom HL valine  position     File  http   ttfinfo desy de CALICEelog data 2007 33 19 08 2007 08 19T16 29 02 00 jpg       2007 08 19 
170. X10cm  muon combined                                        330024 is junk run    2007 07 01 11 39 Erika Beni Yoshi status  At run 330023   muon rate  60Hz   10cmX10cm trigger  6600 events spill      HCAL status  Slot  FE  Module  Comments  9 1 29 Beside lost settings  HV 10V  5 4 35 VFE to be checked  7 6 2 very high noise all module  ECAL status  9 0 chip 11 broken 3  5 All FE off      After 81 config in 75 min of 300k events  run stop   RELOAD HCAL DAQ SETTING  ramp up module 29 to 77 5V  still bad setting     ramp down    run 330025 started    2007 07 01 13 33 Fab Erika Beni Yost  ge shift  SHIFT movable stage 10cm to  Saleve      CALICE eLogbook   Results for search of  20070701 Page 3 of 22       File  http   ttfinfo desy de CALICEelog data 2007 26 01 07 2007 07 01T13 33 33 02 pdf    2007 07 01 13 39 Fab Erika Beni Yobbid scan    pu AAA AA uper  0     we are running with CMB off from this morning    2007 07 01 15 40 Erika Fab Beni Yost  ge shift  Shifted movable stage 10cm to  Jura  with respect to 0 0  position     2007 07 01 15 41 Fab Trigger map for Drift Chambers    Trigger map for the drift chambers       TOCA TDC channel  CS  Er E E  E E A E       CALICE eLogbook   Results for search of  20070701 Page 4 of 22          2007 07 01 15 42 Erika Fab Beni Yo   cmX10cm scintilator shift    10cmX10cm scintilator position was shifted away from ECAL approximatery 67cm compared with original  position in order to keep safety stage moving     2007 07 01 15 52 Erika Fab Beni Y
171. a We  ka ee kk TO    We       File  http   ttfinfo desy de CALICEelog data 2007 29 22 07 2007 07 22T23 53 02 00 jpg    CALICE eLogbook   Results for search of  20070701 Page 1 of 22       2007 07 17 07 32 current increased in HCAL HV ch 6    current in HCAL HV ch 6 increased from 99uA to 101uA at 6 20    monitoring  p Switch On   Off  Voltage  V  Current bei Set  V  Initial  V     am BA 40 2       MW settings   File actions enable   disable ALL chamels    Se   to ERT 7  enable  disable    e 2 Esper   ae ERE o 7  Be   or Eu y MY Crate 0N Wi Crate OFF  mem un  2811 61 82  wee    Swe  ws j ws   remm all trios  MOSS owe J v3   LY deant     dap Wi values to file   so ALL values  position     le Je be be de  e be he he be do le Lo E IE le    Ly   Li   LI   LA   Ly   LJ   LJ   L4   L      e   10 00 W  oW  W       W        K   W   0    10 00    w          ES  HER  Fam            m      a  E  s  mmr   Ep  EE  E  EX    40 70    10 00       File  http   ttfinfo desy de CALICEelog data 2007 29 17 07 2007 07 17T07 32 02 00 jpg    2007 07 17 08 14 Allister  Imad  new shift  Beni   2007 07 17 08 27 Allister  Imad   Beni    330584 ahcGain  30585 ahcPmLedVcalibScan  30586 ahcCmLedVcalibScan  30587 lahcPmNoise   330588 lahcCmNoise       2007 07 17 10 09 new position    moved stage 12 cm to Jura  ECAL 12 cm to Saleve new stage position is  7  1   cm     HCAL Motor control panel    actual position target position  x  7001 po GOXI  yf      0 oc    O lost X reference point  O lost Y reference poin
172. ack  gt getYhits       hitmap for y layers    2007 03 17 10 52 S  Uozumi et al  tungsten inserted to 2nd config SCECAL    We have inserted 25 tungsten plates into SCECAL 2nd configuration     2007 03 17 10 57 daniel tracking code    i made some short documentation for the tracking code here     http   ttfinfo desy de CALICEelog show jsp dir  Analysis SCECAL amp pos 2007 03 17T10 42 33    2007 03 17 11 44 S  Uozumi 1 GeV run  400493 beam T1L 1Gev direct 2 5V scanned 36  fiber 9 10  stage scan  2007 03 17 11 44 S  Uozumi 1 GeV run  400493 beam T1L 1Gev direct 2 5V scanned 36  fiber 9 10  stage scan  2007 03 17 11 47 S  Uozumi 2 GeV run  400495 beam T1L 2GeV direct 2 5V scanned 36  fiber 9 10  stage scan  2007 03 17 11 47 S  Uozumi 2 GeV run  400495  beam TIL 2GeV direct 2 5V scanned 36  fiber 9 10  stage scan  2007 03 17 13 30 S  Uozumi 2 GeV run  400497 beam T1L 2GeV direct 2 5V scanned 36  fiber 9 10  stage scan  2007 03 17 13 30 S  Uozumi 2 GeV run    400497 beam T1L 2GeV direct 2 5V scanned 36  fiber 9 10  stage scan    CALICE eLogbook   Results for search of  20070301 Page 15 of 76       2007 03 17 14 42 S  Uozumi 3 GeV tuns  400499 beam 3GeV direct scanned  fiber E 10  stage scan  2007 03 17 14 42 S  Uozumi 3 GeV Eun  400499 beam 3GeV direct scanned  fiber E 10  stage scan  2007 03 17 18 14 S  Uozumi 3 GeV tuns  400500 beam 3GeV direct scanned  fiber 10  stage scan  2007 03 17 18 14 S  Uozumi 3 GeV tun   400500 beam 3GeV direct scanned  fiber E 10  stage scan  2007 03 17 
173. age 22 of 22       330739  30740  30741    30743  30744    electron    Run No  Energy  GeV mean  mips  sigma  mips  sigma mip        330733 5 201 0 ER            EE  SII   a  30742  12    OI   pe  Po     18     15   12   12   10         CALICE eLogbook   Results for search of  20070701 Page 1 of 19       2007 07 14 03 38 Hakan Oliver Wenbha Gain  230425  Jahccain        gt  gt                        L1  S o      2007 07 14 04 13 new position  moved stage to x center and 1 cm upwards  0  10 in target position  l  Stag coni Ce  j    HCAL Motor control panel    actual position target position    D   0 00 I ox              0  Y   210 00 po Qi   O lost X reference point       lost Y reference point  O lost 2 reference point  O stage in motion    Qo manual operation get Pos                           OFF set  Bj   au p PRE tous            File  http   ttfinfo desy de CALICEelog data 2007 28 14 07 2007 07 14T04 13 02 00 jpg    2007 07 15 00 24 Valeria Wenbiao  lleetron run       Pb    DAQ rate  70 Hz    2 1k pps in 10x10 trigger coincidence   TRIMO6 reading 0 2 BeamRef  0 167  control room confirmed within limit of 0 3  Problems with elog    2007 07 15 00 53 Valeria Wenbiao  lloey    elog is running without problems again    2007 07 15 01 28 Valeria Wenbiao Glleetron run  30 GeV    Pb       DAQ rate  66 Hz   2 0k pps in 10x10 trigger coincidence    stopped after 249 kev because no beam from CPS for 10 minutes    2007 07 15 02 35 Valeria Wenbiao GilwCMNoise  ahcPMNoise  ahccmLEDVcalib
174. age desiplyed in the  DAQ gui    t control    sendAndRecv    std  string recv ed header 11 bytes    gt daqrequest  lt     message  15 bytes    gt 1183421169 0 0  lt      Number of bytes written   2000020316    2000000000    The errlog Log1183412671 err icalice00 says at its end    RcdWriterBin  close   Closed file data run Run330050 003 bin   RcdWriterBin  open   Opened file data run Run330050 004 bin   RcdIoSkt  write   ERROR Number of bytes to send   56556 for record   RcdHeader  print   Record Time   00 02 38 951 602 Tue Jul 3 2007  Type   5   event  RcdIoSkt  write   ERROR Number of bytes to send   56556 for record   RcdHeader  print   Record Time   00 02 38 976 890 Tue Jul 3 2007  Type   5   event  RcdWriterBin  write   Error writing to file number of bytes written 15320    45220 Success  runner  daquser inc run RunWriterDataMultiFile hh 53  virtual bool  RunWriterDataMultiFile  record const       RcdRecord amp     Assertion   writer   write r   failed     There was not run start at all possible    It looks as if the reason for the failure was that the data00 disk   was full  Using my check and delete scripts I have freed some   1 25 TByte on that disk    After that I could start an emcNoise Run    A start of a combined run  however  was  not  possible    Already the DAQ startup was blocked  obviously by the following problems  runner compiled at Mon Jul 2 18 56 42 2007    Command   runner    CrcConfiguration  trgSlot    CrcLocation  print      Crate number   172   0Oxac
175. ain Peak Main Sigma Mean Comment s  GC E AS SEN E miaen ime mem  330657 228 5k 3125 175 4   oisy  A HO AAA    Moes  330658 228 5k 2531 195 3 Sie  330659 229k 2268 SE  EE ee AE  330660 15 231k 145 6 oisy  AAA  330663 K2 185 4k 127 5 oisy   AA a a E a  330664 10 2221k oisy    23 7k  665   668   669   gt one run    7553K Noisy  layer 19  330669 92 9k Noisy  layer 19    20 07 2007    RunNum Energy Beam event Num Main Peak Main Sigma Mean Comment s  br position  reak wictn  OU  ome  330712 67 8k 1074 101 9 5  oisy                      214k 1075 101 8 S oisy  layer 1  oisy  layer 1  232k 1557 4 Ecal    2472                      p JAM Bd jii p ES    2007 07 21 05 12    330751 lbeamData  10x10 I 10GeV e  520 air air 2 6126  3  3 250k  Events    DAQ rate  96 Hz       2007 07 21 05 58       330752 lbeamData  10x10  12GeV e  524 air air 2 1224  3  3 320k  Events    DAQ rate  96 Hz    2007 07 21 06 59 Mandi  Petr   Nicola    330753 lbeamData  10x10  15GeV e  521 air air 1 3613 35  3 270k  Events       DAQ rate  103 Hz    2007 07 21 07 49 Mangi  Petr     CALICE eLogbook   Results for search of  20070701 Page 7 of 19       Nicola    330754 lbeamData  10x10  18GeV e  522 air air 0 9421  3  3 96k  Events       DAQ rate  90 Hz    2007 07 21 08 05 Morin  Lutz   Kieffer    330755 beamData  10x10  18GeV e  522 air air 0 9421  3   3 185k  Events       DAQ rate  90 Hz    2007 07 21 08 54 Morin  Lutz   Kieffer   2330757   abcPmNoise CE we 20000  2007 07 21 08 57 Morin  Lutz   Kieffer       330758 l
176. ake this into account when you go through the checklist     In principle there is no need to restart the cogri scripts by the shift crew  unless    a  being told by the run coordinator s  or by me   b  after a restart of the caliceserv     The cogri scripts should always be started by  start cp sh  from anywhere on caliceserv    This script contains now a safety check whether there are already  running cogri scripts     If it finds one or more running processes it   a  kills the process es  and eveything belonging to it  them    b  re launches the cogri scripts which resume their work two runs  after the one indicated by the special file  last nnnnnn    This avoids the hanging of the scripts on incomplete copied runs   The necessary repair work will then be done afterwards by me   Cheers     Roman    2007 07 19 22 42 Oliver Allister Hengne       Beam is back  at 22 42  continue taking data     10x10 trigger coincidence  1 3 kpps  DAQ rate  60 Hz    2007 07 19 22 43 Beni raise of current limits    For following HV channels the current limit was raised  2  6  25  26  33  58    W IRSA  F 28   E RIA   OW ARTS   E spe   F geg  Se EA  w 9 238 pow 741378   3 MODA   ia TSA  196 W 01308   173 W 75 558  105 KIDS   e pmpa     e   6 W pie 175 70098  129 MOLA 273 W T9M0A   7  OF    146 725 W ad od 160 W ud         fi ae W    rer UP DON   LAV ert   em  uv Ww OFF CAT    LLL LL       S He Crate UN EN Crate  FF  ss   e 7 Ee      recover    Kit ay    R   I  o4   7 19    Wi settings   Filo acti
177. aliceonOcalicesei  dmoise  During current run some noise effect showed up for short time  Most probably it originates from  layer 14  module 16      CALICE eLogbook   Results for search of  20070701 Page 15 of 20    rat MCAL   TCM   Teacsa  Paescieo   Dietei30   CrosChechs   Statue   Setings      Wesgenia   Poses   Leos   nc            I  f       mp     us  u   fa            D  R    File  http   ttfinfo desy de CALICEelog data 2007 31 31 07 2007 07 31T01 05 03 00 gif    2007 07 31 01 05 caliceon calicehda 6b 2   daghoturrent in HCAL channel   CALICE user     This mysterious picture suddenly appeared in logbook ten times  Strange thing happenes on night  shifts           2  2  S  F  11  i 2        21 50 2200 22730 23700  0 2007 SO  OP  2007 SO  07  2007 50 07  2007       File  http   ttfinfo desy de CALICEelog data 2007 31 31 07 2007 07 31T01 05 14 00 jpg    2007 07 31 01 14 HCAL noise    I see a unusual noise in the corresponding FE  Reloading DAC settings  but no effect on module  current     CALICE eLogbook   Results for search of  20070701 Page 16 of 20       hHcalPedRMSPerChannel_slot17       3  E  E  ki     e  co  a  uo  z   74  vc  co  n   channel number  File  http   ttfinfo desy de CALICEelog data 2007 31 31 07 2007 07 31T01 14 02 00 gif  2007 07 31 01 17 caliceonOcalicehcalsc02 cern ch   CALICE user   331199 DAC reload  2007 07 31 01 17 EE cern ch  wee user        i     Stopped after 38k ev     2007 07 31 01 21 HCAL noise    After reloading DAC settings the noise loo
178. alized that we see a lot of pions tagged  as electrons  After some tests we found that the calculated threshold value of 0 94 bar  is too high for 18 GeV  We lowered it to 0 8 bar and we see a much better electron pion  separation  We will keep an eye on this and make a new pressure table        2007 07 22 11 35 Morin  Brient   Lutz       COLL10 positive side went to 4 5mm during run  could be reset  trigger rate 2 8K  daq rate 77 Hz    2007 07 22 12 41 Morin  Brient   Lutz       2007 07 22 12 50 Morin  Brient   Lutz       330804 lbeamData  10x10  12GeV 99K  E events    Noise inside hcal at slot 17 fe 0 modulel    2007 07 22 13 10 Morin  Brient     CALICE eLogbook   Results for search of  20070701 Page 15 of 19       Lutz    330806 beamData  10x10  12GeV pi  524 4mm Air 1 88  6 0 150K  Pb events       2007 07 22 13 46 Brient Morin Lutz beamData       330807 beamData  10x10  10GeV pi  520 4mm Air 2 3  6 0 65  Pb Kevent s    50 50 e pi change to beam absorber 8 mm lead    2007 07 22 14 00 Brient Morin Lutz       330808 beamData  10x10  10GeV pi  520 8mm Air 2 3  6 0 251K  Pb events  2007 07 22 15 21 Brient Morin Lutz beamData    330809 beamData  10x10  8GeV pi  525 8mm Air 2 7  6 0 253k  Pb Events    trigger rate 800  DAQ rate 44 Hz       2007 07 22 16 50 Hengne Oliver Mangi    BhcGain Od  SR LL  B30818                               PheemredvealibSean                     LL  530817                 INE  LLL   HA    330819  test run option  e 10000  v 197  random Trigger    test 
179. all modules    monitoring    safety circuit  printed circuit boards in the middle  plates with connectors left and right side      HH                   Sas  a  aj  CIC  CIAO CEO CUA  i 7s CEA CEA CEA IE  61 71        File  http   ttfinfo desy de CALICEelog data doc electronics 2007 06 14T16 41 30 00 paf    _2007 06 1 7 12 23 Erika module  amp  cable checks    EE HAB 5 coherent noise       CALICE eLogbook   Results for search of  20070601 Page 3 of 19    eecht  A cable broken  HAB 6  coherent noise   B cable broken   B cable broken    HAB 4 coherent noise  HAB 2 coherent noise    DEE EE        lt     lt     ee A cable Broken    STNE broken         SE SILS  2    HAB 6 coherent noise  k    PRA S ceshsrent noise  PESE s   coherent noise    This test was done with the  DESY  test beam DAQ and a protype CRC in slot 12           2007 06 17 16 32 Erika repair work  AHCAL LV HV distribution to modules checked  works fine   All LV HV cables repaired    All installed modules are powered and tested   8 modules at CERN   4 modules at DESY are missing     A total of 15 SCSI cables have to be exchanged for the AHCAL     2007 06 20 09 53  gt   Fwd  caliceserv   Icio and root installed     Dear Colleagues    I have installed root  v5 14 00f  and lcio  v01 08 02  on caliceserv   root has been configured in the following way      configure   disable rfio   disable castor   disable dcache   disable gfal    The machine can be freely used   I will announce times when I need it for test purposes   
180. amHoldScan    30044  30045  2   2     30046 20x20  30047  30048 20x20 ku                  1    330043  spill sync enabled  Split beamRun disabled  Stopped by pressing  START ECAL SAFETY RUN  Button          330044  ahcPmLedHoldScan  spill sync disabled  Split beamRun disabled  extracted Hold PM Led values    got reasonable values for 19 modules only  for modules 1 16 18  20 and 34  extraction of Hold PM led values did not work     pressed  START ECAL SAFETY RUN  Button before run 330045   330045  ahcCmLedHoldScan  spill sync disabled  Split beamRun disabled   extracted Hold CM Led values    Comparision with those already in AHC cfg  Beni put the new ones in    Values were almost all consistent  There were 10 bad modules  where no   proper hold values could be determined from this run    Two of them  Modul 24  CMB 30  and Modul 36  CMB 33  did not answer in mainCMB panel   The other bad Modules are  1  2  10  12  13  14  20 and 34    CALICE eLogbook   Results for search of  20070701 Page 9 of 22       330046  spill sync enabled  Split beamRun enabled    run 330046 stopped by Beni for ahcGain  330047  ahcGain run  Beni     330048  spill sync enabled  Split beamRun enabled  H6B EXPTO1  02  Rates   10 10 6700 and 20 20 7000    Everything looks fine on the monitoring  leaving to run overnight       S 2007 07 03 03 17 Roman Run 330050 stopped on DAQ error   Full Disk   Further  Problems  Dear Experts     The Run330050 stopped on a DAQ Error     In the following the tail of the mess
181. ame     CALICE eLogbook   Results for search of  20070701 Page 6 of 20       The noise looks goog again    mm 2007 07 30 02 25 Roman Noisy layer 9 fixed   The noise in layer 9  central part  has been fixed by replugging the power cables   After the intervention the layer 1 shows a slightly enhanced noise     Roman    2007 07 30 02 26 Roman  Beni 10GeV pion    331145 lbeamData  10x10  10Gev  oi  520 8mm air 2 0 20     Pb 120k       Revived CMBs   beginning of run  I did some test with large VCalib signals before  Seems this can  switch the boards off     2007 07 30 03 00 Roman  Beni       331146 lbeamData  10x10  8GeV  oi  525 8mm air 2 3 20     Pb junk  331147 lbeamData  10x10  8GeV pi  525 8mm air ES 20     Pb junk  R    TRIMO6 adjustment runs  TRIMO6 is  4 5 A now       2007 07 30 03 02 Roman  Beni 8 GeV pion    331148 lbeamData  10x10  8GeV pi  525 8mm air 2 3 20     Pb 156k          10x10 trigger   560 spill  DAQ   35 Hz    The low rate comes from a reduced rate on the main target T4  The user of T10 asked for a reduced  rate and this reduces the rate on T4  too  But the controlroom promised to do check if  they can improve the situation without disturbing the others     We get now 660 triggers spill  DAQ   40 Hz        2007 07 30 04 07 Beni temperature development    Actual temperature development in the detector  The records starts after the CMB mapping was    corrected   overall temperature development    28  27 5  27  26 5    26    A A AA A RAS          25 5   i  24 0
182. an  9x9 points    400531     entire detector scan  9x9 points    400532     entire detector scan  9x9 points    400533     entire detector scan  9x9 points    400535     one strip scan  3x10 points    400536     one strip scan  3x10 points   may be stopped before finished  400537 one strip scan  3x10 points   may be stopped before finished    end of precise scan                                            MIP calibration with 3rd configuration  36 points scanned   400543   00544   00545   00546   00547   00548   00549   00550   00551     stopped at 34th point  data available until 33th point  end of MIP calibration with 3rd config                                        4  4  4  4  4  4  4  4    energy and position scan with 3rd configuration   00650 400651  1GeV 36 9 10 points    00653 400654  2GeV 36 9 10 points    00655 400656  3GeV 36 9 10 points    00657  4GeV 36 9 points    00772  400773  4GeV 36 9 10 points    00765  5GeV 36 points    00774  400775  5GeV 36 9 10 points    00647 400648  6GeV 36 9 10 points    end of energy scan with 3rd config                               4  4  4  4  4  4  4  4    3rd configuration precise scan with tungsten                    00658  9x9 points     00659  9x9 points   00767  9x9 points   00768  9x9 points   00769  9x9 points    9x9 points     00778  00779  00780  00781    5x20 points    4  4   4   4   4  400770  4    4    4 5x20 points   4                   00777  5x20 points                5x20 points  5x20 points       endjot  precise 
183. and it vorks  Problema is  sol   created by spsop on cwo ccc a7lc       Logbook entry   CALICEelog data 2007 29 22 07       22 07 2007 a i    Y MN Hengne Oliver Manqi Beni  17 30  Problems with Beam steering  the Program Cesar NetBeans does not allow anymore changes to  the beam  setup         it always throws exceptions and doesn t change anything    gt  called control room   gt  control room called expert   gt  expert suggest to call other  expert     gt  other expert is not reachable   gt  control room calls other other expert   gt    other other experts suggests to call oracle support   gt  oracle support doesn t answer        other expert called control room     other expert called expert to delete some space    gt  problem solved    REMARK  We do not know on what settings the system got stuck  We checked the scalers in  the computer  room and we have 2k events in 10x10  So we do not worry about the detector        SEE ALSO  entry from 19 01    19 of 28 27 09 2007 11 39    CALICE eLogbook http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log          Logbook entry   CALICEelog data 2007 29 22 07       Ey BEN Beni Geneva temperature development       Logbook entry   CALICEelog data 2007 29 22 07    Ed O Beni temperature development    20 of 28 27 09 2007 11 39    CALICE eLogbook    21 of 28       http   ttfinfo desy de elog results jsp docroot  export web htdocs  zlog        overall temperature development    Logbook entry   CALICEelog data 2007 29 21
184. arch of  20070701 Page 6 of 20                      It SUN d z   E Es X it fc i es    hari mmm d Bai Mr arr eae SR W    BI   E 250  E B         200    150    100     50         SS T 1 rU I T na my  adu  5    1   test  pra   1   1    40 21 00 2320 21340  JF 2007 12 07 2007     12 07 2007     12 07 20    gt  File  htp  ftinto desy de CALICEelog data 2007 28 12 07 2007 07 12T21 26 02 00 Jpg H    2007 07 12 21 34 Marie Cruz  Tao  electron run  Nanda    330377 lbeamData  10x10  15GeV e  536 air 6mm 1 35 10 kev  Pb loeam  gone       600 pps in 10x10 trigger coincidence  DAQ rate  40 Hz    A 2007 07 12 21 49 Marie Cruz  Tao  bugs  Nanda    Bug in online monitoring panel   HCAL pedestals in slot 19 are overwriting the LED panel    Negative pedestalvalues for slot5 FE 7B   module 32    2007 07 12 21 55 Marie Cruz  Tao  electron run  Nanda    PRA Pep Pp PP fe P E UL  Pb       Still bad beam conditions   200pps in 10x10 trigger coincidence  DAQ rate  20Hz    stopped to get an access in order to exchange a TCMT HAB          2007 07 12 22 48 Marie Cruz  Tao  key access    Nanda  Stopped run and asked for a key access to exchange broken TCMT HAB   Had to switch off HCAL  amp  TCMT voltage   ECAL is up  ECAL safety run is ongoing    2007 07 12 23 51 New shift crew    Valeria  Trygve  Hakan    2007 07 13 00 14 Key access finished    Key access finished  Powering up HCAL     CALICE eLogbook   Results for search of  20070701 Page 7 of 20       2007 07 13 00 16 Kurt  Sasha  TCMT ADJUSTMENT  Vis
185. are missing    2007 07 25 03 58 Robert  Riccardo       Run stopped to tune the beam profile  Y at  80     2007 07 25 04 08 Robert  Riccardo  pee mae          2007 07 25 04 32 Robert  Riccardo    Channel 10 again noisy when running at 15GeV e   At lower energies only ch 19    CALICE eLogbook   Results for search of  20070701 Page 11 of 20       was noisy        pa f   f    Ln EE  Er ke Zeien Een men gm e e mees    vult Sl pari ero    Sen  i y ESO IA  RS POCO been E       collin kt    eri aip       Sayar rimar  File  http   ttfinfo desy de CALICEelog data 2007 30 25 07 2007 07 25T04 32 29 00 ps          2007 07 25 04 40 Robert  Riccardo Message from Control Room  SPS long MD    No beam for physics from 25 7 8 00 until 26 7 8 00       2007 07 25 05 22 Robert  Riccardo  330948 fac  2007 07 25 05 23 Robert  Riccardo  Pp AAA AA AA    events    2007 07 25 06 06 Robert  Riccardo HCAL pedestals still shifted    They show no change during the all night    CALICE eLogbook   Results for search of  20070701 Page 12 of 20       Pea Pac oan rel       L  E     File  http   ttfinfo desy de CALICEelog data 2007 30 25 07 2007 07 25T06 06 58 00 ps       2007 07 25 06 57 Robert  Riccardo  AAA AAA A AAA  events  2007 07 25 08 03 Robert  Riccardo As announced  no more beam for today  2007 07 25 14 59 Vasiliy evidence of ECAL noise    comparison of two 10 GeV electron runs from last and this year   excess of hits and therefore energy observed in the ECAL energy sum     CALICE eLogbook   Results for 
186. arget scan with John Maidment  Schaetzel    best rate at target position  4 3mm    2007 05 04 15 00 Riccardo Fabbri Modules 25 32 34 35 in Testbeam       Wi   OG   Hovenent   250 Lus mp       ISEG EHO 8605 Detail No  1          Wi rat  Righ Voltage    all Mon   ati oer            recover    KILL WN     recover all triss f       monitoring Switoh On   Off commots       Weltege  V  Current  bal Set  V  bestial  V     41 10    10 00 Eh Module 254 at     44 40    10 00    73 10    10 00    76 30    10 00  40 30 19 00    10 00       Experts  Maja  LY export peral   Bota to SCIT       10 00       10 00          File  http   ttfinfo desy de CALICEelog data 2007 18 04 05 2007 05 04T 15 00 28 00 png    2007 05 04 15 22 Riccardo Fabbri MIP calibration modules 25 32 34 35    1000 pedestal events   50000 events per position  no LEDs connected   cfg StagePositions dat    stopped while scanning position x  240  y 180 for the second time          2007 05 05 17 21 Beni switched off test beam    Seems  somebody  Sebastian   decided that we have enough data in the last run  DAQ was off and MKR  had switched off our testbeam  after they received a call from one of us     Therefore  I closed the beam shutter and switched off the magnet to save power     2007 05 07 18 51 Beni  Sebastian 1st switch on    CALICE eLogbook   Results for search of  20070501       2007 05 07 18 53       File  http   ttfinfo desy de CALICEelog data 2007 19 07 05 2007 05 07T18 51 07 00 gif    1st switch on    currents of
187. atb ea web cern ch ab div atb ea documentation Target T4 T4Wobbling3 pdf    2007 01 29 14 11 E  Garutti sketches of the tb areas  http   www pp rhul ac uk  calice fab TestBeams   2007 01 29 18 50 E  Garutti  M  Beam Tuning manual for SPS H6  Groll       File  http   ttfinfo desy de CALICEelog data doc CERN docu 2007 01 29T18 50 39 00 ps    
188. beamData  10x10  10GeV e  520 air air 0 76 L3 ES 20K  Events    trigger rate 9 8K  I have load the wrong file  so it is a 10 GeV run with wrong Cherenkov presure  0 76 instead of  2 10 for 10 GeV     2007 07 21 09 04 Morin  Lutz   Kieffer       330759 lbeamData  10x10  10GeV e  520 air air 0 76  3  3 70K  Events    trigger rate 9 8K  I have load the wrong file  so it is a 10 GeV run with wrong Cherenkov presure  0 76 instead of  2 10 for 10 GeV     A 2007 07 21 09 06 Beni high current  The current HV channel 50  layer 25B  raised for 1 1 mA since the 2007 07 19 22 43  From  the current history this raise happen more than 3 hours ago     Nobody noticed it  as the HV panel did not start to blink red     Additinally there is a small rise in channel 58     I checked the noise signal of the module with the big current increase  Except that there are two  dead channels I cannot see any problems  We continue data taking without changes  I will  raise the current limit for HV channel 50  amp  58     EDIT    While editing the file with the current limits I realized that the limit of HV channel 50 was  raised between my ELOG entry from 19th and now  I cannot find any documentation of this   and most probably the person that did this change also forgot to reopen the HV panel   Therfore  the additional raise did not trigger the red light        NW 354 a SSR3A    1 mus  w 39854  w 70358  d    D le MEC 102 E DITA   on 16 98 w 378 8  Sep       120 WoA CS w 35 185 X Era      100  Rm g 55 188
189. calibration run for 2nd  direct fiber  configuration before putting  tungsten     2007 03 16 09 15 daniel temperature data to 16 03 07 and plotting script    I attach temperature data for this range   suspected bad measurements  due to some noise   are commented out with         these are taken from caliceonG8flcscecalsc   temperature ana output summaryTemp txt  I also post the script to plot the data     from caliceon8  flcscecalsc   temperature ana plotAllTempChan C   format of the data is    date  amp  temperature    voltages read on 16 channels  temperatures on the 16 channels    these are the positions of the 16 sensors   tl   AddEntry  graphs  layer3 side       t1  gt AddEnt ry  graphs  layer3 top     tl   AddEntry  graphs  layer  side     t1  gt AddEnt ry  graphs  layer 7 top        layerll side      layerll top      plastic support      steel plate      layerl7 side       t1  gt AddEnt ry  graphs  t1  gt AddEnt ry  graphs  t1  gt AddEnt ry  graphs  t1  gt AddEnt ry  graphs                                       0 JAU04SWNRAO                t1  gt AddEnt ry  graphs                CALICE eLogbook   Results for search of  20070301 Page 10 of 76       tl   AddEntry  graphs  9  tl   AddEntry  graphs  10  t1  gt AddEnt ry  graphs  11       layerl7 top             tl   AddEntry  graphs  12                 layer21 side      layer21 top      layer25 side      layer25 top      air in box       air outside box       t1  gt AddEntry  graphs  13  tl   AddEntry  graphs  14  tl   Add
190. can                                   RSS    MIP calibration after tungsten removal with 3rd configuration                 400784   400785   400786   end of MIP calibration                                    precise scan after tungsten removal with 3rd configuration       400788  3x10 one strip precise scan   400789  3x10 one strip precise scan   400790  3x10 one strip precise scan   400791  3x10 one strip precise scan   400792  3x10 one strip precise scan     CALICE eLogbook   Results for search of  20070301 Page 22 of 76       400793  3x10 one strip precise scan   400794  3x10 one strip precise scan   400796  9x9 precise entire detector scan  400797  9x9 precise entire detector scan                       400798  9x9 precise entire detector scan     9x9 precise entire detector scan                       400799   400800  9x9 precise entire detector scan  400801  9x9 precise entire detector scan  400802  9x9 precise entire detector scan    end    Of precise sedan  222 3334 a SS 24    2007 03 19 18 59 S  Uozumi dead channels in 2nd configuration    Summary of dead channels in 2nd configuration   Layer3 strip9   Layerl4 strip8   Layerl6 stripl   Layerl8 strip10   Layer21 strip17    2007 03 20 00 06 S  Uozumi MIP calibration runs    Run between 400519 400527 are MIP calibration runs     2007 03 20 00 11 S  Uozumi precise scan    400529 MIP run T1L 3GeV direct 2 5V scanned 9x9  fiber precise  scan       2007 03 20 00 11 S  Uozumi precise scan    400529 MIP run TIL 3GeV direct 2 5V
191. can for one strip with 2nd configuration without tungsten  but not  sure all the points are scanned   It may be stopped on the way before it was finished     2007 03 20 18 46 S H Chang et al  Configuration is switched to 3rd Configuration KNU    fiber     We have switched from 2nd Configuration  Direct fiber  to 3rd Configuration  KNU fiber       2007 03 20 19 00 S  Uozumi et al  comissioning run for 3rd configuration    comissioninHl1L 3GeV KNU fiber  2 5V scanned 36   run points  detector  scan for  lcomissionin    2007 03 20 19 00 S  Uozumi et al  comissioning run for 3rd configuration    200538 MIP run TTL 3GeV KNU fiber  2 5V scanned 36  points  detector  scan for  comissioninb    2007 03 20 20 12 S  Uozumi comissioning run for 3rd configuration             comissioning 1L 3GeV KNU fiber  2 5V scanned 36   run points  detector  scan for  comissioninb    HV for KNU module is tuned from  1 2V above Overvoltage 2   Just see the MIP peak position  so QUEE after lst point    2007 03 20 20 12 S  Uozumi comissioning run for 3rd configuration   400539  comissioningiL GeV  NU fiber fp 5v    scanned pe       CALICE eLogbook   Results for search of  20070301 Page 25 of 76    points  detector  scan for  comissioninp    comissioning run for 3rd configuration           KNU fiber  2 5V   15  42  comissioninp  position  run  of strip 1  on X layer  S  Uozumi et al  comissioning run for 3rd configuration  KNU fiber  2 5V   15  42  comissioninp  position  run  of strip 1  on X layer  2007
192. ccc a71lc     vhen the physicist from H6 tried to load a beam file in CESARhe got the folloving error  nessages  Annotation  Exception occurred in Request Processor  org  springframevork jdbc UncategorizedSQLException  Hibernate operation  Could not execute JDBC  batch update  uncategorized SOLException for SOL derer into LOGBOOK  NAME  LOGDATE  VHO     WHY  WHAT  LOGTIME  ID  values                        SOL state  72000   error   code  1536   ORA 01536  space quota exceeded for tablespace  DATAO1     nested exception is java sql BatchUpdateException  OR   01536  space quota exceeded for  tablespace    DATAO1    created by spsop on cwo ccc 871c    ve call Vito and he suggests to call mister Roman Gorbonosov  Roman instructed us to reboot cs ccr spseal  This didn t solve the problen  created by spsop on cwo ccc a7lc  After ve told Roman that in the error messsgge there vas a line    space quota exceeded for    tablespace  DATAO1   he suggested to call Chris roderick to increase the quota so that ve can  vork until tomorrow    created by spsop on cwo ccc a71c  Ve cannot reach Chris and call R  Billen He suggested to call Oracle support  created by spsop on cwo ccc a7lc    Oracle support is not ansvering   created by spsop on cwo ccc 871c  Chris called and he vill call Roman to know vhat he can erase in oprder to liberate sone space  because he is unable to increase the quosta   created by spsop on cwo ccc a7lc  Chris called again to let us try again to send a file vith cesar 
193. cessary    All the macros used to make these plots are found at   flcscecalana  home satoru analysis    Note  this plot also have wrong z axis scale     CALICE eLogbook   Results for search of  20070301 Page 67 of 76       CO  o  o    600    Measured energy  MIPs        File  http   ttfinfo desy de CALICEelog data 2007 12 25 03 2007 03 25T23 30 04 00 jpg    2007 03 26 00 51 S  Uozumi    If the bump is really come from non uniformity  I think we should not plot X and Y layers in a same  Shower profile plot     Then I split the shower profile plot into two  namely for only X layers and Y layers    This is a plot only for x layers    In this plot I apply  inter module  calibration factor to the response of direct readout  module    The inter module factor is set to be 0 8  this number is just my guess to have smooth shower  curve       CALICE eLogbook   Results for search of  20070301 Page 68 of 76       X layers    O   e    I  e    NO  e       Measured energy  MIPs     Se    10 20  Layer ID    File  http   ttfinfo desy de CALICEelog data 2007 12 25 03 2007 03 26T00 51 53 00 jpg    2007 03 26 01 00 S  Uozumi  Same plot for Y layers  with inter module factor of 0 8   So  in fact  the bump is not a bump        It is difficult to say this curve is smooth or not smooth   I think we have to overlay expected shower curve from simulation or simple calculation     CALICE eLogbook   Results for search of  20070301 Page 69 of 76       Y layers    O   e    Measured energy  MIPs   NO A     
194. changed from 5A to  2A   TRIMO6 is changed from  2 5A to  1 5A   COLL01 05 06 10 11       8mm  increased from 5mm since rates are less than lkpps   new trigger rate  4 9 kpps     amp  2007 07 24 02 56 Roman cogri scripts restarted   gt  copy to lyon    Dear Colleagues    I have restarted the cogri scripts  The output is sent to lyon since   at desy there is always the risk to stumble over two or three very   Slow dcache doors which even caused a several hours delay this afternoon     Copying has resumed at 330871 and once more at 330882 after the switch  from data01     data00  see elog entry by Beni      Control of cogri scripts handed back to shiftcrew   Cheers     Roman    A 2007 07 24 03 10 Allister  Robert  noisy ECAL layer 17    CALICE eLogbook   Results for search of  20070701 Page 2 of 20       Riccardo    In run 330886 we see a noisy layer 17  aside from the previously known noisy layers 10 and 19        LJ     mm D sl     TE d   L  n          sU AEREE    CH E       I     de ANM  dere geen AA           don   UT n    E eg a  es E     m       J         La  gt       LE   x E   E slo   DI E eeu       MI Lier     a art                         m m das om oom m om     moo som coma moo moo m    T  Hn E ra ras  File  http   ttfinfo desy de CALICEelog data 2007 30 24 07 2007 07 24T03 10 42 01 ps   2007 07 24 03 50 Allister  Robert  electron run    Riccardo    330887 beamData  10x10  18GeV 522 i i 0 80  6 0 beam  position  tuning   O    eamData  10x10  18GeV i i 0 80 beam  cent
195. cise detector scan  MIP run direct 2 5V scanned  fiber scan of  one strip   3x10points  2007 03 16 14 52 1 strip precise scan    MIP run direct 2 5V scanned  fiber scan of  one strip   3x10points    2007 03 16 14 52    MIP run          1 strip precise scan    scanned  scan of  one strip   3x10points       2007 03 16 16 07 1 strip precise scan  MIP run direct 2 5V scanned    fiber scan of  one strip   3x10points    2007 03 16 16 07 1 strip precise scan    MIP run direct 2 5V scanned    fiber scan of  one strip   3x10points       2007 03 17 09 07 S  Uozumi    Runs 400478 400483 are one strip scan   2007 03 17 09 09 S  Uozumi precise detector scan    400484 MIP run T1L 3GeV direct 2 5V scanned precise  fiber scan  9x9    2007 03 17 09 09 S  Uozumi precise detector scan    400484 MIP run T1L 3GeV direct 2  5V scanned precise  fiber scan  9x9          2007 03 17 09 09 S  Uozumi    Runs 400484 400488 are over night runs for precise detector scan with 2nd configuration     2007 03 17 09 10 S  Uozumi 1 strip precise scan    400491 MIP run TIL 3GeV direct 2 scanned precise  fiber scan of  one strip   3x10points       2007 03 17 09 10 S  Uozumi 1 strip precise scan    400491 MIP run TLE 3GeV direct 2 0 scanned precise  fiber scan of  one strip   3x10points       CALICE eLogbook   Results for search of  20070301 Page 13 of 76       2007 03 17 10 23 S  Uozumi 1 GeV run    400492  beam bg 1Gev direct 2 5V scanned 36  fiber 9 10  stage scan       2007 03 17 10 23 S  Uozumi 1 GeV run  40049
196. cked the files  which are already   converted  The last one converted is Run330427 xxx slcio   The latest binary file transfered   to the GRID is Run330429 xxx bin   We sent a mail to Roman to check the conversion to   slcio           2007 07 14 08 16 Mandi  Allister  new shift  Beni  A 2007 07 14 08 22 Mangi  Beni TRIMO6    Cannot steer TRIMO6  It tells standby   For the current run  330431  we leave it as it is  For the next run we will try to reactivate it     EDIT  for 20 GeV  file 533  TRIMO6 works again          2007 07 14 08 56 Beni  Allister  Beam Run Information 330432  2007 07 14 09 56 Ce Beni  Beam Run  di cd R   E E E E PSP P  2007 07 14 11 09 Allister  Beni  Beam Run  Mangi       trigger rate 7k    2007 07 14 11 12 Allister  Beni  option at Run330434  swtich to pion run  Mangi  Collimator 1 closed to 4 5cm to lower the trigger Rate  2007 07 14 11 56 Allister  Beni  Beam Run  Mangt       trigger rate 9     2007 07 14 12 44 Allister  Beni  Beam Run  Mangi    Pb       trigger rate 5 5 k particle per spill    2007 07 14 12 49 Allister  Beni  Guassian Fit for electron run 330420 330433  Mangi    Performe Gaussian fit to the main energy peak in ECAL     Beam Energy  GeV               E  MIP  Sigma  MIP  Sigma E    EE A E88 a  EE  15  B96  897  B99  768    709  ES         EET ET  5         EEEE 526    917    X     en  ox    En  Lo  oe    CO  oe    SEIN IEN IN ES IN  of   i nj   E  ae  al aef el aef del dei ae          P mw opepsroreo e        E    n  g  SES TA  GW  oO a
197. current increases beyond 500muA     CALICE eLogbook   Results for search of  20070701 Page 13 of 25       2007 07 10 23 25 Mary Cruz Hengne Niels    By performed a gaussian fit on the peak of the hHcalEnergy histo we obtain     Ru             Bnergy Particle   Mean zit value T                             330325 ZE ER ECHT PS  330326 ELOGIA    330327 O                           enr E                 330378 ET5Gev                 ei                         DIT mp E                330330 D  er                     dei me IO       2007 07 10 23 38 Mary Cruz Hengne Niels    Beam gone  so we take a gain run       E 2007 07 10 23 44 Beni AHC cfg issues   I corrected some CMB mapping errors in AHC cfg    AHC CERN07 beamhold 10x10 cfg  AHC CERNO7 beamhold 20x20 cfg and AHC CERN0O7 beamhold 100x100 cfg  were checked  the CMB mapping should be right in all three now     There are still wrong hold values in some of the files   new AHC CERNO07 beamhold 10x10 cfg below  REMARK  Due to the full disk  no AHC cfg changes were possible  During changes the original    AHC CERNO7 beamhold 10x10 cfg  was lost for the same reason but could be recovered from a backup file     CALICE eLogbook   Results for search of  20070701 Page 14 of 25       File  http   ttfinfo desy de CALICEelog data 2007 28 10 07 2007 07 10T23 44 35 00 cfg    2007 07 11 00 06 Kurt Wenbiao TryBeam Data    330332 lbeamData  10x10  10Gev  oi  520 4mm air 2 6064 200k  Pb Events       2007 07 11 00 47 Kurt Wenbiao Tryd  eam data  p
198. d back as values for HAB 0        HAB 6 is read back as HAB 5  Calling this  scrambled numbers  in the table below     A different problem with FE6  values are read back seemingly for the correct HAB but the DAC values  have changed  Example HAB 9  written 40  read back  32     DAQ feature request  automatic comparison of written and read values and a clear error message when  the two are different                                                                                Slot number   3  CRC Component   6   FE6  Label   1   0x01   0b00000001   write  AhcVfeStartUpData   print     Verification data   2863289685   0xaaaa5555   0b10101010 10101010 01010101 0101010  HAB 0 DACs   Se 139 238 143 1389 129 129 1221339 117 139 1129 128 133  132  1339 122  128  HAB 1 DACs   22 l22 1285 129 129 133 192 1259 132 129  40  15 122 133 129 133 122 129  HAB 2 DACs   38  115 7 122  133 CT LC X07 X17 X117 112 107 X07 107 2  A220  9T  5l  HAB 3 DACs   07 117 2 107 107 117 107 412 112 212 46  66 92 25 107 107 412 114  HAB 4 DACs   339 148 153 122 117 163 30  81  Bt 76 122 133 112 11v 122 122 153 117  HAB 5 DACs   12168  56 97 SI  38 138 1423 168 153 163 128 117 117 3423 82   76 87  HAB 6 DACs   232 148 128  10 70 128 122  128 117112012261 Vl  OP 20  133  133  133  143  HAB 7 DACs   33  128 T 143 117 122 Ae 82   71 61 1389 128 122 112 107 112 112  107  HAB 8 DACs   12 112  92  97   81 107 102 107 307 112 07 107 112  112 2 66 46 51  HAB 9 DACs   12 112 107 117 117 112 107 107 112 107 46 97 76 
199. d plot shows it has been increasing steadily lately     Channel 73 is also behaving strangly  but without crossing thresholds               2  2  c cr m  i A    ub    200    150      100       ny O OR A qn   Dt T na  H  01120 01 50 02320 02150  1 07  2007 11 07  2007 11 07 2007 11 07 2007    File  http   ttfinfo desy de CALICEelog data 2007 28 11 07 2007 07 11T03 12 43 00 jpg                2007 07 11 04 41 Roman More files deleted from data01  I have removed files of runs 220000 220238 from data01 data debmJan06 run    Looks like these are files of runs taken while testing the hcal modules   All the files are present in the desy dcache     This freed   170 GByte   Remaining Free Space  590 GByte     2007 07 11 05 45 Kurt TCMT update  New DAC files  DAC tcmt1 11072007 dat and DAC tcmt2 11072007 dat  uploaded to calice00 and  links  DAC tcmtl dat and DAC tcmt2 dat  updated  DAC settings were increased by  0 1mV for  approximately  30 SiPMs     Values should take effect on next DAQ shutdown and startup     2007 07 11 07 30 Trygve HCAL Channels 73 and 74    Channel 73  green  seems now to be stable  Channel 74 is still going up and down  but the current  has been below threshold most of the time     CALICE eLogbook   Results for search of  20070701 Page 2 of 20             0223 ch E      Sai                                            A       A EL 5 mau Ee fleegen  WRC Le PARA AA    02300 04300 Ui z     19700  11 07  2007 11 07 2007 11 07  2007 11 07  200  e  mpmtfinto desy de CA
200. d the reference    20  A  This happened already in the past  Unfortunately the control room is still to busy  to fix the problems caused by the compensator fault from this morning  We will call them  again in 30 min     2007 08 19 12 55 Jamie  Pavel  e   Beni    350260 junk  as for 350261 but with the SPS and magnets  all over the shop       350261 beamData  10x10 H50GeV e  624 ade 3mm 0 04 110k   pa prerr PE pp De  D  D  Ke  350262 lbeamData  10x10  40GeV e  620 air 3mm 0 1 20deg   peg cc O a y eeu   350263 beamData  10x10 H30GeV e  621 air 3mm 0 3 100k   pM DES PP e IL esee   2007 08 19 13 29 Jamie  Pavel  CESAR problem   Beni    The beam steering GUI froze  We killed and restarted     2007 08 19 14 15 Jamie  Pavel  e   Beni       m 2007 08 19 14 15 Jamie  Pavel  GRID problems   Beni  There are general grid problems  All grid based processes are stopped  We called Roman  and he will  monitor the problem and annoy Experts if necessary     Addendum RP  22 09 CEST   The error happened due to today s expiration of the  lfc host certificate  The renewed certificate did not propagate fully through the  system  DESY it reacted on my mail during sunday afternoon        2007 08 19 15 50 Jamie  Pavel   Beni    350266 beamData  10x10  20GeV e  unknown air 6mm broken 21k    Pb Cherenkov  unknown   beam off  after  some time       CALICE eLogbook   Results for search of  20070801 Page 6 of 22       The Cherenkov could not be set  Called control room  They have a look into the problem   
201. dvealibsean             L TTS  BD30536          ahePmNoise           Fe 50000          MEN  530537          esa SSCS  SSO YJ  While waiting for the beam  more calibration runs        2007 07 16 14 25 Allister  Imad muon run    Split beamRun in 250kEvts  but beam stopped at 219Kevts    2007 07 16 14 30 Allister  Imad new beam files    Edda Gschwendtner created new beam files     H6B 515   35GeV beam  but she said the quality is not so good     H6B 518   45GeV beam    2007 07 16 15 14 Allister  Imad calibration runs  30539 ahcGain    2007 07 16 15 41 Allister  Imad muon run    Split beamRun in 250kEvts    2007 07 16 16 28 Manqi  Wanbiao  muon run  Nanda  D30541 Joeambata  io0x100  806ev   u             55 p bir      0 09 EUR 200kEvts    2007 07 16 17 12 Manqi  Wanbiao  remove muon trigger  Nanda  removed the first 1xlm trigger  in front of ECAL and HCAL  from the beamline  2007 07 16 17 31 Beni switched DAQ to 10x10 trigger  DAQ setup for 10x10 trigger    changed AHC cfg link to AHC CERNO07 beamhold 10x10 cfg    changed sequenceDelay in calice00 EmcConfiguration hh to 13    build runner on calice00    DAQ shutDown and startUp  2007 07 16 17 34 Mangi  Wanbiao  gain run  Nanda  D30453 regar YT   2007 07 16 18 02 Mango  Wanbiao  testing new beam file  Nanda       Pb    CALICE eLogbook   Results for search of  20070701 Page 13 of 19    lo  4  4    330550 Ee Ee pesce j 250 kevt    testing Eddas new beam files  see double peak in DC  two types of particles    330545   double structur
202. e 17 of 24       Run 330192 0 Event 1040 COAL Wis  27 Energy  43 7912 mio    E HCAL Hits  519 Energy  3712 12 mips  Time  18 13 58 988 982 Sat Jul 72007   rca Hits  48 Energy  121 267 mips    E   4  Im H f   H Sams Fe  44  HW om E  MES itl  d    i   gm  moa   Hat a WT    4 ie E  ei   li   ix mes     4 D y ine      nee               IEN    L    PE               File  http   ttfinfo desy de CALICEelog data 2007 27 07 07 2007 07 07T21 26 37 00 eps    2007 07 07 21 42 Francois  Sacha   Jean Yves       Pb    run stopped to test the changes in ECAL     2007 07 07 21 49 Francois  Sacha   Jean Yves   Marcel  Roman    Se  Pb       problem with layers 15 and 17 of the ECAL resolved by echanging the corresponding power cables     CALICE eLogbook   Results for search of  20070701 Page 18 of 24       Eesha    8  Q        es Er          S       feru                                umm  w   i Gen       AS    mm    H 1  a  LS              i    fs  E Fa   E       mu ME  DUM Ip qeu          ess  Ed  am  Lj  P  mm  hEcalHitsPerLayer       Pm  n I Dd D  I  a n a acm        d Le   a Ae Ser DE   a Ps E    D r m we E  i   mut    8    a adt    thes rem   gu ET  ww  uo X H Oe  law ni oe  hEcalEnengyPerLayer     A  t  E     T EZ  V    i  o f   d     ha  Em   BE    mm     t      Matem E  E ma   p  mum     t   A TEs  EE   E    I r 5  me    sn   Sg SS  Li 7 T  M     i ao  ee  r E zi     Pa mim  File  http   ttfinfo desy de CALICEelog data 2007 27 07 07 2007 07 07T21  49 49 01 eps  2007 07 07 21 59 Francoi
203. e 6GeV beam file did not work anymore     We had order of 20 particles per spill almost independent of the collimator settings  A look into  the wire chambers showed that the beam got lost somewhere in the vertical plane  After  optimizing the rate in our experimental counter by changing trim01 and trim02 we now  get 5 7k events per spill     We checked the energy measured by the detectors and it is within the fit errors the same as last  night  Therfore  we will run for the rest of the night with the changed trim settings     changed settings for this run compared to beam file     TRIMO01   1A  TRIMO2    6A  COLLO1   COLLO5   COLL10     15    After this optimization  the beam is already centerd in x and y without using TRIMO5 and TRIMO6     2007 07 30 08 00 Erika    23    stopped run 331153 with 220kev    TCMT channels 3 4 and few HCAL channels have lost DAQ settings  reloaded them   the morning rate  6GeV is still 5500 pps             2007 07 30 08 12 Erika  Ariane       short    just checking that all was fine with the reloaded DAC settings    2007 07 30 08 20 caliceon 2calicehediarfi  easurrent thresholds   CALICE user     CALICE eLogbook   Results for search of  20070701 Page 9 of 20       change thresholds for 3 channels corresponding to modules  15A  18A  31A    94  RIES 101 W SA  wa E esa H    187 mosa          8  Ep 150 W MIN   14    RSR   i  25 w 30338    i21 aA as 114 W IRSA    wr SIMA     i01 W  a we i44 W a28 276 68jMe    106 jw 30 96 169 W aj278   o 70138  4    
204. e Martin Cattuatibration runs  2007 07 28 13 33 Ariane Martin Catkhkmlterun       331052 lbeamData  10x10  10Gev pi  520 8mm Air 2 0 10deg  Pb junk    run cancelled  beam off in YX  adjusted both trim 6 and trim 5  new settings saved to beam ref      2007 07 28 13 44 Ariane  Martin Catlpirid8 GeV    Pb          DAQ rate  60 Hz  1400 pps    2007 07 28 14 18 Ariane Martin Catljarikeagain       AAA qe ape es AA e  2007 07 28 14 28 Ariane  Martin Catlpirid2 GeV    Bb       DAQ rate  60 Hz  1200 pps    unstable beam reported  data seems fine though  100k events    2007 07 28 15 05 Ariane Martin Cathpirine GeV    Pb       DAQ rate  60 Hz  1300 pps    CALICE eLogbook   Results for search of  20070701 Page 8 of 16       beam still unstable    2007 07 28 15 11 increase current thresholds  found few channels on too high current  all about 10uA above threshold  maybe linked to higher  temperature     Please close HV pannel every 8 hours and open it again to ake sure it refreshes the threshold  check  Modified thresolds for modules  15 A B  16B 19A  29A    y ima       p      101 MONTA le 186 WOA  emm ien wp    zs 1098    120   310A   ae we 43 2288 200 m 60348    ms   w IMIA   199  Re  seo w 399A    me y 38036 i70 Mag  01 waosa    RS     sf ela  oo  ofl sls alu   gt  all Ole    368 Id eu e 214 w 74 978    N MA 2 Sa w TSROA    W snap   79 w 01308 1 W HER    EOE 77 x5 TA 36  Tas  f 103 wonse             Bd    w 338       167        X X 4 X 4 X W X X aa X  X a          LY Ww OFF GLY LV OFF  
205. e for module 1 was taken instead and redrilled correctly   Therefore a new plate has to be prepared and we have a delay     Unfortunately the drilling machine brake down  Spare has to be ordered and we do not know when it  will be delivered     Nevertheless we try to finish assembling  and testing  Friday 22 06 07     cassettes work in progress  1  2  cassettes waiting for spare SiPMs  24  36  cassettes ready for transport  Soc  status of production  Ze  cassette annem b  ng HAL   FLC              module assembling steps staff  4 5 t 1 D 9  fibres           3  SPMs FRA plate cabin tempsensors cover plate no of       base frame ties pets       noord oko NNNNE    Mor  ke de vi D Daid F1  06 06 07    File  http   ttfinfo desy de CALICEelog data 2007 23 06 06 2007 06 06T 17 51  23 00 pdf    2007 06 11 08 00 Erika 2007 installation  START OF 2007 INSTALLATION    2007 06 14 10 38 Doerte David assembling status  Modules   1 and 2    Good news  The broken drilling machine was repaired already on Monday though production is going  on continuously     cassettes work in progress  l2  cassettes waiting for spare SiPMs  24  36  cassettes at CERN  354438    CALICE eLogbook   Results for search of  20070601 Page 2 of 19    status of production  Mow  cassette osse bing HAL  FLC       roll te vt O Dod  F1  14 06 07    File  http   ttfinfo desy de CALICEelog data 2007 24 14 06 2007 06 14T10 38 59 00 paf    2007 06 14 16 41 Doerte David slow control PS 40 modules  power supply    LV and HV for 
206. e last two days   1  Regulator tests  The regulator tests could not be performed since the  emergency  power supply  does not deliver 7 V which would be needed to supply the regulators and the ASIC  chips  This test had to be postponed until a decent power supply is available  already  ordered by the LLR colleagues    2  Dead    layers  During the last two days we observed two dead layers in the online histogramming   It looks as if layer 18 central and layer 28 bottom are not delivering signals   Still no reason to panic  It might be that these layers are not fully supplied after    the intervention of Christophe de la Taille and me from yesterday     Since Marcel and me will install now bottom layers tomorrow evening we can use this  opportunity to have a closer look at the problem     Cheers     Roman          2007 07 03 17 51 Roman cogri scripts    Dear Colleagues     after the restart of the caliceserv this morning  the cogri scripts needed to be  restarted     Since then everything works fine    This is what the shift crew should do in that case   1  Restart of cogri scripts   The restart of the cogri script should  from  now on be done by     gt  start_cp sh  anywhere on caliceserv     Remark     Do this only after you ve seen that none of the scripts is running   gt  ps aux   grep cp     If in doubt call expert    2  Setting of network parameters  After a restart of the caliceserv machine the network parameters have to be set     a  Try to find a expert or root user o
207. e tails on the low energy side  are prominent  Thus the sigmas are not very reliable              2007 07 13 07 44 Hakan Valeria Trypien run  50 GeV  pepe ese xcu me qe A qe   amp  2007 07 13 08 00 Mandi  Erika  shift start  Beni  2007 07 13 08 57 Mandi  Beni 80 GeV pion  Pb  2007 07 13 09 33 Mandi  Beni 100 GeV pion  Pb  2007 07 13 10 41 caliceon 2calicehawseqfs  iipmh   CALICE user   Moved the stage position 6cm to Jura   60 0 in the input mask  The read position shows a minus     It has to be crosschecked if this is intended     This means that looking in the direction of the beam the detectors moved to the left  Therefore   the beam impact point looking on the front face of the decors moved to the right     How this translates to your favorite coordinate system you have to figure out yourself   HCAL Motor control panel Y        actual position target position    500   x  9 b ox       e fs ar   500    O lost X reference point  O lost Y reference point    O lost 2 reference point    O stage in motion  o manual operation get Pos   jns   fast   pulsezfree   extern STOP   re  fz8   839 906 073  905    OFF set     D  900 E  290       File  http   ttfinfo desy de CALICEelog data 2007 28 13 07 2007 07 13T 10 41  02 00 jpg       S 2007 07 13 11 32 Beni beam settings readout    For runs 330393 and 330394 we tried to get the beam conditions into the readout  Some values were  read back but then the runner crashed with an assertion     Paul is informed     2007 07 13 11 45 Mangi  Ben
208. e to the coarse structure without the 3x3 tiles the offset of the whole puzzle is smaller than  before  That is good but now the produced wedges are undersized  New wedges were made     The colleagues reprogrammed the measuring and drilling machine   Tests for ready cassettes were started   cassettes work in progress  32   31    cassettes ready  24   30  l  23    CALICE eLogbook   Results for search of  20070301 Page 5 of 76       status of production  11cw07 cassette assem bling HAL  FLC       module assembling steps statt  1 3 4 s D 7 8 9  8     baseframe _ tes apen   PRA plate       pa  s      cabing   fibres  tempsensors cover plate no of persons          sousi koat  dn vt D  Daid F1  14 0307    File  http   ttfinfo desy de CALICEelog data 2007 11 14 03 2007 03 14T14 25 46 00 paf                      2007 03 14 16 57 S  Uozumi LED run  p SS GEN GE ae  run for  inter calibr  tion  2007 03 14 16 57 S  Uozumi LED run  400440 CmLed LED fiber direct  2 5V CmLed  run for  A 13444 88a  2007 03 14 17 01 S  Uozumi LED run  400441 PmLed LED fiber direct  2 5V PmLed  run for  pq p E AO ls  2007 03 14 17 01 S  Uozumi LED run  MEE a dd MEE QNNM      run for  inter calibr  tion  2007 03 14 17 02 S  Uozumi LED run   00142        wmWoise Pedestal      JEiber directp 5v          1 T     2007 03 14 17 02 S  Uozumi LED run   00142        mWoise Pedestal   fiber directp 5v         I T     2007 03 14 17 02 S  Uozumi LED run   00143      pmNoise Pedestal                     foe directp 5v     
209. e trigger and DWC cabeling was done at CERN     NIM CRATE setup  24 102006    usen usan  are    ho o   gt  0  D o  o o  e o  e o  o o    o   o          E  e  e   rm  o  o  o  o        9o0o0 oo0             DEE  d  MM am ten m mhi ene are rnt    d ELI  NENNEN 4 ne Sup stase         mmm n vu  musem ze  ftm  4 3  aries  H WoW ata  H NOH Een    mior e o a  D rr s edie 71 6 22   cima ped aral ca     a iaaa cali Vio nima waia 3  D arena sl cade E   ite cat MPa reem     D rara sl cde AS    File  http   ttfinfo desy de CALICEelog data doc electronics 2007 02 16T12 10 44 00 pdf    la 2007 02 16 12 40 Marius CRC Trigger backplane cabeling    CALICE eLogbook   Results for search of  20070201 Page 3 of 21    mi 3 SN  PPRERELE  i  L    BEA REGGE  m n ad    M9IA31VH2 40 42v8  zr e    ala       S on Ee ba H  dT  d 035553  SE SI H 2  ERE e ee ee eR  SSE  amp  tk   Bee Ze  EET 3  gs 3   E E       File  http   ttfinfo desy de CALICEelog data doc electronics 2007 02 16T12 40 30 00 pdf     a 2007 02 16 14 24 Marius LVDS NIM converter    Thanks to Matt Warren    Buid dew 2919  u02 SQATTLLWIN       File  http   ttfinfo desy de CALICEelog data doc electronics 2007 02 16T 14 24 28 00 pdf    2007 02 20 16 52 Sebastian Intercalibration runs  Schaetzel    Module 24 in FREI  Module 26 in FE6    CALICE eLogbook   Results for search of  20070201 Page 4 of 21       2007 02 21 16 59 Doerte David assembling status  Assembling is going on   For cassette 30 we have to use a spare and not complete prepared FR4 
210. e vanished  now move beam to centre of DC   gt  change trim06  7 5 to  3 5A  330546   better  but change trim06 further to 1 0A   330547   still not centred   gt  change trim06 to 3 0A   330548   change tim06 to 8 0A   330549   change trim06 to 15 0A   330550   10x10 trigger rate  1 5 kpps   DAQ rate  62 Hz          open collimator 2 3 8 to  8 after 97000 evt   too much  go back to  6 0 after 100 000 evt   now we get 7 kpps in the 10x10 coincidence  and 64 Hz in the DAQ    2007 07 16 19 21 Mango  Wanbiao  testing new beam file  Nanda       Pb    testing Eddas new beam files II  already opened collimator 2 388 to  10  looks good  just leave it running    10x10 trigger coincidence  8 2 kpps   DAQ rate  90 Hz    2007 07 16 20 05 Mangi  Wanbiao  calibration runs  Nanda    B30552          hecmweise            3   B30553             hcPmNoi se                              AAA         B30556          hedain foo                        2007 07 16 20 38 Mangi  Wanbiao  Gaussian Fit for Current Pion Runs  Nanda    By performing a Gaussian fit to the main energy peak in HCAL we extract the following values     Beam Energy  GeV  Mean E  MIP  Sigma  MIP  Sigma E             TO   40  EE  87     CE   200    5304 90 pei pa         p os        50  53031606 pi   pn20      eo    90    5305319                732 fT E   2007 07 16 21 15 Mangqi  Wanbiao  pion run  Nanda       Pb    5G  10x10 trigger coincidence  8 2 kpps  DAQ rate  85 Hz    2007 07 16 21 56 Mango  Wanbiao  Gaussian Fit for Current 
211. eLogbook   Results for search of  20070701 Page 8 of 19    MATO          I  j                      e                                he GLO RUN Pen ad 15200 WR       File  http   ttfinfo desy de CALICEelog data 2007 29 16 07 2007 07 16T00 45 39 03 jpg    3 43 lbeamtune   20k  events    2007 07 16 01 45 Hakan Oliver Valep amp  50 GeV  lbeamtune       50GeV  10k    3305 lbeamData  10x10 pi                       3  4  6    DCs not aligned  Changing TRIMO6 current from  8 3A to  9 7A   330513   Still not aligned  Changing current to  7A    330514   Still some problems  Changing to  5 0A      330515   Problems  Now  3 3A    330516   Aligned now        CALICE eLogbook   Results for search of  20070701 Page 9 of 19       CAL   HEAL   veut  Traiter   Paguwi   Dispin30   Creech   toi   Sume     Toco toci Chanter       gt   E               111114161    ebetisit  KLLLLLZLL        D    Gel      te  1    if  r8 NOB 20S  TUTTI    Irma    silali  ISSESESSER         ir  uia ler sde H  n     H  n    oe m  11114141    MAMAR           File  http   ttfinfo desy de CALICEelog data 2007 29 16 07 2007 07 16T01 45 04 02 jpg    2007 07 16 02 42 Hakan Oliver Valepk 80GeV    330517 beamData  10x10  80GeV pi  510 4mm Air  A  3  3 lbeamtune   Pb 9k spill   rate  330518 beamData  10x10  80GeV pi  510 4mm Air  A T3  43 lbeamtune   Pb 9k spill   rate    Pb    Spill rate was too high  so set collimators 2 and 3 from     2 0mm to     1 5mm  DCs not aligned  changing TRIMO6 from 13 3A to 7 3A   330518   Still n
212. ector scan    400798  9x9 precise entire detector scan    400799  9x9 precise entire detector scan    400800  9x9 precise entire detector scan    400801  9x9 precise entire detector scan    400802  9x9 precise entire detector scan     end ot Precise Scan  eee RRA AA RRA    CALICE eLogbook   Results for search of  20070201 Page 1 of 21       2007 02 02 13 24 E  Garutti H6 beam manual       File  http   ttfinfo desy de CALICEelog data doc CERN docu 2007 02 02T13 24 52 00 paf    2007 02 13 14 42 Doerte David assembling status  In January of this year the new tiles for cassettes 29 to 32 and spare tiles for previous cassettes  arrived at DESY and Sasha started with the tile tests at DESY     Assembling of the cassettes restarted end of January  when first Natasha and then Andrey came to    DESY again   After some complications in the beginning  assembling is now business as usual     But with cassette no  31 we will start with the coarse structure of tile layout  That is why we  have to make some changes in some of the procedures at that time     cassettes work in progress  27   29    cassettes ready  24   26  1   23    CALICE eLogbook   Results for search of  20070201 Page 2 of 21       status of production  ow  cassette osse Peng HAL  FLC       A 0 ere NMENEN    sousi de vi D Daid F1  1000 07  File  http   ttfinfo desy de CALICEelog data 2007 07 13 02 2007 02 13T14 42 06 01 pdf  la 2007 02 16 12 10 Marius NIM crate   cabeling  The pictures are trying to explain how the NIM crat
213. ed base voltage   80V   400631     CmLed to see lpe peak around strip 15     For Layerl9 stripl4                 400633     CmLed for inter calibration   400634     PmLed for inter calibration   400636     CmLed to see lpe peak  can see also on strip 15    400637     Same with 400636 but with  1V increased base voltage   80V   400639     CmLed to see lpe peak around strip 15    Pedestal run   400640     PmNoise with LED off   400641     CmNoise with LED off    end  of LED PUNS  2552522522022 222205424    energy and position scan with 3rd configuration eee niii eet uelis  400650 400651  1GeV 36 9 10 points    400653 400654  2GeV 36 9 10 points    400655 400656  3GeV 36 9 10 points    400657  4GeV 36 9 points    400772  400773  4GeV 36 9 10 points    400765  5GeV 36 points    400774  400775  5GeV 36 9 10 points    400647 400648  6GeV 36 9 10 points    end of  energy scan with 3rd config eieiei     LED runs with  3rd configuration during no beam period                                   HoldScan again to check whether current holdvalue is OK for KNU module or not   400664     CmLedHoldScan  400665     PmLedHoldScan    found that current hold values are also fine for KNU module     For Layerl2 stripl14 15        400666     CmLed for inter calibration  400667     PmLed for inter calibration  400670     CmLed to see lpe peak   For Layerl0 strip14 15    400672     CmLed for inter calibration  400673     PmLed for inter calibration  400675     CmLed to see lpe peak   For Layer8 strip1
214. eference point  O stage in motion    O manus  operation get Pos                Jess 1 Fast V pulsesfree   extern STOP 1 reft   410 215410 001    OFF set     Kl   900 H Le PING  to elog       File  http   ttfinfo desy de CALICEelog data 2007 34 21 08 2007 08 21T21 39 03 00 jpg    2007 08 21 21 39 Daniel  pi  runs  Anne Marie  Petr    350369 lbeamData  10x10  60GeV Git 610 4mm air 0 04 150k 30  Pb degs       350370 beamData  10x10  80GeV Git 611 4mm air 0 04 152k 30  Pb degs    350371       350372 lbeamData  10x10  50GeV Git 612 4mm air 0 04 154k 30  Pb degs    2007 08 21 22 07 Beni changed CM VCALIB    Reduced the CM Vcalib value for module 25    2007 08 21 22 23 raised current limit    raised current limit of HV channels 53  amp  63    CALICE eLogbook   Results for search of  20070801 Page 19 of 22    ap MITA an Ww UA  e SSA     8 W100 145 MANTE 1471 w 36 538    Taas  Kam a2  Sne    ap ow HOB so w SSIES 27 w base  275      60 w 7058  1030 NM  7 ES c    325 WIA    56 WIMA   35 wn    105 Mane iia w 30188    gaan 4X a a    118 M poa    120 w aoa    409 RIA  iss w 74 978  e TIROA    427  42218  a w ORA    w 44228 w 75 558    101 PASA  93 w IRA w Tas    WS    NER ge gia ass oh T         I A        pum  E e He Crate ON   HY Crate  FF  o zen EE        recover    KILL gl AJ H4 jJ w7 ene    Ww  j    Wi settings   File actione enable   disable  LL chaenele m cm    Nna d   aj  i ver t N j LV detai  sw   ue   on enable   diable   recover all trips  31 wl wg J  LV dta    158   315 4579   
215. electron Runs  Nanda    CALICE eLogbook   Results for search of  20070701 Page 14 of 19       By performing a Gaussian fit to the main energy peak in ECAL       2007 07 16 22 10 new position    moved stage 6cm in direction of Saleve  ECAL was moved 6cm in opposite direction to compensate the  stage movement    Stern   cut j    HCAL Motor control panel    actual position target position    po ox        o   SW ari  Oireina je    Q lost Y reference point   Q lost 2 reference point   Q stage in motion   O manual operation get Pos       ims 1 Fast V pulsesfree   extern STOP   ref 24   950 195950  018                         OF set     RES a           File  http   ttfinfo desy de CALICEelog data 2007 29 16 07 2007 07 16T22 10 02 00 jpg    2007 07 16 22 14 Mangqi  Wanbiao  pion run  Nanda    Pb       test new stage position   ECAL centred and HCAL 6cm off towards Saleve    10x10 trigger coincidence  7 3 kpps  DAQ rate  90 Hz    A 2007 07 16 22 30 Beni high trigger rate    The very high trigger rate comes from an increased rate on our target T4  Called the control room   They are optimizing losses and promised to reduce the rate on the target     2007 07 16 23 00 Mangi  Wanbiao  pion run  Nanda    Pb       10x10 trigger coincidence  6 8 kpps  DAQ rate  80 Hz    2007 07 16 23 41 Mangqi  Wanbiao  pion run  Manoa    330560 beamData  10x10  40GeV 4m 200k  Pb events       Very high trigger rate       had to close SE T SEP E to 5  collimator 10 11 to 10    Now 10x10 trigger coincidence  9 
216. ell ZU     m  WIR Re D    e e eu    No  rR  nm  ojo    N  VO          SEN   CESS ety SS  ej 3  afa u n r o O  Q   a  E   0      p  D  e  p  D  Fh  p      0   Bie BI a n  a  e r a  oje N          N     o  Iii         N          w  N  FIFI     m EN ajej w  Of FR  SB  wt Fr N    w  o  N  BD    3 of 28 27 09 2007 11 39    CALICE eLogbook    4 of 28    http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log          Logbook entry   CALICEelog data 2007 30 23 07    23 07 2007  19 25       Nicola  Hengne  Mangi       30874 ESO REA ES Sa pz Eis  ip 53 E  HO    Efficiency of electron detection  99 4     Logbook entry   CALICEelog data 2007 30 23 07    23 07 2007  19 11       Nicola  Hengne  Mangi       EA e GES Es    Trigger rate is 4 49 kpps    Logbook entry   CALICEelog data 2007 30 23 07  23 07 2007  19 08    330870 achCmVcalibScan    330871 achPmVcalibScan      Nicola  Hengne  Mangi    Logbook entry   CALICEelog data 2007 30 23 07  23 07 2007  19 08    330872 beambata testrun junk       Nicola  Hengne  Mangi    Logbook entry   CALICEelog data 2007 30 23 07    23 07 2007  17 56    330868  beamDataltestrun junk       Nicola  Hengne  Mangi       330869 beanbacal10x10  8GeV e  525 air air 2 53  6 0    Collimator 1 5 6 10 7 set at   8mm  Collimator 2 3 8 set at    3mm       electron run    electron run    136k    27 09 2007 11 39    CALICE eLogbook http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log       Trigger rate is 4 53 kpps       The che
217. embe  rs of the Calice Collaboration and others to whom permission has been given     TITLE   Preliminary study of the combined response of the CALICE calorimeter prototypes to pion beams    CALICE eLogbook   Results for search of  20070601 Page 8 of 19       File  http   ttfinfo desy de CALICEelog data Analysis AHCAL Hadronic shower analysis 2007 06 25T18 07 15 00 pdf    4 2007 06 25 19 57 Beni DAQ fibres are routed wrongly  ATTENTION     During ECAL cabeling work we realized that the fibres reading the two VME crates are routed on a  wrong path     If we would do the plate stack shift now this fibres  would get ripped off     2007 06 26 00 50  gt  Electronic Test of Ecal    Today Jean Charles  Akli and R P  have made the electronic installation  of the SiW Ecal        Verification of power supply cables and proper  and creative   arrangement of power supplies     Routing of  new  optical fibre for slow control  the one which was  installed lead to crashes   of the slow control program        Electronic Tests    a  Noise was very sensitive to the actual voltage of the  5 V Power  supply  After proper adjustment    CALICE eLogbook   Results for search of  20070601 Page 9 of 19       the noise fell from 1000 ADC counts to 6 ADC counts  Is this sensitivity a reason for our noise problems which we have from  time to time     All  observable  Channels between 6  90  of the FE  and 14  10  of the  FE  ADC counts    b  Connectors of lower slabes were installed in the wrong order   
218. embled modules no  1 and  2    In the following drawing you can see the areas of positions in modules 3   38 and 1   2     And now  good luck for the run period     position areas of the 5 temperature sensors  module 3  38 module 1  2 cmo no    in HCAL tile geometry fine and coarse  haf    A    and  E     2  T  5  5  i    ros  cable   connectors    Et       s          Cebindex xy x j ee    File  http   ttfinfo desy de CALICEelog data 2007 27 03 07 2007 07 03T14 09 15 00 paf    2007 07 03 14 22 Erika HCAL pedestals   EEE Jheemveise   EEN   EE EC EE  2007 07 03 14 44 Yoshi Francois Eummron run  stage moved in X and Y    230062 fbeambata  20x20   mu combined       Eis    _______    2007 07 03 14 46 Yoshi Beni Erika Haeferenceisettimmfor y position    14 00 14 40 Movable stage is shifted most to  Heaven  position in order to set reference  for y position     2007 07 03 15 07 Beni Yoshi FrancdMdvalpie stage shift  Movable stage is shifted to  x y    10   20   2007 07 03 15 08 Beni Yoshi Francdigldugean    330063 peamHoldScan 10x10 nu Teombimea          poa  _      CALICE eLogbook   Results for search of  20070701 Page 12 of 22       Beam is lost after  40min    200k events are OK     2007 07 03 16 05 Michele Alexandeti 8dMGalib run  330064 ahccmLedvcafibscan Jno beam    si  lt  T     T T        330065        JehcPmLedVCaliibScan      no beam          TT        2007 07 03 16 48 Roman Ecal Status  Dear Colleagues   I would like to summarize observations actions on the Ecal during th
219. emeng rate     i  HCAL Motor control panel  actual position   target position   x m   E ex   v  ss    fs ori  awk Kerg eg  O lost Y reference point  O lost 2 reference point  Ous onis  O ant seen pl        fais V Fast I pulseztres   extern STOP   refs24   043 506 049  2    OFF set    fi    ou H 23 PANEL to deg          File  http   ttfinfo desy de CALICEelog data 2007 29 21 07 2007 07 21T12 22 02 00 jpg    A 2007 07 21 12 33 Beni VETO ampitude readout    LATE ENTRY  On 20 07 2007 I tried to debug the VETO amplitude readout     MAPPING CHANGE  For the debugging the VETO was moved to slot 12 FE 4 chip 0 chan 18    Following tasks were performed     added delay  63ns   new holdscan  amp  holdvalue  6 ticks   increase of PMT high voltage  exchange of polarity transformer    Now the signal looks better  We can distinguish a pedestal from the MIP peak  but still the  separation  is poor  Further work seems to be neccessary     2007 07 21 12 50 Beni  530768                                      shcemLedvcalibSean SSS CCS    330766 ahcPmLedVcalibScan  M     o   330767  ahcGain SSCS id       2007 07 21 12 57 Morin  Beni     CALICE eLogbook   Results for search of  20070701 Page 9 of 19       Kieffer    Ap a A Te PRE  Pb       position checked  looks ox  Cherenkov pressure for 35 GeV optimised to 0 48 bar instead 0 5 bar    2007 07 21 13 03 position check    position in ECAL   3  3             CECAF 4 4 4       MS ieaie rightew     les       nHcaBarycesterSuera XY E    pet er LE        
220. engo  sy ECAL slot 19    Some chip in slot 19 seem to be very noisy early in a run  here run 330774  but then gradually  reduces   to the same level as the other chips  The circled part in the upper graph is actually   much higher in the earlier part of the run     CALICE eLogbook   Results for search of  20070701 Page 11 of 19    2g     i  o     a  5   3  a    PedRMSAverPerChip       File  http   ttfinfo desy de CALICEelog data 2007 29 21 07 2007 07 21T17 08 00 00 png    2007 07 21 17 15 Oliver Allister Henpien run       continuation of previous run to add more events after the beam resumed     A 2007 07 21 17 27 Oliver Allister Hengo  sy ECAL slot 17  Same as the earlier observation for slot 19  Some chips start twice as noisy as others then  gradually    goes down later in the run        2  E   5  E  H  4  Ed  z at  D       tl  n   AE  2     a 10 20 30 40 50 60 70 80  90  chip number  File  http   ttfinfo desy de CALICEelog data 2007 29 21 07 2007 07 21T17 27 28 00 png  2007 07 21 17 32 Oliver Allister Henpien run    APRA Pere p TIT  Pb    10x10 trigger rate  2 71kpps  DAQ rate        2007 07 21 18 28 Oliver Allister Henpien run    Pb    10x10 trigger rate  6kpps  DAQ rate  90 Hz             2007 07 21 18 38 Roman Cogri Scripts restarted   gt  Copy to desy    Dear Colleagues     I have restarted the cogri scripts  The files will be copied to desy again     CALICE eLogbook   Results for search of  20070701 Page 12 of 19       Replication cc lyon   gt  desy for binaries of r
221. ents             2007 07 06 09 09 Erika ECAL hold value  One word of explanation for the ECAL hold issue   Following the concern of Paul we have checked the ECAL hold setting for the runs we are taking with  100x100 m  Thanks to Goetz last year the ECAL hold was left at the right value to start   this year     so the current value used in all the runs with 100x100 is 2 ticks which   is in agreement with what we expect  An hold scan has been taken to cross check this   value  but in principle all the runs we are collecting are good for ECAL calibration     la 2007 07 06 10 39 Erika change wobbling settings    run 330130 stopped with low stat     change of wobbling settings from  120 GeV to  80 GeV    CALICE eLogbook   Results for search of  20070701 Page 5 of 24       no data for 30 min     2007 07 06 12 07 Sven Changes in the eLogbook  1   the button in the top area of the eLogbook is changed to  insert new CALICE run info   with the corresponding functionality in the background  2    COMBINED run  has been added as a new keyword in the eLogbook  3   it s possible again to print directly the SlowControl panels and trends into the eLogbook    2007 07 06 12 25    330132 ahcPmLedVcalibScan  330133 ahcCmLedVcalibScan    2007 07 06 14 23    14 02 No beam  RF Problem in PS  2007 07 06 14 36 Fab Erika received new beam files  Received new beam file from beam expert     ile name  H6B 510    H6B 512  H6B 514       Loaded 80 GeV file  H6B 510   Collimator C8 changed to select muons   C
222. ents rising    SiPM currents in Ch 6 and Ch 58 rising slowly  Ch 6  100uA  Ch 58  201uA    CALICE eLogbook   Results for search of  20070701 Page 15 of 22       MN chorals    WA   35 WA 103 w SA 300 W 325A    w SSA     5 W 18 98 143 w MIB 149  Rep     yr 361330    120 W J104   au w SoA    ae  EG   Re 101 X108 ou 356  SRA 60 WIA aw MISA    evo  A 100 w 218 es w 9996    AA       317 w 331204  aye   310   113 w 4008     115 RS    i 114 W IRGA   364 AA  E 143 W 2B zas m osma     ns    190  KS      ssi 2954  153 W ajere  09 701388  151 M SBA   16 P   A    3a   a  EI f    339 W 41 214 175 w MSA i27 Ww NNA    150 W ojas 201 w 74 978  152 W 900A   w TSIRIA    aza  W   n8  ao Ww ORA    2 NW ND   3751 428 196 W Upon y 70568    ISA   38 mW 453A 1058 W IBIA 4 w T as    EZ    w 46238 37 Mae    79338 100      100 MA 153 W ATE4A RSC 76    se e ag e oo A 0   MEC M ramp UP DOWN Law or mU HY OFF   muer     guo     o A e    s       recover  KILL wy   HAS SS 839    ed j    Wi settings   File actions enable   disable ALL chenels ii E a  Over u W3 we wa Getai  Bi  wa   um   ewble   dimble   recover all trips   i             duae Wi values to file   dap AL values  position             File  http   ttfinfo desy de CALICEelog data 2007 29 19 07 2007 07 19T18 36 02 00 jpg     amp  2007 07 19 18 47 poeschl cogri scripts to be stopped    Hi there     since the desy mass storage had some problems I need to stop and  restart the cogri scripts now     I will keep them under control until further notice   C
223. er of the Hcal  Mangi    Hcal layer correponding to slot 17 is noisy  observed in Run330824 and 330823  Might be layer 14    ron MCA    rcur   manchar  Parkes   cogi   Cowchecia   stia   za   Fee Pose len   iunc    se  vo Les   vno   wt m mn            File  http   ttfinfo desy  de CALICEelog data 2007 29 22  07 2007 07 22T23 28 23 01 png     amp  2007 07 22 23 50 caliceonOcalicehealeo  ftdaroiti81 is increasing   CALICE user     current in ch 31 is increasing  it is  slightly  beyond limit    Switch On   Off  Set  V   Initial  V     win 20 00 2 00  ein    22 07 2007 22 07  2007 22 07 2007 23 07  200    178 74 9 0 2            zrs29900 2       Trend to stoy    SE Or ken wa L      BS 3 41 1  0 2   10 00 W Rang Feeler Ir    pa arene WY settings   file actions     renable   disable ALL chaerels       10 00 M Rue SE za       LET s   twp   EMT     eae   aset      39 B 43 6 90 2             10 00   r Gm Mm i ETA 2 mum      10 00 W S   oz o c LM Crate  EF        10 00   Ree   oF  I Pi B 43 1  0 2  xu we   TR amd rur  MA EXC AOS tri  _    JWE        10 sat      10 00 W ive   oFF Ji  P2 B 22 4  0 2  dap W volus te file   due il values  position                   File  http   ttfinfo desy de CALICEelog data 2007 29 22 07 2007 07 22T23 50 03 00 jpg    2007 07 22 23 53 raised current limits    for following channels the current limit was raised  26  28  31  28 is back with high current     CALICE eLogbook   Results for search of  20070701 Page 19 of 19    LEE Lhe fall B     Ka Foon  A ak
224. er trigger bits are only for offline analysis    2007 06 30 11 46 Erika HV mapping    B                        dH      vec up Ieft                      1869                        1      Ol       B             LLL CES EH  B  DEER EL   5     ME CZF             2007 06 30 14 09 Marius Pedestal first look into Run 350012  The plot shows the distribution of the pedestal RMS      in total 6809 channels contribute to the histogram  the requirement is   mean    10  amp  amp  RMS    20 channels to reject dead channels and electronic noise      MPV is around 33 channels  which is ok      number of channels in the tail   gt 70 ch   is 252  where 109 belongs to module 1  amp  2  and 86 ch  show long discharge  The other 166 ch  have mostly broader pedestals     No signal at all                slot 5 fe 5   module 34  chips 0 3  first half  no signal     slot 5 fe 6   module 33  chips 0 3  first half  no signal   one dead channel in second half  chip  8  chan 15      slot 7 fe 2   module 36  chips 8 11  second half  negative pedestal  quite high      slot 17 fe 3   module 13  chip 6 chan 2  negative pedestal     slot 19 fe 0   module 9  chip 0 5  first half   negative pedestal  quite high    Channels with VFE noise only  SiPM not working     slot 5 fe 0   module 38  3 channels off     slot 5 fe 3   module 37  3 channels off     slot 5 fe 4   module 35  5 channels off     slot 5 fe 5   module 34  3 channels off     slot 5 fe 6   module 33  3 channels off     slot 5 fe 7   module 32  4 cha
225. eration get Pos     iras i slow 1 pulsesfree   extern STOP 1 refsd     270342 270 082       OF see    E  900 E 1290       File  http   ttfinfo desy de CALICEelog data 2007 30 26 07 2007 07 26T11 12 02 00 jpg    2007 07 26 11 17 Erika start data taking  system in correct position  beam centered on the ECAL front fact at the real 0 0  we have been taking data sofar with the beam shifted  1cm higher on the ECAL face to avoid shower  leakage to the un instrumented bottom   Now with almost fully equipped ECAL we go to the geometrical 0 0  we will always refer to this point as 0 0 since this lable is meant for analysis purpouses   The correct stage position can be retreived from the binary   lcio files    2007 07 26 11 26 Erika beamData    330960 beamData  10x10   35GeV 515 4mm good a  Pb run   k ev       2007 07 26 12 06 Erika Fabrizio MarttieamData       330961 beamData  10x10   45GeV  oi  518 4mm Air 0 1 good  Pb run   200k ev    2000 pps  DAQ rate  70 Hz  the Veto counter is OFF  power off     2007 07 26 12 58 Erika Fabrizio MattieamData    330962 beamData  10x10   80GeV  oi  512 4mm nae  Pb run   k ev       Veto back on    2007 07 26 13 48 Erika Fabrizio Martieam cancelled       330963 lbeamData  10x10  25GeV e  535 Air 6mm 0 46 stopped  Pb 25000  events    DAQ rate  60 Hz  2000 pps    beam stopped after 10 minutes    CORRECTION  by Allister  Yoshi  Daniel on Aug 16  2007   Run 330963 is actually as follows   confirmed  with paper log and hEcalEnergy gaussian fit        330963
226. ered  at DCs   206  kEvts    330888  330889 beamData  10x10  18Gev 522 i i   50  pee ue B pee Eve    TRIMO5  3A   gt   1A   gt   2A   TRIMO6   3A   gt   1A   gt  0A   Trigger rate  2 2 kpps    daq rate  72 Hz       Stopped run 330888 at 206kEvts because of sudden drop in trigger rates  Reason not known   Opened collimators for 330889  COLL01 05 06 10 11      13mm    2007 07 24 05 11 Allister  Robert   Riccardo   330890  beamData Host     20Gev   e         523   eir fJeir          0 688  56 0  250kkEventls  TRIMO6 changed to  1 0A    2007 07 24 05 56 Allister  Robert     CALICE eLogbook   Results for search of  20070701 Page 3 of 20       Riccardo    330891  beambata  t0x10   256ev Je  23  fir  fix    pan  60  PP  Akkvts  TRIMO5 changed to  0 5A    Collimators slightly reduced  COLL01 05 06 10 11      5mm  trigger rate   5   kpps  daq rate   85 Hz          2007 07 24 06 17 Allister  Riccardo  stability of ECAL response  updated    Robert  Plot of ECAL response  obtained by gaussian fit on hEcalEnergy  vs run number  updated with recent  electron runs  The error bars are actually gaussian sigma        Stability of ECAL Response     Fit Mean  mips        200 300 400 500 600 700 800 900  Run Number    File  http   ttfinfo desy de CALICEelog data 2007 30 24 07 2007 07 24T06 17 06 00 png          2007 07 24 06 32 Allister  Robert     Riccardo  HCAL Voltage channels 1 1A  2 31B  33 17A and 1 26A became  red many times during this whole shift     The channel 2 31B is lately blinking 
227. ev           File  http   ttfinfo desy de CALICEelog data 2007 16 20 04 2007 04 20T17 03 18 01 png    CALICE eLogbook   Results for search of  20070401 Page 11 of 14       2007 04 20 17 08 Sebastian normal noise in 27B    Schaetzel    The pedestal RMS is normal   30 ADC ch   also for module 27B  FE2B Channels 540 648 in online  histogram   for which the SC shows a large current of 0 5mA      3 HstCrcNolseCrnieiSioti2Tot EL AN  Em Ces yaw Gest jape Clan    Hun DIE El E As RU e IFC  Man      NT    d i  Pao   PAI ws       File  http   ttfinfo desy de CALICEelog data 2007 16 20 04 2007 04 20T17 08 43 00 png    2007 04 20 17 41 Sebastian MIP calibration modules 25 27  Schaetzel    cfg StagePositions dat  50000 beam events per position   1000 pedestal  no CMB LEDs     runEnd while scanning position 221   210 390   all 216 positions scanned and repeating  from beginning     2007 04 21 20 37 Erika check currents  currents of module 25  27 same as yeaterday   0 09  0 11  0 08  0 47        2007 04 24 17 46 Beni FE 5 6  amp  7 are working    The problem with the PIN readout was not related to the FEs  Most probably the SCSI cable is the  source of the strange signals     In addition the PIN HAB was found to have some missing soldering connections     CALICE eLogbook   Results for search of  20070401 Page 12 of 14       Cable and HAB are replaced now     A 2007 04 24 17 50 Beni high current in module 27 B  amp  bad baseboard module    connection  Sebastian and me tried to find the source
228. f  20070601 Page 15 of 19       File  http   tttinfo desy de CALICEelog data 2007 26 29 06 2007 06 29T 19 19 42 00 gif    2007 06 29 19 20 G Mavromanolakis    Run 330008 0 Event 1020 Pe Msn tr n    HCAL Hits  54 E 54 7943 mi  Time  06 45 56 071 972 Fri Jun 29 2007 id ZE  Tomit Event info       a ig y sinu Ze  K airi ft  SE VUE BEER    EES  WA    T m  eq   1 E   E sj Y       e  K                            File  http   ttfinfo desy de CALICEelog data 2007 26 29 06 2007 06 29T19 20 44 00 gif  2007 06 29 19 22 Portal online  Dear Colleagues   the portal enabling e g  remote observation and the piloting of the  ongoing testbeam at cern  is available now   This tool has been developed by Sven Kartensen from DESY   You re invited to have a look at    https   calice portal cern ch    You need to know the password of the caliceon account on the control  room computers     Thanks to CERN IT to open the gate in the cern firewall     Cheers   Roman  2007 06 29 20 00 Beni gui debug runs   330011 WE  330002 WE  2007 06 30 10 58 Beni debug and coherent noise  2007 06 30 11 33 Erika trigger logic scheme    NIM LVDS converter connections to VME back plane       CALICE eLogbook   Results for search of  20070601    VDS_input Backplane_input    Page 16 of 19    po ox trig  E     LE     ein signal        the positions of the 4 veto counters are given w r t  the incoming beam  looking downstream        the two 10x10 triggers are used in coincidence  internally in the DAQ  as main trigger  all oth
229. f  the DAQ would have gone to a slow monitoring run until the operator  Started run 330101 manually  I would guess the run was fine   Paul    2007 07 05 19 45 Francois Jean Yveni amp as naferika Fab    330101 beamData 100x100 mu  Combined HCAL TCMT  voltages 0 0       330102 beamData 100x100 mu  Combined HCAL TCMT  voltages 0 0    CALICE eLogbook   Results for search of  20070701 Page 22 of 22       run 330101 250k events  run 330102 140k events    the trigger s delay of the ECAL may not be perfect   gt  to be checked  possible offset in ECAL mip calibration collected with these runs          2007 07 05 19 49 SE Yven Rasinaferika Fab    Keen    Keen      Keen    pM EI       change hold values for 100x100 trigger according to Marius analysis   see log entry in analysis folder    run 3300103 can have different MIP position w r t  previous runs   due to this change  Please use this run  and next ones  for calibration    run 330104 250k events  run 330105 250k events  run 330106 250k events  run 330107 248k events    the trigger s delay of the ECAL may not be perfect   gt  to be checked  possible offset in ECAL mip calibration collected with these runs      2007 07 05 22 09 PaulD   Looking at the run 330100  I see the ECAL slot 5  was not being read out  even though it was enabled    Fully enabled  which is probably not intentional    I don t know when this error restarted     Because the runs since are being taken without the  histogramming enabled  it is not easy to check if  t
230. f L8s15    CALICE eLogbook   Results for search of  20070301 Page 42 of 76            2007 03 23 18 41 S  Uozumi LED run  400679 CmLed LED KNU fiber 2 5V for  ii B DL O GO Euer  of L8s15  2007 03 23 18 41 S  Uozumi LED run    qe o3 d  AAA AL OA  inter calibr  tion  of L8s15  2007 03 23 18 41 S  Uozumi LED run  RE E  inter calibr  tion  of L8s15  2007 03 23 18 42 S  Uozumi LED run  400682 CmLed LED KNU fiber 2 5V to  peak at  L8s15  2007 03 23 18 42 S  Uozumi LED run  TEIG  see lpe  peak at  L8s15  2007 03 23 18 42 S  Uozumi LED run  400683 CmLed KNU fiber  AE A We eak d  peak at  L8s15  2007 03 23 18 42 S  Uozumi LED run   l L    por Lb OE  See lpe  peak at  L8s15  2007 03 23 18 43 S  Uozumi LED run  o por Lb E  see lpe  peak at  L8s14  2007 03 23 18 43 S  Uozumi  400685 CmLed KNU fiber   TT T TFI 4 eak    peak at  L8s14  2007 03 23 18 44 S  Uozumi LED run         E ETA oa  inter calibr  tion  of L4S15  2007 03 23 18 44 S  Uozumi LED run    400689 CmLed LED KNU fiber GEN for  inter calibr  tion  of L4S15    CALICE eLogbook   Results for search of  20070301 Page 43 of 76                 2007 03 23 18 44 S  Uozumi LED run  400690 PmLed LED KNU fiber 2 5V for  mL PL LOL L Fet  lof L4S15  2007 03 23 18 44 S  Uozumi LED run  KNU fiber S for    ii D Fe  of L4S15  2007 03 23 18 45 S  Uozumi LED run  Pee MEL  see lpe  peak at  L4514 15  2007 03 23 18 45 S  Uozumi LED run  DOTT TIE GE  see lpe  peak at  L4914 15  2007 03 23 18 46 S  Uozumi LED run  p que qq  E  inter calibr  tion  of 
231. finfo desy de CALICEelog data 2007 30 27 07 2007 07 27T13 57 48 00 ps    2007 07 27 14 00 Erika Fabrizio Martir   GeV  change wobblings to  10 GeV    2007 07 27 14 08 Erika Fabrizio Martiax10 pi     II  LL oli odi Dor om I     Pb    DAQ rate  90 Hz  6000 pps    2007 07 27 14 34 Ariane Martin 10x10 pi        Pb    DAQ rate  60 Hz  3600 pps    collimeters opened after 50k events  Cherenkov efficiency below 50     2007 07 27 15 20 Ariane  Martin 10x10 pi        Pb    CALICE eLogbook   Results for search of  20070701 Page 4 of 16       DAQ rate  80 Hz  3700 pps    trying to see the efficiency of the Cherenkov  Cherenkov efficiency turns out worse    2007 07 27 15 24 Ariane Martin 10x10 pi     Pb       DAQ rate  80 Hz  2000 pps    Lowered the pressure under the default 1 85  Cherenkov is working fine  Switching the absorber to 8mm Pb    2007 07 27 16 19 Lukas  Alex 10x10 pi     Pb       changed collimation  C1 C5 C6 C10 C11  from   10 to   12  150k ev     2007 07 27 16 59 Lukas  Alex 10x10 e      C1 C5 C6 C10 and Cll narrowed to  5 mm    200k ev     2007 07 27 17 38 Lukas  Alex 10x10 e      C1 C5 C6 C10 and Cll narrowed to mm     Rate   6  7k ev     Beam X position offset to  X    gt    Compensated by changing TRIM6 from  1 to  5      Aborted after 40k ev        It    2007 07 27 18 05 Lukas  Alex Junk run    B31014   peambata  roxi0    pee Je     pa Riz Pie Pe oo DEE  2007 07 27 18 06 Lukas  Alex 10x10 e        C1 C5 C6 C10 and C11 narrowed to  3 mm     2007 07 27 18 43 Lukas  Ale
232. for DQM  is rebooted because it had stuck   Not yet launched data copy script at this moment     2007 07 03 12 02    330054 beampata   20 20 Ru              EcAr RER  330055 peambata      0x20 Iu             JTEcAz   tow T                1 9        IO    good runs for muon calibration    ECAL still off     2007 07 03 12 02 Erika    ECAL os off   run taken with movable stage moving in x direction  possible higher noise due to the motor     otherwise x scan of the HCAL at y 0        2007 07 03 12 22 Yoshi Beni Erika Haeferencei settimmfor x position    CALICE eLogbook   Results for search of  20070701 Page 11 of 22       Movable stage reference was lost during test of stage moving in this morning  before 330054      therefore the reference setting was searched again by moving stage most to left  most  to Jura  where gives reference definition for x position     2007 07 03 13 04 Yoshi Francois Euglehy Scan   330058 beamHoldScan ox10 fu combined   ________pios s gt    _______      Average event rate  94Hz   81 configuration   300k events    2007 07 03 13 42 Yoshi Francois junk run   230059   oeamData    20x20        combined       r10  3         _____   _____    Just after run start  beam is lost    gt  Junk run   2007 07 03 14 09 Doerte David assembling status    All modules are well done and also installed at CERN   Though the  assembling status  is finished     But one addition about temperature sensors    The positions of the temperature sensors were changed in the last reass
233. for inter calibration   400412 PmLed for inter calibration   400413 CmLedHoldScan for confirmation   400414 PmLedHoldScan for confirmation   400416 CmLed to see lpe peak  100k ev    400417 CmLed to see lpe peak  100k ev  with over voltage  1V 3 5V    400421 CmLed to see lpe peak  100k ev  but slightly higher light to see adjacent strips    end of LED r  ns                                                                                                                                   between 1st and 2nd round  the drift chamber is tuned to improve efficiency     lst configuration  fiber direct  2st round                             CALICE eLogbook   Results for search of  20070301 Page 20 of 76       Each run contains data with 55 points   36 points for center of each strip of X and Y layers   9 points for central part scan   1 points    x 10 times for central injection      1 points have 10k events of beam data and 1k events of pedestal   1 GeV     400422  400423  2 GeV     400425  400426  3 GeV     400427  400428  4 GeV     400429  400438  5 GeV     400432  400437  6 GeV     400434  400435    end of Ist configuration 2nd EOund   2    3329 9322392924222 293242323   LED run after lst config 2nd round to see gain of Layer 6 strip 14                  Pedestal runs   400442     CmNoise    400443     PmNoise    for Layer6 Stripl4     400440     CmLed for inter calibration   400441 PmLed for inter calibration   400445     CmLed to see lpe peak   400446 CmLed to see lpe peak around
234. for search of  20070701 Page 13 of 22    297 Rass  148 F ose    113 W IRBA    143  199 MIRTA  167 W ajere    ne   150 F 3A Ea  ER 145 W ezo   xm  32 e  LETT RA     sis we 148  SSES  ou  s9 W ORA 15    74 ae 196 vna      7 mea 105 W IpiA as     ew  105 w sepie 96 Mapa 178 07098   20 w  149 W Njaa 12 275 Rip EM E K       i   WAppe ze   E ss ps 6 ov    vum   mm                       Hi Crate ON HN Crate  FF     E   a          recover    KILL Hy   AS A  187 1 839     HY settings   file actions     enable   disable iLL chaenela   Y ca A rper         gt     a     i    23  Ej    d  ch  dk  iif    d  IF  351    SEKR        d  ae    1058    ii    BEGG             y            l                File  http   ttfinfo desy de CALICEelog data 2007 29 19 07 2007 07 19T10 12 02 00 jpg    2007 07 19 11 10    330669  beambata   10x10   ece Je          525  air Jair peo     po          530668      beampata  10x10       8Gev  e EE Jar Jair      2 39     po         530665      beampata i0xi0      8G6ev Je         1325     Jair Jair     12 399 po            same configuration  160K event    same configuration   Stopped after 83 kevts because of beam stop  col 1 2 3 5 set to  3 3  trigger rate 4 58  24K events  The second layer is noisy since the night shift  kd e ES    ACA  HEAL   ren   rencias  partes   Dugiaya0   Comchacha   Sen sating    Eemer   peana   vom   wc            e        emper          fu aes 3 Rosende  X Menos A CALCE tu he cacon caer    File  http   ttfinfo desy de CALICEelog data 2
235. ge 8 of 20       checked signal in run 330922   gt  good pedestal MIP separation    2007 07 24 17 43 Beni trigger changed to 20x20  changed AHC cfg link to AHC_CERNO7_beamhold_20x20 cfg  set ECAL sequence delay to 10 ticks    build runner  restarted DAQ       2007 07 24 17 54 Nicola  Hengne   Mangi  at 81000 events the CMB are switched off    does the pedestal change  No  the pedestal doesn t move  at all     change hold file to 20x20 trigger   3 ticks earlier than 10x10  for all detectors    2007 07 24 18 10 Nicola  Hengne   Mangi       bad Vcalib values in the AHC cfg file for 20x20 trigger  many CMB show no light in this run  changed Vcalib settings for the next run    2007 07 24 18 55 Nicola  Hengne   Mangi    2007 07 24 19 30       330932 lpbeamData  10GeV mu  556 air air 3 02 250k  events  2007 07 24 20 14    2007 07 24 20 18    2007 07 24 21 10             330935 lbeamData  10GeV mu  556 air air 3 02 130k  events    2007 07 25 00 00    2007 07 25 00 13 Nicola hengne Anormal in Hcal pedestal  mangi  Another anormal corresponding to slot 5 appeared   the third block become quite negative  Appear  from run 330935     CALICE eLogbook   Results for search of  20070701 Page 9 of 20       ECA PAL   mower   teacher   Partes   Dopiny20   Cross Chen   tits   zeg   Response Pesestes lag Jus    wm Log   oc Leg Log Log Les            Serge  ber age rd    hHcalnergyVs Time               Mes  gt  r e  File  http   ttfinfo desy de CALICEelog data 2007 30 25 07 2007 07 25T00 13 10 00 p
236. gen BESAS  Clg 0  Ara 1550055  Db 529  36709 E  Murder  L    HsiLabelAdc    Entries 2818048                i Mean D  E 1 Rus D     Lu  0 5  D   0 5     1L    i poli   30000 20000 10000 0    10000 20000 30000  ADC counts    File  http   ttfinfo desy de CALICEelog data 2007 19 11 05 2007 05 11T16 08 02 00 ps    2007 05 11 16 16 Sebastian pedestal events   LED events  Schaetzel    this is an offline histogram  ahcBinHst   huge RMS is a result of sampling of pedestal events together with LED events    run type is ahcPmLed  Is that meant to include pedestals     CALICE eLogbook   Results for search of  20070501         5037  3s    Entries  Mean  RMS    Cm  Be f             Slot 5  FEO  Chip 11  Chan 17            SEROO7    File  http   ttfinfo desy de CALICEelog data 2007 19 11 05 2007 05 11T16 16 01 00 ps    2007 05 14 09 21 Sebastian high currents in modules    Schaetzel  LED holes are closed with tape  Module 4 has CMB attached  cover closed     1000 2000 3000 4000 5000 6000 7000 S000 9000    Page 6 of 9    CALICE eLogbook   Results for search of  20070501 Page 7 of 9                          monitoring    Switch On   OFF comments  Vokaoe V Curent  uA  Set bi intial fv        mma       PSEUD  CMB o                       Im 10 01  amp  Ramd OFF U H 1 cus on     paa 1 0 aor qun UD  dl SSES gree T   wor 10 01  amp  Rand OFFI broken a   10 01 0 Baal OFR a    10 01 Bane OFF a   Sg GEES res db  ferm WHK  orri  nr gr   aor SES E d  IN a2  R fame OFF   a   Ke SI i  mw HELP Se   nae T j
237. gt  Can   be solved by   switching the connectors or rejumpering the Crc Board  might be the  easier solution     Correction RP  No rejumpering but switching from left to right connectors  on Crc Board    Therefore no measurement of the noise in these layers     Cheers     Roman          2007 06 26 11 49  gt  No computing in control room   26 6 07 20 00h   22 00    Dear Colleagues     due to the movement of the control room computers from the actual  control room to the cooled cabinet opposite to it       The calice computers won t be available today 26 6 07 between 20 00h  and 22 00h     Please logoff well in advance  The computers will be shutdown on 20 00h  without further notice     I will announce via the elog when the computers will be available again   Cheers   Roman  2007 06 27 04 30 Computers up again  All computers in the control room are up again     Beni and Roman    2007 06 27 21 34 Alexander  Noise on SiPM  Sebastian    There is noise on the SiPM signal line   The attached scope picture shows       purple  Tcalib taken on CMB connector at module    green  SiPM analogue HAB outlet  LEMO     Picture is average over 128 events  randomly shared by scope triggered  on Tcalib and on noise     CALICE eLogbook   Results for search of  20070601       Tek Run    ade dej    2007 06 27 22 34    Page 10 of 19            ds  aes A  64 0mV      1 18V   b A  22 1ps        22 2us     Ch4 Mean   E 33 7mV     i    M  4 00gs A Chi Y 76 0mV   200mV Chal i    mV O 27 Jun 2007   ii 9 
238. gue out of  Module 1  HAB 7  3 different HABs show the same signal    signal does not change in shape when dimming the LED    hold 35 ticks       File  http   ttfinfo desy de CALICEelog data 2007 26 01 07 2007 07 01T22 39 59 00 png    2007 07 01 22 49 S  Schaetzel    same setup but hold 100 ticks  signal goes through funny shape transformation during hold scan    we skip this HAB for the LED tuning    CALICE eLogbook   Results for search of  20070701 Page 6 of 22          TUE    File  http   ttfinfo desy de CALICEelog data 2007 26 01  07 2007 07 01T22 49 11 00 png    2007 07 01 23 14 S  Schaetzel    the nightly calicehcaldaq01 crash happened at 23h    enabled hyperthreading in BIOS for a change    2007 07 01 23 49 S  Schaetzel DAC files are not synchronised    LED tuning was done on calicehcaldaq01  using the DAC value files in   home caliceon cfg ahcal    These are not the files on the raid disk which  are used by calice02     Example  Module 2  tuned yesterday to the file DACvalue mod2 A68 0 B65 0 txt   on calice02  the file DACvalue mod2 A77 0 B79 9 txt is used    For Module 1 the files were also different  I copied the calice02  file to calicehcaldaq01 and used it for the tuning     hcaldaq01 has the raid disks in  etc fstab but they are not mounted      Why   the daq01 files are from June 17    files in  srv calice data00 cfg ahcal from June 28    4  rw r  r   1 caliceon 3145 1799 Jun 28 20 39 DACvalue modl A68 1 B64 7 txt  4  rw r  r   1 caliceon 3145 1768 Jun 28 20 40 
239. h di T  E 4c y J  H  amp c A ai En 4  f ats    i  z ac IT DI a  amp   i   d d      nc 5 E d il  P am 4 A  ap He ia ch Hi 1  1b ie 4 Li  A WC a en d z  H am d a  it fio id F   v   S  12 T5 D TE A B  il id c r m Lr  amp   ni nd  D   a  24 L   C Uhinak BA      3  34 Ha a i F  15  c   im na f  15 Sp D ED a4  T  3 23     Fi EI  H     d no   F     ar 23 C B a uF p  qr GE  Pe        af ma F CH ra 5  3B o VS    1  ap ap e 10 n En H  d ru o F 1  20 IT c za  5A  i B  28 15 He F 3  21 mc   1 uF i  31 S    HD   i  n HE E EI ti T  t    amp  n5 ne F     21 mc iz m di 3  n    d Ho   5  du ae   r id a     H Sp E 7     5 ME ch L     1  ze Sp ri 8 D  28 IR  S   LEA a 4  mi ao HG F E  i me Z   rad  a Fi  aT Bu up   5  J   me A CH a 5  iE ap He F     28 zi KI nam  a D    2007 07 06 01 21    File  http   ttfinfo desy de CALICEelog data 2007 27 06 07 2007 07 06T00 37 20 01 pdf    Alex  Yoshi  AM muon run  HCAL TCMT voltage scan    Page 1 of 24    CALICE eLogbook   Results for search of  20070701 Page 2 of 24    330114 beamData 100x100 mu  combined HCAL amp TCMT  250k events  nominal 0 3V    330113 is junk run       330114  ECAL  layer 0 is back to normal at the beginning of the run   Average event rate in Run is  80Hz     HCAL distribution looks weird with 2 peaks  see below   Should be noise peak  when events have no muon  happen quite often in the display                     T      Rime ach   e E E  I T     a  B    cb     T ETE E E   pue pa mtem  hifcalEnengyPertayer mm dina    E D mm   s    A p
240. hHcalLedPerLayer  Beni  for CMB checking reference  run 330596    hHcalLedPerLayer       5 10 18 20 28 30    t 40  ayer number    File  http   ttfinfo desy de CALICEelog data 2007 29 17 07 2007 07 17T 14 41  22 00 png    2007 07 17 15 20 Allister  Imad  calibration runs  Beni   2007 07 17 15 46 Allister  Imad  calibration runs  Beni   2007 07 17 16 59 Valeria  Mangi  LED check  Nanda    For run 330600 we don t see LED light in layers   1 3 4 5 6 10  14 15 16 17 18 19 20 28 29  33  34 35  36 37  38  just as in the previous runs     CALICE eLogbook   Results for search of  20070701 Page 4 of 22       A 2007 07 17 18 31 caliceon calicehcaisn  atgamych   CALICE user     around 16 45 the current in HV channel 58   module 2b was rising from 60 to 80   HV channel 57   module 2a remained stable               2  ccr ai         200      160       I   I I   I I   I I   I  15700 16 00 17700 15 04   H  200r LF OF 2007 17 07 2007 17 07  2007    File  http   ttfinfo desy de CALICEelog data 2007 29 17 07 2007 07 17T18 31 02 00 jpg       2007 07 17 18 47 Nanda Mangi ValdriBD check   adjusted vcalib settings in AHC cfg  performed DAQ shutDown  DAQ startUp  CMB manual re init      CMB check    since some channels had too much light  and 5 still did not show any signal        still no signal on online moitoring plots  Run330604     gt  seem not to work at all       now below 16 000 ADC counts  before above 20 000 ADC counts    A 2007 07 17 20 02 caliceonOcaliceheisingrcerrentil    CALICE user   
241. harged up for a long time      2007 07 02 17 00 Alex All CMBs tuned now    All CMBs are mounted and tuned now  see photo below        File  http   ttfinfo desy de CALICEelog data 2007 27 02 07 2007 07 03T21 48 52 00 jpg    2007 07 02 17 29 A  M  Magnan resuming muon runs    E A E Sl PA  run       330037 beamData 10x10 mu  combined stage 0 10 good but  run low DAQ rate    strange feature in the DAQ   trigger rate not in coincidence with spill one every other spill     ECAL looks weird  no muons in event display whereas HCAL and TCMT see some  and HCAL noisy     Stopped the run  restart everything from scratch for run 330037   Explanation  beam was stopped from the beginning of the line as experiment upstream was in access       And as ECAL bottom slabs are not in the event display yet so we see only noise       2007 07 02 18 34 A  M Magnan Playing with DAQ rate    run       run    good for HCAL  only bottom row in beam  last 10 layers     DAQ modifs  disable ECAL stage control socket because DAQ couldn t communicate with it   Change trigConfig hh to solve DAQ trigger rate not in coincidence with spill  miss one spill every  other spill     CALICE eLogbook   Results for search of  20070701 Page 8 of 22            ped in spill enabled  ped in spill rate 10 Hz   Now for 1 spill over 5 7  DAQ still at 0 Hz   compatible with config change every 10 000 events   So average trigger rate is 80 Hz now  was 60 with ped in spill rate disabled    330038 stopped to change ped in spill rate 
242. he error is still present  It is not normally a    good idea to disable the histogramming   Paul    2007 07 05 22 43 Francois Jean Yves Sasha    330109 ahcPmNoise 100x100 mu  Combined HCAL TCMT  voltages 0 0    20 000 events stored    2007 07 05 22 44 Francois Jean Yves Sasha  330110 ahcGain 100x100 Imu    Combined HCAL TCMT  voltages 0 0    2007 07 06 00 18 Alex  Yoshi  AM beamHoldScan    330111 lbeamHoldScan   100x100 mu  combined nominal garbage  voltages       330112 lbeamHoldScan 100x100 mu  combined nominal  voltages    330111 was started with wrong DAQ settings     330112 scalers  H6B_EXPT01 1 8E 03 pps  10cmX10cm   H6B EXPTO2 4 0E 03 pps  20cmX20cm   H6B_EXPT03 1 4E 04 pps  H6B_EXPT04 4 2E 01 particles spill    ECAL noisy layers   1   15 and  17   Average event rate in Run is  90Hz     Run330112 stopped at 81 configurations and collected 300k events    caliceserv local time is two hours late     CALICE eLogbook   Results for search of  20070601 Page 1 of 19       2007 06 06 17 51 Doerte David assembling status    All cassettes are done once  But       Modules   24 and 36 have many bad SiPMs which should be exchange  The very last spare SiPMs will  come from Moscow 17 06 07 and will be changed Monday 18 06 07     New modules   1 and 2    All tiles are already tested at DESY   2 New FRA plates were prepared with standard holes for mirrors     We started with cassette 2  Due to a misadjustment all holes for SiPMs were shifted though we could  not use that plate  The plat
243. heers     Roman    2007 07 19 19 14 Oliver Allister Henpien run    330683 fbeamData   10x10 10Gev pi 520 8mm Pb air 2 09 O   3 195k  Events    2007 07 19 19 37 Oliver Allister Hen uassian Fit for electron and pion runs    gaussian fits for hHcalEnergy for pion runs    and for hEcalEnergy for electron runs  A A POP    E AAA  pu PA TARA ee  E    E  0 23  as    Sigma  mips        CT    2        v  H  Q  3  D     3  m  IO    lr  ops  tal Do Of o  PR  al Sy cof    lolo  e  e Jo  SEAS ES 19 2  TG Oy oy       I oof oo   ul Of AP  o  FI oo   wo  Oo  FY  oo  ol  00        2007 07 19 20 16 Oliver Allister Hengalibration runs  Beam stopped due to power failure  so take lots of calibration data  BhccmLedvcalibScan SSS                          S       BhcPmLedVcalibScan     E  ENTES    ahcGain      junk  accidentally clicked  Start Run       ehcGain OOOO     CALICE eLogbook   Results for search of  20070701 Page 16 of 22    30651 A A  30692 ahcPmLedVcalibScan HA    m      B     al   HA    30693 Reload Hcal DAC settings  30694 ahcCmLedVcalibScan             2007 07 19 22 06 Roman Cogri scripts restarted    Dear Colleagues     I have restarted the cogri scripts  The data will be shipped to desy as before since  during my work the desy dcache became alive again     The control of the scripts has been handed back to the shift crew   Due to the interruption copying and conversion will be behind  I hope that  it will have kept up with the data taking until the 930h meeting tomorrow     Please t
244. hnu    Replaced 6 HABs on TCMT 1 baseboard to correct problem with unstable current and voltage     2007 07 13 00 23 Valeria Hakan Trygveun 15 GeV       ii MEN ot ee am ge l W    Pb I    Fab  collimators c2  c3  c8 opened to allow more pps   new settings  c2   c3   c8       8  J1 2     2007 07 13 00 35 Fab TCMT 1 unstable    An intervention was needed this afternoon to look at an unstable channel in the TCMT  TCMT 1     see previous entry by Kurt     Things to do when on shift  monitor the current in TCMT 1  is corrent flactuates     10 around the  reference value  46 uA   just note it in the elog   paper log   but do not act  If fluctuation  is above these limits  turn channel off from slow control panel  go from HV main to HV  61 80 and click button  off  corresponding to TCMT 1   channel 77 39A   If there is a  sudden jump of the current to    100uA  switch off TCMT 1 immediately     Please note  intervene on this particular channel only if it is the only one that flactuates  i e   goes from green to red   if most of channels flactuate  just stop run and reload HCAL  settings  Once the setting have been reloaded and the panel is back to being green  then  re start the run     2007 07 13 01 32 Hakan Valeria Trygveun 20 GeV    330381 lbeamData  10x10  20GeV e  533 Air 6mm  Pb  M IN pl  Pb    2007 07 13 02 44 Hakan Valeria Trygveun 25 GeV    330383 lbeamData  10x10  25GeV e  535 Air 6mm 0 5 250k  Pb events          TRIMO6 off nominal value during first 5k events    Seems that
245. holdtime   with KNU       2007 03 21 21 57 S H Chang Tungsten was inserted in the 3rd Configuration     Tungsten was inserted in the 3rd Configuration     2007 03 21 22 00 S  Uozumi 3 GeV run  400643 beam TLL 3GeV KNU fiber  2 5V scanned 36  2007 03 21 22 00 S  Uozumi 3 GeV run  400643 beam T1L 3GeV KNU fiber  2 5V scanned 36  9 10  PUT lee eI E E   2007 03 21 22 03 S  Uozumi 3 GeV run  400644 beam TIL 3GeV KNU fiber  2 5V scanned 36  scan  2007 03 21 22 03 S  Uozumi 3 GeV run  400644 beam T1L 3GeV KNU fiber  2 5V scanned 36  scan  2007 03 21 22 05 S  Uozumi 6 GeV run  pepe we eee ee    9 10  scan  2007 03 21 22 05 S  Uozumi 6 GeV run  AAA AAA  eae  9 10  2007 03 22 09 52 S  Uozumi 6 GeV run  400648 beam T1L 6GeV KNU fiber  2 5V scanned 36  9 10  qe cup YE AAA AAA   2007 03 22 09 52 S  Uozumi 6 GeV run  pp qc ge queen 7   9 10  2007 03 22 09 53 S  Uozumi 1 GeV run       400649   em n          iGev         ENS fiber E 5v     scanned     Bs  7     CALICE eLogbook   Results for search of  20070301 Page 37 of 76    9 10  scan    starts with no good  positron  beam       2007 03 22 09 53 S  Uozumi 1 GeV run    400649 beam TIL 1Gev KNU fiber  2 5V scanned 36  9 10  scan    starts with no good  positron  beam  Stopped  Useless data     2007 03 22 10 08 S  Uozumi slides for AHCAL meeting today    This is the slides for status report of SCECAL BT at DESY AHCAL meeting        File  http   ttfinfo desy de CALICEelog data 2007 12 21 03 2007 03 22T10 08 28 01 ppt    2007 03 22 12 06 S  Uo
246. howing an increase in total energy for 400 to 900 MIPS  then it goes suddenly to 50   MIPS per event  and stays relatively flat     Roman tried to look if the power supply not in slow control was doing weird things  but he saw  nothing    special   Roman and Marcel will try to track that down tomorrow when installing the new slabs     Time  00 10 58 397 335 Wed Jul 4 2007    Run 330079 0 Event 5070 Hits  160 Energy  688 775 mips    eee tt n n ng                            8 J  File  http   ttfinfo desy de CALICEelog data 2007 27 04 07 2007 07 04T02 11 00 03 eps    2007 07 04 03 22 Anne Marie  Heng fte  Abhweird behaviour in first 100s  How can we put 2 plots in one entry      Plot below associated to last entry        hEcalEnergyVsTime  Entries 150000  Mean 1158  508 8  57 98       o 20 40 60 80 100 120 140 160 180 200  time elapsed  sec   File  http   ttfinfo desy de CALICEelog data 2007 27 04 07 2007 07 04T03 22 09 02 eps    2007 07 04 03 25 Anne Marie  HengmaMihale 0 5    CALICE eLogbook   Results for search of  20070701 Page 20 of 22       330081 Praise     ICAL nominal  0  5  TOM  nomina                     combined       BCAL nominal  0 5  TCMT nomina  Some observations on HCAL from the monitor plots  330083    a  Looks noisy in display  probably mip value needs updating     b  on avg  pedestal seems to be  10 to  15 counts  C  Led per layer shows   ve value for 31 lt layer lt 34  1st 10 layers inoperational       2007 07 04 04 17 Anne Marie  Heng has  isetting back to 
247. i 100 GeV pion    Pb       2007 07 13 12 19 Mangi  Beni 80GeV pion    CALICE eLogbook   Results for search of  20070701 Page 10 of 20    Pb changing  Conditions            run stopped  as there will be no beam for at least one hour     amp  2007 07 13 13 27 Mangi  Beni no beam  Damper problem  access in BA2   no beam    more news at 14 30        updates        14 00  access in BA2 at 14 30   no beam  more news at 15 00    2007 07 13 13 35 Mandi  Beni calibration runs    330397 ahcGain spill  sync off   330399 iahcPmLedVgapib can  sync off   330400 ahcCmLedVgapib   can  sync off    330401 jahcPmNoise e 50000  spill  sync off   330402 lahcCmNoisq e 50000  spill  sync off          2007 07 13 14 08 Beni DAQ speed test run  peer  2007 07 13 14 18 Mango  Beni Gaussian Fit Result for Pion runs    By performing a Gaussian fit to the main energy peak in HCAL we extract the following values     Beam Energy  GeV  EE E  MIP  Sigma  MIP  Sigma E    EE E392 E38 ERE E  T   ee E E E             stage shift after run 330393  For data file 330395 and 330396  the main peak is less significant  comparing to 330393 and 330392       2007 07 13 15 01 Beni CMB pickup noise   pedestal shift    switched SiPM HV to 10 0V    330403 ahcGain spill sync SiPM  off HV off   pickup  noise    pedestal  shift run    2007 07 13 15 42 Beni PM LED holdscan    There is still a mess in the AHC cfg files VCALIB settings and for the holdvalues of the LED modes   We have a lot of missing PM holds     330404 ahcPmLedHds
248. iane    Pb       4300 pps  DAQ rate  85 Hz  150k events    CALICE eLogbook   Results for search of  20070701 Page 5 of 20       2007 07 29 12 08 Beni  Martin  e  25 GeV  Ariane  pee eee AA A ATA AAA   5100 pps    DAQ rate  83 Hz  200k events    beam went off for a few seconds    2007 07 29 12 50 Beni  Martin  junk run  Ariane       junk    beam slightly off  new run started    2007 07 29 12 56 Beni  Martin  e  18 GeV  Ariane  po A AA A AE che  5100 pps    DAQ rate  85 Hz  200k events    changed momentum collimators according to general to obtain acceptable pps rate    2007 07 29 13 51 Beni  Martin  e  8 GeV  Ariane  peg emp ye cue cu e 7  1700 pps    DAQ rate  75 Hz  200k events    2007 07 30 00 00 Roman  Beni new shift    2007 07 30 00 15 Roman  Beni calibration    31137 BheGain Od  31138 ahcCnLedvealibscan  ER  31139 ahcPmiedVcalibScan HA   Fe 20900    51130 ahcCmNorse SCS  FOO  31141 BhcPmNoise Fe 20000       2007 07 30 01 12 Roman  Beni 10 GeV pion  root RI D Lo PP f    Pb junk  Pb junk    runs for TRIMO5 adjustment  10x10 trigger   1 4k spill  DAQ   60Hz       2007 07 30 01 37 Roman  Beni 10 GeV pion       Pb    run stopped to check on noisy ECAL layer 9    2007 07 30 01 39 Roman  Beni access to check on ECAL noise      A 2007 07 30 01 57 Roman  Beni no more checklist  We run out of CALICE TESTBEAM 2007 checklist sheets     E 2007 07 30 02 02 Beni new PIN baseboard  I replaced the PIN diode readout HBAB  that got damaged during the rotation  The PIN HABs are still  the s
249. iber 2 5V to  peak  2007 03 21 21 39 S  Uozumi LED run  400618 CmLed LED KNU fiber 2 5V for  Pops OE 8M ll nm  2007 03 21 21 39 S  Uozumi LED run  400618 CmLed LED KNU fiber 2 5V for  pee Xe ce zc  A Jh    0 deiieiepn  2007 03 21 21 39 S  Uozumi LED run  400619 PmLed LED KNU fiber 2 5V for  eee listez estirjeion  2007 03 21 21 39 S  Uozumi LED run  400619 PmLed LED KNU fiber 2 5V for  LEE listezestirjeton  2007 03 21 21 39 S  Uozumi LED run    400621 CmLed LED KNU fiber 2 5V to  see lpe  peak    CALICE eLogbook   Results for search of  20070301 Page 34 of 76    2007 03 21 21 39 SS Uozumi LED run  400621 CmLed KNU fiber  peak  2007 03 21 21 40 ms Uozumi LED run  400623 CmLed KNU fiber  peak  2007 03 21 21 40 SE Uozumi LED run  400623 CmLed KNU fiber  peak  2007 03 21 21 40 S  Uozumi LED run  400625 CmLed LED KNU fiber 2 5V for  A A A  ar    listezestirjeton  2007 03 21 21 40 S  Uozumi LED run  400625 CmLed LED KNU fiber 2 5V for  AAA EE AAA E as  2007 03 21 21 40 S  Uozumi LED run  400626 PmLed LED KNU fiber 2 5V for  pee AA AA E ege Loo needed  2007 03 21 21 40 S  Uozumi LED run  AAA qu eee gh ee  inter calibr  tion  2007 03 21 21 41 SE Uozumi LED run  400628 CmLed KNU fiber  peak  2007 03 21 21 41 xi Uozumi LED run  400628 CmLed KNU fiber  peak  2007 03 21 21 42 ES Uozumi LED run  400629 CmLed KNU fiber  peak  2007 03 21 21 42 E Uozumi LED run  400629 CmLed KNU fiber  peak  2007 03 21 21 42 5 Uozumi LED run    400631 CmLed KNU fiber  see lpe  peak    2007 03 21 21 42 S  Uozu
250. ich are buffered  and hence  need to be read after the spill      CALICE eLogbook   Results for search of  20070701 Page 3 of 20       Run 330247  Cig 5  Aes 18 5  Ext 213120315700  Clg record rato             A   Eb  7  ion 128  ep 188    Time Iech  Ron 330247  Cig 5  Atg 18 8  Eet 21315780 15 790  Cig record data rate    Data rate  Mitytexis     10 d   e      w     1 0  m       150  Time  sec     File  http   ttfinfo desy de CALICEelog data 2007 28 12 07 2007 07 12T11 30 45 00 jpg          2007 07 12 12 50 Beni HCAL switched off    For the debugging of one ADC line ASIC in module 1 the HCAL  ASICs  amp  CMB  was switched off     2007 07 12 12 58 Marcel power cables changed  changed power cables for bottom layers 16 and 17   Noise for buttom layer 17 dropped from 12 to 10 ADC   ECAL turned off after tests     2007 07 12 13 10 Marcel ECAL stage position wrong  The display in the stage control program shows X 4950  while the ruler says X 4750   Apparently the control program has been restarted without having edited the config file   This last movement seems not to be recorded in the e log           2007 07 12 13 30 Beni HCAL powered up again  HV on   CMB on     amp  2007 07 12 13 36 Beni slot 7 FE 6 is bad    CHANGE IN HCAL MAPPING  AHC cfg     The bad read chip in module 1 could be a broken FE     slot 7 FE 6  SERO08  shows too high noise on one ADC channel without anything connected to it  Also  the noise of another ADC seemed to low     CALICE eLogbook   Results for search 
251. igger rate  6 05 kpps  DAQ rate  85 Hz    Logbook entry   CALICEelog data 2007 30 23 07  23 07 2007  13 48    330860 set trim06 to  5A  wrong direction  330861 set trim06 to  5A  right direction but too much       Beni  Petr   Laurent    330863 set trim06 to  2 5 and it s good       Logbook entry   CALICEelog data 2007 30 23 07  23 07 2007  13 27    330859 beamData 10x10  8GeV e  525 air air 2 563  6 0     Events    x was not centered in drift chamber       Beni  Petr   Laurent 8Gev electron run          Logbook entry   CALICEelog data 2007 30 23 07  23 07 2007    13 19 petr   laurent    E  GeV   mean  mips  sigma  mips   sigma mean      EE    232 46 19 8  183 42 22 95             E  GeV     15    pa      27 09 2007 11 39    CALICE eLogbook http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log       Logbook entry   CALICEelog data 2007 30 23 07  23 07 2007       E 1259  petr   laurent calibration runs    330854 ahcCmNoise 20000  330855 ahcPmNoise 20000    330856 ahcCmLedVcalibScan EE       330857 ahcPmLedVcalibScan EE       Logbook entry   CALICEelog data 2007 30 23 07      23 07 2007  zy ILo petr   laurent    Logbook entry   CALICEelog data 2007 30 23 07    E Tr E petr   laurent test of 6 GeV runs    330851 beampata 10x10  6GeVpi  p23    air 2 5  6 0 test run only few events  330851 beamData 10x10  6GeV e   023lairlair 2 5  6 0  run only few events    we can do 6GeV      pion run has about 30  pions   electron run has a ratio 6 1 electron pions  less tha
252. inally will remove them middle of  September     Cheers     Roman    CALICE eLogbook   Results for search of  20070701 Page 1 of 20       2007 07 29 00 13 Roman  Nanda pi  run    Pb       10x10 trigger rate  1 4 kpps  DAQ rate  60Hz  150kevt    2007 07 29 01 00 Roman  Nanda e  run       closed collimators 1 5 6 10 11 from 12 to 5mm  10x10 trigger rate  4 1 kpps   DAQ rate  80Hz   150kevt    2007 07 29 01 26 Roman  Nanda e  run    10x10 trigger rate  10 9 kpps   DAQ rate  90Hz   beam not centred in DC  x mean at  100     gt  change trim06 from  5 to  15A  331078   wrong direction again        gt  adjust trim06 to  5A   331079   now nicely centred beam   have to reduce collimators 1 5 6 10 11 from 6 to 5mm  then 10x10 trigger rate  6 8kpps  150kevt   DAQ rate  80Hz       2007 07 29 02 17 Roman  Nanda e  run    junk       good    not centred in DC    gt  adjust trim05 from  4 to  6A  amp  trim06 from  2 5 to  10 5A    10x10 trigger rate  8 6kpps  DAQ rate  80Hz  150kevt    2007 07 29 02 47 Roman  Nanda e  run    10x10 8 8kpps       trigger rate  8   DAQ rate  80Hz  150kevt    2007 07 29 03 19 Roman  Nanda pi  run    Pb    10x10 trigger rate  3 2kpps  DAQ rate  70Hz  150kevt       2007 07 29 03 52 Roman  Nanda pi  run       Pb    10x10 trigger rate  4 1kpps  DAQ rate  70Hz  150kevt    2007 07 29 07 57 Beni  Martin  pi  12 GeV  Ariane    PD       CALICE eLogbook   Results for search of  20070701 Page 2 of 20       1800 pps  DAQ rate  60 Hz    previous beam settings running contino
253. ints precise scan for X layer strip 5 and 14   end of 2nd configuration energy   position scan                           2nd configuration MIP calibration after removing tungsten                  400519   400520   400521   400522   400523   400524   400525   400526   400527    end of MIP calibration    Precise scan with 2nd configuration without tungsten    400529 entire detector scan  9x9 points    400530 entire detector scan  9x9 points    400531 entire detector scan  9x9 points    400532 entire detector scan  9x9 points    400533 entire detector scan  9x9 points    400535 one strip scan  3x10 points    400536 one strip scan  3x10 points   may be stopped before finished  400537 one strip scan  3x10 points   may be stopped before finished    end oft Precise  Scan     1222221822    MIP calibration with 3rd configuration  36 points scanned     4       4  4  4  4  4  4  4  4    LED runs with 3rd configuration before    00543  00544  00545  00546  00547  00548  00549  00550  00551        Stopped at 34th point   end of MIP calibration with 3rd config    Pedestal runs                   data             400555 CmNoise with LED on   400556 PmNoise with LED on   400557 CmNoise with LED off   400558 PmNoise with LED off   For Layer20 stripl5    400553 CmLed for inter calibration   400554 PmLed for inter calibration   400562 CmLed to see lpe peak   400564 CmLed to see lpe peak around strip 15   For Layer21 stripl5    400567 CmLed for inter calibration   400568 PmLed for inter calibrati
254. ired    A 2007 07 26 00 35 Beni wrong runnumbers    Today  I had to change the data link to the non full disk on calice02  Unfortunately  I forgot to  update runNumber txt  Therfore  we took two runs with already existing runnumbers  After  consulting Roman I changed the name of the files to the runnumbers these should have  been     We have to check if these runs pass conversion and processing without problems     2007 07 26 00 39 Beni Switched HV off          2007 07 26 08 01 Erika   many CMB off  try with Manual re init  2007 07 26 08 59 Erika  2007 07 26 09 09 Erika Fabrizio Mar  hcCmLedVcalibScan  2007 07 26 09 13 Erika Fabrizio MarahcGain   330956 ahcGain  2007 07 26 10 41 Erika  Fabrizio Martest run       Pb    DAQ rate  80 Hz    2007 07 26 11 00 Erika Fabrizio Martest run       Pb    2007 07 26 11 04 Erika Fabrizio Martest run       Pb    checking the disconnection of the drift chamber channel nearest to the ECAL   drift chamber no 2  in online monitoring program    2007 07 26 11 12 caliceon 2calicehawseqis  ipmh   CALICE user     Moved stage lcm up to move beam to geometrical center of ECAL  We used a displaced position as  center  beferore  as the ECAL was missing some bottom slabs     CALICE eLogbook   Results for search of  20070701 Page 16 of 20          HCAL Motor CU panel    actual position   target position    x   710 10 be CR  Y   19 9 Lo   i  Yi    O lost X reference point    lost Y reference point    O lost 2 reference point    Q stage in motion  O moral op
255. isted for example in run 330037       Bad channel  layer 7  strip 4  pedestal 948 807     63 6077       Bad channel  layer 7  strip 5  pedestal 820 662     289 902       Bad channel  layer 5  strip 19  pedestal 559 559     124 605       Bad channel  layer 5  strip 15  pedestal 896 465     61 5737       Bad channel  layer 1  strip 9  pedestal 981 226     73 8279       Bad channel  layer 16  strip 16  pedestal 879 609     118 966       ECAL looks really noisy on the monitoring  and unstable noise  See file below and in next entry     CALICE eLogbook   Results for search of  20070701 Page 16 of 22        ITT    A TS  aHEREEEEEE    ei i        marp mih PAGO IT  H  E          ae py    W  Ww La  had Wi    File  http   ttfinfo desy de CALICEelog data 2007 27 03 07 2007 07 04T00 09 32 06 eps  2007 07 04 01 04 Anne Marie Hengiee4Misboking like french Alps    In run 330073  ECAL energy distribution looks bad   because of layer 0 noisy periodically     CALICE eLogbook   Results for search of  20070701 Page 17 of 22       hEcaHita     d Ee       a  1  el T T                   a E     M     hEcalEnerngyPerLayer  i wb  Fest  be  m    m  de    xa    E C4    amp        AL  kbabkEELPAGR  SRMRGREAGE EGOUGRFCRAGER OR E Gb moa mi arme IBI i diki b     8 a kl     Qm ior  File  http   ttfinfo desy de CALICEelog data 2007 27 04 07 2007 07 04T01  04 24 00 eps  2007 07 04 01 07 Anne Marie HengiwaljBig El CAL voltages      40 0  TCMT    330076 lbeamData 20x20 combined HCAL TCMT  nominal 0 3V       ANN
256. it better  especially in the bottom part where V and T calib signals are connected    after this action all CMB but 2 are working  Missing are    layer 28  mod 1  slot fe 12 1   layer 35  mod 34  slot fe 5 5   both are working when setting Vcalib from CANBUS and not from the DAQ    we set a low value from CANBUS steering  110   2 157 V   28000 DAC settings from the DAQ   and test a gain run    Pedestal stability  we have exchanged the SCSI cable for slot 5 1 B  ex 5 7 B  module 32 pos 33 we need to see if this  has improved the pedestal stability over the next few runs    2007 07 25 20 22 Beni  Erika   350068  wrong runNumber txt should be 350087    note  layer 28 and 35 set to CANBUS controlled amplitude for LED  please check if anything is visible in the LED spectra   maybe light not sufficient  slot fe 7 4 sees no light module 2   also 9 6  module 25  has very low light  enough       2007 07 25 20 34 Beni  Erika       CALICE eLogbook   Results for search of  20070701 Page 15 of 20       A 2007 07 25 22 45 Beni SR loss    I found a lot of modules with lost shift registers  Reloaded the SR     2007 07 26 00 31 Jean Charles  New slabs installed  Roman  Marcel      New slabs  bottom layers 6 11  installed  noise level is very good    rotation tested succesfully     regulators tests performed but rejected due to higher noise level    New power supply  5V  installed     back in ECAL slow control        ECAL is off  fans are running        Position is 47 5    crew is hungry and t
257. it chisq fit NDof    CALICE eLogbook   Results for search of  20070301 Page 65 of 76       File  http   ttfinfo desy de CALICEelog data 2007 12 25 03 2007 03 25T23 26 57 00 dat    2007 03 25 23 28 S  Uozumi    Then I look into 6 GeV central injection run   First I look longitudinal shower profile to see the calibration constants are reasonable or not     Attached plot shows the profile  sorry no error bars   there are some bumps especially on layer 6     There is another bump at the border of fiber and direct readout modules    Note  I made a mistake about scale of vertical axis  sorry  the vertical axis is not a measured  energy  actually it is measyred energy summed over all events    That s why the measured energy is so huge       CALICE eLogbook   Results for search of  20070301 Page 66 of 76       x10     2000    1000    Measured energy in a layer  MIPs     0 10 20  Layer ID    File  http   ttfinfo desy de CALICEelog data 2007 12 25 03 2007 03 25T23 28 00 00 jpg    2007 03 25 23 30 S  Uozumi    To see what is happening  I make a response map for the 6 GeV center injection run    As you can see the bump in the shower profile comes from high response at Layer 6 strip 14   I checked the MIP fit for this strip  but looks OK    I suspect this may be caused by the non uniformity of the strip    If the shower hits the region very close to MPPC  it gives larger response    But still not sure  if it s true  we should see same effect on other layers      Anyway investigation is ne
258. k   Results for search of  20070701 Page 7 of 20          wey mor   ower    We Crave ON  HY Crate OFF    recover    KILL Hy   wiwa j w2z Im    HY settings   File actions  ele   disable ML chaenels   w2 ws  wel    SE   oa  an    we   disable   gene all tne  el  geg Hi values to file   dae HL vie  position         File  http   ttfinfo desy de CALICEelog data 2007 30 24 07 2007 07 24T14 43 01 00 jpg                   2007 07 24 15 24 Hengne ECal Linearity Check  Low Energy         for different stage position     0 0     0 43      3  3      3  3      6 0      6 0       linearity of stage position  0 0       6 0    6 0  seems to be better than others        Since Run330717  Run330719  Run330720  which correspond to energy 15GeV and 18GeV at stage  position     0   3   were not online Analysed  these two points on the plot of stage position  0 0     are left blanx     no error considered since they are neglectable    ECal Linearity Check  Low Energ       Stage Position    0 0    0  3     3  3     3  3     6 0     6 0        File  http   ttfinfo desy de CALICEelog data 2007 30 24 07 2007 07 24T15 24 03 00 png    2007 07 24 15 51 Beni  Petr   Laurent  Guassian Fit for recent electron runs    CHE SCENE          igna mipsy             isns7nesn 8   3085 5       2007 07 24 16 00 Beni VETO ampitude readout    Identified the bug in the veto amplitude readout setup   updated values for    HV   1600 V   hold   20 ticks  10x10 trigger     CALICE eLogbook   Results for search of  20070701 Pa
259. kkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkzk    WARNING  ECAL slot 17 was not booted correctly and hence completely missing  for runs 330135 to 330142 inclusive  All these runs are junk for the ECAL        Paul   2007 07 06 19 33 Fab Erika Francoisnuon run  beam tuned  good run     DAQ rate  88 Hz    10x10 rate 3700 pps  100x100 rate 15000 pps    Note  values of collimators and quadrupoles changed wrt beam reference file   see previous posting for run 330141     stop the run with 132 000 events    2007 07 06 19 35 Fab Francois NEW WAY TO INSERT NEW RUN INFO    Please from now on use the    insert new CALICE run info  button on top of the WEB page  to insert a new run in the elog     You have to provide the following information     CALICE eLogbook   Results for search of  20070701 Page 7 of 24       Run Number  33xxxx    Run Type  beamData  emcNoise       Trigger   Run Option  10x10  20x20  100x100    Energy  e g   80 Gev  etc     Beam type  e g  muon electron pion    File Name  just insert file number  e g  for  H6B 510  just insert  510    Absorber  e g  air  Pb  etc     Target  e g  air  etc     Cherenkov status  e g  off  on   if on  write pressure    Comment  e g  good run     There is also a Comment 2  box  however this at present only allows you to write one line   To add more comments just save your entry and then re edit it from the main browser page     2007 07 06 19 51 Fab Erika Francois  key access to reboot CRC in slot 17 ECAL  It is working now    2007 07 06 19
260. ks different  even if there is no effect on the module  current     hHcalPedRMSPerChannel slot17       PedRMSPerChannel       00  channel number    File  http   ttfinfo desy de CALICEelog data 2007 31 31 07 2007 07 31T01 21 02 00 gif    2007 07 31 01 21 Beni  Lukas    Pb         JK pps   Stopped after 10k events    Adjusting position  TRMO5 set to  3 0    2007 07 31 01 30 Beni  Lukas    Pb       2007 07 31 02 13 Beni  bukas    E     2k pps  E because of beam adjustments    TRMO5 adjusted to  2 5       2007 07 31 02 19 Beni  bukas    E       2007 07 31 03 01 Beni  Lukas    CALICE eLogbook   Results for search of  20070701 Page 17 of 20       E    1 8k pps   Aborted because of adjustments    TRM06   20  TRMO5    3  2007 07 31 03 06 Beni  SES    s        Aborted  further X adjustments  TRM       2007 07 31 03 11 Beni  Lukas  pote ge ee  Pb   1 8k pps    200k ev   2007 07 31 03 54 Beni  Lukas  popa tr q E cp ges 1   Pb   1 8k pps   Aborted after 10k ev   Adjusted TRM05    5 and TRM06   30  2007 07 31 04 04 Beni  SORS       s    1 8k pps   beam off after 52k ev    2007 07 31 04 12 caliceon caliceseest ts for electron runs    fitted energies for electron runs                            2500      E  2000    140  e  Cs 720 25 30 35 40 48 80 UAE  No M D X    beam setting  GeV  beam setting  GeV     resolution  E L  t  B    E  a 2  65  H  6 D  ss    e  SES  s    45           A A A A A E EE    File  http   ttfinfo desy de CALICEelog data 2007 31 31 07 2007 07 31T04 12 02 00 gif    2007 
261. l  ahcGain  Anne Marie  Petr  FEIER EEE EE  2007 08 19 23 38 Daniel  80 GeV pion      Anne Marie  Petr    350282 lbeamData  10x10  80GeV Git 611 4mm air unknown 100 k    Pb 20deg    too low rate    CALICE eLogbook   Results for search of  20070801 Page 11 of 22       2007 08 20 0 00 Beni they care a lot        This is a copy of the SPS elog entry concerning our Cherenkov problem     More info on the problem with the BXCET  reported by previous shift crew at 15 26   I could  find out that there is a process called BXCET  but this process is running at three machines   v1n438  vln472 and V3n351   We can t find any info which machine is responsable for  the concerning problem  We could try a wreboot  but don t know what side effect it will    have for possible other users  We leave the problem as is for the moment and hope the  BI group can solve the problem in the morning  In the meantime the H6 beamline can still  use the Cherenkov detector  however without the possibility to change settings          2007 08 20 00 52 Satoru  Allister  Drop of trigger rate and HCAL energy plot  Hakan    We notice a drop of trigger rate for run 350283 after a loss of beam              E ue E  E ue E     1006       a P    aee     E  Amb  36  fee elapsed  aet   File  http   ttfinfo desy de CALICEelog data 2007 34 20 08 2007 08 20T00 52 24 00 png  2007 08 20 01 16 Hakan Allister Sat6 GeV pions  350284 Beam 10x10  60GeV Git 610 4mm Air 0 5 20deg   Data Pb 160k ev   2007 08 20 02 03 Satoru  Allister 
262. l and recover DORIS  data useless after 20h45 because of switch over to positrons  good data for 13 tiles       PIA and DESY II Wed Jan 17 08 28 16 2007  Currents  m    Final Energy DESY II  GoVJ  5 16    Or Op uo dud    dq  Oe m0 UO NY O0        8  Time  h     DESY II Energy     PIA e  Cur   18 PIA e  Cur   10 DESY II Curent  File  http   ttfinfo desy de CALICEelog data 2007 03 16 01 2007 01 16T17 53 14 02 gif    2007 01 17 17 47 Sebastian mip scan  Schaetzel   202270 k ehcBeamStagescan_  v 13  3 GeV fe    flchcaldaq01  home caliceon cfg StagePositions innercore flipped 20070117 12h45 dat       2007 01 22 11 12 Niels Position of temperature sensors  Hi all     I made this sketch with the closest cells to the five in module temperature sensors   Hope it helps    Niels  Update Feb 1  2007    The previously posted picture contained an error   thanks to Sebastian Schaetzl for catching the  wrong chip numbers in the lowest row of the middle panel  Plot is corrected  please download    again     Cheers   Niels    CALICE eLogbook   Results for search of  20070101 Page 4 of 5       CALICE eLogbook   Results for search of  20070101 Page 5 of 5       File  http   ttfinfo desy de CALICEelog data doc HCALproduction 2007 01 22T11 12 11 02 jpg  2007 01 29 14 10 E  Garutti MWPC manual  Here you will find a User Manual for the CERN Delay Wire Chambers   http   sl web cern ch SL Publications bi98 023 pdf  2007 01 29 14 11 E  Garutti beam orbits  Beam orbits on the T4 target    http   ab div 
263. le     date   dat   During the august period of serious data taking the DAC files have been not modified   Therefore only one file per module is available     CALICE eLogbook   Results for search of  20070201 Page 11 of 21          File  http   ttfinfo desy de CALICEelog data doc CERN docu 2007 02 27T09 03 13 02 pdf    la 2007 02 27 09 30 Marius AHC cfg files for october run    enclosed all used AHC cfg files during the october run period  more information in the next entry    CALICE eLogbook   Results for search of  20070201 Page 12 of 21       File  http   ttfinfo desy de CALICEelog data doc CERN docu 2007 02 27T09 48 03 00 zip     4 2007 02 27 09 39 Marius AHC cfg files for october run period    this files gives a overview about the used AHC cfg files during the october run period        a small comment in the first line gives an idea why a new file was needed     the first files  in which not all 23 modules are listed  because they were not yet installed are  Skipped  but they are in the zip file     all files  where only the TCMT frontends were activated a left out  but they are in the zip file  in the entry before    CALICE eLogbook   Results for search of  20070201 Page 13 of 21          File  http   ttfinfo desy de CALICEelog data doc CERN docu 2007 02 27T09 39 04 01 paf    2007 02 27 09 52 Marius DAC files for october run period        enclosed all used DAC files during the october run period    CALICE eLogbook   Results for search of  20070201 Page 14 of 21       Fi
264. le  http   ttfinfo desy de CALICEelog data doc CERN docu 2007 02 27T10 12 32 00 zip     4 2007 02 27 09 53 Marius DAC files for october run period  DAC files for october      gives a overview about the HV settings       shows which DAC files for which module have been used  latest version        some comments    The mentioned DAC files are in the zip file in the entry below this one   In this file all DAC files during the october run period are stored    not only the latest ones     A new DAC file have been used from the date  which is specified in the filename  on   DAC module     date   dat    During the october period of serious data taking the DAC files have been modified  to reduce for  example long discharge   Therefore in the zip file are sometimes more than one DAC file per module     The reason for the not properly working Module01 can be found here  I have messed up the DAC file  for Module01    it is identically to Module02  which leads to the conclusion that I did  a mistake during copying     CALICE eLogbook   Results for search of  20070201 Page 15 of 21          File  http   ttfinfo desy de CALICEelog data doc CERN docu 2007 02 27T09 53 46 00 pdf    4 2007 02 27 10 00 Marius SlowControl HV configuration files    enclosed all used SlowConfiguration for the HV during both run periods    CALICE eLogbook   Results for search of  20070201 Page 16 of 21       File  http   ttfinfo desy de CALICEelog data doc CERN docu 2007 02 27T20 27 30 00 zip       2007 02 27 14 02 
265. led to control room and ask to tune one parameter to be optimum to  the positron beam  and we now get higher rate   300Hz with 3GeV  even with the 6GeV positron  in DESY II    Now we return to the beam data taking    Thanks Erika  Beni      2007 03 23 18 36 S  Uozumi LED run       CALICE eLogbook   Results for search of  20070301 Page 41 of 76                2007 03 23 18 36 S  Uozumi LED run  400666 CmLed LED KNU fiber 2 5V for  A PA A AA Ts  2007 03 23 18 37 S  Uozumi LED run  KNU fiber 3 for  inter calibr  tion  of L12S15  2007 03 23 18 37 S  Uozumi LED run  400667 PmLed LED KNU fiber 2 5V for  eee E listezestiejeion  2007 03 23 18 38 S  Uozumi LED run  400670 CmLed LED KNU fiber 2 5V to  see lpe  peak at  L12514 15  2007 03 23 18 38 S  Uozumi LED run  400670 CmLed LED KNU fiber 2 5V to  see lpe  peak at  EMEN E ug E 15  2007 03 23 18 39 S  Uozumi LED run  KNU fiber Drs SV  for  inter calibr  tion  ETE ih  see  2007 03 23 18 39 S  Uozumi LED run  400672 CmLed LED KNU fiber GV for  inter calibr  tion  p eq um AAA A  2007 03 23 18 39 S  Uozumi LED run  KNU fiber 2 5V for  inter calibr  tion  ESTATE TA  2007 03 23 18 39 S  Uozumi LED run  KNU fiber   for  inter calibr  tion  ESA E      2007 03 23 18 40 S  Uozumi LED run  400675 CmLed LED KNU fiber 2 5V to  see lpe  peak at  L10s15  2007 03 23 18 40 S  Uozumi LED run  400675 CmLed LED KNU fiber 2 5V to  see lpe  peak at  L10s15  2007 03 23 18 41 S  Uozumi LED run       400679 CmLed LED KNU fiber 2 5V for  inter calibr  tion  o
266. m  0123456789101 012345678 931011          File  http   ttfinfo desy de CALICEelog data 2007 27 03 07 2007 07 03T 18 37 49 05 png    2007 07 03 19 01 Kurt Francis TCMT debuggin run    330066 ahcCmLedNois   internal test  run to  verify  faulty HAB  in TCMT    2007 07 03 21 10 Kurt Francis    BY aa  sd  330088     ER Cd Cd ER a                          2330069  ahcGain X  LedTrigger Jno beam  car            2007 07 03 22 37 Michele AlexandertiashdeamData    330070 beamData 20x20 mu  combined  10  20 155k  good run  ruri       330072 beamData 20x20 mu  combined  10  10 155k  good run    2007 07 03 23 27 Sasha  Vishnu TCMT movement  Move TCMT another 30 cm in x      gt  Jura   2007 07 03 23 57 Fab check list during shifts    Please run check list  at least  twice per shift   In the time between the check lists  please check that both HCAL and ECAL are on while we take data     If not on  and there are no obvious resons why they should be off     please turn on and make an  entry in the elog     2007 07 04 00 07 Fab    Movable stage back to the  0 0  position   Start programme for the night  voltage scan of HCAL  muon beam      CALICE eLogbook   Results for search of  20070701 Page 15 of 22       2007 07 04 00 09 Anne Marie Hengriiratishiseries of beamData runs that will scan in SiPM  HVs    8 PL   330074 ahcPmNoise  LED combined nominal  ECAL HCAL    40 in x  for TCMT   330075 ahcGain combined nominal  ECAL HCAL    40 in x  for TCMT    Monitoring list new bad channels  weren t l
267. mi LED run    CALICE eLogbook   Results for search of  20070301 Page 35 of 76    400631 CmLed KNU fiber  see lpe  peak       2007 03 21 21 43 S  Uozumi LED run  pee quee cip oso cm EE  inter calibr  tion  2007 03 21 21 43 S  Uozumi LED run  poer quee AAA O Cererem  inter calibr  tion  2007 03 21 21 44 S  Uozumi LED run   o ee Nest com  AAA qe  Je a  inter calibr  tion  2007 03 21 21 44 S  Uozumi LED run  pe upee ge can ARA EL a  inter calibr  tion  2007 03 21 21 45 SS Uozumi  400636 CmLed deg T  peak  2007 03 21 21 45 SC Uozumi LED run  400636 CmLed KNU fiber  peak  2007 03 21 21 45 S  Uozumi LED run       see lpe  peak  2007 03 21 21 45 Imm Uozumi  400637 CmLed deg      peak  SS 03 21 21 46 S  Uozumi pedestal run    PmNoise pedestal KNU fiber   DV Noise    Run with  200k events       2007 03 21 21 46 S  Uozumi pedestal run  400640 PmNoise pedestal KNU fiber 2 5V Noise  Run with  200k events  2007 03 21 21 47 S  Uozumi pedestal run  400641 CmNoise pedestal KNU fiber 2 5V Noise  Run with  200k events    2007 03 21 21 47 S  Uozumi pedestal run    400641 CmNoise pedestal KNU fiber 275V Noise  Run with  200k events             CALICE eLogbook   Results for search of  20070301 Page 36 of 76       2007 03 21 21 48 S  Uozumi pedestal run    400642 sceBeamHoldlEd  n 3GeV KNU fiber  2 5v  25 20  holdscan  detector  to check  center holdtime           2007 03 21 21 48 S  Uozumi pedestal run    400642 sceBeamHoldlE    n 3GeV KNU fiber  2 5v  25 20  holdscan  detector  to check   center 
268. mi uniformity of direct readout module      and this shows uniformity of light yield with a direct readout strip     CALICE eLogbook   Results for search of  20070301 Page 59 of 76    200    responsemap  Entries 10000    15073  Mu      Meanx 24 03      Meany 41 11            w    O RMSx 6421   100 RMSy 13 54    50     50 MIMI AN  60 40 20 0    File  http   ttfinfo desy de CALICEelog data 2007 12 25 03 2007 03 25T12 34 11 00 jpg    2007 03 25 15 19 S  Uozumi MIP calibration  KNU fiber  2 5V scanned 36  points  scan for  IP  calibration    MIP calibration  KNU fiber  2 5V scanned 36   points  scan for   IP  calibration       2007 03 25 15 19       MIP calibration  KNU fiber  2 5V scanned 36  points  scan for    2007 03 25 15 19       IP  calibration    2007 03 25 15 19 S  Uozumi MIP calibration  KNU fiber  2 5V scanned 36  points  scan for    IP  calibration       2007 03 25 15 20 S  Uozumi    Switched to one strip precise scanning     2007 03 25 15 22 S  Uozumi MIP run   400788  beam  MIL  BGev   KNU fiber p  scammed     Pxi9     7       CALICE eLogbook   Results for search of  20070301 Page 60 of 76    one strip  precise  scan    2007 03 25 15 22 S  Uozumi MIP run          400788 beam TIL 3GeV KNU fiber  2 5V scanned 3x10  one strip  precise  scan    2007 03 25 19 36 S  Uozumi    Runs 400788 400794 are precise one strip scan     2007 03 25 19 43 S  Uozumi over voltage for fiber module changed  The DAC setting for fiber readout module has been changed from  0 4V to  1 2V to di
269. minutels  during   this run  at 65k  events        330283 beamData  10x10   50GeV 0 1178bar 0   3 202k  Es events    2007 07 10 06 27  gt  cogrid scripts restarted   OUTPUT SENT TO CC IN2P3  Lyon    Dear Colleagues     I have started the cogri scritps  and give them back under shift crew  control taking into account  the lines below     As the srm dcache turned out to be pretty unstable tonight and it is  likely that there will   be more interruption during tomorrow due to tests on the problem at  desy  all data binary and converted files   are currently copied to cc in2p3 at lyon  I have modified the scripts  accordingly and first tests   turned out to be successful     This is true for runs  gt   330279    All runs until then are available at desy  however the conversion was  blocked by this and another problem with the   ressource broker we have used at desy  I have also switched to another  ressource broker which seem to work   better     Runs betwen 330257 and 330278 have not been converted so far as well as  14 runs earlier which got stucked   in the  Scheduled  state at epgce2 ph bham ac uk   Unfortunately  the  problem of today prevented me from   curing the first one by resending the jobs     CALICE eLogbook   Results for search of  20070701 Page 9 of 25       Transfer rates into lyon are with  10 15 MByte s satisfactory  such  that there is no big worry concerning speed     At least everything looks much healthier now than earlier tonight   Cheers    Roman   P S   T
270. mm  Pb  330582 beamData  10x10  100GeV fpi  516 4mm Air N A spill  Pb rate  6 5k                CALICE eLogbook   Results for search of  20070701 Page 16 of 19    330583 lbeamData  10x10  100GeV   pi  4mm spill  Pb rate  6 5k    DCs offset  changed TRIMO6 from    Still offset  changing to  8 0A  DCs aligned  Beam stopped for several minutes  will continue taking remaining events when it s back       2007 07 17 10 16 Allister  Imad  pion run  Beni    330589 lbeamData  10x10   50GeV pi  512 4mmPb air n a  6 0  position   ok     352kEvts    2007 07 17 12 13 Allister  Imad  40 GeV pions  Beni          Pb    Changed COLLO1 to   8 8  to reduce rate to 10k per spill  Saved this change to beam file              2007 07 17 12 57 Allister  Imad  45 GeV pions  Beni  2007 07 17 13 55 Allister  Imad  calibration runs  Beni  330593 ahcGain  330594 junk run          2007 07 17 14 26 Beni 35 GeV pion    WW SS Le  Pb       Pb    This run was taken directly after the ahcGain run 330593   Some CMBs are missing in the online monitoring    Performed manual re init  and status check of CMBs   330596   no change of the CMB situatuion    2007 07 17 15 40 Allister  Imad  pion run  CMB check  Beni    Pb       just to check if there s a change in hHcalLedPerLayer  no change found     2007 07 17 16 41 Nanda Mangi Valgsipn run       Pb    trigger rate  100 kpps      close collimators  done by Fabrizio    all collimators set to 5    trigger rate now  8 3 kpps  DAQ rate  90 Hz    2007 07 17 17 47 Valeria  
271. ms that copying has resumed successfully     It looks as if the  slow  dcache doors at desy sometimes block transfers  completely  The problem with the bandwidth differences is under heavy  investigation by desy it     Control over cogri scripts is handed back to shift crew        Roman   2007 07 30 22 30 Alex Catherine calibration runs  231187 Jenccain oT E E E         l1 I  No beam      2007 07 30 23 01 Alex Catherine    88 lahcCmLedVyt xibScan  15GeV 20 deg   ro tel   funk    lbeamData  10x10  15GeV 20 deg   pecu qe  PA PO p     Pb    331192 Reload       p o 0   0         pim               CALICE eLogbook   Results for search of  20070701 Page 14 of 20    HCAL DAC current   Settings in mod  3A  leds  still  off    20 deg   junk     331193 beamData  10x10  15GeV  GER  SE    LEDs  but 29 and 35  are ok now  intervention in the CMB panel was necessary        2007 07 30 23 58 caliceon calicehdaN6b 2 dighoturrent in HCAL channel  Layer 3A   CALICE user            as lada dr al              AAA              A m s Dp    21 30 22300 eet au 23100  Hi  2007 30  DE  2007 30 07  2007 30  OF  2007    File  http    ttfinfo desy  deCALICEelog data 2007 31 30  07 2007 07 30T23 58 45 00 jpg       2007 07 31 00 03 Alex Catherine 20 GeV e  run  pugne que ee A AA A A  Pb positioning     2007 07 31 00 12 Alex Catherine 20 GeV pion run    Pb    2007 07 31 00 28 Beni  Lukas    Pb          331198 lbeamData  10x10  20GeV e  533 Air 6mm 0 7 20 deg   Pb stopped  after  169k ev     2007 07 31 01 05 c
272. mu   mu   u     p    ahcGain 100x100   80GeV m i i    configuratlions       HCAL HV  set to  0 3V  file CERN_030707_ 0 3V txt    CALICE eLogbook   Results for search of  20070701 Page 11 of 24             2007 07 07 09 08 Kurt TCMT DACs   DAC change in TCMT 1 file chip 4 channel 1 value changed from 70 to 255 to turn OFF SiPM   This was an attempt to turn off a noisy channel that may be causing layer 7 to have an  incorrectly high energy     2007 07 07 09 17 Kurt Cristina Michiid  ankas  tip for  80 GeV pions    Loaded file H6B 510    Collimators closed  a bit  because rate is too high        Good rate of pions   10x10 7500 pps  20x20 8500 pps  100x100 970 pps    Started beam tuning run using above configuration   DAQ changed for 10x10 trigger    2007 07 07 09 42 Fab Erika Cristinalbfamnaltinipgnyun pions  80 GeV    330171 lbeamData  10x10  80Gev pi  510 air air tuning  beam for  pion run    Movable stage in  0  position       potential problem with layers 15 and 17 of ECAL   the number of hits ok  but the energy nearly 0    HCAL chip slot 7 chip 76 drifted to negative values    2007 07 07 09 53 Fab Erika Cristina Michal Evgeny       Movable stage in  0 0  position  trigger rate 10x10  11k  DAQ rate  70 Hz    2007 07 07 10 11 Kurt Cristina Mich amp l  arikesealp   50 GeV pions    Loaded file H6B 512    Collimators setting not loaded from beam file  only magnets    Settings for collimators modified from previous file        Some electron contamination is visible from the on li
273. n 1000 particles per spill                   Logbook entry   CALICEelog data 2007 30 23 07      23 07 2007  Si 1000 petr   laurent       4mm    330850 beamData 10x10 10GeV pi  520 5  air 1 8635  6 0 235k Events       DAQ rate  90 Hz    Logbook entry   CALICEelog data 2007 30 23 07      23 07 2007  A 09 17 petr   laurent          Logbook entry   CALICEelog data 2007 30 23 07       3 23 07 2007  y 08 25    330848 beamData 10x10  15GeV pi  521 en ds 0 98  6 0 207K events    Logbook entry   CALICEelog data 2007 30 23 07      23 07 2007  m 08 22          7 of 28 27 09 2007 11 39    CALICE eLogbook    8 of 28    http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log       4mm  Pb  Run stopped because we have to reload DAC setting  slot fe 17 0 B lost setting  second time     330846 beamData 10x10  15GeVpi  521 air 0 98  6 0 35K events       Logbook entry   CALICEelog data 2007 30 23 07    23 07 2007  08 21          TRIMO5 5A   gt 0 5A   gt  1A   gt  3A    Logbook entry   CALICEelog data 2007 30 23 07    23 07 2007  08 20    same as yeasterday slot fe 17 0 lost settings   lowered current limit for HV ch  28       Erika reload dac settings    recovered after reload    3 A 40 7   0 2    Bras  e      Drax  28 40 6  0 2  Eras  mg ras  in ras    Or on     Breas       _          Sisi    MA a  DE 20  5     Sg B 41 1 90 290 5         Gm      po Hose   Mr a 1 AER  oe  EU OU ie   ceo gin wm    E    We 207 2007 23 07 2007 23 07 2007 23 07 2007    REG    Bra      Dr an    Dr
274. n caliceserv  Don t call him her in the middle of the  night        b  The root user should then set the network parameters by     gt  cd  scratch caliceon bin       setNetwork serv sh    Cheers     Roman     amp  2007 07 03 17 59 Roman cogri scripts stopped    The cogri scripts have been stopped for test purposes     CALICE eLogbook   Results for search of  20070701 Page 13 of 22       2007 07 03 17 59 Evgeny HCAL temperature    HCAL temperature after one night of cooling down with fans  temp taken at about 11 00 am    T1 5 module sensors   T6 7 CMB top  bot    e 30    28  26  24    22       Layer number       zo   e 30 e 3  28 28 36  34  26 26  32  24 24 30  22 20 ew 2  Layer number  E  File  http   ttfinfo desy de CALICEelog data 2007 27 03 07 2007 07 03T17 59 28 00 gif  2007 07 03 18 37 Alex  Beni CMB fuses  LEDs    After changing fuses for CMBs 30 and 33 we found three other CMBs with dead fuses  8 9 and 12   We exanged all burned fuses with new 3 A fuses  After turning CMB power off an on  all LEDs were    turned off  During CMB tuning some LEDs were turned off by intention   The information which LEDs were off was lost  but could be reconstructed from  http   www desy de  sschaetz hcal cmbstatus cmbstatus  xhtml    Current LED settings are shown in the picture below     CALICE eLogbook   Results for search of  20070701 Page 14 of 22    malncMHB VZ pri  on elceheateco2     CMB Main 123 45 6789 10 11 12 13 14 15 16 17 18 19 20 21 22 2    m Pulser LED ensble disable     z
275. n the  readout                 Tp mum eene ON Hd O erre nene AN npe pros ven rompen eren pe prenne em emen erem  14 00 20 00 00 00 08 00 14 00 20 00 40 00 08 00 14 00 20 00 02 00 08 00 14 00 20 00  2   07 2007 06 07 2007 09 07 2007 09 07  2007 09 07  2007 09 07 2007 10 07 2007 10 07 2007 10 07 2007 10 07 2007 11 07 2007 11 07 2007 11 07 2007 11 07 2007 12 07           File  http   ttfinfo desy de CALICEelog data 2007 28 11 07 2007 07 11T21 29 02 00 jpg    CALICE eLogbook   Results for search of  20070701 Page 25 of 25       ls  e Pos 35   Module 34   CMB 37  PEL to eon    Tene  07 10 2007 06 40 08 44 813    ue                  117 000 07 10 2007 06 49 08 4 813                  117 000 07 10 2007 06 40 08 AH  813         117 000   07 10 2007 06 40 08 ez                  9  cic n di s    117 000 c   ezztaz2007 06 40 66 as                  File  http   ttfinfo desy de CALICEelog data 2007 28 11 07 2007 07 11T21 31 02 00 jpg       2007 07 11 21 39 Beni HV off    We switch the HV to 10 V over night  as there is nobody who could recognise a SR loss  LV and CMB  stay powered to keep the temperature distribution stable    CALICE eLogbook   Results for search of  20070701    la 2007 07 06 00 37    Roman Ecal Mappin    Please find an updated Ecal Mapping in the attachment         Lampa P C  d cone d Sep     ef Kar la       c 13 Ek  uf A  p ls 4 D     HE i al H4 B  1 Ve 4     1 ac E E ir 5  T n d 1  A 56 E a   1  3 1D H 4  i ac iB EI En fi  il Ke y 1  z Se 3 a ps  2  i iD d x  E 16 14 R
276. n the local disk     Current run 202402 stopped at configuration 516   I will restart the copy scrip  and check back later      y 2007 04 29 14 34 Beni continued stage scan    While the cleanup of the data is still ongoing  I continued the stage scan at position 170     202103 ahobeamStagepeame electrons                  IL            TA     positions file  StagePos remaining 29 04 07 dat  run stopped after all positions scanned more than once    CALICE eLogbook   Results for search of  20070301 Page 1 of 76             2007 03 13 00 09 Roman Converter Script started    Dear Colleagues    I have restarted the converter script on ilc 1log01    The conversion has successfully started with run 400370 and is now happily   waiting for new runs  The script can be found under  scratch convscript ilcFileCopy pl  To check whether the converter script is running     a  logon to ilc 1og01  afs password   b  ps aux   grep perl    Output should look similar to that    caliceon 7986 0 4 0 5 4320 2604 pts 0 S 00 07 0 02  usr bin perl  w   scratch convscript ilcFileCopy pl    caliceon 9157 0 0 0 0 1616 468 pts 0 S 00 17 0 00 grep perl    Note that these files are not registered automatically to the grid and hence  end up in the desy internal dcache which should speed up the access to these  files from inside desy  I will register them later such they   will be also availble via the grid     Runs up to Run400369 are now being processed in the grid  and should be ready by tomorrow morning     
277. nd HCAL VME     2007 07 07 18 32  30187  peambata   JUNK        lo  330188 d dd  50GeV    DS d        off 20000   evts   off  70 K  P odes                A    e  530 6mm air  Pb  530 6mm air  Pb       CALICE eLogbook   Results for search of  20070701 Page 15 of 24             2007 07 07 18 35 Fab Erika Paul  ECAL slot 5 is now being read out from runs 330188 onwards  2007 07 07 19 22 Francois  Sacha   Jean Yves    Pb    Pb    DAQ rate 80 Hz   10x10 2400 pps   20x20 3200 pps   100x100 out of the beam line       2007 07 07 20 15 Francois  Sacha   Jean Yves    Pb       DAQ rate 65 Hz   10x10 7200 pps   20x20 7700 pps   100x100 out of the beam line    no beam after 90k events    Here s one picture of the drift chambers    CALICE eLogbook   Results for search of  20070701 Page 16 of 24       j  itil    pitied    Track Xo IU Lach     DT   i              M      PERE     i  3      if  E  E   j  4    T    LL NE          ii       EERE    ET         ATrackiz SS ES       E wd   E x    bai E         a  egen  anm  Hri     TERMS               File  http   ttfinfo desy de CALICEelog data 2007 27 07 07 2007 07 07T20 15 33 07 eps    2007 07 07 21 15 Roman Marcel changed power cables of layer 15 and 17  Changed power cables of layers 15  15 7  and 17  17 0    This reduced the noise in both of them  plus the one of layer one  15 0  to a normal level     2007 07 07 21 26 Francois  Sacha     Jean Yves  3d picture of one event of run 330192      CALICE eLogbook   Results for search of  20070701 Pag
278. ne monitoring   Try adding an absorber to remove it  Run 330172 stopped    Added 4 mm lead absorber  Started run 330173   Electron contamination disappeared  We are happy with the settings now     2007 07 07 10 24 Cristina Michal EvBeram tuning  zo GeV pions       CALICE eLogbook   Results for search of  20070701 Page 12 of 24    for 50    GeV pions       ECAL position shifted compared with HCAL     30mm in x   35mm in y    ECAL layers 1 2  15  17 noisy   large nb of hits  high energy for first two layers  no energy for  1S9  317     Notice the triangular shape of the HCAL shower barycentre   BCAL  iscan   porer  tinker  Patet Gesten  Cresent   Suis   Sera    Recp  ca     Peseta Derrojectora      hEcalYX      H m        mum    dd         AR gum  TUTTI    E M Wee        File  http   ttfinfo desy de CALICEelog data 2007 27 07 07 2007 07 07T 10 24 20 04 jpg       2007 07 07 10 27 Niels AHCAL DAC update    There now is a script available on calice00 to automatically retrieve the latest version of AHCAL  DAC files and updating all internal links on the DAQ system     Usage     log on to calice00 and change directory to   online      execute  updateDACfiles ch  lt yyyymmdd gt    with the tag yyyymmdd being the date by  convention    done       The new files will be stored in the directory   online cfg ahcal yyyymmdd    thus old files are  kept for later reference     2007 07 07 11 05 Cristina Michal Evbeam tuning  40 GeV pions    330174 lbeamData  10x10  40Gev pi  514 4mm air off t
279. ng          2007 07 25 00 30 Fabbri  Kieffer   No beam for us    330938 beamData    8GeV 525 air air 2 56 238K  events    From 00 20 we have no beam  At the control room they are investigating  the reason of this        Run 330938 interrupted to take an ECAL safety run     2007 07 25 01 14 Robert  Riccardo    ahcCmNoise O    o       529908  BhcPmNoise                             femo       shcCmLedVcalibScan AAA  ghcPmLedVcalibScan E    We start a calibration because there is still no beam           2007 07 25 01 40 Robert  Riccardo Beam is back  The ahcGain calibration is skipped  to restart normal data taking  2007 07 25 01 47 Robert  Riccardo  poe Pt je DS  qp  t            ems    events  2007 07 25 01 59 Robert  Riccardo  Guassian Fit for recent electron runs    SUE mean  mips        sigma  mips  sigma mean  3        PLUS combined  pi   and e    Ran                      E GevVj                 eanimips  sima  MES  sigma mean   330949 AA AAA i2                    A 2007 07 25 02 20 Robert  Riccardo Same pedestal shift as in 8GeV e  run    CALICE eLogbook   Results for search of  20070701 Page 10 of 20         hHclPecPerChanrel siet       Ped Parla perspi    Vechta Pacta nme    a E xy hb m B H         File  http   ttfinfo desy de CALICEelog data 2007 30 25 07 2007 07 25T02 20 23 00 ps       2007 07 25 02 47 Robert  Riccardo  pee eee A A  events  2007 07 25 03 55 Robert  Riccardo  pe pesce cur peer uem E   events   amp  2007 07 25 03 56 Robert  Riccardo Sheets for checklist 
280. nnels off     slot 7 fe 0   module 31  4 channels off     slot 7 fe 3   module 24  2 channels off     slot 7 fe 4   module 01  39 channels off     slot 7 fe 6   module 02  29 channels off     slot 7 fe 7   module 30  1 channel off     slot 9 fe 1   module 29  3 channels off     slot 9 fe 3   module 27  completely off     slot 9 fe 4   module 26  1 channel off     slot 9 fe 7   module 23  3 channels off     slot 15 fe 2   module 21  1 channel off     slot 17 fe 1   module 15  2nd half completely off     slot 17 fe 2   module 14  1 channel off     slot 17 fe 3   module 13  2 channels off     slot 17 fe 5   module 12  3 channels off     slot 17 fe 6   module 11  2 channels off     slot 19 fe 0   module 09  1 channel off     slot 19 fe 1   module 08  1 channel off     slot 19 fe 3   module 06  2 channels off     slot 19 fe 5   module 05  4 channels off     slot 19 fe 6   module 04  8 channels off                   CALICE eLogbook   Results for search of  20070601         slot 19 fe 7      module    03  12 channels off      gt  some modules   modules half have to be recovered    gt  the list of dead is channels is not ok  but module 1  amp  2 have a lot of dead channels and  in addition long discharge as well    p s     Page 17 of 19    detailed list of long dischaged and dead sipm channels has been sent to the AHCAL experts at    CERN    Hentries       O   ro    2007 06 30 14 59    1  format   2  for AHCAL  Hold CM    1  for TCMT  M Vcalib PM  1  for PIN  2  for PMI   5 0 AHCAL8 
281. nominal value    330086 lpbeamData 20x20 Combined HCAL TCMT 0 0  voltages 0 5  TCM        330087 PmNoise Combined HCAL TCMT 0 0  voltages 0 5  CMT 40 0     330089 beamData 20x20 Combined nominal 0 0   pore pedem Pao 40 0    330090 beamData 20x20 mu  Combined nominal 0 0  TCMT   330091 beamData 20x20 mu  Combined nominal 0 0   EAS 1  nm    27 5 0       ECAL  first 120s  12000 events  still high noise in layer 0         8     g  CT peitages 0 5     ter 40 0   Ld  g                Muons rates  6 2k and 6 7k resp  for 10 10 and 20 20    2007 07 04 07 30 Anne Marie  HengieaWiehisinift  ECAL  HCAL and TCMT shutdowned at 07 30 am   Beam gone at 7 33am     Good guess        2007 07 04 12 00 Beni LED holdscans  ahcCmLedVcalibScan    ahcPmLedVcalibScan        330093 AO    330094 LLL  EhePmBoldScan   330095 LLL JehecmBo1dScan    2007 07 04 17 08 Fab Michele Change in the voltage of DC2 and DC3    The drift chamber s voltage has been increased to 2850 V  for all three drift chambers    The threshold of the discriminators have also changed  from 200mV     100mV      amp  2007 07 05 04 54 Marcel and Ecal ready to rumble    Roman  Hi     the three new slabs were installed in layers 17 12 bottom     After recabling all channels were alive     Noise situation satisfactory    Details during tomorrow     Since we ve needed to move the Ecal table  the Ecal might need to  be aligned once more  It is now roughly in the center  fine tunig needed     Ecal is ready to take data   Currently the Ecal i
282. nts  2007 07 20 07 50 Mandi  Petr  8Gev electron run  Nicola    330712 Iox10   86e  e 132  Jair  air     12 9818    beam interuption    2007 07 20 08 21 kieffer  Fabbri   Morin       330713 lbeamData   10x10  8GeV e  525 air air 2 99 0 43 235K  events    same configuratin as previous run    2007 07 20 09 35 kieffer  Fabbri   Morin    330714 lbeamData   10x10  10GeV e  520 air air 2 06 0 43 173K  events  6    trigger at 6 6K       2007 07 20 14 19 kieffer  Fabbri   Morin       330722 lbeamData   10x10  20GeV 523 air air 0 76 0 43 130k  Events    trigger rate 9 8K    CALICE eLogbook   Results for search of  20070701 Page 18 of 22       many pion contamination    2007 07 20 16 18 kieffer  Fabbri   Morin    330728 Garbage Run    2007 07 20 16 19 kieffer  Fabbri   Morin   2007 07 20 17 40 kieffer  Fabbri   Morin    2007 07 20 17 54 Oliver Allister Henglectron run    330732 lbeamData  10x10  25GeV e  526 air air 0 5 0  3 92k  Events    330731 Garbage Run       10x10 trigger coincidence  6 5 kpps  DAQ rate  90 Hz    stopped run due to large pion contamination  n electron n pion   1 4  set COL03 and COL08 to    2mm    2007 07 20 18 17 Oliver Allister Henglectron run    330733 lbeamData  10x10 I 25GeV e  526 air air 0 5 0  3 263k  Events       10x10 trigger coincidence  6 0 kpps  DAQ rate  80 Hz    still large pion contribution  but Cherenkov seems to work fine    2007 07 20 19 20 Beni new position  moved to  3 3   10 10 stage position   Singe cota Gel i       HCAL Motor control panel
283. ny Cristina Michal    CALICE eLogbook   Results for search of  20070701 Page 14 of 24       ul MN     ft FP Ef E  Pb    junk RUN    2007 07 07 13 25 Evgeny Cristina MEGAIL re centered on the beam  2cm shift     stage shifted from x 4750 to 4950    2007 07 07 13 29 Evgeny Cristina Mbfh amp leV e  run    330179 lbeamData  10x10    50GeV e  530 6mm air off 250K  Pb evts       DAQ rate 88Hz  HCAL voltage   3V    2007 07 07 13 56 Evgeny Cristina MEQHAL shifted again by 2 cm   beam centered      stage position shifted from 4950 to 5150    2007 07 07 14 11 Evgeny Cristina M amp h amp leV e  run    330180 lbeamData  10x10     50GeV e  530 6mm air off 144K  Pb evt s       DAQ rate 88Hz  HCAL voltage   3V  run stopped since the beam stopped    2007 07 07 14 45 Evgeny Cristina MidiePmNoise run    330181 ahcPmNoisel0x10 off off 530 6mm  Pb  off off 530    ahcPmNoisel0x10 6mm    Pb  ahcCmNoisej10x10 off off 530 6mm   Pb  IE  ahcCmNoisel0x10 off off 530 6mm   Pb  IE    2007 07 07 17 36 Kurt Modification of TCMT DAC file    Modified DAC tcmtl1 07072007 dat DAC file to turn off chip 0 channel 13  value changed from 116 to  255  which seems to be causing a incorrect high MIP energy reading in layer 7  Chip 4  channel 1 which was turn off earlier by mistake has been turned back on  value changed  back from 255 to 70             amp  2007 07 07 18 30 Erika Fab Paul fix of slot 5    ECAL slot 5 has been fixed by disconnecting and reconnecting multiple times the trigger connection  between ECAL a
284. o       mainly pions in beam   10x10 trig rate   3000 Hz   100x100 trig rate   2200 Hz   DAQ rate 42 Hz   low intensity beam due to booster problems       CALICE eLogbook   Results for search of  20070701 Page 6 of 24       2007 07 06 18 52 Francois HCal Temperature s longitudinal profile  These data were taken during the runs 330104 to 330107    HCal Temperature    s longitudinal Profile  ro T 972 om O T4 WTS    Average  26                  12345678291011121314151617181920212223242526272829 303132333435 3637 38   Comments  HCAL Layer    data taken between 21h00 and 21h30 on july  5  2007      sensors T2 to Ts indicated 117 00   C  not shown in the graph and not taken into account for    che average values  for layers 28 and 35      emer Ts indicated also 117 00  C for layers 19 and 31     File  http   ttfinfo desy de CALICEelog data 2007 27 06 07 2007 07 06T18 52 00 02 paf    2007 07 06 19 03 Fab Erika beam back  still tuning    Beam back  Still continue the tuning to obtain a good  wide  muon beam     B30131         eam data    Tio0xi09     so8ev         Ru                FEexbinsd     F3 58v     1                1    Beam file loaded  H6 510   Collimator 1   20  JAW1  20  JAW2   Collimator 2   5  JAW1  5  JAW2   Collimator 5   20  JAW1  20  JAW2   Collimator 6   20  JAW1  20  JAW2   Collimator 8   20  JAW1    15  JAW2   No absorber    012  015  016  0 current  DAQ rate  80 Hz    10x10 coincidence 3600 pps  100x100 13500 pps    wFkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkk
285. o 13 0A  better  but not good   330496  gt  changed trim 06 to 10 0A  still off   330497    change trim06 to 5 0A   much better  but still not perfect  330498    change trim06 to 2 0A   how far can one be off    330499    Now trim06 is changed to 0 0A    2007 07 15 21 26 George  Mangi  electron run  Nanda    Pb    Pb    10x10 trigger rate  2 0 k pps  DAQ rate  60 Hz       again beam off in the DC  330501    change trim06 from 0 0A to  3 0A    2007 07 15 22 22 George  Mangi  electron run  Nanda    beamData  10x1    p                10x10 trigger rate  1 4 k pps  DAQ rate  50 Hz                again beam off in the DC   330503    change trim06 from  3 0A to  8 0A   330504    oops  overdone     Back to  5 0A   330505    wrong direction    gt  change to  3 0A   330506  gt  much better  but still  lcm off in ECAL   gt  set trim06 to  6 0A    2007 07 15 23 21 George  Mandi   Nanda    Pb  Pb    10x10 trigger rate  1 4 k pps  DAQ rate  50 Hz       330508  gt again  have to tune trim06  change from 15 to 12A    2007 07 16 00 45 Hakan Oliver Valep amp  40GeV  330509 beamData  10x10 I 40GeV i  514 4mm Air N A  3  3 10k  eee qum i  qe   ee ee EEN ee d  330510 beamData  10x10  40GeV pi  514 4mm Air N A  3   3 5k  Eee AAA AE uu eo cae d  330511 beamData  10x10 I A40GeV pi  514 4mm Air N A  3   3 200k  AAA AE A que qp cu fems    DCs not aligned  changing TRIMO06 current from  6 7A to  0 7A   330510  DCs still not aligned  Changing current to  3 7A   330511  DCs now aligned           CALICE 
286. o Erika and Fab  these are also my two pences for the 930h  meeting to which I won t attend today     Addendum  I just saw that the first lcio file created in the  lyon based chain  came  back successfully from the grid     2007 07 10 07 33 Kurt  Wenbiao  ahcGain  Trygve   2007 07 10 07 33 Kurt  Wenbiao  BeamData  Trygve    Amm i 0 2016   0 43  Pb    E    pi i  beamData  10x10  40GeV pi  514 4mm air 0 2016 0  Pb    problem          2007 07 10 09 10 Erika calibration run   330287 ahcPmLedVcalibScan no light in modules 19   12  344 25 2T    230288     J ahcrmed  1000   Jj          230289  ahccmed         e1000                     30291  j ahcembedVcalibScan TT   2007 07 10 09 11 Beni  Michal  e runs   Cristina    DAQ rate 28 Hz       330292 run stopped   beam off    2007 07 10 11 08 change wobbling settings  change wobbling settings from  80 to  10 GeV    2007 07 10 12 27 Beni  Michal  calib run  Cristina  30302     hcGain  oL    30301 ahcCmLedVcalibScan  30300 ahcPmLedVcalibScan         2007 07 10 13 16 Beni  Michal     CALICE eLogbook   Results for search of  20070701 Page 10 of 25          Cristina  330303 lbeamData  10x10  10Gev pi  533 air air off 0  3 beam  tunning  2007 07 10 13 27 Beni  Michal  electron run  Cristina   230304  beamData  10x10     10GeV  Ki      533  it Rit fott  0 3  180K    2007 07 10 14 18 Beni stage movement    new position is  0 0   EH    HCAL Motor control panel    actual position target position 4  x  00 D GOXI  y  om p coyif    O lost X reference
287. o the latest of the startUps     If everything is OK  you should see the following    CrcReadout  ctor   PCI card 0 crate Oxac slot 5 found alive  CrcReadout  ctor   PCI card 0 crate Oxac slot 7 found alive  CrcReadout  ctor   PCI card 0 crate Oxac slot 9 found alive  CrcReadout  ctor   PCI card 0 crate Oxac slot 12 found alive  CrcReadout  ctor   PCI card 0 crate Oxac slot 15 found alive  CrcReadout  ctor   PCI card 0 crate Oxac slot 17 found alive  CrcReadout  ctor   PCI card 0 crate Oxac slot 19 found alive    BmlCaen767Readout  ctor   PCI card 0  VME base address 0xcae00000 found alive  BmlCaen767Readout  ctor   PCI card 0  VME base address 0xcael0000 found alive  2007 07 08 00 44 Alex  Tao  change collimator settings  Nanda  Fab    since the rate was very high we changed the collimator settings   collimatorl reference changed to  5 1  J1   5 1  J2   Now we have  10000 particles per spill in the 10x10 trigger    2007 07 09 00 34 Kurt  Tao  Nanda pion fune     330235 lbeamData  10x10  100GeV   pi  516 250 k  ES Events        llk pps in 10x10 trigger coincidence  Average data taking rate   90 Hz    2007 07 09 01 23 Kurt  Tao  Nanda gain run    330236 JaheGain TT  2007 07 09 01 24 Kurt  Tao  Nanda noise run    230237   ehePmNoise TOK events    2007 07 09 01 49 Kurt  Tao  Nanda pion run    ES        llk pps in 10x10 trigger coincidence  Average data taking rate   90 Hz    2007 07 09 02 40 pion run       330239 lbeamData  10x10   50GeV pi  9135 air 4mm Off 250k  Pb Events     
288. obbid scan   230029  BeamMoldScan i0cmxX10cm__  muon combined                  tx yv    10cmf0                         Average event rate  60Hz    Scaler  H6B_EXPTO1 6 0E 03  10cmX10cm trigger   H6B EXPTO2 6 7E 03  20cmX20cm trigger   H6B EXPTO3 3 6E 03    2007 07 01 16 11 HV increase    Increased HV for 100x100 muon counter in the back of the TCMT from 2100 to 2500 V     2007 07 01 18 46 stage shift  Shifted movable stage back to x 0 position and vertically  up 10 cm  to  Heaven      2007 07 01 18 50 Erika change of AHC cgf    found a mistake in AHC cgf  modules 1 and 2 positions were inverted                   corrected back  module 2 is not anymore noisy  Marius was right    wrong DAC settings      new AHC cfg  5 0 AHCAL8 38 38 36 38 34 32 35 200 28000 31000  5 3 AHCAL8 37 37 40 37 33 32 35 59 28000 31000  5 4 AHCAL8 35 36 39 35 32 32 34 59 28000 31000  5 5 AHCAL8 34 35 37 34 34 32 34 200 28000 31000  5 6 AHCAL8 33 34 28 33 32 34 35 200 28000 31000  5 7 AHCAL8 32 33 35 32 32 34 35 200 28000 31000  1   7 0 AHCAL8 31 32 34 31 31 34 35 200 28000 31000  9 0 AHCAL8 36310 0 99150 0 00  7 3 AHCAL 24300 0 O 15000000  7 4 AHCAL 2 290 0 O 15 00 0000 swapped  7 6 AHCAL 1 280 0 O 15 00 0000 swapped  7 7 AHCAL 30 27 23 30 27 32 34 59 28140 31000  1    9 1 AHCAL 29 26 26 29 26 32 34 59 28140 31000  9 2 AHCAL 28 25 10 28 25 32 35 59 42500 46000  9 3 AHCAL 27 24 38 27 24 32 34 59 27900 31000  9 4 AHCAL 26 23 22 26 23 32 33 59 28080 31000  9 6 AHCAL 25 22 16 25 22 32 31 59 41800 46000  9 7
289. of  20070701 Page 4 of 20       I moved module 1 to slot 12 7  Unfortunately the CMB doesn t work on this module  Therfore  we have  to wait for beam data before we can be sure     2007 07 12 16 25 Nanda  Tao  start up  Marie Cruz    Searched and closed area H6a and H6b   Power up ECAL  HCAL  TCMT  amp  CMBs   Expect beam back in 5 minutes     2007 07 12 16 59 Erika DAQ problems    Communication to ECAL stage control laptop failed again     Removed it from the DAQ socketing    2007 07 12 17 02 Marie Cruz  Tao  calibration runs  Nanda   2330362   ahcPmLedVqalibScan             o                   J JJ  J          2330363 jahcCmLedVdalibScan                                             J JJ     ZI      2330362  jJanccain                                       l  L     2330365   ahcemNoisqd e 20000                                    Jg    J         330366   JehePmNoisq e 20000        0 E 0 p _  _       330367    fahcCmledVdalibScan      0 E 0 p  p _   p    330368   JshePmbLedVdalibScan    0 p  0    d    d        Still no beam  so checking CMBs with LED light  found dead CMBs in slot FE module  15 3 20  17 5 12  5 5 34  7 4 2  12 7 1       2007 07 12 19 54 Marie Cruz  Tao  beam tuning  Nanda    WW ak    B NM mM       NM 8    1800 pps in 10x10 trigger coincidence  DAQ rate  75 Hz    Finally beam is back   Started beam tuning          2007 07 12 20 29 Fab Marcel Romakcal Stage Moved    THe Ecal stage has been moved using  local control  to  49 5 cm  i e  centered on the beam     Fab 
290. og data 2007 28 1 1 07 2007 07 11T20 41  01 00 jpg        2007 07 11 20 51 Beni bug in automatic pindiode mapping to database  Detected an error in the automatic datbase writing of the PIN mapping  I believe  that this is fixed  now  Together with the cleanup of the AHC cfg PIN mapping errors yesterday the actual   values should be reliable     2007 07 11 21 06 problematic CMB    Module 1   CMB 14  Pulser Settings  Anel DAC  Width WE  Ae Veslib  Width  St  J52     Jael DAC  Width  351  352  pNidth DAC  c5 D     fel Veslib  Width DRC po Gelee    y Enable LEDs  9012345 672825 101    rfmplitude DAC  ly A  500 y    n ESCH K   121217 000 cB   317 000   tO 1317 000     TS 1117 000 Y    Tupper S   3 833  C Wi upper LED   gt  496 y  L   e manual RE INIT  Y calib_a UT   p 001 V WE upper LED   7 455 v    rOIP Switch SC    feel at   MOM LED  7 5 v   Ia vm u P  30 0  REF 1 299y  5 235 v   20 018 V 12348678  WR upper LES  7 432 V Wcal tbat 051   0 000 y    PREL to elos         File  http   ttfinfo desy de CALICEelog data 2007 28 1 1 07 2007 07 11T21  06 02 00 jpg    2007 07 11 21 09 problematic CMB    CALICE eLogbook   Results for search of  20070701 Page 20 of 25    Module 20   CMB 31     Pulser Settings  F el DC  M   DC cum e    po se  Ji  Vealib  Width  351  J52  M    Jael DAC  a  351  352  Width DAC    Ke  fee  Vealib  Width DRC   y Enable LEDS  0123456789101       UENRA O FXI 3827 96  Mip7   C   ln ae    Tapper H   7 522  e Mie upper LED   7467  y  A  a manual RE INIT  V calib at UT 
291. ogbook   Results for search of  20070701 Page 3 of 19       playing  beam    35K events accumulated                2007 07 20 16 01 kieffer  Fabbri   Morin  330726 Calibration  ahcCmLedVcalibScan  330727 Calibration  ahcPmLedVcalibScan        2007 07 20 16 29 Beni change in DaqConfiguration hh    There was a wrong setting for the number of configurations in the ahcPmLedVcalibScan  Number of  configurations is now 32 instead of 65    2007 07 21 00 14 Mandi  Petr   Nicola    330742 lbeamData  10x10  12Gev pi  524 4mm air 2 08 Spo 230k  Pb Events       continue from 330741 run due to lost beam    2007 07 21 01 05 Mangi  Petr   Nicola       330743 lbeamData  10x10  10Gev pi  520 4mm air 2 5815  37 ts 250k  Pb Events    DAQ rate  85 Hz    Some electron component in ECAL energy Spectrum  Try thick absorber next run    2007 07 21 02 04 Mandi  Petr   Nicola    Pb       2007 07 21 03 40 Manai  Petr  calibration runs  Nicola   2007 07 21 03 42 Mangi  Petr  calibration runs  Nicola   2007 07 21 03 49 Mandi  Petr  calibration runs  Nicola   2007 07 21 03 57 Mangi  Petr  calibration runs  Nicola   2007 07 21 04 24 Mandi  Petr   Nicola    330750 lbeamData  10x10  8GeV e  525 air air 2 9816  3   3 250k  Events       DAQ rate  97 Hz    2007 07 21 04 31 Mangi Summary of Current Electron Runs    Here is a summary of all the current electron run with statistic    100 k is given     07 07 2007    CALICE eLogbook   Results for search of  20070701 Page 4 of 19       Layer 2 keeps noisy  low hit ra
292. on   400570 CmLed to see lpe peak   400572 CmLed to see lpe peak around strip 15   For Layer23 stripl5    400577 CmLed for inter calibration   400578 PmLed for inter calibration   400580 CmLed to see lpe peak   400582 CmLed to see lpe peak around strip 15   For Layer22 stripl5    400586 CmLed for inter calibration   400587 PmLed for inter calibration   400590 CmLed to see lpe peak   400592 CmLed to see lpe peak around strip 15   For Layerl4 stripl5    400594 CmLed for inter calibration   400595 PmLed for inter calibration    available until 33th point    CALICE eLogbook   Results for search of  20070301 Page 74 of 76       400597     CmLed to see lpe peak  400599     CmLed to see lpe peak around strip 15     For Layerl5 stripl5     400602     CmLed for inter calibration   400603     PmLed for inter calibration   400605     CmLed to see lpe peak   400607     CmLed to see lpe peak around strip 15     For Layerl6 stripl5        400609     CmLed for inter calibration   400610     PmLed for inter calibration   400612     CmLed to see lpe peak   400614     CmLed to see lpe peak around strip 15     For Layerl7 stripl5        400618   CmLed for inter calibration   400619     PmLed for inter calibration   400621     CmLed to see lpe peak   400623     CmLed to see lpe peak around strip 15        For Layerl8 stripl5           400625   CmLed for inter calibration   400626     PmLed for inter calibration   400628     CmLed to see lpe peak   400629     Same with 400628 but with  1V increas
293. ons     penable   disable PLL channels Stee RE 0 eet    sw   Lom   um  j enable f disable          S     gt  Zeg Hi values to file dump FLL values  position   File  http   ttfinto desy de CALICEelog data 2007 29 19 07 2007 07 19T22 43 02 00 jpg                      CALICE eLogbook   Results for search of  20070701 Page 17 of 22       2007 07 19 23 32 Oliver Allister Hengne       330696 fbeamData   10x10  12Gev pi  524 8mm Pb air 2 10 O   3 200k  Events    10x10 trigger coincidence  1 0 kpps  DAQ rate  54 Hz    A 2007 07 19 23 37 Oliver Allister HenBIMET readout power supply failure    One channel of the PMT test board readout LV power failed   We replaced the power supply   Nevertheless  the readout shows no resonable signal     2007 07 20 02 08 Mangi  Petr  12 GeV pion run  Nicola    330702 lbeamData   10x10  12GeV pi  524 8mm Pb air 2 10 0 43 80k  Events    Continued from run 330696       2007 07 20 02 30 Mandi  Petr   Nicola       330703 lbeamData   10x10  15GeV pi  521 8mm Pb air 1 3552  0  3 250k  Events    change colimator 2 3 8   6 6  triger rate 1 5k    2007 07 20 03 39 Mandi  Petr  18 GeV pion run  Nicola    330704 lbeamData   10x10  18GeV pi  522 8mm Pb air 0 9421  0  3 250k  Events    change colimators 2   4 4  and 3   6 6   triger rate 2 6k       2007 07 20 04 35 Mandi  Petr  20 GeV pion run  Nicola       330705 lbeamData   10x10  20GeV pi  523 8mm Pb air 0 7607  0  3 250k  Events    triger rate 1 7k       2007 07 20 05 36 Mangi  Petr  25 GeV pion run  Nicola  Eve
294. ot aligned  Now 4 3A    330519   Aligned now        2007 07 16 03 49 Hakan Oliver ValeRecision test    WW  Pb    M    Pb    ECAL successfully aligned        CALICE eLogbook   Results for search of  20070701 Page 10 of 19       ECAL Iren   vow   tancias   Paco   Coaupescgl Cresscheci   Sees   Samos    Response   renger CoPranctons         er  1911464      ais A ER          20              File  http   ttfinfo desy de CALICEelog data 2007 29 16 07 2007 07 16T03 49 19 03 jpg    2007 07 16 03 58 new position     10  10  corresponds to the ECAL center    HCAL Motor control panel    actual position target position    710 00  ho  xi   10 06 FT sec    O lost X reference point    Q lost Y reference point       O lost 2 reference point    O we in mrin  O eroi eration set Pos     fea Y Fast 1 gulsesfree   extern STOP   ott   690 2  880 000       r Offset  CES Hi zm PANEL to dea     File  http   ttfinfo desy de CALICEelog data 2007 29 16 07 2007 07 16T03 58 02 00 jpg        2007 07 16 04 47 Beni changed to muons    Following actions were performed to switch to muon data taking mode     raise  amp  connect front 100x100 scintillator   change AHC cfg link to AHC_CERN07_beamhold_100x100 cfg   change in EmcConfiguration hh sequenceDelay to 1   new beamsettings  file 555  absorber  amp  target   air    adjusted collimator settings to 10k events in 100x100 coincidence COLLO2 COLLO3   4 4   COLLO5  COLLO6   20  20     left beamdump closed    2007 07 16 05 06 Beni Valeria muon run data quality  
295. ots 17 and 19 monitor  YA 00 06  Oliver Allister Hengne  shot 1     noisy ECAL slots 17 and 19 monitor  shot 1              28 of 28 27 09 2007 11 39    CALICE eLogbook   Results for search of  20070701 Page 1 of 19    2007 07 20 00 53 Mandi  Petr  ahcCMNoise Run  Nicola  EEE SEC E  2007 07 20 00 55 Mandi  Petr   Nicola  beam lost  linac cooling problem  2007 07 20 00 56 Mandi  Petr  ahcPMNoise  Nicola   330698   a EFEC E  2007 07 20 00 59 Mangi  Petr  ahcCMLedVScan run  Nicola  30699 ENS EEES ESE E  2007 07 20 01 05 Mandi  Petr  ahcPMLedVScan run  Nicola   330700   BhcePmLedVealibScan o  2007 07 20 01 22 Mandi  Petr  ahcGain run  Nicola   330701 EEES E  2007 07 20 06 26 Mandi  Petr  ahcCmNoise  Nicola   330707 Perras TI     2007 07 20 06 29 Mangi  Petr  ahcPmNoise  Nicola   330708 Peris TI     2007 07 20 06 32 Mandi  Petr  ahcCMLedVScan run  Nicola   230709 SS   BhecmbedVcalibScan o E  2007 07 20 06 39 Mandi  Petr  ahcPmLedVScan run  Nicola   230710 SS ehcPmedvcalibScan LI OE  2007 07 20 07 00 Mandi  Petr  ahcGain  Nicola  FEIER faci    2007 07 20 08 37 Morin  Fabbri Noise in layer 19 of the Ecal  The layer 19 of the Ecal is much more noisy than yesterday morning   The ratio of the number of its hits with the neighbouring channels ones  oscillates sizably           celis hit       H H 10 15 20      ayer number  File  http   ttfinfo desy de CALICEelog data 2007 29 20 07 2007 07 20T08 37 49 00 png    2007 07 20 10 05 kieffer  Fabbri     CALICE eLogbook   Results for search of 
296. p e p O PS DE D  send   Events    Pb Itransitio  Events  AAA A A E PA qe   Pb Events    2007 07 11 01 44 Kurt Wenbiao Tryd3eam Data    330335 lbeamData  10x10  25GeV e  526 air air 0 4715 40k  Events          330336 beamData  10x10  25GeV e  526 6mm air 0 4715 12k  Pb Events  moved in  absorber  to  increase  e  pi   ratio    CALICE eLogbook   Results for search of  20070701 Page 15 of 25       330337 beamData  10x10  25GeV 0 4715 180k  Events    removed absorber e  beam highly contamianated by pions   5pions lelectron   lost beam for about 10minutes HCAL slow control output 34 58 started to flash red          EODEM 0208 _KurtWenbiao Trygeam Data  Po pee Er CS eS  p       23     fair     Bir Jo  7eo7_    garbage    EE ge p  Iw que m Ce EE  events  e  beam contaminated by pions   1 to 1 ratio   2007 07 11 03 48 Kurt Beam Data  AAA MENGE GEN beens   events  e  beam contaminated by pions   1 to 1 ratio   2007 07 11 04 36 Kurt Beam Data    330342 lbeamData  10x10  15GeV e  521 air air in 10k  pr pp Pp E esah beer     330343 lbeamData  10x10  15GeV 1 3613 200k  AAA AE AA aC       e  beam contaminated by pions but better than previous runs 2 e  to 1 pion        2007 07 11 05 23 Kurt Beam Data  330344 lbeamData  10x10  12GeV e  524 air air in 25k  pa pon peee po pr Dt send    dens    330345 lbeamData  10x10  12GeV 2 1100 182k  ARA  APP dee  r     e  beam contaminated by pions but better than previous runs 6 e  to 1 pion        2007 07 11 06 03 Kurt Beam Data  330346 lbeamData  10x10
297. pasd  n  settings  were  wrong in  E cfg          330405 ahcPmLedHgs gSct  n  sync off    CALICE eLogbook   Results for search of  20070701 Page 11 of 20       2007 07 13 16 02 Marie Cruz  Tao  pion run    330406 Kee ee    Nanda       DAQ rate  50 Hz     1000 pps in 10x10 trigger coincidence  but not very stable    stage at 60 0  beam gone again  need another SPS access   We will take some calibration data in the mean time     2007 07 13 16 55 Marie Cruz  Tao  calibration runs  Nanda    330407 EncPmLedveal psean      J     T Y    L               L                 330408 BhcCmLedVcal bScan                  LL         TL        L            L                  330409 ieheGain o poo po  oo  LLL   330410 ehcemNoise LI ITT A       la 2007 07 13 18 15 PaulD Beam line data    The CERN beam line data have been changed this year and now no longer  include the  RP  values  which I think stands for  Radiation Protection     This is why the DAQ data reader was not able to correctly parse the char  string sent  Sven and I were not told about this before now       have removed the section of code which interprets these data and the  interaction now seems fine  The RP values will now all be set to zero   will leave this enabled in the DAQ        n case there are further problems  then to remove the beam data readout  from the DAQ  comment the line       define BML SLOW DATA    Gets beam information  in src runnerDefine icc on calice00 and recompile the DAQ  i e        build runner    2007 07
298. per  W Crate  N Wi Crate OFF  recover  KILL Wl M1 1 wa   7   wio    RS SM    dep M values to file dup AL values  position     See    e   el  See   ee   ee   ee    ee   Sc   LEX  E  see   Se   See   e   ee   fae    fee    el       I3  8 8    i          File  http   ttfinfo desy de CALICEelog data 2007 34 22 08 2007 08 22T02 25 04 00 jpg    2007 08 22 03 20 Satoru Hakan  Allister  350388    Waiting for beam to come back        2007 08 22 03 32 Satoru Hakan Allister    350389 beamData  10x10  20GeV e  523 air air 0 5 105k  30 degs    2007 08 22 03 53 Satoru Hakan Allister  350390 beamData  10x10  12GeV pi  524 8mm air 4 5 101k 30  OT PTT pep I Emo E  P  Be    350391 beamData  10x10  8GeV loi  525 8mm air 1 5 0 0 k 30  AE ger qo aee qus ve A A TES      2007 08 22 04 47 Satoru Hakan Allister    B pp          ETUR eee ERO pm He MERC BRUM  em Ppa pep    2007 08 22 08 00 end of 2007 running   CERN  THE END       8 ttt       the last shift says goodbye    see you all again   FERMILAB    CALICE eLogbook   Results for search of  20070801 Page 21 of 22       File  http   ttfinfo desy de CALICEelog data 2007 34 22 08 2007 08 22T09 05 26 01 jpg    2007 08 22 08 42 current   10V       mn MITA 07  Ww Sjt5A  o 1098 0 Sp        w 310A F IRSA   o w moe 1 F axs  0 FIMA W WIA  4 wie W aja  p    0 XA    ane    LR es dead    1 MiplA   0 MW 298    o w    ox 392      o re    Cm eae    fH settings   file actions enable   disable iLL channels    9t   umb   UNT      ele disable             File  http   
299. plate because 1 prepared plate  was drilled wrong last year  So additional to the holes for the current SiPMS all holes  for the mirrors have to be drilled too     cassettes work in progress  30          29  28  cassettes ready  24   27  1   23  status of production OScw07 cassette 098a b  ng MAL FLO  module assemb  ng steps staff  1 2 3 4 5 D 7 8 9     base frame ties SONS FRA plate pads cabing fibres  tempsensors cover plate no of persons       amour ic de we    Deed F1  210207    File  http   ttfinfo desy de CALICEelog data 2007 08 21  02 2007 02 21T 16 59 48 00 pdf     4 2007 02 26 19 43 Marius ASIC positions before coherent noise update   This gives a list of the HAB position before exchanging them with the modified HABs to reduce the  coherent noise    HAB production number from the label of the HAB vs module and chip number    CALICE eLogbook   Results for search of  20070201 Page 5 of 21          File  http   ttfinfo desy de CALICEelog data doc CERN docu 2007 02 26T19 43 35 00 pdf    la 2007 02 26 19 50 Marius Holdscan results for august and october    Holdscan results for august and october vs slot fe or module    CALICE eLogbook   Results for search of  20070201 Page 6 of 21       File  http   ttfinfo desy de CALICEelog data doc CERN docu 2007 02 26T19 50 15 00 pdf    la 2007 02 26 20 19 Marius AHCAL readout  DAQ  map 8 power connection map    CALICE eLogbook   Results for search of  20070201 Page 7 of 21          File  http   ttfinfo desy de CALICEelog data doc CER
300. points entire detector    400474  81 points entire detector    400475  81 points entire detector    400477  30 points  precise scan for X layer strip 5   400478  30 points  precise scan for X layer strip 5   400479  30 points  precise scan for X layer strip 5   400480  30 points  precise scan for X layer strip 5   400481  30 points  precise scan for X layer strip 5   400482  30 points  precise scan for X layer strip 5   400483  30 points  precise scan for X layer strip 5  stopped at 7th point   400484  81 points entire detector    400485  81 points entire detector    400486  81 points entire detector    400487  81 points entire detector    400488  81 points entire detector    400491  30 points  precise scan for X layer strip 5              end ert precise scan ee 9e    2nd configuration energy and position scan                     CALICE eLogbook   Results for search of  20070301 Page 73 of 76          400492  400493  1GeV 36 9 10 points    400495  400497  2GeV 36 9 10 points    400499  400500  3GeV 36 9 10 points    400502  4GeV 36 9 10 points    400508  4GeV 36 9 5 points    400512  4GeV 90k central injection    400509  400510  5GeV 36 9 10 points    400504  400505  400506  6GeV 36 9410 points    400513  3GeV 3x10 points precise scan for X layer strip 5    400514  3GeV 5x20 points precise scan for X layer strip 5 and 14   400515  3GeV 5x20 points precise scan for X layer strip 5 and 14   400516  3GeV 5x20 points precise scan for X layer strip 5 and 14   400517  3GeV 5x20 po
301. possibly replaced during next access    2007 07 28 21 05 Erika Roman Beni1Mek0_pkas    Pb    ECAL tripped because of high current    2007 07 28 21 21 Beni Alex Lukas 10x10 pi        Pb     ECAL tripping again   CMBs were off  since when is unknown    Stopped because of technical difficulties after  113k ev     CALICE eLogbook   Results for search of  20070701 Page 10 of 16       2007 07 28 22 00 Beni stage angle measurement    The picture shows the angles and the measured distances to calculate the angle     results   in degree        Following formulas were used to calculate the angles   tan  alpha     b a  d  tan beta    c d    10 deg position      b a    19 4    0 4 cm  e   19 4    0 2 cm  d   110 0    0 2 cm    20 deg position           a  16 7    0 2 cm  b  56 1    0 2 cm  gc 39 5     0 2  cm  d   110 0    0 2 cm  File  http   ttfinfo desy de CALICEelog data 2007 30 28 07 2007 07 29T00 37 32 00 gif  2007 07 28 22 33 Beni Alex Lukas  2007 07 28 22 58 Beni Alex Lukas 10x10 pi        Pb    Stopped afer 133k because of lost beam     2007 07 28 23 48 Alex Lukas       2007 07 29 04 39 Roman  Nanda  e  pi  run  beamData fo pe e  pi  pe BRE Air   20deg    peambata   oxi0 fee       7pi  023 Rir ER zdeg       331086  331087  331088       CALICE eLogbook   Results for search of  20070701 Page 11 of 16       ES mm  rr    0x10 trigger rate  8 8kpps  Un rate  75Hz  Split beam run in 250 kevt    NOTE  Good data seems to be run 331088    Allister    2007 07 29 13 34 Beni  Martin  junk r
302. pped at 7th point   400484  81 points entire detector    400485  81 points entire detector    400486  81 points entire detector    400487  81 points entire detector    400488  81 points entire detector    400491  30 points  precise scan for X layer strip 5     end of precise scan                              2nd configuration energy and position scan                   400492  400493  1GeV 36 9 10 points   400495  400497  2GeV 36 9 10 points     CALICE eLogbook   Results for search of  20070301 Page 21 of 76       400499  400500  3GeV 36 9 10 points   400502  4GeV 36 9 10 points    400508  4GeV 36 9 5 points    400512  4GeV 90k central injection   400509  400510  5GeV 36 9 10 points   400504  400505  400506  6GeV 36 9 10 points   4   4   4   4   4       00513  3GeV 3x10 points precise scan for X layer strip 5   00514  3GeV 5x20 points precise scan for X layer strip 5 and 14   00515  3GeV 5x20 points precise scan for X layer strip 5 and 14   00516  3GeV 5x20 points precise scan for X layer strip 5 and 14   00517  3GeV 5x20 points precise scan for X layer strip 5 and 14   end of 2nd configuration energy   position scan Een    2nd configuration MIP calibration after removing tungsten                  400519  400520  400521  400522  400523  4  4  4  4       00524  00525  00526  00527  end of MIP calibration                              Precise scan with 2nd configuration without tungsten                   400529     entire detector scan  9x9 points    400530     entire detector sc
303. pps  100x100 rate 16700 pps    CALICE eLogbook   Results for search of  20070701 Page 9 of 24       100k events    2007 07 06 22 27 Francois Sacha Jean Yves   230148    ahcPmNoisq100x100     80GeV    m       5910         fir      fair ote    DAQ rate  92 Hz   10x10 rate 8800 pps  20x20 rate 11500 pps  100x100 rate 16700 pps    23k events    no run 4330147   What we ve done   after run330146       click  stop run  button       change run type to ahcPmNoise      click on  start run     e Ck Ck Ck Ck CK CK CK CK CK CK CK CK CK CC CC CC CK CK CK CK CC CC CC CK CC CC CC CK CC CC CC CC CC CC CC CC CC CC CK CK CK CK CC CK CC CK CK CK CK CK CK Ck kk kc kckckckckck    You were running a  sequence  of runs of type beamData which would continue   indefinitely unless interrupted  Hitting  stop run  stops the run but not the   Sequence  for the sequence  use the sequenceEnd command  Hence  the DAQ went   to the next run in the sequence  which was 330147  Then  clicking  start run    stopped this run almost immediately and started yet another  330148  See  Log1183744318 0ut icalice00   for details  Just doing the  start run  is all that is needed     Run 330147 contains just the runStart and runEnd records and so is junk     If you want the previous run to finish at the right time  e g  after the  Specified number of events  rather than immediately  then use the non interrupt  feature when you  start run  the run to follow it  This is the  i option of   the command runStart which should al
304. quite stably           2007 07 24 06 42 Allister  Robert  calibration data  Riccardo  330892   JaheGain     A  330893      JhePmLedVcalibScan           330894 SE TEE ESE                   2007 07 24 07 50 Allister  Robert   Riccardo  330897 beamData  10x10  25 e  526 air air 0 42 138K  events  2007 07 24 08 24 Allister  Robert   Riccardo    check  new beam  position    beamData  20GeV check  new beam  position                   check  new beam  position       330900 beamData  20GeV check  new beam  position    ui pepe E ai    not for physics  after moving to 0 0 we check the beamspot    CALICE eLogbook   Results for search of  20070701 Page 4 of 20    330902  peambata  0x10 Peceve  P3 gt      Bir pir Bo 70  _     beam is out in X  we changed trim6 0   gt   2       330903 fpeamData Jr0x10 pec po E3 bir bir po po   1    ie to que in ECAL RE    this corresponds to  10  10  in stage coordinates          File  http   ttfinto desy de CALICEelog data 2007 30 24 07 2007 07 24T10 09 02 00 jpg       ECAL position  0 0     CALICE eLogbook   Results for search of  20070701 Page 5 of 20         hEcal Y X         E 10 Hd 14 16    x  cm     UEM L   DR  EEN    E 10 Hd H fon    E d   tt 14 415    y  om     File  http   ttfinfo desy de CALICEelog data 2007 30 24 07 2007 07 24T 10 36 02 00 gif    2007 07 24 10 48 Beni  Petr    Laurent    closed x acceptance  optimized Quad 9 amp 10    optimized Quad 11612   13614   15816    removed Quad  de optimization 15 amp 16       RESULT    old values for 
305. r 2 5V LED  run to see  lpe peak  2007 03 21 21 15 S  Uozumi LED run  400562 CmLed LED KNU fiber 2 5V LED  run to see  lpe peak  2007 03 21 21 16 S  Uozumi LED run  400564 CmLed LED KNU fiber 2 5V LED  run to see  lpe peak  2007 03 21 21 16 S  Uozumi LED run  400564 CmLed LED KNU fiber 2 5V LED  run to see  lpe peak  2007 03 21 21 16 S  Uozumi LED run  500567 CmLed LED KNU fiber 25V for  inter calibr  tion  2007 03 21 21 16 S  Uozumi LED run    500567 CmLed LED KNU fiber 2 5V for  inter calibr  tion    CALICE eLogbook   Results for search of  20070301 Page 30 of 76       2007 03 21 21 17 S  Uozumi LED run  400568 PmLed LED KNU fiber 25V for  Pp O A AL ss  2007 03 21 21 17 S  Uozumi LED run  400568 PmLed LED KNU fiber 2 5V for  ob E TI ER E a  2007 03 21 21 19 S  Uozumi LED run  400570 CmLed LED KNU fiber 2 5V LED  lpe peak  2007 03 21 21 19 S  Uozumi LED run  400570 CmLed LED KNU fiber 2 5V LED  lpe peak  2007 03 21 21 25 S  Uozumi LED run  400572 CmLed LED KNU fiber 2 5V LED  lpe peak  2007 03 21 21 25 S  Uozumi LED run  400572 CmLed LED KNU fiber 2 5V LED  lpe peak  2007 03 21 21 26 S  Uozumi LED run  ehr eee cre  inter calibr  tion  2007 03 21 21 26 S  Uozumi LED run  E A A LPRA ee  inter calibr  tion  2007 03 21 21 26 S  Uozumi LED run  400578 PmLed LED KNU fiber 2 5V for  eee A Ap E LE LE o lintez catibreion  2007 03 21 21 26 S  Uozumi LED run  400578 PmLed LED KNU fiber 2 5V for  pem amer quie o AO quee rs  2007 03 21 21 27 S  Uozumi LED run  400580 CmLed LED KNU fiber 
306. r configurations    From these two files one can deduce how the trigger was configured  at which time and for which run    I could have made it a bit more sophisticated but it should be ok      The first one is attached to this entry  the other one i will attach to the  next entry           Cheers   Roman   File  http   ttfinfo desy de CALICEelog data 2007 12 21 03 2007 03 21T00 40 55 00 txt  EF 2007 03 21 00 50 Roman     and the second one    here we go    CALICE eLogbook   Results for search of  20070301 Page 27 of 76       File  http   ttfinfo desy de CALICEelog data 2007 12 21 03 2007 03 21T00 50 00 00 txt    2007 03 21 08 29 S  Uozumi overnight MIP calibration runs    Runs between 400543 400550 are overnight MIP calibration run with 3rd configuration   Run 400551 is stopped at 34th point  but data until 33th points are usable     2007 03 21 18 29 S  Uozumi et al  Thanks     Thanks  Roman     2007 03 21 18 30 S  Uozumi et al     The DESY II is down from this morning to afternoon  just running LED runs       2007 03 21 19 41 saori linearity and resolution of 2nd configuration  direct fiber  2nd configuration   preliminary results of linearity and energy resolution    Linearity looks good   Energy resolution is 12 43    It has approached the G4 simulation result  12 3       CALICE eLogbook   Results for search of  20070301 Page 28 of 76          File  http   ttfinfo desy de CALICEelog data 2007 12 21 03 2007 03 21T19 41 10 00 paf    2007 03 21 21 11 S  Uozumi LED run   4005
307. r number  File  http   ttfinfo desy de CALICEelog data 2007 34 21 08 2007 08 21T07 53 39 00 png  2007 08 21 08 49 Jamie  Pavel  e     Beni  beamData  10x10 H15GeV e  625 air 6mm  6 0 110k 30  pum  ppp fe D DU Bs  Pb degs  Pb degs  lbeamData  10x10 H30GeV e  621 air 3mm 0 30  6 0 100k 30  ape pop pp fe Pf   ke  lbeamData  10x10  40GeV e  620 air 3mm 0 10  6 0 100k 30  pepe qux SN GEET  lbeamData  10x10 H50GeV e  624 air 3mm 0 04  6 0 107k 30  popa AAA    2007 08 21 13 31 Jamie  Pavel  pi   Beni    P    PS shutdown for machine development at 15h30                        6 0 12k 30  e ME   6  0 100k 30  INE  T6 0 165k 30  IINE EE  T6 0 50k 30  SE       2007 08 21 16 39 Daniel  calibration runs    Anne Marie  Petr    350358 ahcGain setting up manual steering of  module 2    8503602 een                 CELL   550361                     LL BhePmNoise SS  550387                                  BheemvealibSeam         MU  550363                     LL bheemVealibSean SSS OOOO SOOSOSOS    SO       2007 08 21 20 01 Daniel  pi  runs  Anne Marie  Petr    350364 lbeamData  10x10 H50GeV  oit 612 4mm air 0 04  6 0 85k 30  Pb degs       CALICE eLogbook   Results for search of  20070801 Page 18 of 22       350365 junk    Pb  Pb    e       Pb    2007 08 21 21 39 caliceonOcaliceheescGtage position   CALICE user        Ld l A    HCAL Motor control panel      actual position target position    ELO  x Les Yoo mx  Y Tas 10 55 e    O lost X reference point    Qo lost Y reference point      lost 2 r
308. reased  over voltage    LED run    KNU fiber   same  with  previous run  but with  increased  over voltage          2007 03 23 19 00 S  Uozumi       2007 03 23 19 00 S  Uozumi LED run    CmLed KNU fiber       2007 03 23 19 00 S  Uozumi LED run    CmLed KNU fiber    2007 03 23 19 01    CmLed KNU fiber 25V    2007 03 23 19 01 S  Uozumi LED run    CmLed KNU fiber 2 5V    2007 03 23 19 02 LED run    CmLed KNU fiber 3 5V  with  previous run  but with  increased  lover voltage               2007 03 23 19 02 S  Uozumi LED run    400739 CmLed LED KNU fiber 35V same  with       CALICE eLogbook   Results for search of  20070301 Page 49 of 76       previous run  but with  increased  Over voltage  2007 03 23 19 03 SC Uozumi LED run  400740 CmLed KNU fiber  see lpe  peak at  L19s15  2007 03 23 19 03 A Uozumi LED run    400740 CmLed KNU fiber  see lpe  peak at  L19s15    2007 03 23 19 03 S  Uozumi LED run    KNU fiber 3 5V same  with  previous run  but with  increased  over voltage    2007 03 23 19 03 S  Uozumi LED run    LED KNU fiber 3 5V same  with  previous run  but with  increased    2007 03 23 19 04 S  Uozumi LED run    l NI i MN    2007 03 23 19 04 S  Uozumi LED run    l    i EN    2007 03 23 19 05 3  Uozumi LED run    400743 CmLed KNU fiber  see lpe  peak at  L20515    2007 03 23 19 05 ES Uozumi LED run    400743 CmLed KNU fiber  see lpe  peak at  L20s15                         2007 03 23 19 05 S  Uozumi LED run    CALICE eLogbook   Results for search of  20070301 Page 50 of 76    
309. rectly compare  light yield between fiber readout and extrusion modules   DAC files have suffix of LYcompare     2007 03 25 19 45 S  Uozumi MIP run    400796 beam T1L 3GeV IKNU fiber A scanned 9x9  precise  detector  scan           2007 03 25 19 45 S  Uozumi    400796 beam T1L 3GeV IKNU fiber 4 scanned 9x9  precise  detector  scan    2007 03 25 23 14 S  Uozumi tuning MIP selection and fit for calibration constant   ve tried to optimize MIP selection and fitting to obtain calibration constant   n addition to the Daniel s cuts  I apply following additional cuts     T1R T2R  have signal consistent with 1 MIP  V1U  V1D  V2D have signal consistent with pedestal  Other strips in a same layer with the targer strip have signal consistent with pedestal    Latter 2 cuts reduce upper tails of MIP distribution  which may come from shower events    Such large upper tail by shower events are significant for downstream layers   Even without tungsten  positron sometimes occurs small EM shower in MIP runs      One note for MIP events    Even with such MIP selection  there are some events consistent with pedestal    Those events are observed especially at the first part layers  layer 1 2 3       think it may happen since for those layers we can not require signal on upper stream  layers    f we want to kill those pedestal MIP events  I think tracking information is necessary   For now I limit the fitting region to avoid such events        Then I tried to use several types of functions to find 
310. renkov works very well  We observe 98  efficiency electron detection     Logbook entry   CALICEelog data 2007 30 23 07    23 07 2007  ay    PIT ee Erika HCAL linearity  pi     17 10   interesting the residual from linearity when normalizing to the 8 GeV point  Without  correction   for saturation effects in SiPM the non linearity of the HCAL response has a max at 25  GeV of 10               1    75 100  beam E GeV                   75 100  beam E GeV       Logbook entry   CALICEelog data 2007 30 23 07  z 23 07 2007  Si 14 07    330864 lahcCmNoi se  e 20000  330865  ahcPmNoi se  e 20000    Beni  Petr   Laurent    EES  psose7janccain              Logbook entry   CALICEelog data 2007 30 23 07    23 07 2007  ze       oro e Oliver stability of HCAL response  14 02  I plotted the mean of the gaussian fit of hHcalEnergy for pion runs against the run  number    i e  time   The error bars are actually the gaussian sigma  The corresponding plot for  the ECAL       can be found in the elog on the 20 07 2007  The plot has been done using the code made  by Allister for the ECAL  thank you      5 of 28 27 09 2007 11 39    CALICE eLogbook    6 of 28    http   ttfinfo desy de elog results jsp docroot  export web htdocs  zlog        Stability of HCAL Response    N  a  o  o       2  E       c  D  o  z     u       Run Number    Logbook entry   CALICEelog data 2007 30 23 07    23 07 2007  13 54       Beni  Petr   Laurent    330863 beamData 10x10  8GeV 525 air air 2 556 SS  70K Events          tr
311. ror bars are actually the gaussian sigma     CALICE eLogbook   Results for search of  20070701 Page 21 of 22       Stability of ECAL Response    Fit Mean  mips        200 300 400 500 600 700  Run Number    File  http   ttfinfo desy de CALICEelog data 2007 29 20 07 2007 07 20T23 26 20 00 png          2007 07 20 23 32 Oliver Allister HenB GAL Response    The plot shows the response of the HCAL in MIPs for pions beams of different energies  The HCAL  response is given by the mean of a gaussian fit of hEcalEnergy spectrum  The error bars  indicate the standard deviation of this fit     At the moment there are four different scans for different days  ranging from July 13th 2007 to  July 19th 2007  displayed in the plot     HCAL response    2400  2200         070719          070717   2000         070716  1800 E  ors  1600  1400  1200  1000  800  600  400  200  0    visible Energy   MIP       0 20 40 60 80 100  beam energy   GeV    File  http   ttfinfo desy de CALICEelog data 2007 29 20 07 2007 07 20T23 32 32 01 png    2007 07 20 23 52 Oliver Allister Henpien run    330741 beamData  10x10 I 12GeV pi  524 4mm air 2 08 bon F3 90k  Pb Events    10x10 trigger coincidence  3 8 kpps  DAQ rate  75 Hz       2007 07 20 23 54 Oliver Allister HenGuaBbialaFiefbkelegtron and pion runs    gaussian fits for hHcalEnergy for pion runs  and for hEcalEnergy for electron runs    pions ll  Run No Energy  GeV  ES ei sigma mip        330737  330738       CALICE eLogbook   Results for search of  20070701 P
312. rresponds to the switch on of the   detector for data taking  Before this  the HCAL had some time to cool down  After this    it was practically on continously     overall temperature development       27    26 5  26  25 5  25  24 5    24          23 5       05 07 06 07 07 07 08 07 09 07 10 07    time  File  http   ttfinfo desy de CALICEelog data 2007 28 11 07 2007 07 11T11 53 46 00 gif    2007 07 11 12 00 ECAL off    CALICE eLogbook   Results for search of  20070701 Page 17 of 25       due to Paul exercise with the DAQ we turn the ECAL off    2007 07 11 12 02 Beni temperature development in Geneva    Genf  Daten aus Genf  Schweiz       Terperatur C C3  28    26  24    22  20  15  16  14    to  S    1          04 07  06 07  07 07  08 07  09 07  14 07  11 07    w  S    Legende      Durchechnitt     Maximum     Minimum    Zeitraum  04 07 2007   11 07 2007       File  http   ttfinfo desy de CALICEelog data 2007 28 11 07 2007 07 11T12 02 30 00 png    2007 07 11 14 40 Beni reversed fibere swap  optical fibres are back to standard configuration    2007 07 11 16 25 Beni DAQ optical cable    The wrongly routed fibre cables were re routed  Now the HCAL absorber plates shift should be  possible     2007 07 11 20 14 Nanda calibration runs    330356 JahcCmNoisd e 50000  330357 lahcPmNoisd e 50000       30359  330380 fancGain  __     high current in module 25b HV channel 50  469 uA    Eo  o o    LL          p uoa qe AAA  330358    ahcCmVcalibScan Eo  o o    TENIS  Eoo  oo oo  SL S  A     
313. run  but with  increased  lover voltage          2007 03 23 19 11 S  Uozumi LED run    CmLed KNU fiber 3 5V same  with  previous run  but with  increased    2007 03 23 19 11 LED run    CmLed KNU fiber 2 5V    LED run          2007 03 23 19 11                   i o MEE  2007 03 23 19 12 S  Uozumi pedestal run  400756 CmNoise pedestal KNU fiber 25V CmNoise  with LED on  2007 03 23 19 12 S  Uozumi pedestal run  400756 CmNoise pedestal KNU fiber SEN CmNoise  with LED on  2007 03 23 19 12 S  Uozumi pedestal run  400757 PmNoise pedestal KNU fiber 2 5V PmNoise  with LED on  2007 03 23 19 12 S  Uozumi pedestal run  400757 PmNoise pedestal KNU fiber 2 5V PmNoise  with LED on  2007 03 23 19 13 S  Uozumi pedestal run    400759 PmNoise pedestal KNU fiber 25V CmNoise  with LED  off       2007 03 23 19 13 S  Uozumi pedestal run    400759 CmNoise pedestal KNU fiber ERE CmNoise  with LED  off    2007 03 23 19 14 S  Uozumi LED run    400760 CmNoise pedestal KNU fiber 2 5V CmNoise  with LED  Off          CALICE eLogbook   Results for search of  20070301 Page 53 of 76          2007 03 23 19 14 S  Uozumi LED run  with LED  Off  2007 03 23 19 15 S  Uozumi 5 GeV run  pq ue em  9 10  Scan   2007 03 23 19 15 S  Uozumi 5 GeV run    400765 beam TIL 5Gev KNU fiber  2 5V scanned 36  points  scan    with positron beam  DAQ rate   40Hz    2007 03 23 21 23 S  Uozumi 3 GeV run    KNU fiber E  scanned    detector  scan          with positron beam  DAQ rate is  120Hz           2007 03 23 21 23 3 GeV run    KN
314. run option  e 10000  v 197  random Trigger     test run option  e 10000  v 197  random Trigger                2007 07 22 17 16 caliceon 2calicehd iph  sureeniin Ch 28   CALICE user     current in Ch 28 increased by approx  500uA during the last half an hour     sonitoriog Switch On   or cossents  Voltage  V  Current  uA Set  V   Initial  V      RE A 40 7 40 2     uie   IERCH    LEI  doe      2  ic g  n d x  10 00    10 00  10 00  10 00  10 00    10 00    i   8  8  8  3 8      10 00    10 00    10 00    BDONDONSSUSOSUIHEGOSNOGON    10 00    Trend to elos  3 e Y    10 00    i    10 00  HY settings   File actions erable   disable ALL charnels    n Sw   wow   EMT     ele   _ dile         10 00    10 00  Experts    rere rrr  W Crate N W Crate OFF  recover  KILL Wl A J wa 1 WwI j wis j    recover all trips  3 J we J m3 J LY otan   dee W values to File   _ de REL values  position       10 00    10 00    10 00    10 00          15 lalala lalalalala lalala    10 00          File  http   ttfinfo desy de CALICEelog data 2007 29 22 07 2007 07 22T17 16 01 00 jpg          2007 07 22 19 52 caliceonOcalicehdalrr  sstegnctirrent in ch 26   CALICE user     The current of channel 26 has increased during the last two hours by approx  3uA  It is now beyond  limit     CALICE eLogbook   Results for search of  20070701 Page 16 of 19    VARAS nu A     SCH      41 80    10 00   mje Bian auazen     IE 10 00 M  ee ffl os o een      10 00 Rest D               ERP zi pordle   disable AL chwnels    se   wm
315. s  Sacha     Jean Yves    Pb       10x10 6300 pps  20x20 6600 pps  100x100 out of the beam line    2007 07 07 22 37 Francois  Sasha   Jean Yves    Pb       10x10 4600 pps  20x20 4800 pps  100x100 out of the beam line    2007 07 07 22 42 Francois Sacha Jean Yves    CALICE eLogbook   Results for search of  20070701 Page 19 of 24       run 330194  shower shape of an electron in ECAL    Time  19 49 45 297 225 Sat Jul 7 2007    Run 3301940 Event 1160 Hits  427 Energy  6352 97 mips    e     e   e e    99     e       49999  e   4     mg     e     EE  e         d  13   3      ti  H      o  01      n  o           ege g  s             File  http   ttfinfo desy de CALICEelog data 2007 27 07 07 2007 07 07T22 42 28 00 eps    2007 07 07 23 34 Francois  Sasha   Jean Yves       Pb    10x10 10700 pps   20x20 10900 pps   100x100 out of the beam line  stopped after 254k events    2007 07 08 00 33 Alex  Tao  Nanda pion run       Pb    10x10 trigger coincidence   12000 events per spill  Data taking rate   70 Hz   stage at 0 0 position   1x1 out of beamline   no cerenkov available    220k events    2007 07 08 01 23 Alex  Tao  Nanda no beam  stopped run 330198 after 220k events  because beam is down  beam should be back in a few minutes    2007 07 08 01 31 Alex  Tao  Nanda electron run    Pb    10x10 trigger coincidence  3800 particles per spill  data taking rate   80 Hz   stage at 0 0 position   checked CRCs and found all alive for AHCAL and ECAL    stopped after  250k events    CALICE eLogbook   
316. s 24 26 28 30 installed  Schaetzel    mae TEL as       Trigger  3x3cm2 new from Petr Smirnov  HV  CAEN channels 20 and 21 3000V each  PMT voltage is 1500V   LVDS lines 10 and 11 on VME backplane  use  v 26 for coincidence   there is no hardware  NIM  coincidence     hold from run 202364  ahcBeamHoldScan  v 26     Voltages and currents as of 13 April 2007  13h30 using  DACvalue mod24 A68 4 B75 2 txt   DACvalue mod26 A37 9 B40 8 txt  DACvalue mod28 A82 2 B77 5 txt  DACvalue mod30 A79 9 B78 3 txt    CALICE eLogbook   Results for search of  20070401 Page 2 of 14       Vision  1  1_Maln_HCAL pni E Sry x    He  OG   Movement   Bid  QOL peral             W ieit    High Voltage  ali Me ON   oll WY OFF    monitoring Switoh On   Off  Weltage  V  Current  A  Set  V  Destin  V     10 00   Ree    10 00 Feel  1 10 00 e   CECR  10 00 Feel    12 00 Seel  10 00 Seel                10 00 Ree  La b  de 208 77 27 syed a j  Bote to KII             File  http   ttfinfo desy de CALICEelog data 2007 15 12 04 2007 04 12T16 19 33 04 png    2007 04 13 14 30 Sebastian MIP calibration  Schaetzel    50000 events per positions   run stopped at position x 105 y  15 on Saturday  14 April at 12h40   beam was gone  shutter closed  presumably around 23h48 when Linac failed   Linac Freigabe gone      cfg StagePositions dat   3 rounds of 216 positions     no LED light  CMBs off    2007 04 14 13 09 Sebastian MIP calibration modules 24 26 28 30  cont d   Schaetzel    like 202365 but with cfg StagePositions202366 da
317. s amp log       z 23 07 2007      20 56 Nicola  Hengne  Mangi       EE E EISE  O E ZE    Trigger rate  2 2 K    From the drift chambers we observe that the alignement is not symmetrical around 0    we  stop the   run  The collimator settings was WRONG    this run is NOT good for analysis    Logbook entry   CALICEelog data 2007 30 23 07    20 06 Nicola  Hengne  Mangi          EE    collimator set up changed in order to improve the trigger rate   collimator 1 5 6 10 11    6 mm  collimator 2 3 8   4 mm    Cherenkov efficiency 99     The beam looks not perfectly centered  A shift of  60 in the y  and  76 in the x    Logbook entry   CALICEelog data 2007 30 23 07  2 23 07 2007         19 54 Nicola  Hengne  Manqi    A En An        JUNK RUN        Logbook entry   CALICEelog data 2007 30 23 07      23 07 2007  D 19 26    During temperature analysis I realized that the HCAL slowcontrol data records in the  binary do not  match the data displayed in the HCAL slowcontrol for several modules   Additionally  for module 20 the records is missing  The values from module 24 are  displayed twice in the record for module 18  amp  24  The record of module 20 contains  the values of module 18  Details are shown in the table below        slow readout mapping error    Comparing the values displayed in the slowcontrol with the records in the binary file  I  found follwing  situation     D  E  D  a  o  B  pan  D   lt    0   B  B  D  u  3  D  ct  Q  D    reads data from module CMB  1    LA Ry RY  db  
318. s switched off     Marcel and Roman    2007 07 05 08 56 Bernard    ECAL stand position tuned   correspondance with front plate of HCAL done in a more reliable way than before    i e  could be reproduced with the same precision by different operators     centered  within  0 5mm   center position set to X 4750  as before   defined as beam X 0 position   HW range is from 5627  JURA side  to 3292  SALEVE side    SW range will be set to 5620  3600  current  5700  3600   the HW limit on the JURA side is 1mm close to the  mobile  stage hitting the ECAL stage control  box    m 2007 07 05 11 36 Beni CRC repair    Matt fixed the broken ADC in FE 2 on CRC SEROO7  This gives us one spare FE in slot 5     CALICE eLogbook   Results for search of  20070701 Page 21 of 22       A 2007 07 05 12 08 Beni HCAL HV supply    We have some lose connections on the HCAL HV power supply  Careful adjustment of the connectors  is necessary        2007 07 05 12 11 Beni HCAL stage    During Svens test this morning we lost the Y reference  I checked afterwards and found the driver  of the Y motor in error mode  status 5   Most probably this caused the loss of the reference     We performed a reference movment  Y is referenced again and at the upper Y limit  I will wait till  the motor cooled down before I move the stage to 0 0 again     2007 07 05 15 49 Erika    30096 AhcPmLedHoldScan no light in 17 0 17 5 15 4 7 6  30097 AhcPmLedHoldScan increased light    hopefully with enough light on all modules         
319. scan    CALICE eLogbook   Results for search of  20070301 Page 39 of 76       2007 03 22 15 18 S  Uozumi 3 GeV run  400655 beam T1L 3GeV KNU fiber  2 5V scanned 36  9 10  pa A AA TAE  2007 03 22 15 30 S  Uozumi 3 GeV run  400655 beam T1L 3GeV KNU fiber  2 5V scanned 36  9 10  i i cE E  ac aS EES  2007 03 22 15 30 S  Uozumi 3 GeV run  pp wc TE  9 10  2007 03 22 16 43 S  Uozumi 3 GeV run  400656 beam T1L 3GeV KNU fiber  2 5V scanned 36  9 10  AE AS PE EHNEN  2007 03 22 16 43 S  Uozumi 3 GeV run  400656 beam TIL 3GeV KNU fiber  2 5V scanned 36  9 10  pq AA AAA  2007 03 22 17 36 S  Uozumi 4 GeV run  pq ue gr qe e 4p dee  l  9 10  2007 03 22 17 36 S  Uozumi 4 GeV run    400657 beam T1L AGeV KNU fiber  2 5V scanned 36  9 10  scan    Stopped at 45th points  so central injection is missed     2007 03 22 23 07 S  Uozumi    After 7 pm there are only positron beams in DESY II   I called to control room several times to have electron beam  but they said there are some problems  on acclerator and after solving the problem they will work on HERA and DORIS injection     So  we may not have the beam overnight         2007 03 22 23 14 S  Uozumi    Called again to control room after their shift change   and they could give the 7GeV electron beam now  but due to linac problem  intensity  number of  bunches   is not high   They will give us the high intensity beam as soon as the linac problem is solved     2007 03 23 07 31 S  Uozumi 3 GeV run    Eventually during almost of last night we had just
320. scanned 36  calibration direct points  scan for  IP  calibration       CALICE eLogbook   Results for search of  20070301 Page 8 of 76       2007 03 14 23 47 S  Uozumi MIP calibration runs    2007 03 14 23 56  Runs between 400453   400457 are MIP calibration run without tungsten   Run 400458 is useless     2007 03 14 23 59 S  Uozumi precise detector scan    400459 MIP run T1L 3GeV fiber 2 5V scanned precise  direct scan  9x9    2007 03 14 23 59 S  Uozumi precise detector scan    400459 MIP run T1L 3GeV fiber 2 5V scanned precise  direct scan  9x9           amp  2007 03 15 00 03 Roman Poeschl Stage positions in database    Dear Colleagues    due to a bug in the steering file the stage positions were not automatically filled  up to run 400450  This has been corrected and for runs  gt  400450 they should be  immediately available     Runs smaller than this 400450 are currently processed in the grid where for most of the  runs the db is already updated     As an example I have read out the db for run 400368  and found     Moving stage position x mm   34 8  Moving stage position y mm   783 2  The jobs should be finished by tomorrow morning  Try to access the data for any desired run    and contact me in case of problems     Sorry for the inconvenience        Cheers   Roman  2007 03 15 07 53 S  Uozumi precise detector scan  GE EUN LN Laud LANE Ge  precise scan points   2007 03 15 07 53 x Uozumi precise detector scan       precise scan  9x9 points     2007 03 15 10 09 S  Uozumi preci
321. scrambled numbers  all zeros   all zeros   all zeros  scrambled numbers  all zeros  scrambled numbers  all zeros   wrong values    wrong values       A EE  Bi zeros Cd  PA EE  BB S zeros CY  58 EE  B EE  WA perembled numbers A   8            S es Cd   88 ong values                         LL        CALICE eLogbook   Results for search of  20070401 Page 1 of 14       2007 04 12 09 31 Doerte David assembling status    The tiles for the last modules 37   38 are on the way to DESY   Tiles for the other modules 33  34 and 36 are ready to be tested at DESY     In the assembling of the cassettes we have a problem  The type of the micro coax cable from CERN  which we use for the SiPM signals was changed  Now it is difficult to adjust the stripping  machine for the new cable  One reason is that we still do not have a data sheet at the   moment  Therefore we cannot finish the modules but we prepare them till the step of cabling     cassettes work in progress  35       32  cassettes ready  1     31  status of production  15cw 07 Cassette amar bing HAL   FLC  module assembing steps staff  1 2 3 4 5 LJ 7 8 9   S Boso frome tios   SR _FR4 piste ges cabing____fibres  tempsensors cover plate no of persons             K   TT   oir purgas   NENNEN    sous  Senet ds vt Q Daid F1  12 04 07    File  http   ttfinfo desy de CALICEelog data 2007 15 12 04 2007 04 12T09 31 48 00 paf    2007 04 12 12 01 Doerte David  The tiles for the last modules are already at DESY   2007 04 12 16 19 Sebastian Module
322. scratch software i386 extern root lib libCore so  1  5 Oxf7ac7ddd in TUnixSystem  StackTrace     from  scratch software i386 extern root lib libCore so  1  6 Oxf7ac5e26 in TUnixSystem  DispatchSignals     from  scratch software i386 extern root lib libCore so  1  7 Oxf7ac3e4d in SigHandler     from  scratch software i386 extern root lib libCore so  1  8 Oxf7aca345 in sighandler     from  scratch software i386 extern root lib libCore so  1  9   signal handler called    2  10 Oxf5fefafl in TFitEditor  UpdateGUI     from  scratch software i386 extern root lib libFitPanel so  1  11 0xf60031f5 in GGFitPanel 270 0 15          from  scratch software i386 extern root lib libFitPanel so  1  12 Oxf734da8b in Cint  G__CallFunc  Execute     from  scratch software i386 extern root lib libCint so       1  13 0xf79ac504 in TQConnection  ExecuteMethod     from  scratch software i386 extern root lib libCore so  1  14 0xf79b0233 in TQObject  Emit     from  scratch software i386 extern root lib libCore so  1  15 Oxf6d0b94e in TPad  RangeAxisChanged    from  scratch software i386 extern root lib libGpad so  1  16 Oxf6d38d1f in TPad  RangeAxis     from  scratch software i386 extern root lib libGpad so  1  17 0xf604a08c in THistPainter  RecalculateRange     from  scratch software i386 extern root lib libHistPainter so  1  18 0xf603c45c in THistPainter  PaintFrame     from  scratch software i386 extern root lib libHistPainter so  1  19 0xf6034abd in THistPainter  Paint     from  scratch software i3
323. se detector scan    400462 MIP run TIG 3GeV fiber 2 5V scanned precise  direct scan  9x9    2007 03 15 10 09 S  Uozumi precise detector scan    400462 MIP run T1L 3GeV fiber 25V scanned precise  direct scan  9x9          2007 03 15 14 05 S  Uozumi 1 strip precise scan    400464 MIP run TIL 3GeV fiber 265V scanned precise  direct scan of  strip 5   of X  layers   strip   3x10points       CALICE eLogbook   Results for search of  20070301 Page 9 of 76       2007 03 15 14 05 S  Uozumi 1 strip precise scan    400464 MIP run bg 3GeV fiber 25V scanned precise  direct scan of  one strip   3x10points       2007 03 15 17 10 S  Uozumi et al  configuration changed  We have switched from the 1st configuration  fiber direct combination  to the 2nd configuration   direct fiber    Then proceed to the comissioning run to check dead channels       2007 03 15 18 34 S  Uozumi comissioning run for 2nd configuration       MIP run fiber 2 5V scanned entire  direct detector  scan for  comissioninp  2007 03 15 18 34 S  Uozumi comissioning run for 2nd configuration    MIP run fiber 2 5V scanned entire  direct detector  scan for  comissioninp            S  Uozumi       2007 03 16 09 12 calibration run without tungsten    direct 2 5V scanned   fiber i i  tungsten   36    2007 03 16 09 12 calibration run without tungsten    400468  beam TLG 3GeV direct 2 5V scanned  fiber i i  tungsten   36  points        2007 03 16 09 12 S  Uozumi MIP calibration runs for 2nd configuration    Runs 400468 400471 are MIP 
324. se voltage   80V   400752     CmLed to see lpe peak around strip 15    For Layer23 stripl5                       400753     CmLed to see lpe peak  400754     Same with 400753 but with  1V increased base voltage   80V   400755     CmLed to see lpe peak around strip 15    Pedestal runs    400756 CmNoise with LED on  400757 PmNoise with LED on  400759 PmNoise with LED off  400760 CmNoise with LED off       end of LED runs           a a a a    3rd configuration precise scan with tungsten                    400658  9x9 points   400659  9x9 points   400767  9x9 points   400768  9x9 points        CALICE eLogbook   Results for search of  20070301 Page 76 of 76       400769  9x9 points   400770  9x9 points   400777  5x20 points                 400778  5x20 points     5x20 points                 400779  400780  5x20 points  400781  5x20 points    end of precise scan                                        i 2030220 iioii    MIP calibration after tungsten removal with 3rd configuration                      400784   400785   400786   end of MIP calibration                                  precise scan after tungsten removal with 3rd configuration  400788  3x10 one strip precise scan    400789  3x10 one strip precise scan    400790  3x10 one strip precise scan    400791  3x10 one strip precise scan    400792  3x10 one strip precise scan    400793  3x10 one strip precise scan    400794  3x10 one strip precise scan    400796  9x9 precise entire detector scan    400797  9x9 precise entire det
325. search of  20070701 Page 13 of 20              Electron 10 GeV  old and new    ECAL GeVs    File  http   ttfinfo desy de CALICEelog data 2007 30 25 07 2007 07 25T14 59 40 00 paf    2007 07 25 15 02 Vasiliy evidence of ECAL noise    the ECAL noisy slabs contribute to the small spikes on the right of the 10 GeV peak in the spectrum   but cannot explain the overall shift between the new and old run  More coherent noise             tego um        ud Larisa  150 200 250 300 350       Number of hits    LJ       Electron 10 GeV old and new       N    ECAL HITs    CALICE eLogbook   Results for search of  20070701 Page 14 of 20       File  http   ttfinfo desy de CALICEelog data 2007 30 25 07 2007 07 25T15 02 04 00 paf    2007 07 25 15 09 Vasiliy  this shift is visible already at the MIP scale   also note the shift at the MIP peak is smaller than the shift at the 10 GeV peak  i e  shift  proportional  to energy  not offset     MIPs    Electron 10 GeV old and new    ECAL MIPs       200 400 600 800 1000 1200 1400 1600 1800          File  http   ttfinfo desy de CALICEelog data 2007 30 25 07 2007 07 25T 15 09 43 00 pdf     4 2007 07 25 20 20 Beni Erika HCAL expert report  CMB work    during machine day we have opened the CMB side of the HCAL and      reinserted CMB in layer position 10  module 12  slot fe 17 5 which now works fine     fixed better connections of all CMBs  some found not fully screwed in       extracted CMB in layer 18  module 20  slot fe 15 3 for inspection and try to fix 
326. sed    2007 03 23 18 56 S  Uozumi LED run    CmLed KNU fiber    2007 03 23 18 56             see lpe  peak in  neighboring  strip of  L8515  but  pedestal too    2007 03 23 18 56 LED run    a  2007 03 23 18 56    2007 03 23 18 57 S  Uozumi LED run    CmLed KNU fiber S same  with  previous run  but with  increased  lover voltage          2007 03 23 18 57 S  Uozumi LED run    KNU fiber   same  with  previous run  but with  increased  lover voltage       CALICE eLogbook   Results for search of  20070301 Page 47 of 76       2007 03 23 18 57 S  Uozumi LED run    400731 CmLed LED KNU fiber 2 5V to   see lpe  peak in  neighboring  strip of  L15515  but  pedestal too  wide         2007 03 23 18 57    ul   13    2007 03 23 18 58 LED run    E  2007 03 23 18 58    2007 03 23 18 58 LED run    CmLed KNU fiber z   with  previous run  but with  increased  over voltage   2007 03 23 18 58   KNU fiber 3 same   with  previous run  but with  increased  lover voltage   2007 03 23 18 59   400734 CmLed LED KNU fiber 25V    but    pedestal too  wide     2007 03 23 18 59 S  Uozumi    400734 CmLed LED KNU fiber 25V to  see lpe  peak in  neighboring  strip of  L16515                2007 03 23 19 00 S  Uozumi LED run       400735 CmLed LED KNU fiber 2 5V to  see lpe    CALICE eLogbook   Results for search of  20070301 Page 48 of 76    peak at  L16515          2007 03 23 19 00 S  Uozumi LED run    iln MENE          2007 03 23 19 00 S  Uozumi LED run    CmLed KNU fiber 7  with  previous run  but with  inc
327. sestsss       Logbook entry   CALICEelog data 2007 30 23 07          23 07 2007 e         PV Tp Allister  Robert  Riccardo  Beni pion run  03 51   330836 beamData 10x10  25GeV pi  526 SC air 0 4115  6 0 250kEvents       10x10 trigger rate  3 6 kpps  DAQ rate  80 Hz    Logbook entry   CALICEelog data 2007 30 23 07    23 07 2007    03 46 Allister  Robert  Riccardo  Beni       330835 beamData 10x10  8GeV e  526  air air 0 4115  6 0 junk  short run for position check              15kEvents  Logbook entry   CALICEelog data 2007 30 23 07  BEN Allister  Robert  Riccardo  Beni position check    27 09 2007 11 39    CALICE eLogbook    12 of 28    http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log       position in ECAL   6 0        Logbook entry   CALICEelog data 2007 30 23 07  23 07 2007  03 29    330834 ahcPmLedVcalibScan    Allister  Robert  Riccardo  Beni calibration data    Logbook entry   CALICEelog data 2007 30 23 07    23 07 2007  03 26       Allister  Robert  Riccardo  Beni position check       E GG 4mm _   short run for position check     Logbook entry   CALICEelog data 2007 30 23 07       23 07 2007 caliceon G calicehcalsc02 cern ch  CALICE  03 17 user     moved to   6 0   in stage coordinates this is   50  10   Actually  the difference in the  stage coordinate   system is 13 cm in respect to the last position  But we used this position before and   for a 12 cm movement  checkrun 330829  we are more at  5 than at  6 in the ECAL  coordinate   system   
328. sh the contact with the stage itself     the  set  light is blinking on the slow control panel    the two lights inn the box on the floor of the control room are not blinking anymore    Since the connection with the DAQ is established  we left enabled the ECAL stage socket in the DAQ  configuration  The position that the program reads is the one in the configuration file   x 4750  y 0            2007 07 28 20 00 Erika Roman Beni Alex Lukas  331064 ahcPmNoise  2007 07 28 20 17 new zero position    We moved the x axis to go back to the ECAL center after turning to 20deg angle     E    a om  o eem iJ    HCAL Motor control panel          actual position target position      00  x  S88 E TT    dE 20 Gei   500       O lost X reference point   O lost Y reference point   O lost 2 reference point   le BE   O manual operation get Pos          feas 1 fast V pulsesfree   extern STOP I refs24   534 314 534 914    OFF set     E 900 H La PING  to elog       File  http   ttfinfo desy de CALICEelog data 2007 30 28 07 2007 07 28T20 17 03 00 jpg       2007 07 28 20 26 Erika Roman BenidMek0_pkas  331065 lbeamData  10x10   20GeV pi  523 4mm Air 0 5 20deg  Pb  eee Pap E E ET E EST  2007 07 28 20 40 Erika Roman BenidWMek0_pkas    Pb     Stopped after 40k ev to reload HCAL DAC settings    TRIM6 adjusted from  1 to  5    2007 07 28 20 51 Erika Roman Beni Alex Lukas  2007 07 28 20 53 Erika  Beni PIN HBAB removed    PIN HBAB 12 2 A removed after damage during rotation   Board is being investigated and 
329. so be available in the run control GUI    or should be added if it isn t      Paul    wkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkkk  Thank you for these explanations       Francois    2007 07 06 22 33 Francois Sacha Jean Yves   230149   ahcGain  100x100  F80cev__ j m       510         Bir         fir      off      DAQ rate  96 Hz  10x10 rate 8800 pps    20x20 rate 11500 pps  100x100 rate 16700 pps    165k events    2007 07 06 23 01 Francois Sacha Jean Yves    Lun       HCal configuration file changed to CERN_030707_ 0 7V txt  DAQ rate  83 Hz   10x10 rate 8800 pps   20x20 rate 11500 pps   100x100 rate 16600 pps    250k events    2007 07 06 23 47 Francois Sacha Jean Yves    A a cda  run       SEHEN cdd eae  run    DAQ rate  99 Hz   10x10 rate 8900 pps  20x20 rate 11700 pps  100x100 rate 17100 pps    250k events for each    CALICE eLogbook   Results for search of  20070701 Page 10 of 24       2007 07 07 01 21 Yoshi Tao Alex         DM    CERN 030707  0 7V txt    au  o w  C   wl c w w  o  o O O  mn mn m mn  g gl mn mn   ot  o gl Ed       2007 07 07 03 15 Yoshi Tao Alex muon run       E          Im  Im  Im  im       B B       CERN  030707  nominalV txt                         Sometimes Magnet QUAD11 status differs a little     0 5  from BeamRef    Average event rate in Run  100Hz    2007 07 07 05 40 Yoshi Tao Alex beamData run    i  k    CERN_030707_ 0 3V txt                                  2007 07 07 08 22 Evgeny Cristina Michal    M GE    
330. sp docroot  export web htdocs amp log          AA ad ah a rh ah    KE   has per event        db s5EBESUSS       Logbook entry   CALICEelog data 2007 30 23 07    23 07 2007  05 03       Allister  Robert  Riccardo    trigger rate  3 3 kpps       Logbook entry   CALICEelog data 2007 30 23 07    23 07 2007  04 58       Allister  Robert  Riccardo       Changed TRIMO6 reference current from  3 333A to  2A and see how the beam position  changes   Correction seems to work fine        Logbook entry   CALICEelog data 2007 30 23 07    23 07 2007  04 47    330837 beamData 10x10  20GeV bi  523         Allister  Robert  Riccardo    mm    b air 0 682  6 0 52kEvents          10x10 trigger rate  3 36 kpps  DAQ rate  84 Hz  stopped to correct the beam position  not centered on DCs     Logbook entry   CALICEelog data 2007 30 23 07       Alister Robert Riccardo Fit result of HCAL pion energy  04 10 spectrum    27 09 2007 11 39    CALICE eLogbook    11 of 28     gr    http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log       Run   E  GeV   mean  mips   sigma  mips   sigma mean      330836 25 616 69 11 2  18             Logbook entry   CALICEelog data 2007 30 23 07    23 07 2007  03 52    After the bad x beam position in the last 8GeV run I wanted to retune the 8GeV file  But  after reloading   the beam file the position is ok  I expect that the effect comes from some magnet  failure    We should keep an eye on this        Beni 8GeV x position check    pssst cesses      sese
331. t       O lost Z reference point   Q stage in motion   O manual operation get Pos    ee  fast  puse free   extern STOP   ref 24   929385 829985       OF set   B 900 NM 1290 PANEL to eLog  File  http   ttfinfo desy de CALICEelog data 2007 29 17 07 2007 07 17T 10 09 01  00 jpg  2007 07 17 12 32 Beni temperature development    Avergae temperature of the HCAL from the first record in each run since 4th of July    CALICE eLogbook   Results for search of  20070701 Page 2 of 22       overall temperature development    28  27  26  25    24    TLT UU UNA AUS AZUL     a  i E          07 07 14 07  time  File  http   ttfinfo desy de CALICEelog data 2007 29 17 07 2007 07 17T12 32 18 00 gif     amp  2007 07 17 12 36 Beni outside temperature  For reference the Geneva temperature for the same time periode   Genf       Daten aus Genf  Schweiz  Terperatur C C3   32   30    25    26    i N    t                07  Ou    D  s         12 07  13 07   4 97  15 07  16 07  17 07    08 07  09 07  10    06 07  07 07       L1  Ss    04  14    Legende      Durchschnitt   Maximum     Minimum  Zeitraum  04 07 2007   17 07 2007    File  http   ttfinfo desy de CALICEelog data 2007 29 17 07 2007 07 17T12 36 08 00 png    2007 07 17 13 16 Allister  Imad  Guassian Fit for pion runs 330558 330582  Beni    HCAL energy gaussian fits for pion runs where the ECAL is centered and the HCAL is 6mm off towards  Saleve     X   z    igna Mean 4   L 222       r3  A  LO        en  en    5 87    JJ 00  VO     ON    7 48    EI 
332. t    EDIT  beam gone in night to Sunday  machine problems  Linac and DESY 2    run stopped on Monday  16 April at 8h00 at position  240  120  only 64 positions scanned     Page 3 of 14    CALICE eLogbook   Results for search of  20070401    ful    guam i aa ti   an  o A Ol   mt md tel Lt       aa  B wa kp ZG 0 SE GG DE a ki EL o E S pP E 0 EE DE BL MH FH EL Ob       D D  i Eu  t L  t   El  k   E  F F    D E  A ES  E F  B Ep  Lion  8 1530 Also   eg Ja  pang    JOKE PERSO sp aly scm 0 Ausg PUT id    CALICE eLogbook   Results for search of  20070401 Page 4 of 14       File  http   ttfinfo desy de CALICEelog data 2007 15 13 04 2007 04 16T08 09 44 02 ps    2007 04 16 10 35 Sebastian increased current in module 28B and 30B  Schaetzel    current  mA  13 April current  mA  16 April   16              ISEG EHO 8605 Detail No           LADO    High Voltage  all Me ON   oll MN OFF                monitoring Switch On   Off cosmata  j Sat  V  bestial  V     65 60   10 00   Re  or Module 244 68 5    75 40    10 00 Ramp  CFF If  eate 20 75    38 20    10 00 e  GFF  EES  10 00 Ree  oF Si esie 2 dun  10 00   Bee  La   Module 300 90 1V  10 00   Bee  a    gen Srat  2 1   10 00   Bee siile 29 82 44  i a fine stunt OFF wawa    10 00 Raw dule 208 77 77 pela veltager  y  E T rA  tem eet ien daa vo cnt    NA    Voltage  V  Current Ia           Experts       File  http   ttfinfo desy de CALICEelog data 2007 16 16 04 2007 04 16T10 35 15 00 png    2007 04 16 10 47 Sebastian MIP calibration modules 24 26
333. t good  enough     An another red image under strip 5 implies the optical light cross talk exists     CALICE eLogbook   Results for search of  20070301 Page 18 of 76       trackYPos       trackXPos    File  http   ttfinfo desy de CALICEelog data 2007 12 19 03 2007 03 19T14 56 23 00 jpg    2007 03 19 15 13 saori etal  Z position    http   azusa shinshu u ac jp  saori desy2007 z position pdf    2007 03 19 15 18 S  Uozumi New version of ntuples  I ve put very preliminary version of new root tree file under the directory   flcscecalana  home anauser newtrees vl 0     Some more information will follow     2007 03 19 16 10 S  Uozumi 3 GeV run    direct 2 5V scanned 5x20  fiber points         2007 03 19 16 10    direct 2 5V scanned  fiber    2007 03 19 17 08 MIP calibration    MIP direct 2  5V scanned  calibration fiber  calibration       CALICE eLogbook   Results for search of  20070301 Page 19 of 76       2007 03 19 17 08 S  Uozumi MIP calibration    MIP TIL 3GeV direct 25V scanned detector  calibration fiber scan  36  points   for MIP  Calibration       2007 03 19 17 08 saori et al  tungsten removed    we removed the tungsten plates from the SCECAL   2007 03 19 18 09 S  Uozumi Run summary  Again I summarize all the useful runs here     lst configuration  fiber direct  lst round                           MIP calibration before putting tungsten  400262          detector scan  36 points     1GeV     400309  point0 35    2GeV 400308  point0 35    3GeV     400305  point0 35    4GeV   
334. t if nhits  x y     4 and the chisq contribution of that hit is    500       assumes the same drift velocity in all chambers  not true  since the applied voltages are not  equal  1 700  1 625  1575  1650 V in chambers 0  gt 3 respectively     all of the above can and should be improved     CALICE eLogbook   Results for search of  20070301 Page 14 of 76       here is a code snippet of how to get the track information from your code     1  include  lt StraightTrackImpl h gt        const std  vector lt std  string gt   collections   evt  gt getCollectionNames      for  int i20  i lt collections  gt size    i       if  collections  gt at  i    driftchamber_track      cout  lt  lt   found track      LCCollection   tracks   evt  gt getCollection  driftchamber_track     if  tracks   getNumberOfElements    1     cout  lt  lt   WARNING  found    lt  lt  tracks   getNumberOfElements    lt  lt   tracks   lt  lt  endl     else    StraightTrackImpl   theTrack    StraightTrackImpl    tracks  gt getElementAt  0    float chisqx   theTrack   getChisqX       chisq for x hits  float chisqy   theTrack   getChisqY       chisq for y hits  float x   theTrack  gt getX       X at calo front face  float y   theTrack  gt getY        Y at calo front face  float xslope   theTrack  gt getXslope       X slope of track  float yslope   theTrack  gt getYslope       Y slope of track  std  hitset lt 4 gt  xhits   theTrack  gt getXhits       hitmap for x layers layer 0  gt 3    std  hitset lt 4 gt  yhits   theTr
335. t seems startSkt was not successful  there   I would guess after the crash a  panic  would be needed on all three PCs   maybe this was not done   In such a case  it is better to log onto calice01 and  calice02 and do the startSkt by hand  as opposed to using the GUI  as the lock file  error is then seen explicitly and it becomes clearer that a panic is needed   Cheers    Paul    2007 07 03 08 30 Erika bug  fixing     one comment to the problem  observation  pointed out by Roman    the DAQ should be started only from the GUI and not manually from the shell  if not otherwise specified  When not starting from the GUI all socketing  has to be started by hand before  I think this was the problem of Roman   Concerning the update of run number I ll have to investigate     2007 07 03 08 45 Erika 20x20 hold scan    0 0    my guess was yesterday evening that the 20x20 comes about 12 ns later w r t  the 10x10 trigger  So I ve set all hold values 2 ticks earlier  would be nice to confirm this trend at least for few  tiles  thank you     2007 07 03 09 52 Francois Eugeny Yesbn run   230052 fbeambata  20x20 fru combined    _____p0___ _______      start muon run    2007 07 03 10 26 Beni Francois Eugengl  asiri   230053    beambata    20x20     mu  combined            Wi         150K events  run terminated because DAQ computer hanging     For safety  the ECAL is switched off  DAQ computer needs to be rebooted     2007 07 03 10 44 Beni Francois Eugealy    selnv reboot     10 15 caliceserv  PC 
336. tDown  amp  startUp    2007 07 18 14 40 Beni CMB check run   230635   foeamData Fv 197  NA ___  JN A      JN A            N A  Va      A    2007 07 18 15 18 CMB test    this is the result of the CMB test    hscallLadParl ayer  ie ERE       5 10 15 20 25 30 i   E  yer number    File  http   ttfinfo desy de CALICEelog data 2007 29 18 07 2007 07 18T15 18 02 00 gif    2007 07 18 15 18 Beni gain calibration         330636 ahcGain stopped  after a  few  configurat ions  for slow  control  work    2007 07 18 16 00    330640 ahcGain                2007 07 18 20 00 Beni summary of todays CMB  slow control work    Sven improved the CMB low voltage display in the slow control  see earlier entries    On Sven s request I bridged the voltage sensing of the CMB power supplies  This improves the  switch on behaviour and should make a fault like we had over the last days much more  unlikely  But this also reduces the voltage at the CMB for approximately 0 7 V  There  is a test gain run  330640  to check if this influences the CMB operation or not    Sven  amp  me checked the mapping of the HBAB slow control readout   Nanda  amp  me took out the CMB from layer 10  which did not show light at all  for further  error diagnosis   I tested the TCalib  amp  VCalib signals at the outlet of the module in layer 10 and these  were ok  Therfore  the error has to be in the CMB itself     2007 07 18 20 38 Beni  Oliver pi  8GeV    330641 lbeamData  10x10  8GeV pi  525 4mm air 2 98 250k  Pb events       10x
337. te  88 Hz    Logbook entry   CALICEelog data 2007 29 22 07       27 09 2007 11 39    CALICE eLogbook http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log       Ep aa Ricardo Nicola Kieffer    Events    DAQ rate  77 Hz       Logbook entry   CALICEelog data 2007 29 22 07      22 07 2007       noisy ECAL slots 17 and 19 monitor  Ep 0000     Oliver Allister Hengne  shot 5     noisy ECAL slots 17 and 19 monitor  shot 5            14 DR       edt Maier  Due al       Logbook entry   CALICEelog data 2007 29 22 07      22 07 2007 e   noisy ECAL slots 17 and 19 monitor  A mm Oliver Allister Hengne  shot 4     noisy ECAL slots 17 and 19 monitor  shot 4           26 of 28 27 09 2007 11 39    CALICE eLogbook http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log       fw NM     wa   n Du ene a ntu       Logbook entry   CALICEelog data 2007 29 22 07      22 07 2007 1   noisy ECAL slots 17 and 19 monitor  Ey 0008 Oliver Allister Hengne  shot 3     noisy ECAL slots 17 and 19 monitor  shot 3              Logbook entry   CALICEelog data 2007 29 22 07      22 07 2007 e i noisy ECAL slots 17 and 19 monitor  Ey 0007 Oliver Allister Hengne  shot 2     noisy ECAL slots 17 and 19 monitor  shot 2           27 of 28 27 09 2007 11 39    CALICE eLogbook http   ttfinfo desy de elog results jsp docroot  export web htdocs amp log       D     u  E  5  DH  4  3  D    Pudor    151 91       Logbook entry   CALICEelog data 2007 29 22 07      22 07 2007     noisy ECAL sl
338. the MIP peak position   As a couclusion of a short study  it is found that fitting with assymetric gaussian within a certain  region gives good fit  small chisquare    believe we can find better function than the asymmetric gaussian  then we can fit wider region  and reduce statistical uncertainty  but for now I adopt this function   will attach some plots for MIP fits   This is for a fiber readout strip        CALICE eLogbook   Results for search of  20070301    Layer 4 Strip 5  MIPEvent2_3_4  Emnes 6255  Y Mean 3245  200 qus 1042    Underfiow o   Overfiow 0   E ind 25 92 134   1 50 Prob 0 8009  21983140   2252 4 1    48 32   244  70 77 1 264    100    50       200 400 600  ADC counts    File  http   ttfinfo desy de CALICEelog data 2007 12 25 03 2007 03 25T23 14 43 00 jpg    200 0    2007 03 25 23 21 S  Uozumi    And this is a fit result for a direct readout strip     Layer 15 Strip 5    MIPEvent2 14 4    400    x  indt 29 07   22  0 1429    Prob   po 4477269  p    p2   p3    300    95 65   2 53  28 58   1 52  52 09   1 57    200    100       200 400 600  ADC counts    File  http   ttfinfo desy de CALICEelog data 2007 12 25 03 2007 03 25T23 21 08 00 jpg    200 0       Page 61 of 76    CALICE eLogbook   Results for search of  20070301 Page 62 of 76    2007 03 25 23 23 S  Uozumi    This is a mapping of the obtained calibration constant for lst configuration   MIP calibration runs after tungsten removal are used for this result     1st configuration after tungsten removal      
339. theless  To be investigated tomorrow     2007 04 27 09 24 Sebastian no currents in Modules 25A and 29B  Schaetzel                Nigh Voltage  all Mc ON   dt DEF    recove  EU Ww   fecto all trips    sonitoritg Switoh On   Off comata  Weltage  V  Current bal Set  V  bestial  V       0 00 1 10     10 00   Rise  LI   Module 25  41 14     19 00 Fee  LA   Module 268 44 44   10 00 Ree  WEE   19 00 Ree  e  este 20 en   10 00 Seel or Module 29 73 0   10 00 Rew  La   Wodule   10 00   fae  or   Fodule D      10 00 Ree  OFF  ff moore siz 0 TEE   Bota to CIE   Lm a                     File  http   ttfinfo desy de CALICEelog data 2007 17 27 04 2007 04 27T09 24 06 00 png    m 2007 04 27 12 05 Beni current problem    After disconnecting and reconnecting the relevant LV HV cables the two module halfes 25A  amp  29B  show a reasonable current again  These connectors do not have screws to fix  If the problems  happens again  we should replace the powerlines      y 2007 04 27 12 08 Beni stage scan restarted    202402 ahcBeamStagefPoar 6 e  3 GeV modules scan  25 27 29 31    I restarted the stage scan from the beginning  as the signal was gone for module half 25A       S 2007 04 29 13 22 Beni disk full again    CALICE eLogbook   Results for search of  20070401 Page 14 of 14       I just checked the status of the test beam DAQ from remote     The disk is full again  It looks like the copy script is off since quite some time  Therefore  the  clean script could not remove some of the older files o
340. tings   J Ampl DAC  Width DAC  J Ampl Vcali  Width  451  352     3 Ampl DAC  Width  J51  J52    Ruth Dae    7  AB  C Ampl cal Wan pac   1  v  j2882V         2216V    Enable LEDs  0123456789105       121728913 1131 27 710      14 120 076  c   TS  28 520      T_upper_CMB   30 333  c  V  calib at  U041  0 001 v  Power 12V   13 421 Y  7 517 V  REF  1 235V f 235 Y PIN_10V_bias  9098 Y  VIDA upper LED  7 4532 v W calib at UOST   0 001 v       File  http   ttfinfo desy de CALICEelog data 2007 29 18 07 2007 07 18T03 11 02 00 jpg    2007 07 18 03 17 caliceon 2caliceheulsliti  mal shift duties   CALICE user     Due to the problems with the CMBs following tasks should be performed in addition to usual shift  duties     Check for every beam run if the LED ampilude changed in respect to the reference plot Nada  made the last day  If you recognize a change check first if the CMB has the right supply  voltage  see last enttry  and then RE INIT and check the CMB  There exist two buttons  to performe this actions in the slowcontrol CMB main panel  which is shown in the picture  below  Any difference from the reference should be documented in the logbooks  ELOG amp paperlog     Performe the CMB power check before and after each LED calibration run  ahcGain   ahcPmLedVcalibScan    amp  ahcCmLedVcalib scan   Both positive and negativ results have to be documented in  the logboobks    CALICE eLogbook   Results for search of  20070701 Page 7 of 22    CME Main 1234506789 10 11 12 13 15 16 17 18 18 
341. tio layer 15  17   large low energy tail    Energy Beam event NumMain Peak ain Sigma Mean Comments  position Peak Width    08 07 2007  Longitudinal profile better than those of 07 07 2007   Only non significant noisy layer exists    RunNum Energy Beam event Num Main Peak Main Sigma Mean Comment s  pene AC ti  n reak emm  PRA  omens  330199 128k 5969 314   Slightly  noisy layer  us E ET  330203 40 230k 4775 305 5 6 4 Ecal  noisy 1  a Pp    E  330210 40 228 6k 4792 300 6 3  Ecal  noisy 1  12  330212 30 75 8k 3686 221 6 0 Ecal  noisy 1    330213 30 85 2k 3684 221 6 0   212   213   gt  one  file  beam  position  changed  330217 20 230 8k 2515 159 3 6 3  Slightly  12  30 239 2k 2510 162 2 t m  ML p OS OS GO LE  2    40 230k 4760 4  Ecal  noisy 1  12  Ecal  noisy 1  12                    09 07 2007 10 07 2007  Most runs for this 2 days are mu runs or pion runs    RunNum Energy Beam event Num Main Peak ain Sigma Mean Comment s  pen E Pron mt positions Deae uam inn pome    330270 181 2k 5887 294   Noisy  MAA es    330778 B5         NETAS   593             187              7            A  330293 B5      EEE 2 75 1557 CEA DEE    11 07 2007  Many hadrons in the beam     Energy Beam event Num Main Peak Main Sigma Mean Comment s  AAA EA fr AAA  caes 168 4  Ecal       2516                      13 07 2007    Energy Beam event NumMain Peak Main Sigma Mean Comments  position Peak Width    330380 E      ED 1917 E34               Tr 93        ETT       CALICE eLogbook   Results for search of  2
342. to 100 Hz for 330039  Doesn t change anything     2007 07 02 19 37 A  Kaplan Too high current in HV channels  red blinking   There is a to high current in the following modules   Position  26A 26B 27A 27B 28A  Module  29A 29B 30A 30B 1A  Current  274 300 520 545 1054        2007 07 02 19 40 Erika for HCAL analizers    the last muon runs taken at position 0 10 should all be fine and can be used to determine mip  calibrations     you will see a difference in the number of pedestals in spill  would be interesting to know the  rate of pedestals in spill for runs 330038 and 330039  Thank you     2007 07 02 19 47 Erika  2007 07 02 22 27 Alexander beamHoldScan ahcPm Cm LedHoldScan runs    330043 beamHoldScan u   30043 S  30045  30048    30017  30048  30049    30050       330043  spill sync enabled  Split beamRun disabled  stopped by pressing  START ECAL SAFETY RUN  Button       330044  spill sync disabled  Split beamRun disabled    pressed  START ECAL SAFETY RUN  Button before run 330045  330045  spill sync disabled  Split beamRun disabled    330046  spill sync enabled  Split beamRun enabled       CMB 30  Modul 24  and CMB 33  Modul 36  do not answer in mainCMB panel   run 330046 stopped by Beni for ahcGain  330047  ahcGain run  Beni     330048  spill sync enabled  Split beamRun enabled  H6B EXPTO1  02  Rates   10 10 6700 and 20 20 7000    Everything looks fine on the monitoring  leaving to run overnight     2007 07 02 22 27 Alexander beamHoldScan ahcPm Cm LedHoldScan runs  30043 be
343. to see lpe peak   400386     CmLed to see lpe peak with over voltage  1V  3 5V     for Layer8 Strip15     400393     CmLed for inter calibration   400396     PmLed for inter calibration   400394     CmHoldScan   400395     PmHoldScan   400401     CmLed to see lpe peak  100k ev    400402     CmLed to see lpe peak  100k ev  but with slightly higher light to see adjacent strips  400419     CmLed to see lpe peak  100k ev    400420     CmLed to see lpe peak  with over voltage  1V  3 5V     for Layerl4 Strip13                       400411   CmLED for inter calibration   400412 PmLed for inter calibration   400413 CmLedHoldScan for confirmation   400414 PmLedHoldScan for confirmation   400416 CmLed to see lpe peak  100k ev    400417 CmLed to see lpe peak  100k ev  with over voltage  1V 3 5V    400421 CmLed to see lpe peak  100k ev  but slightly higher light to see adjacent strips    end of LED rung EEN    between 1st and 2nd round  the drift chamber is tuned to improve efficiency     CALICE eLogbook   Results for search of  20070301 Page 72 of 76       lst configuration  fiber direct  2st round                             Each run contains data with 55 points   36 points for center of each strip of X and Y layers   9 points for central part scan   1 points    x 10 times for central injection      1 points have 10k events of beam data and 1k events of pedestal   1 GeV     400422  400423  2 GeV     400425  400426  3 GeV     400427  400428  4 GeV     400429  400438  5 GeV     400432  4
344. trongly  supported and gave great helps to the SCECAL  beam test     2007 03 30 23 49 S  Uozumi run summary again  Since I could not modify the old entries in e log from Japan  I put another updated run summary  here     Now all the LED runs are added     lst configuration  fiber direct  lst round                           MIP calibration before putting tungsten  400262    detector scan  36 points     1GeV     400309  point0 35    2GeV 400308  point0 35    3GeV     400305  point0 35    4GeV     400302  point0 35    5Gev 400296  point0 35   400310  point0 35   400311  point0 35   400314  point0 35   6GeV 400285 400291  point0 35     center part scan  9 points             1GeV     400309  point36 44   2GeV 400308  point36 44   3GeV     400305  point36 44   4GeV     400302 point36 44   400313  point36 44   5GeV 400296  point36 44   400310 point36 44   400311  point36 44   400314  point36 44   6GeV 400264 Run400285 400291  point36 44   Center injection  1 points x many times   1GeV     400317 90k ev   2GeV 400318 90k ev   3GeV     400319 90k ev   4GeV     400320 90k ev   5GeV 400314  point45 54  400321 90k ev   6GeV 400322 90k ev     end  of  dst configuration dst round  2352233332335 233 7372233203920 SR SS SSS    LED run with 1st configuration  only 3 layers                                            Pedestal runs    400382     CmNoise   400398     PmNoise   for Layer7 Strip15    400406     CmLed for inter calibration   400407     PmLed for inter calibration   400385     CmLed 
345. ts for search of  20070701 Page 6 of 19          meng Cem  m jj  HCAL Motor control panel Y    actual position target position    200  x   20 0 p oxi   gt   Y  14 95 35 ei      Q lost X reference point       O ones pix    Q lost 2 reference point  Q stage in motion  O maras  operation get Pos       feas V Fast Y pulsesfeee   extern STOP   refs24   849 506 049  906       OFF set    B     200 E 29  PING  to elos          File  http   ttfinfo desy de CALICEelog data 2007 28 15 07 2007 07 15T18 26 02 00 jpg    2007 07 15 18 45 George  Mangi  finish calibration  Nanda   230489                                        k ehePmnedvealibseaa LI     just to make the set complete     2007 07 15 18 46 George  Manqi  electron run  Nanda       Pb    10x10 trigger rate  3 4k pps  DAQ rate  60 Hz    beam position not centered in the drift chambers    gt  changed magnet current of TRIMO6 to 1A less  now at 17A     2007 07 15 18 58 George  Mangi  electron run  Nanda    10x10 trigger rate  3 4k pps  DAQ rate  60 Hz    BANI NN       beam even further off      change trim06 to 20 0 A  and try again     much better but not perfect      change trim06 to 22 0 A    2007 07 15 19 46 George  Manqi  electron run  Nanda    p ee PP gue ques quem  Pb  pedem ec m d gue jp  Ph  Pb    Pb  0x10 34                      Pb  Pb       CALICE eLogbook   Results for search of  20070701 Page 7 of 19       10x10 trigger rate  2 0 k pps   DAQ rate  60 Hz   again beam off in the DC   330495  gt  change trim06 from 15 0A t
346. ttfinfo desy de CALICEelog data 2007 34 22 08 2007 08 22T08 42 03 00 jpg    2007 08 22 09 00 Beni gain tricks  350398 junk       350399 only modules 1 2 14 25  manually controlling  module 2  14  25 see next table       350400 ahcCmLedHoldScan with manual amplitude setting    CALICE eLogbook   Results for search of  20070801 Page 22 of 22       new holds extracted for module 14  amp  25    350401  350402    350403  350404       2007 08 22 15 00 Beni problems over problem    SlowControl problems  amp  DAQ problems after switching off and remounting module 1 CMB    2007 08 22 17 00 Beni running again    After several powerdown  powerup  slowcontrol twiggeling  restaring DAQ  going directly on calice02  I can steer the DAQ  amp  detector again    2007 08 22 18 00 Beni module 1 conclusions    After checking all signals of module one   The module delivers a wrong TCalib signal  Instead of a differential signal I see one line grounded   The CMB does not accept this signal as trigger     There is not enough time left for further debugging and fixing  We have to live without a    calibration  for this module for now           2007 08 23 11 21 Roman Poeschl Transfers finished    As far as I see all primary files  i e  binaries and related log and slw files  have been shipped out from the control room disk array     Therefore I need no further access to the disk array    In principle the computers can be switched off now but as said earlier   we may leave them simply running until we f
347. un  Ariane    junk       x tracker off  run junked    2007 07 29 14 37 Beni  Martin  junk run  Ariane       331104 lbeamData  10x10  10 GeV   pi  520 8mm Air 2 0 20deg  Pb junk    y tracker off    2007 07 29 14 45 Beni  Martin  pi  10 GeV  Ariane       Pb    1200 pps  DAQ rate  60 Hz  150k events    2007 07 29 14 55 caliceon 2calicesdzC  amp drmolise    We welcome back the ECAL square event    Comment  Marcel     This is not what is referred to as square event  It shows only a pedestal shift in one layer    For now an event has been called  square event  when  nearly  all the border cells of one wafer  show a hit     CALICE eLogbook   Results for search of  20070701 Page 12 of 16       Time  12 49 23 471 761 Sun Jul 29 2007  Hits  274 Energy  350 079 mips    Run 331105 0 Event 14970       File  http   ttfinfo desy de CALICEelog data 2007 30 29 07 2007 07 29T14 55 02 00 png    2007 07 29 15 12 caliceon calicehediark   cai threshold   CALICE user     change curernt threshold for 2 channels corresponding to modules 18B and 29A         92 ITA   100   w SMITA   311 W sesa  an Ww 18 98   150 ANA 143 RSE     120  Raps    34 w Siia 113 W IRSA  i01   3108   wo we     i43 2288  Fima   so SA   199 MORA    105 w 961198   159 w apra   1 XA ma e seoa   149 W poa      x10 Ws   i116 w 40808 143 W ess    112 34  sn een   17 SES    an  Em   370 maae   3147 m fee  145 FA 39 m AzA   a 9904  125 Kaap   T   w 4428    197 XI      103 patsa   e RES   103 wN Ipa f      OSB ow 45208  se m apie   2    
348. uning  Pb beam file  for 40  GeV pi        DAQ average rate  73 Hz  120k evts    CALICE eLogbook   Results for search of  20070701 Page 13 of 24       2007 07 07 11 29 Cristina Michal Evgeny    330174 lbeamData  10x10  40Gev pi  516 4mm air off tuning  Pb beam file  for 40  GeV pi        DAQ average rate  73 Hz  120k evts    2007 07 07 11 36 Evgeny Cristina Mh  emEsttalb abi 00 GeV pions  Loaded file H6B 516    Collimators setting not loaded from beam file  only magnets    Settings for collimators modified from previous file        Absorber  4mm Pb  Target  air    10x10  10800 pps  20x20  11500 pps  100x100  1450 pps       2007 07 07 11 51 Evgeny Cristina Mix amp arh tuning   100 GeV pions       IILLPLLE E  Pb  Stopped after 80K events  2007 07 07 12 41 Fab Erika beam tuning   50 GeV electrons    extracted 50 GeV e  beam file from 80 GeV pi   File name H6B 530  extracted from H6B 510     Opened position collimators  Cl  C10  C5   20  20    h position  C11  C6   20  20    v position  Opened momentum collimators  C2   4  4    h momentum  C3 C8   4  4    v momentum    Changed quadrupoles   Q6   gt  130   Q7       70   Q8   gt    45   Q11   gt   140   012   gt  50 6    Beam satisfactory  Started run 330176    2007 07 07 13 07 Evgeny Cristina Michal    8 NIE MAA  PH       Stopped after 90K events since shower partially leaking from the ECAL side    2007 07 07 13 20 Evgeny Cristina Michal       330177 lbeamData  10x10  50GeV e  530 6mm air off  Pb  junk RUN    2007 07 07 13 24 Evge
349. uns 330725   330775 also  started     Control of scripts handed back to shift crew   Cheers     Roman    2007 07 21 19 24 Oliver Allister Henpien run    330778 lbeamData  10x10   8GeV pi  525 4mm air 3 0  3   3  cherenko  E pressure  changing   h 0000 e table    stopped wit v when cherenkov sta  ratio pi e   7 2  insert 8mm Pb absorber to try and improve it       2007 07 21 19 28 Oliver Allister Henpi  n run    Pb       10x10 trigger rate  1 0kpps  DAQ rate  55Hz    pi e   20 1 3         2007 07 21 19 57 Roman Run 330394  Hi there    I am bit puzzled    It seems that we have never taken a run 330394     It doesn t appear neither in the logout of the cogri scripts  nor in the checks performed before deleting files     In the elog I have found that there were DAQ Problems for runs 330393 and 330394      but somehow the latter was never taken     Maybe it s just for the record     Roman  2007 07 21 20 48 Oliver Allister Hengalibration runs  2007 07 21 20 58 Oliver Allister Hengnergy resolutions pion runs    21 07 2007 Pion Runs    Energy  GeV  lenergy mean  mip  width of mean  mip   Sigma Mean    349 1    E19     eE  EST       2007 07 21 20 59 Oliver Allister Henghange of beam setting    Moved absorber to air  in preparation for electron runs     Reduced collimators to decrease 10x10 trigger rate  New collimator settings   COLL1   COLL5   COLL6   COLL10   COLL11      5  old value was    15     2007 07 21 21 14 Oliver Allister Henglectron run    4 96kpps    10x10 trigger rate   D
350. ure how long we can have the beam  the DESY II may back to the positron beam mode  very soon    DAQ rate  130 Hz with 4 GeV beam     2007 03 24 13 43 S  Uozumi 5 GeV run  400774 beam T1L 5GeV KNU fiber  2 5v scanned 36  9 10  AAA AAA adig CN NN  2007 03 24 13 43 S  Uozumi 5 GeV run  um Hu c r itl  9 10  2007 03 24 17 23 S  Uozumi 5 GeV run    KNU fiber  2 5V scanned 36  9 10  position         2007 03 24 17 23 5 GeV run    T    3 GeV run    KNU fiber  2 5V scanned  points    S  Uozumi    S  Uozumi    ue       2007 03 24 18 43       CALICE eLogbook   Results for search of  20070301 Page 55 of 76    14 to see  position  dependence    2007 03 24 18 43 S  Uozumi 3 GeV run       400777 beam T1L 3GeV KNU fiber  2 5V scanned 5x20   points  precise  scan for  IX layer  strips 5  14 to see  position  dependence    2007 03 25 10 10 S  Uozumi overnight runs    Runs 400777 400781 are 5x20 precise scan overnight   Run 400782 was stopped around 50th point        2007 03 25 10 49 S Chang etal  3rd config  removed tunsten  We have removed tungsten plates from detector   2007 03 25 10 51 S Chang etal   3rd config   Responce of KNU modules   Calibration  Run   3 7V   Changing overvoltage  3 7V  gave improvement    55 4   for  2 9V  it was 41 4      But  KNU Modules response is still lower than Kuraray fiber coupling modules     55 4     153 2   276 4   100    ek informations of data KREKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKEKEK    Calibration Run   400543  2nd Calibation Run for KNU module   Run
351. usly until now  331086 87 88 89        2007 07 29 08 20 caliceonOcaliceseroisyECIAL layer    noisy layer 9 in run 331090                                            hEcalHitsPerLayer    EE LEE EEN BEEN Seel Mo  a d I i A F H     0 5 10 15 20   ayer Amber  e hEcalEnergyPerLayer  a H H  E eeoa E     gt    BP 44 Lip  o  5 12 AAA  10  A o O NN ANN   MB o Pe ei Di AO LIB   4  ayer number  File  http   ttfinfo desy de CALICEelog data 2007 30 29 07 2007 07 29T08 20 02 00 gif  2007 07 29 08 40 Beni  Martin  calibration runs  Ariane  331091 ahcGain  2007 07 29 08 43 Beni  Martin  pi  10 GeV  Ariane    331092 lbeamData  10x10  10Gev  oi  520 8mm Air 20  Pb  beam off    1400 pps  DAQ rate  60 Hz  150k events       2007 07 29 09 50 Beni  Martin  junk run  Ariane    331093 lbeamData  10x10   8GeV pi  525 8mm Air 2 3 20deg   Pb junk       X tracker slightly off  adjusting current  run stopped    2007 07 29 09 59 Beni  Martin  pi  8 GeV  Ariane    Pb       low rate  changing collimeter position according to general rule  1200 pps    DAQ rate  60 Hz  150k events      2007 07 29 10 11 Beni  Martin  GUI crashed    CALICE eLogbook   Results for search of  20070701 Page 3 of 20       Ariane  The online monitor GUI crashed  You can still press the buttons and zoom in the windows  But the  displayed data is not updated anymore  The error output is attached     This has already happened a few times  namely while rapidly starting and stopping beams           Break     segmentation violation          
352. ut with  1V increased base voltage   80V   400728     CmLed to see lpe peak around strip 15  pedestal too wide  difficult      For Layerl5 stripl5     400729     CmLed to see lpe peak  400730     Same with 400729 but with  1V increased base voltage   80V   400731     CmLed to see lpe peak around strip 15    For Layerl  6 stripl5     400732     CmLed to see lpe peak  400733     Same with 400732 but with  1V increased base voltage   80V   400734     CmLed to see lpe peak around strip 15    For Layerl8 stripl5     400735     CmLed to see lpe peak  400736     Same with 400736 but with  1V increased base voltage   80V   400737     CmLed to see lpe peak around strip 15    For Layerl7 stripl5     400738     CmLed to see lpe peak   400739     Same with 400738 but with  1V increased base voltage   80V    For Layerl9 stripl5    400740     CmLed to see lpe peak   400741     Same with 400740 but with  1V increased base voltage   80V    400742     CmLed to see lpe peak around strip 15  pedestal too wide  difficult      For Layer20 stripl5     400743     CmLed to see lpe peak  400744     Same with 400743 but with  1V increased base voltage   80V   400745     CmLed to see lpe peak around strip 15    For Layer22 stripl5        400747     CmLed to see lpe peak  400748     Same with 400747 but with  1V increased base voltage   80V   400749     CmLed to see lpe peak around strip 15    For Layer21 stripl5     400750     CmLed to see lpe peak  400751     Same with 400750 but with  1V increased ba
353. vice the srm dcache desy de will not be available  tomorrow   July 12th from 11 00 until about 12 30     No queues of the CEs at DESY HH will be drained     Regards  Christoph  for the DESY HH admin     2007 07 10 17 47  gt   Fwd   calice vo support  Short Service Interruption  srm dcache desy de     Roman Poeschl wrote   A few words more    The problem with the srm dcache doess not mean that users cannot  continue to use it     It might only happen that access attemps will fail and one has to retry   Why this is to some extend acceptable   for users we cannot use it for the current primary management of the  data  i e  permant copying and processing     CALICE eLogbook   Results for search of  20070701 Page 12 of 25       So the lines below affect only the actions performed with the caliceserv  in the control room  but it remains true that  the currently data are stored in lyon  binaries and converted files         Cheers   Roman  2007 07 10 18 44 Mary Cruz Hengne Niels  P  I  m p  D M E       Pb evts   2007 07 10 19 07 high current    during beam tuning some currents had jumps or increased slowly  Maybe effect of the higher number  of particles per spill       sonitoriog Switch On   Off  Voltage  V  Current  un  Set EV  Initial  V     10 00  10 00  10 00  10 00  10 00  10 00  10 00  10 00  10 00  10 00  10 00 Doo a Y  10 00  HY settings   File actions enable   disable AL chamnels  SWE Lom   er   enable   disable      b  ebelebe e    10 00       10 00    I3 18  ae  zry    E 
354. x No beam  31016    Beam stopped before we could start another run           2007 07 27 19 38 Lukas  Alex Junk run  331017 ahcGain  2007 07 27 19 40 Lukas  Alex 10x10 e        Stopped after 30k ev because of collimators adjustment     2007 07 27 20 22 Lukas  Alex 10x10 e         C1 C5 C6 C10 and C11 narrowed to  4 mm  This run doesn t exhibit excentricity problems observed  in previous run      Pressure in Cerenkov counter was stuck at 0 79 and didn t want to go to prescribed value  Finally  it reached 0 5 after  29 minutes of run     200k ev        The beam went off and on for many short periods of time during this run   http   elogbook cern ch eLogbook eLogbook jsp shiftId 1007833    CALICE eLogbook   Results for search of  20070701 Page 5 of 16          2007 07 27 21 14 Lukas  Alex 10x10 pi   Pb    Because of different size of slits defined in file 3  collimators C1 C5 and C6 were adjusted  to  7 mm to be the same width as C10 and Cll    190k ev   2007 07 27 21 53 Lukas  Alex 10x10 pi        Pb    Aborted after 26k ev to adjust beam in Y axis     2007 07 27 22 04 Lukas  Alex 10x10 pi     Bb    Aborted after 14k ev because of further adjustment of beam in Y  now from 3 A to 1 A     2007 07 27 22 09 Lukas  Alex 10x10 pi   331024 lbeamData  10x10  15GeV pi  521 4mm Air 10deg  Bose eee ipee Ier reus ape opp n pies pe  2007 07 27 23 41 Lukas  Alex 10x10 pi     Pb       2007 07 28 00 24 Nanda  Roman pi  run    Pb       Miserable rate  5pps        opened collimators 1 5 6 8 10811
355. zumi 1 GeV run    400650  beam TIL 1Gev KNU fiber  2 5V scanned 36  9 10  scan    2007 03 22 12 06 S  Uozumi 1 GeV run    CALICE eLogbook   Results for search of  20070301 Page 38 of 76    400650  beam TIG 1Gev KNU fiber  2 5V scanned 36  9 10  scan       2007 03 22 13 45 S  Uozumi 1 GeV run  400651  beam TIL 1Gev KNU fiber  2 5V scanned 36  9 10  fe a p c pipe vg  2007 03 22 13 45 S  Uozumi 1 GeV run  400651 beam T1L 1Gev KNU fiber  2 5V scanned 36  9 10  pq gp wet e x   2007 03 22 13 46 S  Uozumi 2 GeV run  400653  beam TIL 2GeV KNU fiber  2 5V scanned 36  9 10  pap ETE TA e sete  3   2007 03 22 13 46 S  Uozumi 2 GeV run  400653 beam T1L 2GeV KNU fiber  2 5V scanned 36  9 10  pre E AAA TEA   2007 03 22 13 46 S  Uozumi Just to remenber in Japan    Here is a table of magnet current to set the beam energy    attached to the magnet control PC     Energy  GeV  Magnet current  A        1 00 37 40  1 20 44 88  1 40 52 36  1 60 59 84  1 80 67 32  2 00 74 80  2 20 82 28  2 40 89 76  2 60 97 24  2 80 04 72  3 00 12 20  3 20 19 68  3 40 27 16  3 60 34 64  3 80 42 12  4 00 49 60  4 20 57 08  4 40 64 56  4 60 72 04  4 80 79 52  5 00 87 00  5 20 94 48  5 40 201 96  5 60 209 44  5 80 216 92  6 00 224 40    But since magnet control system can accept only integer current values   we use following currents for our runs     Energy  GeV  Magnet current  A   3  75  112  150  187  224    CDs GAbAt3    2007 03 22 15 18 S  Uozumi 2 GeV run    400654 beam T1L 2GeV KNU fiber  2 5V scanned 36  9 10  
    
Download Pdf Manuals
 
 
    
Related Search
    
Related Contents
Super Multi DVD Writer  ENVI Services Engine User Guide - Exelis Visual Information Solutions  Powerocks Deck  Télécharger  BASE ELEVES 1er degré Manuel utilisateur de la Mairie  SLW1 Weather Station Installation and Operating Instructions  Tecumseh AJA7441AXAXW Technical Data Sheet  Manual ST-412  EUROLITE LED Strobe with Cable/Plug User Manual  Jules Verne    Copyright © All rights reserved. 
   Failed to retrieve file